Die Pakete

0.1 Package descriptions

0.1.1 A: Linux Base System (You need it!)

aaa base S.u.S.E. Linux base package

This package installs several important configuration files. Central scripts like SuSEconfig are also in this package.

Contents: aaa base fillup skel

Copyright:  (c) 1996,97 S.u.S.E. GmbH Fuerth, Germany

Author:

Burchard Steinbild <bs@suse.de>, Florian La Roche <florian@suse.de>, Werner Fink <werner@suse.de>, Martin Scherbaum <maddin@suse.de>

Version:  98.8.12-0

aaa dir S.u.S.E. Linux directory structure

This package installs the directory structure. It also includes the home directories of the system users.

Contents: aaa dir

Copyright:  (c) 1997 S.u.S.E. GmbH Fuerth, Germany
Version:  98.8.13-0

at Job Manager

With this program you can run jobs at a specified time.

Contents: at-3.1.8

Copyright:  (c) 1993 David Parsons, 1993, 1994 Thomas Koenig
Author:  Thomas Koenig <ig25@rz.uni-karlsruhe.de>, David Parsons <orc@pell.chi.il.us>
Version:  3.1.8-11

base Some important GNU packages

This package normally also includes the program ”rmt”, which provides remote tape drive control. Since there is a compatible ’rmt’ in the ’dump’ package we didn’t put ’rmt’ into this package. If you are planning to use the remote tape features provided by tar you have to install the ’dump’ package as well.

Contents: freeramdisk grep-2.0 hdparm-3.3 ktools m4-1.4 sed-2.05 tar-1.11.8 time-1.7 which

Author:

Francois Pinard <pinard@iro.umontreal.ca>, Tom Lord <lord@cygnus.com>, David J. MacKenzie <djm@gnu.ai.mit.edu>, Akiko Matsushita <matusita@sra.co.jp>, Alan Magnuson <awm@osc.edu>, Alexander Lehmann <alex@hal.rhein-main.de>, Amos Shapira <amoss@cs.huji.ac.il>, Andreas Gustafsson <gson@niksula.hut.fi>, Andreas Schwab <schwab@issan.informatik.uni-dortmund.de>, Assar Westerlund <assar@nada.kth.se>, Ben A. Mesander <ben@piglet.cr.usgs.gov>, Bengt Mertensson <bengt@mathematik.uni-bremen.de>, Bernhard Daeubler <Bernhard.Daeubler@physik.uni-ulm.de>, Bjorn R. Bjornsson <brb@falcon.is>, Brendan Kehoe <brendan@cygnus.com>, Erez Zadok <ezk@cs.columbia.edu>, Eric Allman <sendmail@sendmail.ORG>, Eric Backus <ericb@lsid.hp.com>, Geoff Russell <grussell@guest.adelaide.edu.au>, Greg A. Woods <woods@web.apc.org>, Hal Peterson <hrp@pecan.cray.com>, Hoang Uong <hoang@ornews.intel.com>, Ian Lance Taylor <ian@cygnus.com>, Jason Merrill <jason@cygnus.com>, Jim Avera <jima@netcom.com>, Jim Kingdon <kingdon@cygnus.com>, Jim Meyering <meyering@na-net.ornl.gov>, Joel Sherrill <jsherril@uahcs2.cs.uah.edu>, John David Anglin <dave@hiauly1.hia.nrc.ca>, Joseph E. Sacco <jsacco@ssl.com>, Karl Berry <karl@cs.umb.edu>, Karl Vogel <vogelke@c-17igp.wpafb.af.mil>, Kaveh R. Ghazi <ghazi@noc.rutgers.edu>, Kristine Lund <lund@lpnaxp.in2p3.fr>, Krste Asanovic <krste@icsi.berkeley.edu>, Marion Hakanson <hakanson@cse.ogi.edu>, Mark Seiden <mis@seiden.com>, Matthias Rabe <rabe@mathematik.unibielefeld.de>, Michael Fetterman <mafetter@ichips.intel.com>, Michael L. Welcome <welcome@bigbird.llnl.gov>, Mike Lijewski <lijewski@theory.tc.cornell.edu>, Nick S. Kanakakor <skanan@otl.scu.edu>, Nicolas Pioch <pioch@inf.enst.fr>, Noah Friedman <friedman@gnu.ai.mit.edu>, Pete Chown <pete.chown@dale.dircon.co.uk>, Pierre Gaumond <gaumondp@ere.umontreal.ca>, Pierre Mathieu <mathieu@geod.emr.ca>, Robert Bernstein <rocky@panix.com>, Rene’ Seindal <seindal@kb.dk>, Richard Stallman <rms@gnu.ai.mit.edu>, Roland H. Pesch <roland@wrs.com>, Roland McGrath <roland@gnu.ai.mit.edu>, Scott Bartram , Simon Leinen <simon@lia.di.epfl.ch>, Skip Montanaro <skip@automatrix.com>, Stephen Perkins <perkins@cps.msu.edu>, Steve Williamson <willy@uinpla.npl.uiuc.edu>, Thorsten Ohl <ohl@physics.harvard.edu>, Tom Mc-Connell <tmcconne@sedona.intel.com>, Tom Quinn <trq@dionysos.thphys.ox.ac.uk>, Tom Tromey <tromey@cns.caltech.edu>, Ulrich Drepper <drepper@gnu.ai.mit.edu>, Vern Paxson <vern@ee.lbl.gov>, Vivek P. Singhal <singhal@cs.utexas.edu>, Walter Wong <wcw+@cmu.edu>

Version:  98.2.10-15

bash The GNU Bourne-Again Shell

Bash is a sh-compatible command interpreter that executes commands read from standard input or from a file. Bash also incorporates useful features from the Korn and C shells (ksh and csh). Bash is ultimately intended to be a conformant implementation of the IEEE Posix Shell and Tools specification (IEEE Working Group 1003.2).

Contents: bash-2.01.1 readline-2.1

Copyright:  (c) 1989 Free Software Foundation, Inc.
Author:  Chet Ramey <chet@po.cwru.edu>
Version:  2.01.1-15

bash1 The GNU Bourne-Again Shell (older version)

Since the new bash, version 2.x, is somewhat stricter in its syntax checking than version 1.x, we are keeping the older version as bash1 in the distribution.

If you have scripts that don’t satisfy the stricter rules of 2.x and you don’t want to look for the errors, simply replace ”/bin/sh” with ”/bin/bash1” in the first line.

The script will then be run by the old bash shell.

Contents: bash-1.14.7

Copyright:  (c) 1989 Free Software Foundation, Inc.
Author:  Chet Ramey <chet@po.cwru.edu>
Version:  1.14.7-14

bdflush Linux update daemon

Bdflush is used to start the kernel daemon to flush dirty buffers back to disk.

A: Linux Base System (You need it!)

Contents: bdflush-1.5

Author:

Eric Youngdale <eric@aib.com>, Phil Bostley <bostley@cs.colorado.edu>, Daniel Quinlan <quinlan@yggdrasil.com>

Version:  1.5-19

compress Standard Un*x compression program

Author:  Peter Jannesen <peter@ncs.nl>
Version:  4.2.4-13

cpio The GNU cpio backup and archiving utility

This is GNU cpio, a program to manage archives of files. This package also includes mt, a tape drive control program. Cpio copies files into or out of a cpio or tar archive. An archive is a file that contains other files plus information about them, such as their pathname, owner, timestamps, and access permissions. The archive can be another file on the disk, a magnetic tape, or a pipe. This package normally also includes the program ”rmt”, which provides remote tape drive control. Since there is a compatible ’rmt’ in the ’dump’ package we didn’t put ’rmt’ into this package. If you are planning to use the remote tape features provided by cpio you have to install the ’dump’ package as well. Contents: cpio-2.4.2

Copyright:  (c) 1989, 1990, 1991, 1992, 1993 Free Software Foundation, Inc.

Author:

David J. MacKenzie <djm@gnu.ai.mit.edu>, Jim Meyering <meyering@nanet.ornl.gov>

Version:  2.4.2-16

cron cron daemon

Automatically starts programs at specific times. Add new entries with ”crontab -e”. (See ”man 5 crontab” and ”man 1 crontab” for documentation.) Contents: cron-3.0.1

Author:  Paul Vixie <paul@vix.com>
Version:  3.0.1-15

db BSD database

Static and shared libraries for the BSD database routines. Contents: db.1.85

Version:  1.85-11

devs Device files

This package creates special files in the /dev directory, so that your system’s hardware can be accessed.

Author:  Hubert Mantel <mantel@suse.de>
Version:  98.7.26-0

diff GNU diffutils

The GNU diff utilities find differences between files (to make source code patches, for instance). Contents: diffutils-2.7

Copyright:  (c) 1985-94 Free Software Foundation, Inc.

Author:

Paul Eggert <eggert@twinsun.com>, David J. MacKenzie <djm@gnu.ai.mit.edu>, Richard Stallman <rms@gnu.ai.mit.edu>, Roland McGrath <roland@gnu.ai.mit.edu>

Version:  2.7-13

eazy Compatibility links

This package contains some links and enhancements of the /etc files which should make it easier to install software packages from the Internet. A typical example of where you can use this package is the link from /usr/sbin/sendmail to /usr/lib/sendmail.

Version:  98.2.9-5

elvis A vi-clone called elvis

A vi clone, the standard text editor on Unix systems. Contents: elvis-2.0

Copyright:  (c) Steve Kirkendall
Author:  Steve Kirkendall <kirkenda@cs.pdx.edu>
Version:  2.0-14

ext2fs Utilities for the second extended file system

Utilities needed to create and maintain ext2 filesystems under Linux. Included in this package are: chattr, lsattr, mke2fs, mklost+found, tune2fs, e2fsck, and badblocks.

Copyright:  (c) Remy Card, Theodore Ts’o
Author:  Remy Card <card@masi.ibp.fr>, Theodore Ts’o <tytso@mit.edu>
Version:  1.10-21

file determine file type

With ”file” you can obtain information on the file type of a specified file. File type recognition is controlled by the file /etc/magic, which contains the classification criteria. This command is used by apsfilter to permit automatic printing of different file types. Contents: file-3.22

Author:  Mark Moraes <moraes@deshaw.com>, Christos Zoulas <christos@deshaw.com>
Version:  3.22-20

fileutil Basic programs such as cp, mkdir,ln, ls

GNU fileutils: the basic file manipulation programs Contents: fileutils-3.16

Copyright:  (c) Patrick Volkerding
Author:  Jim Meyering <meyering@na-net.ornl.gov>
Version:  3.16-14

find GNU find

This package contains the GNU find, xargs, and locate programs. The find and xargs programs comply with POSIX 1003.2. They also support some additional options, some borrowed from Unix and some unique to GNU. Contents: findutils-4.1

Copyright:  (c) 1990,91,92,93,94 Free Software Foundation, Inc.

Author:

David J. MacKenzie <djm@gnu.ai.mit.edu>, Jan Brittenson <bson@gnu.ai.mit.edu>, Brian L. Matthews <blm@6sceng.UUCP>, Mike Rendell <michael@cs.mun.ca>, James A. Woods <jwoods@adobe.com>, Jay Plett <jay@silence.princeton.nj.us>, Tim Wood <tim@toad.com>

Version:  4.1-15

gawk GNU gawk

GNU gawk. It is upwardly compatible with the System V Release 4 awk. It is almost completely compliant with POSIX 1003.2. Contents: gawk-3.0.3

Copyright:  (c) 1989, 1991 - 1995 Free Software Foundation

Author:

David Trueman <david@cs.dal.ca>, Arnold Robbins <arnold@skeeve.atl.ga.us>, Michal Jaegermann <michal@gorte.phys.ualberta.ca>, Scott Deifik <scottd@amgen.com>, Darrel Hankerson <hankedr@mail.auburn.edu>, Kai Uwe Rommel <rommel@ars.de>, Pat Rankin <rankin@eql.caltech.edu>

Version:  3.0.3-21

gdbm GNU database routines

Static and dynamic library for the GNU database routines. Contents: gdbm-1.7.3

Version:  1.7.3-13

gppshare C++ shared libraries

Needed for dynamically linked C++ programs. The sources are available in the libgpp s package.

Author:  H.J. Lu <hjl@nynexst.com>, Per Bothner <bothner@cygnus.com>
Version:  2.7.2.1-1

groff GNU troff document formatting system.

The groff package provides compatible versions of troff, nroff, eqn, tbl, and other Unix text-formatting utilities. Groff is used to ”compile” man-pages stored in groff/nroff format for different output devices; e.g., for displaying to a screen or in PostScript(tm) format for printing on a PostScript(tm) printer. Most programs store their man-pages in either /usr/man/ or /usr/X11R6/man/.

Contents: groff-1.10

Copyright:  (c) 1989-94 Free Software Foundation, Inc.
Author:  James Clark <jjc@clark.com>
Version:  1.10-17

gzip GNU zip compression utilities

Gzip reduces the size of the named files using Lempel-Ziv coding LZ77. Whenever possible, each file is replaced by one with the extension .gz, while keeping the same ownership modes, access and modification times. Contents: gzip-1.2.4

Copyright:  (c) 1992-1993 Jean-loup Gailly
Author:  Jean-loup Gailly <gzip@prep.ai.mit.edu>
Version:  1.2.4-14

kbd Keyboard and font utilities

Load and save keyboard mappings. Needed if you are not using the US keyboard map. This package also contains utilities for changing your console fonts. If you install this package, YaST will include an extra menu to allow you to choose between the different fonts. This package additionally includes fonts from the kbd fonts.tar.gz package on Sunsite. Contents: kbd-0.93 kbd fonts

Copyright:  (c) 1994 Andries Brouwer
Author:  Andries Brouwer <aeb@cwi.nl>
Version:  0.93-18

kernmod Loadable kernel modules

Most Linux drivers are available as modules, they can be loaded at run-time. So, if you have installed a modular kernel the installation of this package is highly suggested. Please refer to the chapter about the kernel daemon in the accompanying handbook. Contents: VideoteXt-0.6 ibcs-2.0.970513 jazz-2.1d joystick-0.8.0 xosview-1.5.1

Copyright:  (c) Linus Torvalds et al
Author:  Linus Torvalds <torvalds@cs.helsinki.fi>
Version:  2.0.35-3

ldso ld.so, the dynamic linker/loader.

ld.so completes the final process of linking all necessary references to sharable objects and unreferenced symbols in an impure executable, (usually a dynamically linked executable), to produce a runnable file. Nearly all Linux binaries are ”incomplete” in the sense that they require further linking at run time. ld.so’s job is to thus complete the linking process started at compile time. Contents: ld.so-1.9.6

Copyright:  (c) 1995 David Engel, Rickard E. Faith,Yggdrasil Computing Inc.
Version:  1.9.9-6

less less - opposite of more

Less is a pager program similar to more (1), but allows backward as well as forward movement within a file. Also, less does not have to read the entire input file before starting. It is also possible to start an editor at any time from this program. Contents: less-332

Author:  Mark Nudelman <markn@3do.com>
Version:  332-16

libaout Libraries for the old a.out format.

Linux has changed its default binary format from a.out to elf. To be able to run older binary programs, you will need to install these old a.out shared libraries. Contents: libc-4.7.5.bin libc-4.7.5 libc-4.7.6

Author:  H.J. Lu <hjl@nynexst.com>
Version:  4.7.6-2

libz Data Compression Library

zlib 1.0.2 is a general purpose data compression library. All the code is reentrant (thread safe). The data format used by the zlib library is described by RFCs (Request for Comments) 1950 to 1952 in the files ftp://ds.internic.net/rfc/rfc1950.txt (zlib format), rfc1951.txt (deflate format) and rfc1952.txt (gzip format). These documents are also available in other formats from ftp://ftp.uu.net/graphics/png/documents/zlib/zdoc-index.html The sources are contained in the libgr source package Contents: zlib-1.1.1

Copyright:  (c) Rob Hooft, Michael Mauldin, (see netpbm)
Author:  Rob Hooft <hooft@EMBL-Heidelberg.DE>, Michael Mauldin <mlm@cs.cmu.edu>
Version:  1.1.2-2

lilo LILO

The LInux-LOader: LILO boots Linux from your hard drive. It can also boot other operating systems such as MS-DOS and OS/2, and can even boot DOS from the second hard drive. The configuration file is /etc/lilo.conf. Contents: lilo-20

Copyright:  (c) 1992-1997 Werner Almesberger
Author:  Werner Almesberger <almesber@di.epfl.ch>
Version:  20-23

loadlin LOADLIN

LOADLIN.EXE is a loader (running under DOS) for LINUX kernel images which fully supports the kernel’s command line feature. LOADLIN (and BOOTLIN) are the safest ways to boot Linux from your hard disk. If you have DOS (in addition to Linux) on your machine, LOADLIN is the right tool for booting Linux. In contrast to LILO, LOADLIN does not modify the master boot record. Also, with LOADLIN, the physical location of the kernel image is not important. With LILO, an update must be run each time the kernel file is moved on the media (i.e. whenever you recompile the kernel or restore the kernel image from a backup medium). This means even if you have to restore your whole Linux filesystem because of a crash, you can still reboot the system with LOADLIN. All you need is LOADLIN.EXE and a kernel image (zImage) copied to your DOS-partition. Another good reason to install LOADLIN is that you can install a mini emergency-system on a UMSDOS partition (such as MiniLinux). Then, you can always boot this system with LOADLIN and you’ll be able to repair your main Linux system in those rare cases where you can’t boot (i.e. problems with your rc scripts or you simply forgot the root password). If you haven’t yet installed LOADLIN by means of the DOS-Setup, you have to unpack the file lodlin16.zip into your DOS-partition and copy the kernel image to the same place. Assuming your DOS C: partition is mounted as /dos, proceed as follows: cd /dos mkdir loadlin cd loadlin unzip -d /root/lodlin15.zip cp -p /vmlinuz .

Copyright:  (c) 1994,95,96 Hans Lermen
Author:  Hans Lermen <lermen@suse.de>
Version:  1.6a-1

localedb database for ’locale’

Version:  1.99-2

man Program for displaying man pages.

Program for displaying man-pages on the screen or sending them to a printer (using groff). Contents: man db-2.3.10

Author:  G. Wilford <G.Wilford@ee.surrey.ac.uk>
Version:  2.3.10-19

modules Load modules into the kernel

Utilities for the loading of kernel modules. Included are ’insmod’, ’lsmod’, ’rmmod’, ’depmod’, ’modprobe’. The configuration file /etc/conf.modules can be used to pass parameters to the modules; for this reason ’modprobe’ is recommended for loading modules.

Contents: modules-2.1.13 modutils-2.1.85

Author:

Bjorn Ekwall <bj0rn@blox.se> , Jon Tombs <jon@gtex02.us.es> , Bas Laarhoven <bas@vimec.nl>

Version:  2.1.85-11

ncurses New curses libraries

As soon as a text application needs to directly control its output to the screen (e.g. wants to place the cursor at location x y and then write text), ncurses is used. The panel and the forms libraries are included in this package. These new libraries support color, special characters and panels. Contents: ncurses-4.2

Copyright:  (c) 1992-1995 Zeyd M. Ben-Halim, Eric S. Raymond

Author:

Zeyd M. Ben-Halim <zmbenhal@netcom.com> , Eric S. Raymond <esr@thyrsus.com>, Thomas Dickey <dickey@clark.net>, Juergen Pfeifer <Juergen.Pfeifer@t-online.de>

Version:  4.2-5

net tool Important programs for networking.

This package contains essential programs for network administration and maintenance: netstat, hostname, arp, ifconfig, rarp and route. Contents: net-tools-1.32-alpha

Copyright:

(c) 1988-1993 MicroWalt Corporation, 1993-1994 Swansea University Computer Society, 1995-1996 Bernd Eckenfels

Author:

Fred N. van Kempen <waltje@uwalt.nl.mugnet.org>, Bernd Eckenfels <nettools@lina.inka.de>, Fred Baumgarten <dc6iq@insu1.etec.uni-karlsruhe.de>, Peter Tobias <tobias@et-inf.fho-emden.de>

Version:  1.32alpha-17

netcfg Network configuration files in /etc

All of the basic configuration files for the network programs including /etc/inetd.conf, /etc/protocols and /etc/services. These are often used by network routines in the C-library and therefore must be installed for all network programs. Contents: netcfg

Version:  98.8.17-0

nkita A collection of basic network programs.

Attention! Former releases of this package included dip and sliplogin. These programs have moved to package slip. Contents: bootp-DD2.4.3 bootpc.v061 bwnfsd icmpinfo-1.10 linux pcnfsd2 netdate nfs-server-2.2beta29 pidentd-2.7.4 tcp wrappers 7.6 traceroute-1.4a5 ytalk-3.0.3-8bit

Author:

Florian La Roche <florian@suse.de>, Gordon W. Ross <gwr@mc.com>, Jon Peatfield <J.S.Peatfield@damtp.cam.ac.uk>, Fred N. van Kempen <waltje@uwalt.nl.mugnet.org>, Laurent Demailly <dl@hplyot.obspm.fr>, Alexander Bezprozvanny <bez@parsy.spb.su>, Donald Becker <becker@cesdis.gsfc.nasa.gov>, Peter Eriksson <pen@lysator.liu.se>, Sven Goldt <goldt@pegasus.in-berlin.de>, Jan Wedekind <jan@todonix.ping.de>, Wietse Venema <wietse@wzv.win.tue.nl>

Version:  98.7.7-5

nkitb A collection of net-programs like ftp, telnet,...

A collection of network Utilities. Contents: NetKit-0.09 NetKit-B-0.06 linux-ftpd-0.09-shadow portmap 5beta routed-8.10

Author:  Florian La Roche <florian@suse.de>
Version:  98.7.20-4

nvi ’vi’ clone

A freely redistributable replacement for the Berkeley ex and vi text editors. Contents: nvi-1.79

Author:  Keith Bostic <bostic@cs.berkeley.edu>
Version:  1.79-15

pcmcia PCMCIA support for Linux

This package contains the kernel modules for supporting the following PCMCIA cards:

Ethernet: 3Com 3c589, 3c589B, 3c589C, Accton EN2212, EN2216 EtherCard, Allied Telesis Ethernet, AST 1082 Ethernet, CNet CN30BC, CN40BC Ethernet, Datatrek NetCard, Digital EtherWORKS Turbo Ethernet, D-Link DE-650, Edimax Technology Ethernet Combo, EFA InfoExpress SPT EFA 205 10baseT, Eiger Labs EPX-ET10T2 Combo, EP-210 Ethernet, Epson Ethernet, Explorer NE-10000 Ethernet, Farallon Etherwave, Genius ME3000II Ethernet, Grey Cell Ethernet, GVC NIC-2000P Ethernet Combo, Hypertec HyperEnet, IBM CreditCard Ethernet Adapter, IC-Card Ethernet, Infotel IN650ct Ethernet, Katron PE-520 Ethernet, Kingston KNE-PCM/M, LANEED Ethernet, Linksys EtherCard, Longshine Ethernet, Maxtech PCN2000 Ethernet, NDC Instant-Link, Network General “Sniffer”, New Media Ethernet, Novell/National NE4100 InfoMover, PreMax PE-200 Ethernet, Proteon Ethernet, RPTI EP400 Ethernet, SCM Ethernet, Socket Communications Socket EA LAN Adapter, Thomas-Conrad Ethernet, Trust Ethernet Combo, Volktek Ethernet Modem cards: All should work. Memory cards: New Media SRAM, Epson 2MB SRAM, Intel Series 2 and Series 2+ Flash SCSI adapter: Qlogic FastSCSI, New Media Bus Toaster SCSI, Adaptec APA-1460 SlimSCSI The following cards don’t work: Xircom ethernet and ethernet/modem cards, Canon/Compaq PCMCIA floppy drive Under construction: Adaptec/Trantor APA-460 SlimSCSI (David Hinds <dhinds@allegro.stanford.e du>), New Media .WAVjammer (Roger Pao <rpao@paonet.org>), Nat’l Inst DAQ-700 (Tim Raymund <traymund@arlut.utexas.edu>), PreMax SC100 SCSI (W.Kevin Pedigo <pedigo@cyberspace.com>), IBM Token Ring (Steve Kipisz <kipisz@vnet.ibm.com>), AT&T GIS / NCR WaveLAN (Anthony Joseph <adj@pdos.lcs.mit.edu>) Contents: pcmcia-cs-3.0.0

Copyright:  (c) 1994, 1995, 1996 David A. Hinds
Author:  David Hinds <dhinds@hyper.stanford.edu>
Version:  3.0.4-4

perl Perl interpreter

What would Linux/Unix be without Perl ;-) To be able to use perl, you need to generate the system-dependent perl-include files. As root, on an installed system, simply enter: cd /usr/include ; h2ph *.h sys/*.h linux/*.h Contents: perl5.004 04

Author:  Larry Wall, Louis J. LaBash, Jr. <llabash@siue.edu>
Version:  5.004 04-11

ps procps

Utilities for displaying process and memory information. Also includes Steve ”Mr. Bassman” Bryant’s enhanced ’w’. Contents: procinfo-12 procps-1.2.6 psmisc

Author:

Helmut Geyer <Helmut.Geyer@iwr.uni-heidelberg.de>, Charles L. Blake <cblake@bbn.com>, Michael K. Johnson <johnsonm@merengue.oit.unc.edu>, Werner Almesberger <almesber@di.epfl.ch>, Branko Lankester <branko@hacktic.nl>, Robert Nation <nation@rocket.sanders.lockheed.com>, David Engel, Larry Greenfield <greenfie@gauss.rutgers.edu>, Roger Binns, Brian Edmonds, Steve ”Mr. Bassman” Bryant <bassman@muttley.soc.staffs.ac.uk>

Version:  98.8.4-0

rpm rpm - Red Hat Package Manager

rpm (Red Hat Package Manager) is the main tool for managing software packages of the S.u.S.E Linux distribution. rpm can be used to install and remove software packages; with rpm it’s easy to update packages. rpm keep track of all these manipulations in a central database. This way it is possible to get an overview of all installed packages; rpm also supports database queries. Contents: RPM-HOWTO.html RPM-Tips.html rpm-2.4.12

Copyright:  (c) GPL
Author:  Erik Troan <ewt@redhat.com>, Marc Ewing <marc@redhat.com>
Version:  2.4.12-12

sh utils GNU shellutils

The programs in this package include: basename date dirname echo env expr false logname nice hohup pathchk printenv printf sleep stty tee test true who whoami yes tty uname Contents: sh-utils-1.12 sh-utils-1.16

Copyright:  (c) 1985-94 Free Software Foundation, Inc.

Author:

David J. MacKenzie <djm@gnu.ai.mit.edu>, Jim Meyering <meyering@nanet.ornl.gov>

Version:  1.16-13

shadow Shadow password suite

This package is configured to prevent root login from other then local consoles. If you want to enable root login via TCP/IP comment out ”CONSOLE” in /etc/login.defs. Please also have a look at our Support Database (sdb.rpm). Contents: shadow-961025

Version:  961025-15

shlibs The shared C libraries libc.so and libm.so

Shared C libraries - needed to run programs linked with libc and libm libraries.

Author:  H.J. Lu <hjl@gnu.ai.mit.edu>
Version:  5.4.46-1

shlibs6 dynamic libraries (built for glibc)

With this package you should be able to execute binaries that have been build for glibc. The libraries are installed in /usr/i486-linux-glibc/lib and ld.so is configured to use this library-path for glibc-binaries.

Version:  98.8.16-0

syslogd Sysklogd

The sysklogd package implements two system log daemons. The syslogd daemon is the general system logging daemon which is responsible for handling requests for syslog services. This version of syslogd is similar to the standard Berkeley product but with a number of compatible extensions. The klogd daemon runs either standalone or as a client of syslogd. Klogd ’listens’ to kernel log messages, prioritizes them and routes them to either output files or to syslogd. This version of klogd will optionally translate kernel addresses to their symbolic equivalents if provided with a system map. Contents: sysklogd-1.3

Copyright:

(c) 1983, 1988 Regents of the University of California, 1994 Dr. Greg Wettstein, Enjellic Systems Development

Author:

Dr. Greg Wettstein <greg@wind.rmcc.com>, Stephen Tweedie <sct@dcs.ed.ac.uk>, Juha Virtanen <jiivee@hut.fi> , Shane Alderton <shane@scs.apana.org.au>

Version:  1.3-20

sysvinit SysV-Style init

System V style init programs by Miquel van Smoorenburg that controls the booting and shutdown of your system. These support a number of system runlevels, each one associated with a specific set of utilities. For example, normal system runlevel is 2, which starts a getty on virtual consoles tty1 - tty6. Runlevel 3 starts xdm. Runlevel 0 shuts down the system. Please see the individual man pages for: inittab, initscript, halt, init, killproc, killall5, powerd, reboot, runlevel, shutdown, and telinit. Contents: mingetty-0.9.4 killproc-1.2 sysvinit 2.65

Copyright:  (c) 1991-1995 Miquel van Smoorenburg

Author:

Miquel van Smoorenburg <miquels@drinkel.ow.nl>, Florian La Roche <florian@suse.de>

Version:  2.65f-19

termcap The libtermcap

Contents: curses-1.0.0 termcap-2.0.8

Copyright:  (c) 1985, 1987, 1993 Free Software Foundation, Inc.
Author:  H.J. Lu <hjl@gnu.ai.mit.edu>, David J. MacKenzie <djm@gnu.ai.mit.edu>
Version:  2.0.8-15

terminfo Complete /usr/lib/terminfo database

This is the terminfo reference database, maintained by Eric Raymond. This database is the official successor of the 4.4BSD termcap file and contains information about any known terminal. The ncurses library makes use of this database in order to use the terminals correctly. If you just use the Linux console, xterm and VT100, you probably will not need this database - a minimal /usr/lib/terminfo tree for these terminal is already included in the ncurses package.

Author:  Eric S. Raymond <esr@thyrsus.com>
Version:  4.2-5

texinfo GNU texinfo

Texinfo is a documentation system that uses a single source file to produce both on-line information and printed output. Using Texinfo, you can create a printed document with the normal features of a book, including chapters, sections, cross references, and indices. From the same Texinfo source file, you can create a menu-driven, on-line Info file with nodes, menus, cross references, and indices. This package is needed to read the documentation files in /usr/info. Contents: texi2html-1.51 texi2roff-2.0 texinfo-3.12

Copyright:  (c) 1987-95 Free Software Foundation, Inc.

Author:

Brian Fox <bfox@gnu.ai.mit.edu>, Jim Meyering <meyering@nanet.ornl.gov>, Noah Friedman <friedman@prep.ai.mit.edu>, Richard Stallman <rms@gnu.ai.mit.edu>, Robert J. Chassell <bob@gnu.ai.mit.edu>, Daniel Hagerty <hag@gnu.ai.mit.edu>, Roland McGrath <roland@gnu.ai.mit.edu>, David J. MacKenzie <djm@gnu.ai.mit.edu>, Charles Hannum <mycroft@gnu.ai.mit.edu>

Version:  3.12-7

textutil GNU textutil

These are the GNU text file (actually, file contents) processing utilities. Most of these programs have significant advantages over their Unix counterparts, such as greater speed, additional options, and fewer arbitrary limits. The programs in this package are: cat, cksum, comm, csplit, cut, expand, fold, head, join, nl, od, paste, pr, sort, split, sum, tac, tail, tr, unexpand, uniq, and wc. Contents: textutils-1.22

Version:  1.22-14

timezone Time zone descriptions

These are configuration files that describe possible time zones. With YaST you can select an appropriate time zone for your system. Contents: tz97e

Copyright:  (c) 1985-95 Regents of the University of California

Author:

Keith Bostic, Bob Devine, Paul Eggert, Robert Elz, Guy Harris, Mark Horton, John Mackin, Bradley White, <tz@elsie.nci.nih.gov>

Version:  97e-16

umsprogs File system utilities

Miscellaneous utilities which are needed for Jacques Gelinas’ UMSDOS file system. Contents: umsdos progs-0.9.src

Copyright:  (c) 1993 Jacques Gelinas
Author:  Jacques Gelinas <jacques@solucorp.qc.ca>
Version:  0.9-26

util Selected utilities

Selected utilities compiled from Rik Faith’s huge utility collection: agetty arch banner cal chroot chsh clear clock col colcrt colrm column ctrlaltdel dmesg domainname fdformat fdisk frag fsck hexdump ipcrm ipcs kbdrate kill last logger login look mesg mkfs.minix mkswap more mount newgrp passwd pwd rdev renice reset rev script setfdprm setserial setsid setterm simpleinit sln strings swapoff swapon sync syslogd tsort tunelp ul umount update wall whereis write zdump zic zones

Contents: mount-2.6g util-linux-2.6

Author:  Rik Faith <faith@cs.unc.edu>
Version:  2.6-20

vim Vim: Vi IMproved

Vim (Vi Improved) is an almost compatible version of the UNIX editor vi whereby almost every possible command can be performed using only ASCII characters. Only the ’Q’ command is missing (you don’t need it). Many new features have been added: multi level undo, command line history, filename completion, block operations, editing of binary data, etc. Vi is available for the AMIGA, MS-DOS, Windows NT, and various versions of UNIX.

For the SuSE Distribution, Vim is used as /usr/bin/vi.

Contents: vim-4.6

Author:  Bram Moolenaar <mool@oce.nl>
Version:  4.6-13

yast Yet another Setup Tool

YaST is a tool for installing and configuring the S.u.S.E. Linux system. It provides easy administration capabilities. Installation and configuration of the system by means of YaST is explained detailed in the accompanying handbook of S.u.S.E Linux.

Contents: misc-yast-bins sel yast-0.89.2

Copyright:  (c) 1995,96 S.u.S.E. GmbH
Author:  Michael Andres <ma@suse.de>, Thomas Fehr <fehr@suse.de>
Version:  0.97.2-1

compat some libraries in compatibility-versions

Contains libncurses.so.1.9.7a (from the old sources)

libgdbm.so.2 and libdb.so.2 (originals with changed major-release)

Contents: db.1.85 gdbm-1.7.3 ncurses-1.9.7a

Version:  97.11.27-7

oldshlbs Old Version of libc

This package includes older versions of libc, along with the corresponding include-files.

Contents: libc-5.4.7

Copyright:  (c) 1981-95 Regents of the University of California. , Free Software Foundation, Inc.

Author:

H.J. Lu <hjl@gnu.ai.mit.edu>, Alan Cox <Alan.Cox@linux.org>, Adam J. Richter <adam@yggdrasil.com>, Alan Modra <alan@spri.levels.unisa.edu.au>, Andreas Schwab <schwab@issan.informatik.uni-dortmund.de>, Andrew Walker <andy@keo.kvaerner.no>, Andries Brouwer <aeb@cwi.nl>, Charles Lopes <tjarls@infm.ulst.ac.uk>, Chip Salzenberg <chip@fin.uucp>, Christoph Kern <kern@informatik.uni-wuerzburg.de>, Corey Minyard <minyard@wf-rch.cirr.com>, Craig Smith <craig@rezrov.natinst.com>, Daniel Quinlan <quinlan@yggdrasil.com>, Delman Lee <delman@mipg.upenn.edu>, Dominik Kubla <dominik.kubla@unimainz.de>, Dsouza <m.dsouza@mrc-applied-psychology.cambridge.ac.uk>, Eric Youngdale <eric@aib.com>, Peter Eriksson <pen@signum.se>, Florian La Roche <florian@suse.de>, Fred N. Van Kempen <waltje@uwalt.nl.mugnet.org>, Gerhard Koerting <g.koerting@koerting.techem.ruhr-uni-bochum.de>, H. Peter Anvin <hpa@yggdrasil.com>, Harald Koenig <koenig@tat.physik.unituebingen.de>, Igor Sharfmesser <igor@irbis.alma-ata.su>, J. Alan Eldridge <alane@wozzle.linet.org>, Jaakko Hyvatti <jaakko.hyvatti@www.fi>, James H. Cloos Jr. <cloos@rahul.net>, Jeschke <i6210109@ws.rz.tu-bs.de>, Jim Van Zandt <jrv@vanzandt.mv.com>, Jochen Hein <Hein@Informatik.TU-Clausthal.de>, Jochen Kiemes <jochen kiemes@s2.maus.de>, John F. Haugh Ii <jfh@rpp386>, Jonathan Kamens <jik@cam.ov.com>, Karl Hakimian <hakimian@eecs.wsu.edu>, Kaz Sasayama <kaz@lilia.iijnet.or.jp>, Marc Gingold <gingold@amoco.saclay.cea.fr>, Martin Kuck <martin@unifix.de>, Martin Pirker <pirker@eiunix.tuwien.ac.at>, Mathew G. Monroe <mmex@shadowland.pc.cc.cmu.edu>, Matthias Urlichs <urlichs@smurf.noris.de>, Matthias Urlichs <urlichs@test.smurf.noris.de>, Michael Riepe <riepe@ifwsn4.ifw.uni-hannover.de>, Mike Kennedy <kmb1@gate.net>, Mike Meissner <meissner@cygnus.com>, Miquel van Smoorenburg <miquels@drinkel.ow.nl>, Mitchum Dsouza <m.dsouza@mrc-applied-psychology.cambridge.ac.uk>, Nick Holloway <nick.holloway@alfie.demon.co.uk>, Nickolay Saukh <nms@ussr.eu.net>, Nicolai Langfeldt <janl@ifi.uio.no>, Niibe Yutaka <gniibe@mri.co.jp>, Olaf Flebbe <o.flebbe@science-computing.uni-tuebingen.de>, Patrick D’Cruze <pdcruze@orac.iinet.com.au>, Paul Eggert <eggert@twinsun.com>, Per Bothner <bothner@cygnus.com>, Peter Himmler <himmler@sparcy.mayn.sub.de>, Philip Gladstone <philip.gladstone@mail.citicorp.com>, Remy Card <remy.card@masi.ibp.fr>, Rick Sladkey <jrs@world.std.com>, Robert Andrew Ryan <rr2b+@andrew.cmu.edu>, Roland Mcgrath <roland@gnu.ai.mit.edu>, Rolf Rossius <rossius@hrz.tu-chemnitz.de>, Ross Biro <biro@yggdrasil.com>, Sean Mcneil <sean@mcneil.com>, Stephen L Moshier <moshier@world.std.com>, Stephen Tweedie <sct@dcs.ed.ac.uk>, Steve Losen <losen@delphi.com>, Steve M. Robbins <steve@nyongwa.montreal.qc.ca>, Swen Thuemmler <swen@uni-paderborn.de>, Theodore Ts’O <tytso@mit.edu>, Thomas Pundt <pundtt@goedel.uni-muenster.de>, Ulrich Drepper <drepper@gnu.ai.mit.edu>, Uwe Ohse <uwe@tirka.gun.de>, Vadim V. Vlasov <vvv@ms3.inr.ac.ru>, Wes Cowley <wcowley@cftnet.com>,

Version:  97.10.9-14

0.1.2 AP: Programs that don’t need X

a2ps Converts ASCII text into PostScript

Converts ASCII text into PostScript. This feature is used by e.g. apsfilter, in order to pretty-print ASCII text.

Contents: a2ps-V4.3

Author:  Miguel Santana <M.Santana@frgu.bull.fr>
Version:  4.3-14

alien Alien - Perl script to convert packages

Experimental Perl script to convert software packages. The following formats are supported: deb (Debian), rpm (Red Hat), and tgz (Slackware).

AP: Programs that don’t need X

Contents: alien-5.19

Copyright:  (c) GPL

Author:

Joey Hess <joey@kite.ml.org>, Christoph Lameter <clameter@debian.org>, Randoph Chung <rc42@cornell.edu>

Version:  5.19-9

apmd APM daemon

1) Daemon for monitoring battery status via APM

2) Application to query /proc/apm Includes: apm, apmd, xapm, and tailf

Contents: apmd-2.4

Copyright:  (c) 1996 Rickard E. Faith
Author:  Rik Faith <faith@cs.unc.edu>
Version:  2.4-12

aps The ultimate printer filter: apsfilter

This printer filter lets you print arbitrary files on your printer. The file type is recognized automatically. When needed ghostscript is run to print PostScript files on non-PostScript printers. In order to configure the filter, start YaST in your installed system. Choose ”System administration” in the main menu, then ”Integrate hardware into system”, and finally ”Configure printers”. Fill out the mask, and YaST will configure apsfilter for you. The conventional SETUP of apsfilter is sure available as /var/lib/apsfilter/SETUP This revised script uses the program dialog to make the usage more easier for you. The documentation of this version of apsfilter revised by S.u.S.E. is placed in /usr/doc/packages/apsfilter/. Contents: susefilter-1.4 Please note: Ghostscript is required if you will be using apsfilter with a non-PostScript printer.

Copyright:  (c) 1995 Andreas Klemm, Thomas Bueschgens, 1996,97 S.u.S.E.-GmbH

Author:

Andreas  Klemm  <andreas@knobel.gun.de>,  Thomas  Bueschgens

<sledge@hammer.oche.de>, werner@suse.de <werner@suse.de>

Version:  1.5.SuSE-5

arc a packer

ARC is a packer that was used on various operating systems. This package should allow you to unpack *.arc files under Linux. You shouldn’t use it for packing Linux-Software. Use gzip instead.

Contents: arc

Author:  Howard Chu <hyc@umix.cc.umich.edu>
Version:  5.21-13

ash ash shell

NetBSD’s ash (Almquist sh) for Linux is a small (62K - no job control) Bourne-compatible shell. Great for machines with low memory, but does not provide all the extras of shells like bash, tcsh, and zsh. Most shell scripts are run compatible to bash. Note that most scripts written for Linux use some bash-specific syntax. The Slackware setup scripts are a notable exception, since ash is the shell used during installation.

Contents: ash-linux-0.2

Copyright:  (c) 1989 by Kenneth Almquist, 1991 The Regents of the University of California

Author:

Kenneth Almquist, Florian La Roche <florian@suse.de>, Branko Lankester <branko@hacktic.nl>, Sunando Sen <sens@fasecon.econ.nyu.edu>, Arjan de Vet <devet@adv.win.tue.nl>

Version:  0.2-13

autolog Auto logout

This programs automatically logs out ”forgotten” logins.

Contents: autolog-0.31

Copyright:  (c) 1995 Kyle Bateman
Author:  Kyle Bateman <kyle@actarg.com>
Version:  0.31-13

bc GNU bc - command line calculator

bc is an interpreter that supports numbers of arbitrary precision and the interactive execution of statements. There are some similarities in the syntax to the C programming language. A standard math library is available through command line options. When used, the math library is read in before any other input files. bc then reads in all other files from the command line, evaluating their contents. Then, bc reads from standard input (usually the keyboard). The dc program is also included. Dc is a calculator which supports reverse-polish notation and allows unlimited precision arithmetic. Macros can also be defined. Normally, dc reads from standard in but can also read in files specified on the command line. A calculator with reverse-polish notation saves numbers to a stack. Arguments to mathematical operations (operands) are ”pushed” onto the stack until the next operator is read in which ”pops” its arguments off the stack and ”pushes” its results back onto the stack. Contents: bc-1.04

Copyright:  (c) 1991, 1992, 1993, 1994 Free Software Foundation, Inc.
Author:  Philip A. Nelson <phil@cs.wwu.edu>
Version:  1.04-14

buffer Buffering stdin and stdout

This is a program designed to speed up writing tapes on remote tape drives. When this program is put ”in the pipe”, two processes are started. One process reads from standard-in and the other writes to standard-out. Both processes communicate via shared memory.

Copyright:  (c) 1990,1991 Lee McLoughlin
Author:  Lee McLoughlin <L.McLoughlin@doc.ic.ac.uk>
Version:  1.17-12

bzip Program for compression of files

Bzip is a very powerful program for compressing files. Contents: bzip2-0.1pl2

Copyright:  (c) 1996,1997 Julian Seward
Author:  Julian Seward <jseward@acm.org>
Version:  0.1p2-10

cdcons Console CD Player

A small CD Player that runs on the text console. Documentation: /usr/doc/packages/cdcons Contents: cd-console-1.1

Copyright:  (c) 1996, Osku Salerma
Author:  Osku Salerma <osku@iki.fi>
Version:  1.1-18

cdparano cd-audio grabber for problem-drives

This CDDA reader distribution (’cdparanoia’) reads audio from the CDROM directly as data, with no analog step between, and writes the data to a file or pipe as .wav, .aifc or as raw 16 bit linear PCM. Beta release, see: http://www.mit.edu/afs/sipb/user/xiphmont/cdparanoia/index.html

Copyright:  (c) GPL
Author:  Monty <xiphmont@mit.edu>
Version:  03alpha7-0

cdrecord Tool to write CD-Rs

Cdrecord has been implemented with a Philips CDD 521 and is now maintained with a Philips CDD-2000. It should run also on lots of Hardware, see /usr/doc/packages/cdrecord/README Contents: cdrecord-1.6

Copyright:  (c) Copyright 1996 J. Schilling
Author:  Joerg Schilling <joerg@schily.isdn.cs.tu-berlin.de>
Version:  1.6-11

cdwrite Writing CDRs

Program to write data and audio CDs using a CD writer.

Contents: cdwrite-2.0 Be very careful using this program. If you enter the wrong SCSI device, you could irrevocably overwrite a hard disk (including partition table and foreign partitions).

Copyright:  (c) 1995, Adam J. Richter
Author:  Adam J. Richter <adam@yggdrasil.com>
Version:  2.0-19

cdwtools CDW tools

Frontend for various programs used to create CDs

Contents: cdda2wav-0.9beta11 cdwtools-0.93

Copyright:  (c) 1996, Thomas Niederreiter
Author:  Thomas Niederreiter <p7003ad@sunmail.lrz-muenchen.de>
Version:  0.93-18

cfengine GNU cfengine

cfengine is an abstract programming language for system administrators of huge heterogeneous networks. With cfengine, system administrators have an easy and elegant way to maintain complicated networks. For introduction and documentation please look at the files doc/cfengine.tex in TeX mode or at /usr/info/cfengine.info.gz in TeXInfo mode.

Contents: cfengine-1.4.10

Copyright:  (c) 1995 Mark Burgess, Oslo
Author:  Mark Burgess
Version:  1.4.10-11

chipmunk Hardware ’debugging’.

Documentation: /usr/doc/packages/chipmunk.

Contents: chipmunk-1.4

Author:  R.E. Wolff <R.E.Wolff@et.tudelft.nl>, R.D.L. Stout <stout@dutecai.et.tudelft.nl>
Version:  1.4-15

ctags Generate C language tag files for use with vi

Copyright:  (c) 1996 Darren Hiebert
Author:  Darren Hiebert <darren@sirsi.com>
Version:  1.5-12

deco Demos Commander

DECO or Demos Commander is a visual front-end to the UNIX operating system, similar to the Norton Commander. Usually, DECOS displays two fields. In each of these, the contents of a directory is listed. The program is controlled using the keyboard. Use the cursor keys to choose a file, press return to select it; TAB changes between the two fields. Commands can be executed by using the F-keys or the menu. Additional commands can be added by the user. Unfortunately, the sources for this program are not freely available.

Contents: deco-3.5

Copyright:  (c) 1989-1993 Serge Vakulenko
Author:  Serge Vakulenkom <vak@kiae.su>
Version:  3.5-7

dialog Menus and input boxes for shell scripts

This program lets you use menus and dialog boxes in shell scripts.

Contents: dialog-0.62

Author:

Savio Lam <lam836@cs.cuhk.hk>, Alessandro Rubini <rubini@ipvvis.unipv.it>, Florian La Roche <florian@suse.de>, Marc Ewing <marc@redhat.com>

Version:  0.62-19

dosfsck Filesystem Check for DOS Filesystems

dosfsck is the Linux equivalent of PC/MS-DOS’ CHKDSK. It checks the consistency of PC/MS-DOS file systems and optionally tries to repair them. The tests dosfsck performs are described in the man page.

Contents: dosfsck.1 Attention: This Package is still in ALPHA state, i.e. not absolutely error free and usage may result in serious damage of your DOS partition.

Author:  Werner Almesberger <almesber@bernina.ethz.ch>
Version:  beta1-16

duconv Unix <-> DOS Converter

A DOS <--> UNIX converter which even correctly converts German umlauts. Something everybody(!) is looking for..;) Contents: duconv-0.1

Copyright:  (c) 1994 Torsten F”ahrmann, TortiSoft
Author:  Torsten F”ahrmann <tfaehr@aol.com>
Version:  0.1-13

dump Backup program

a Backup program. Contents: dump-0.3 e2fsprogs-1.10

Copyright:  (c) Copyright (c) 1980, 1993 The Regents of the University of California
Author:  Remy Card <Remy.Card@freenix.fr>
Version:  0.3-17

ed standard old un*x line editor

Author:  Andrew Moore <alm@worm.talke.org>
Version:  0.2-13

edy An editor with menus

Simple ASCII editor. Several files can be edited simultaneously. Commands can be accessed via pull-down menus. Contents: edy-2.0

Copyright:  (c) 1993 Markus Bulik
Version:  2.0-13

enscript ASCII to PostScript(tm) converter

Enscript converts ASCII files to PostScript and writes the generated output to a file or sends it directly to the printer. Enscript is configured in /etc/enscript.cfg. Contents: enscript-1.5.5

Author:  Markku Rossi <mtr@iki.fi>
Version:  1.5.5-7

gcal A program for printing calendars

Gcal is a program for printing calendars. Gcal displays a calendar for a month or a year, eternal holiday lists and fixed date lists, in many ways. The program correctly omits the dates that were skipped when the current Gregorian calendar replaced the earlier Julian calendar. Contents: gcal-2.40

Copyright:  (c) GPL
Author:  Thomas Esken <esken@uni-muenster.de>
Version:  2.40-11

gdtmon Administrate a RAID array with a Vortex GDT controller

Version:  1-8

git GNU Interactive Tools

A complete system shell, for those who don’t prefer the command line :-) Contents: git-4.3.15

Copyright:  (c) 1994 Tudor Hulubei & Andrei Pitis
Author:  Tudor Hulubei <tudor@cs.unh.edu>, Andrei Pitis <pink@pub.ro>
Version:  4.3.15-13

glimpse Text search in huge databases

You can index lots of ASCII files very easily. Glimpse creates hash tables which enable you to search (and find!) any given text.

In general glimpse does the same as grep but far more efficient and faster. Contents: glimpse-4.1

Copyright:

(c) Udi Manber and Burra Gopal of the University of Arizona, and Sun Wu of the National Chung-Cheng University, Taiwan.

Author:  Udi Manber <glimpse@cs.arizona.edu>
Version:  4.1-9

gpm General purpose mouse support

This program is based on the ”selection” package, and allows you to use the mouse to cut and paste text from the screen. It also acts as a mouse server for applications running on the Linux console, such as the Midnight Commander and dialog. A second program, gpm-root, provides pull down menus to be called up by pressing Ctrl+mouse-key. NOTE: This program may cause problems when you start X on a system that does not use a serial mouse. If you get an ’unable to open mouse device’ error from X, disable gpm. Contents: gpm-1.13

Copyright:  (c) 1994,1995 Alessandro Rubini
Author:  Alessandro Rubini <rubini@ipvvis.unipv.it>
Version:  1.13-17

gs both Ghostscript with X11 and with VGA features

This gs version is for X11 machines and vga terminal. Currently there is the newer version 5.10 of ghostscript available. Due to a change in the licence of this version we are not allowed to put this version onto this cdrom. Nevertheless this version you will find this version on our FTP server.

Copyright:  (c) 1989-96 Aladdin Enterprises.
Author:  Peter Deutsch <ghost@aladdin.com>
Version:  5.10-0

gs fonto Optional fonts for Ghostscript

Some useful fonts for the PostScript interpreter Ghostscript.

Version:  5.10-0

gs fonts Standard fonts for Ghostscript

The necessary fonts for the PostScript interpreter Ghostscript.

Version:  5.10-0

gs serv Ghostscript without X11 and without VGA features

This gs version should only be installed on print servers without X11. Currently there is the newer version 5.10 of ghostscript available. Due to a change in the licence of this version we are not allowed to put this version onto this cdrom. Nevertheless this version you will find this version on our FTP server.

Copyright:  (c) 1989-96 Aladdin Enterprises.
Author:  Peter Deutsch <ghost@aladdin.com>
Version:  5.10-0

gs vga Ghostscript without X11 but with VGA

This gs version is for X11-less machines and vga terminal or graphic card. The graphic card should be supported by the SVGA lib. Currently there is the newer version 5.10 of ghostscript available. Due to a change in the licence of this version we are not allowed to put this version onto this cdrom. Nevertheless this version you will find this version on our FTP server.

Copyright:  (c) 1989-96 Aladdin Enterprises.
Author:  Peter Deutsch <ghost@aladdin.com>
Version:  5.10-0

gs x11 Ghostscript for X11

If you plan to preview PostScript documents under X, you need to install this. Currently there is the newer version 5.10 of ghostscript available. Due to a change in the licence of this version we are not allowed to put this version onto this cdrom. Nevertheless this version you will find this version on our FTP server.

Copyright:  (c) 1989-96 Aladdin Enterprises.
Author:  Peter Deutsch <ghost@aladdin.com>
Version:  5.10-0

hello A friendly greeting program

The GNU ‘hello’ program produces a familiar, friendly greeting. It allows nonprogrammers to use a classic computer science tool which would otherwise be unavailable to them. Because it is protected by the GNU General Public License, users are free to share and change it. Contents: hello-1.3 Hello, world!

Copyright:  (c) GPL

Author:

The  King  <elvis@graceland.gnu.ai.mit.edu>,  Charles  Han-
num  <mycroft@hill.gnu.ai.mit.edu>,  David  J.  MacKen-
zie  <djm@apple-gunkies.gnu.ai.mit.edu>,  Jan  Brittenson

<bson@kropotkin.gnu.ai.mit.edu>, Michael I Bushnell <mib@geech.gnu.ai.mit.edu>,

Mike  Haertel  <mike@churchy.gnu.ai.mit.edu>,  Noah  Fried-
man  <friedman@nutrimat.gnu.ai.mit.edu>,  Roland  McGrath

<roland@geech.gnu.ai.mit.edu>

Version:  1.3-12

id utils Language-independent identifier database tool

mkid is a simple, fast, high-capacity, language-independent identifier database tool. Actually, the term ‘identifier’ is too limiting -- mkid stores tokens, be the program identifiers of any form, literal numbers, or words of human-readable text. Database queries can be issued from the command-line, or from within emacs, serving as an augmented tags facility. Contents: id-utils-3.2

Copyright:  (c) GPL
Author:  McGary <gkm@gnu.ai.mit.edu>, <Tom.Horsley@mail.hcsc.com>
Version:  3.2-18

idled Idle daemon

This program automatically logs out users who have been idle for some time; this time, of course, can be set by the system administrator. Contents: idled-1.12

Author:  Michael P. Crider <crider@cs.hope.edu>
Version:  1.12-14

ingres A relational database system

This powerful relational database comes with the full sources and its own query language. This query language can be easily integrated into your own C-programs using the Ingres-internal preprocessor. It comes without saying that the system provides concepts of transaction and recovery to ensure data consistency.

Copyright:  (c) The Regents of the University of California 1977
Version:  8.9.4-11

ingrtool Tool package for Ingres

This add-on for Ingres gives you a Tcl/Tk Interface to the good old ingres database. For installation change to /usr/lib/ingrestools-1.0 and call ”Install”. Documentation can be found under /usr/lib/ingrestools-1.0/doc. (German and English) Contents: ingrestools-1.0

Copyright:  (c) 1994 Gert Reinauer
Author:  Gert Reinauer
Version:  1.0-13

inlinem inlinemath

With this package you get inlined assembler functions of the mathematic library under /usr/include/i386/. To use this functions the system needs a math coprocessor (386 combined with 387, 486 with builtin coprocessor, 586). Documentation is stored under /usr/doc/packages/inline-math-1.6/ and the header file math.h and extensions.h under /usr/include/i386/. To include this files you have to patch the file /usr/include/math.h with /usr/doc/packages/inline-math-1.6/math.h-patch.

You can control the usage of /usr/include/i386/ math.h with the gcc option ‘-D NO MATH INLINES’ and the usage of /usr/include/i386/extensions.h with the gcc option ‘-D MATH EXTENSIONS’. Note: If the package is removed the patched version of /usr/include/math.h has to be replaced by the original version. Contents: inlinem-1.9 For this package you really need a math coprocessor!!! After installation you have to patch the header file /usr/include/math.h with /usr/doc/packages/inlinemath-1.6/math.h-patch!

Copyright:  (c) (C) 1995 John C. Bowman
Author:  John C. Bowman <bowman@hagar.ph.utexas.edu>
Version:  1.9-13

isapnp ISA plug and play configuration utility

Two programs - one allows the dumping of resource data and generation of a skeleton configuration file, the other configures ISA PnP hardware using a configuration file. More Info: /usr/doc/packages/isapnp/README.SuSE Contents: isapnp

Copyright:  (c) GNU
Author:  Peter Fox <fox@roestock.demon.co.uk>
Version:  1.10-20

jed JED Editor

JED is an extremely powerful but small emacs programmer’s editor that is extensible in a C-like macro language and is able to perform color syntax highlighting. Among the many features: Emacs, wordstar, EDT emulation; C, Fortran, TeX, text editing modes; Full undo; GNU Emacs compatible info reader, and lots more. 8 bit clean, so you can even edit binary files! Contents: jed097-12 slang0.99-23

Copyright:  (c) 1992, 1993, 1995 John E. Davis
Author:  John E. Davis <davis@amy.tch.harvard.edu>
Version:  097.12-12

joe Joe Text editor

JOE (Joe’s Own Editor) is a freeware ASCII editor for UNIX. Joe is similar to most IBM PC text editors. The keyboard combinations are similar to WordStar[tm] and Turbo C. Some of Joe’s features include: -full support of termcap/terminfo (useful for device independent description to control screen output). -optimizing screen refresh even as GNU Emacs (Joe can even be used at 2400 baud). -easy installation. All features for UNIX integration of the VI: a marked block of text can be filtered through a UNIX command; and wherever JOE accepts a filename parameter the following can also be substituted: !command :to redirect from/to another command >>filename :to append onto an existing file filename,start,size :to edit a part of a file/device

- :to use standard input/output When JOE has several files opened at the same time, each file is displayed in its own window. Additionally, JOE allows for shell-windows whereby the output of the executed commands are saved in a buffer, automatic filename completion (via TAB), help windows, undo-redo, search and replace using regular expressions. Contents: joe2.8

Copyright:  (c) 1992 Joseph H. Allen
Author:  Joseph H. Allen <jhallen@world.std.com>
Version:  2.8-16

jove Jonathan’s Own Version of Emacs

JOVE (Jonathan’s Own Version of Emacs) is a text editor based mainly on the original Emacs editor written by Richard Stallman at M.I.T. Although jove is mostly Emacs compatible there are several differences so that you shouldn’t rely on it. Jove offers you e.g. automatic indentation, multiple view of more than one file, and shorthands in form of keys, words, lines and paragraphs A mostly Emacs compatible text editor, with man pages and the ’teachjove’ tutorial program. If you’d rather have the real thing, the actual Emacs editor/environment is available as disk series ’E’. JOVE is configured by /var/lib/jove/jove.rc. Contents: jove

Copyright:  (c) 1986, 1987, 1988 by Jonathan Payne
Author:  Jonathan Payne <jpayne@sun.com>
Version:  4.16-17

lha packer

Lha is a packer comparable to ZIP (PKZIP), ZOO etc. It has been included for compatibility reasons only. For general archiving purposes use GZIP, because this is standard for Linux. Contents: lha

Copyright:  (c) 1992 Masaru Oki
Author:  Masaru Oki <oki@wbg.telcom.oki.co.jp>
Version:  1.0-14

logsurf Powerfull tool to observe logfiles

This program makes it possible to observe logfiles and to react on special events, that can be defined by regular expressions, with a defined action, e.g. to send a mail.

Copyright:  (c) 1995 by Wolfgang Ley
Author:  Wolfgang Ley <ley@cert.dfn.de>
Version:  1.41-2

lsof Lsof lists information about files opened by processes

An open file may be a regular file, a directory, a block special file, a character special file, an executing text reference, a library, a stream or a network file (Internet socket, NFS file or UNIX domain socket.) A specific file or all the files in a file system may be selected by path. Contents: lsof 4.16

Copyright:  (c) 1994 Purdue Research Foundation, West Lafayette, Indiana
Author:  Victor A. Abell <abe@purdue.edu>
Version:  4.31-6

makewhat makewhatis - create a whatis database

The man system in S.u.S.E. Linux (package man) does not need a whatis database. Nevertheless some manual browser (e.g. tkman) still need this database. For this reason we included this package. Contents: makewhat

Copyright:  (c) 1992 Rickard E. Faith
Author:  Werner Fink <werner@suse.de>
Version:  98.7.24-0

mbedit Mutil platform editor

mbedit is a full screen text editor with macro option, online calculator, command history buffer, hex editor and many other features. It is available for some common operating systems as: MS-DOS, os/9, Unix. Contents: mbedit.7.71

Copyright:  (c) 1992-1997 / Michael Braun
Author:  Michael Braun <braun@hydro2.atlas.de>
Version:  7.71-13

mc Midnight Commander

The Midnight Commander is a Norton Commander clone; a program that manipulates and manages files and directories. It is useful, fast, and has color displays on the Linux console. It also has mouse support if you run the ”gpm” mouse server. This program requires the terminal description files in /usr/lib/terminfo, which are found in ncurses.tgz or terminfo.tgz. You can also use the Midnight Commander under X with your mouse. If you enter ”mc -c”, colors will be used.

Under the Midnight Commander, the screen is divided into four sections, whereby the majority of the screen is covered by two directory panels. The second to last line on the screen is the shell command line; the last line displays the function key assignments. At the very top, the menu list is shown. One of the directories displayed is the current working directory. This is where most of the commands are found. For certain commands, like copy and move, the second directory is used as the target directory.

Contents: mc-4.1.11

Copyright:  (c) 1994, 1995 Miguel de Icaza, Janne Kukonlehto

Author:

Miguel de Icaza <miguel@roxanne.nuclecu.unam.mx>, Janne Kukonlehto <jtklehto@paju.oulu.fi>, Dugan Orlando Porter <dugan@b011.eunet.es>, Radek Doulik <rodo@earn.cvut.cz>, Fred Leeflang <fredl@nebula.ow.nl>, Mauricio Plaza <mok@roxanne.nuclecu.unam.mx>, Jakub Jelinek <jj@sunsite.mff.cuni.cz>

Version:  4.1.11-15

mdutils Utilities for ’Multiple Devices’

You can combine different partitions (including SCSI and IDE) to one partition under Linux. More information can be found under /usr/doc/packages/mdutils an in the man pages for add, mdcreate and mdtab. Please be aware that the author himself considers this software as ALPHA. We at S.u.S.E. have not encountered any problems however.

Contents: md035

Author:  Marc Zyngier <zyngier@ufr-info-p7.ibp.fr>
Version:  0.41-6

mkdosfs tool to format dos partitions

This is a tool to create the filesystem on dos partitions which works similar to mke2fs. Documentation is under /usr/doc/packages/mkdosfs This programm is not fully stable yet! Use at your own risk!!

Copyright:  (c) GPL
Author:  Dave Hudson <dave@humbug.demon.co.uk>
Version:  0.4-12

mkisofs Creating CDs under Linux

Actually, mkisofs is a pre-mastering program that generates an iso9660 filesystem . It takes a snapshot of a given directory tree, and generates a binary image which corresponds to an iso9660 filesystem which can be written to a block device.

Contents: mkisofs-1.11.1 mkisofs-1.11

Copyright:  (c) 1993 Yggdrasil Computing, Incorporated

Author:

Eric Youngdale <eric@aib.com>, H. Peter Anvin <hpa@yggdrasil.com>, John Brezak <brezak@ch.hp.com>, Paul Eggert <eggert@twinsun.com>, Peter van der Veen <peterv@qnx.com>, R.-D. Marzusch <marzusch@odiehh.hanse.de>

Version:  1.12b3-2

mktemp Utility for tempfiles

mktemp is a small utility that interfaces to the mktemp() function call to allow shell scripts and other programs to use files in /tmp safely.

Contents: mktemp-1.4

Copyright:  (c) BSD
Author:  Todd C. Miller <Todd.Miller@courtesan.com>
Version:  1.4-12

mmv Multiple Move

This mv command allows you something like ”mmv ”*.c #1.c.old” ” to perform tasks like mv *.c *.c.old. Documentation: man mmv

Contents: mmv-1.01b

Copyright:  (c) 1989 by Vladimir Lanin
Author:  Vladimir Lanin <lanin@csd2.nyu.edu>
Version:  1.01b-13

mpegutil MPEG (de-)compression utility

Mpeg is an utility for the compression and decompression of MPEG files. MPEG animations can be composed out of single images, and decomposed into single images (in YUV format). Contents: mpeg-utils-1.2.1

Copyright:  (c) 1990-93 Andy C. Hung
Author:  Andy C. Hung <achung@cs.stanford.edu>
Version:  1.2.1-13

mt st mt-st controls magnetic tape drive operation

Although this mt program is tailored for SCSI tape, it can also be used with the QIC-02 driver and hopefully with other Linux tape drivers that use the same ioctls (some of the commands may not work with all drivers). Contents: mt-st-0.4

Copyright:  (c) 1980 The Regents of the University of California
Author:  Kai Makisara <Kai.Makisara@vtt.fi>
Version:  0.4-17

mtools Access files on an MS-DOS filesystem

Allows uncomplicated access to a MS-DOS filesystem on Disk, without mounting it. Included commands for working with MS-DOS Files mdir, mcd, mcopy, mformat. XDF support for OS/2 is also provided. Contents: mtools

Author:  Alain Knaff <Alain.Knaff@imag.fr>, David C. Niemi <niemidc@erols.com>
Version:  3.9.1-5

nc Northern Commander

It is not only the name which is similar to the well-known DOS program. Contents: nc-0.74

Copyright:  (c) 1993-94 Andreas Matthias
Author:  Andreas Matthias <matthias@hrz.uni-kassel.de>
Version:  0.74-15

obst The OBject system of STONE

OBST is the Objekt Management System of the STONE project. It was designed as a tool for saving persistent objects. OBST can be seen as an object-oriented database written in C. Contents: OBST3-4.3

Copyright:  (c) 1992-1994 by Forschungszentrum Informatik (FZI)
Author:  Forschungszentrum Informatik, OBST Projekt <obst@fzi.de>
Version:  3.4.3-11

octave Octave

Octave is a high level programming language. It’s designed for the solution of numeric problems. There is a command line interface supplied. Contents: octave-2.0.9

Copyright:  (c) 1992, 1993, 1994, 1995 John W. Eaton
Author:  John W. Eaton <jwe@bevo.che.wisc.edu>
Version:  2.0.9-16

onyx 4GL-Engine

The main goal in the development of Onyx was to create a tool for the fast development of database applications. The main idea is that all application are build in a similar way. Internal tables are used to store the data, and masks are used to view and change the data Transactions define the connection of the internal data with the rest of the world.

Copyright:  (c) 1993 M.Koehne
Author:  Michael Kraehe <kraehe@bakunin.north.de>
Version:  1.29-14

pdksh Public Domain Korn Shell

A public domain Korn Shell clone. Contents: pdksh-5.2.13

Version:  5.2.13-10

pg datab Initialization Database for PostgreSQL

This package contains the initialization for the PostgreSQL database system. If this initialization is installed, the user ”postgres” is able to setup additional database users; cf. the documentation under /usr/doc/packages/postgres/html/index.html. Please, check the configuration file /var/lib/pgsql/data/pg hba.conf; the manual page pg hba.conf(5) explains the background! Please, install this package only during the first installation or while doing a PostgreSQL ”major number update”! If a database already exists, it will be saved unter /var/lib/pgsql/data.old; please, remove the directory /var/lib/pgsql/data.old, if you do not need it anymore.

Copyright:  (c) No copyright
Version:  6.3.2-1

pg ifa PostgreSQL Interfaces (base)

Client package with major interfaces:

- command line insterface psql

- C and C++ library (shared)

- ecpg and libecpg (shared)

- Tcl/Tk (shared)

- PgAccess (a Tcl/Tk application)

- most of the man pages A useful package if you want to setup a PostgreSQL client machine.

Copyright:  (c) 1994-1997 Regents of the University of California

Author:

Andrew Yu <andrew@CS.Berkeley.edu>, Jolly Chen <jolly@CS.Berkeley.edu> , Constantin Teodorescu <teo@flex.ro>

Version:  6.3.2-1

pg iface PostgreSQL Interfaces

In the meantime, there are a lot ”interfaces” available to access a PostgreSQL database system. These interfaces could be quite different: there are ordinary drivers which could be included in foreign programs (even Windows programs) like ODBC or JDBC, or modules which one will use in script languages (Tcl/Tk, Perl, Pyhton), or ready to use database browsers. Right now, this package contains:

1. A Perl5 module. 2. A JDBC driver (Java).

Copyright:  (c) Different: GPL, Artistic License, 1996-98 by Keith R. Davis, Mutiny Bay Software...

Author:

Peter T Mount <pmount@maidast.demon.co.uk>, Adrian Hall <adrian@hottub.org>, Edmund Mergl <E.Mergl@bawue.de>, Keith R. Davis <keidav@whidbey.com>

Version:  6.3.2-3

pico Small, easy to use editor.

The editor normally used by pine. You will find the Sources in pine.spm.

Version:  3.96-15

postgres PostgreSQL - the Database

DBMS (DataBase Management System) PostgreSQL is the successor of the relational database system INGRES. It has been developed at Berkeley, University of California. The main advantages of the new system are:

- better support of complex objects.

- improved user-side adaption of data types, operators and access methods.

- Integration of active databases (alerts and triggers), conclusions including forward- and backward chaining.

- simplification of dbms code for restoration after a crash.

- Provision of a design which concatenates the advantages of optical disks and workstations which consist of many compound processors and customer-specific VLSI chips.

- Avoidance of too many changes according to relational model.

The query language for this system is SQL. This system offers you a wide range of additional features such as transparent history-system, embedding in your own C-routines, shell scripts and more. Contents: appgen-0.2-alpha dbbrowser-0.9.src postgresql-6.2.1

Copyright:  (c) 1994-1997 Regents of the University of California
Author:  Andrew Yu <andrew@CS.Berkeley.edu>, Jolly Chen <jolly@CS.Berkeley.edu>
Version:  6.3.2-1

ppa PBM to PPA converter

This software creates output using the PPA (printer performance architecture) protocol. This protocol is used by some HP ”Windows-only” printers (so called GDI), including the HP Deskjet 820C series, the HP DeskJet 720 series, and the HP DeskJet 1000 series. Currently only black-and-white printing is supported right now. A short description how to use ppa with apsfilter is given in the directory /usr/doc/packages/ppa/. For full information on the PPA protocol and this software, visit http://www.rpi.edu/˜ normat/technical/ppa/ .

Copyright:  (c) 1998 Tim Norman
Author:  Tim Norman <normat@rpi.edu>
Version:  0.8.4-2

quota Disk Quota System

Allows to assign disk quotas to the individual users of the system. The kernel must be recompiled with disk quota support enabled. Contents: quota-1.70 This program only runs if an kernel with quota support has been installed!

Copyright:  (c) 1980, 1990 Regents of the University of California.
Author:  Marco van Wieringen <mvw@planets.ow.nl>
Version:  1.70-5

recode THE Character Set Converter

This filter converts various character sets. Contents: recode-3.4.1

Copyright:  (c) 1987-94 Free Software Foundation, Inc.
Author:  Francois Pinard <pinard@iro.umontreal.ca>
Version:  3.4.1-13

rman rman

RosettaMan is a filter for UNIX manual pages needed by TkMan. It takes as input man pages formatted for a variety of UNIX flavors (not [tn]roff source) and produces as output a variety of file formats. Also included is http-rman which provides access to the man-pages via a normal web-browser. Contents: http-rman rman-3.0.2a9

Copyright:  (c) T.A. Phelps
Author:  T.A. Phelps <phelps@cs.Berkeley.EDU>
Version:  3.0.2a9-14

rpm2html RPM database info to HTML converter

from the original spec file: rpm2html tries to solve 2 big problems one face when grabbing a RPM package from a mirror on the net and trying to install it:

- it gives more information than just the filename before installing the package. - it tries to solve the dependancy problem by analyzing all the Provides and Requires of the set of RPMs. It shows the cross references by the way of hypertext links. Contents: rpm2html-0.85

Author:  Daniel Veillard <veillard@w3.org>
Version:  0.98-0

rpmfind Automatic RPM Package Update

rpmfind analyze the current state of the system, check a remote database using HTTP to lookup the packages providing the best affinity with your current software base and also list the extra packages that you need to install to solve the missing dependencies.

Rpmfind can also do the transfer once the list of package has been established. home page: http://rufus.w3.org/linux/rpm2html/rpmfind.html

Copyright:  (c) 1998 World Wide Web Consortium
Author:  Daniel Veillard <veillard@w3.org>
Version:  0.98-0

sc Spreadsheet

SC is a spreadsheet program. After invoking up sc a spreadsheet is shown which is arranged into rows and columns of cells. The spreadsheet is empty if the program is called without any parameters. Normally though, the filename parameter is read in by sc as a spreadsheet file. Each cell can contain a number, character string and/or expression (formula) which is based on numerical values, or character strings of other cells. Contents: sc-6.21

Author:  Jeff Buhrt , Mark Weiser , James Gosling
Version:  6.21-14

screen Multi Screen on a VT100/ANSI Terminal

With this program you can take advantage of the multi-tasking abilities of your Linux system by opening several sessions over one terminal. The sessions can also be detached and resumed from another login terminal. Documentation: man screen Contents: screen-3.7.3

Copyright:  (c) 1987-93 Oliver Laumann
Version:  3.7.3-13

scsiinfo Low level SCSI configuration

With this comfortable graphical user interface you may easily make ”disastrous” configurations to your SCSI subsystem. Please be very careful! You may destroy both your hardware and installed system by using this tool! Documentation: /usr/doc/packages/scsiinfo Contents: scsiinfo-1.7 Use with utmost care. Improper configuration can lead to severe hard and software damages!!!

Copyright:  (c) 1993, Yggdrasil Computing, Inc.
Author:  Michael Weller <eowmob@exp-math.uni-essen.de>, Eric Youngdale <eric@aib.com>
Version:  1.7-21

seejpeg An image viewer (libvga)

Seejpeg is Evan Harris’s JPEG viewer which utilizes svgalib and contains limited GIF, PPM and TARGA viewing support. The most useful feature of seejpeg is that it automatically decides on the ”best” video mode to use based on the image being displayed and the video card being used. If the image is larger than the screen the cursor keys can be used to scroll. Contents: seejpeg-1.6.1

Copyright:  (c) 1993, 1994 Evan Harris
Author:  Evan Harris <evan@cs.mu.oz.au>
Version:  1.6.1-23

sharutil GNU shar utilities

shar, uuencode, uudecode and others... Contents: sharutils-4.2c uudeview-0.5.13

Version:  4.2c-21

splitvt Split terminals

splitvt splits a VT100 compatible (Linux console, xterm) terminal‘s screen into two. Contents: splitvt-1.6.3

Copyright:  (c) 1993 Sam Lantinga
Author:  Sam Lantinga <slouken@cs.ucdavis.edu>
Version:  1.6.3-13

su1 su1

Program which allows selected users to run selected commands as root. Look at the config file! (/etc/su1.priv)

Copyright:  (c) GPL
Author:  David I. Bell <dbell@canb.auug.org.au>
Version:  4.2-1

sudo sudo

’sudo’ is a command that allows users to execute some commands as root. The /etc/sudoers file (edited with ’visudo’) specifies which users have access to sudo and which commands they can run. ’sudo’ logs all its activities to syslogd, so the system administrator can keep an eye on things. Sudo asks for the password for initializing an check period of a given time N (where N is defined at installation and is set to 5 minutes by default). Contents: sudo.v1.5.2

Copyright:  (c) 1991 The Root Group, Inc.

Author:

Jeff Nieusma <nieusma@rootgroup.com>, David Hieb <davehieb@rootgroup.com>, Ian McCloghrie <ian@ucsd.edu>

Version:  1.5.2-13

svgatext Extended SVGA text modes

No more place on your screen in text mode? Huge letters on the screen? You can change this by calling SVGATextMode which extends a SVGA text display up to 160x60 characters on good SVGA cards. Before use you have to configure the program by editing the file /etc/TextConfig to the capabilities of your graphics hardware. Furthermore you may create your own modelines for /etc/TextConfig with SVGATextCalc. Broad documentation can be found under /usr/doc/packages/svgatext and the respective man pages. Contents: SVGATextCalc-0.9 SVGATextMode-1.8 The author of SVGATextCalc would like to get a short mail message from you to tell him how you like the program.

Copyright:  (c) 1995,1996 Koen Gadeyne, 1996 Marco Paganini
Author:  Koen Gadeyne <kmg@barco.be>, Marco Paganini <paganini@ism.com.br>
Version:  1.8-10

taper Backup program taper

This is a tape backup and restore program that provides a friendly user interface to allow backing/restoring files to a tape drive. Alternatively, files can be backed up to archive files. Selecting files for backup and restore is very similar to the Mouseless Commander interface and allows easy traversal of directories. Recursively selected directories are supported. Incremental backup and automatic most recent restore are default settings. Both SCSI and floppy tape drives are supported. Information about archives is stored on the hard disk allowing information about archives to be obtained without having to have access to the tape. Contents: taper-6.7.4

Copyright:  (c) 1995 Yusuf Nagree
Author:  Yusuf Nagree <yusuf@mail.bogo.co.uk>
Version:  6.7.4-20

tcsh tcsh

tcsh is an enhanced but completely compatible version of the Berkeley UNIX C shell, csh(1). It is a command language interpreter usable both as an interactive login shell and a shell script command processor. It includes a command-line editor, programmable word completion, spelling correction, a history mechanism, job control, and a C-like syntax. Contents: tcsh-6.07.02

Copyright:  (c) 1980, 1991 The Regents of the University of California
Author:  Christos Zoulas <christos@deshaw.com>, Scott Krotz <krotz@mot.com>
Version:  6.07.02-12

teapot A three dimensional spreadsheet

Teapot is a spread sheet program for UNIX, which uses modern concepts besides being portable and extensible, for instance three dimensional tables and iterative expressions.

It provides various output formats e.g. LaTeX and HTML. There are executables in english, dutch and german available. They are started by typing ”teapot en”, ”teapot nl” or ”teapot de”. Just typing ”teapot” starts the english version. Contents: teapot-1.0

Copyright:  (c) Michael Haardt, 1995, 1996, 1997
Author:  Michael Haardt <michael@cantor.informatik.rwth-aachen.de>
Version:  1.0-5

unarj Pack program

Allows for unpacking of .arj files, which are widely spread in the DOS world. It is included for compatibility reasons to extract packages from DOS. It should not be used for compressing packages under Linux; use gzip which is the standard, instead. Contents: unarj

Copyright:  (c) 1991-93 Robert K Jung
Author:  Robert Jung <robjung@world.std.com>
Version:  2.41a-13

units units conversion utility

Copyright:  (c) 1989, 90, 91, 92, 93, 94 Free Software Foundation, Inc.
Author:  Adrian Mariano <adrian@cam.cornell.edu>
Version:  1.53-13

unzip Decompresser

UnZip is an extraction utility for archives compressed in .zip format (known as ”zip files”). Although highly compatible both with PKWARE’s PKZIP and PKUNZIP utilities for MS-DOS and with Info-ZIP’s own Zip program, our primary objectives have been portability and non-MSDOS functionality. Contents: unzip-5.32

Version:  5.32-8

wdiff display word differences between text files

wdiff compares two files, finding which words have been deleted or added to old file to get new file. A word is anything between whitespace. Xwdiff is an handy x frontend, based on Tcl/Tk.

Copyright:  (c) 1992 FSF
Author:  Francois Pinard <pinard@iro.umontreal.ca>, Hans Lermen <lermen@fgan.de>
Version:  0.5-2

xdelta Binary delta generator and RCS replacement library

Xdelta is a binary delta generator and RCS replacement library. The delta algorithm is faster than GNU diff for text files and works on binary files too. Contents: xdelta-0.14

Copyright:  (c) 1997-98 Josh MacDonald, Paul Eggert
Author:  Josh MacDonald <jmacd@cs.berkeley.edu>
Version:  0.19-2

yamm Yet Another Micro Monitor

similar to the program ”top”. Contents: yamm-2.5.6

Copyright:  (c) 1994, 1995 Riccardo Facchetti
Author:  Riccardo Facchetti <riccardo@cdc8g5.cdc.polimi.it>
Version:  2.5.6-19

ytree YTree - A filemanager like XTree

A (DOS) XTree-clone An example configuration file can be found as /usr/doc/packages/ytree/ytree.conf. You may copy this file to .ytree in your home directory. Contents: ytree-1.50

Copyright:  (c) Werner Bregulla
Author:  Werner Bregulla <werner@frolix.han.de>
Version:  1.63-3

z88 z88

Finite Elements programme with several modules: The FE processor reads the general structure data and the restrictions. Basicly, the Z88 entry files can be generated by a CAD convertor Z88X, by an editor or a text processor system or with a mixed procedure. Z88X outputs the processed structure data and restrictions, calculates the stiffness matrices of the elements, compiles the stiffness matrix of the entire system, scales the equation system, solves the system with an in situ Cholesky method and displays the displacements in ASCII format. This is why there are modules for calculating the stress, for the node strength and a net generator. You can graphically display and plot the net generator’s data. Contents: z88

Copyright:  (c) Dr. Frank Rieg
Author:  <dr.frank.rieg@t-online.de>
Version:  98.1.24-5

zip compression program

Zip 2.0.1 is a compression and file packaging utility. It is compatible with PKZIP 2.04g (Phil Katz ZIP) for MSDOS systems. Contents: zip

Version:  2.0.1-13

ziptool Tool for the Iomega ZIP and JAZ drives

With this tool, configuration of the Iomega ZIP and JAZ drives is possible. Automatic eject and write protection (password too) may be configured. Contents: ziptool

Copyright:  (c) 1997 Mirko Kraft, Grant R. Guenther, Itai Nahson
Author:  Mirko Kraft <Mirko.Kraft@ixos.de>, Grant R. Guenther <grant@torque.net>
Version:  1.1b-18

zoo Pack program

Zoo is a packer based on the Lempel-Ziv algorithm. Lots of files on DOS/AmigaDOS and TOS systems used this packer for their archives. The compression rate of gzip are not reached at all. Thus zoo should only be used for decompressing old archives. Contents: zoo

Copyright:  (c) 1991 Rahul Dhesi

Author:

Ian Phillipps <igp@camcon.co.uk>, J. Brian Waters <jbwaters@bsu-cs.bsu.edu>, Paul Homchick <rutgers!cgh!paul>, Mark Alexander <amdahl!drivax!alexande>, Randal L. Barnes <rlb@skyler.mavd.honeywell.com>

Version:  2.10-13

zsh zsh version

Zsh is a UNIX command interpreter (shell) which of the standard shells most resembles the Korn shell (ksh), although it is not completely compatible. It includes enhancements of many types, notably in the command-line editor, options for customizing its behavior, filename globbing, features to make C-shell (csh) users feel more at home and extra features drawn from tcsh (another ‘custom’ shell). Zsh was written by Paul Falstad. Contents: zsh-3.0.5

Author:  Paul Falstad
Version:  3.0.5-11

gs lib Necessary files for running gs (ghostscript)

This package contains all startup files for calling gs (ghostscript). You will find some useful documentations about few printer drivers in the directory: /usr/share/ghostscript/4.03/doc/ Please read the files use.txt, devices.txt and for the experimental driver cdj850 readme.hp8. Currently there is the newer version 5.10 of ghostscript available. Due to a change in the licence of this version we are not allowed to put this version onto this cdrom. Nevertheless this version you will find this version on our FTP server.

BEO: Extreme Linux (Beowulf)

Contents: gs4.03 hpdj-2.4.tar.gz hp8xxs13.zip gdevhl7x0.c.gz gdev820c-0.1.3.tar.gz

Copyright:  (c) 1989-96 Aladdin Enterprises.
Author:  Peter Deutsch <ghost@aladdin.com>
Version:  5.10-0

0.1.3 BEO: Extreme Linux (Beowulf)

bwatch A cluster performance monitor

bWatch is short for Beowulf Watch. It is a Tcl/Tk script designed to watch the load and memory usage on all of the nodes in a beowulf cluster. bWatch does not require root access, and can be run by any user on the system as long they can execute rsh to all other machines. An example configuration file can be found at /usr/doc/packages/bwatch/bWatchrc.tcl.sample.

Copyright:  (c) GPL
Author:  Jacek Radajewski <jacek@usq.edu.au>
Version:  1.0.2-0

lam LAM (Local Area Multicomputer)

LAM (Local Area Multicomputer) is an MPI programming environment and development system for heterogeneous computers on a network. With LAM, a dedicated cluster or an existing network computing infrastructure can act as one parallel computer solving one problem. LAM features extensive debugging support in the application development cycle and peak performance for production applications. LAM features a full implementation of the MPI (Message Passing Interface) communication standard. The documentation will be installed in /usr/doc/packages/lam/ as lam61.doc.ps and in /usr/doc/packages/lam/mpi-quick-tut/ . The WWW home page of LAM is http://www.mpi.nd.edu/lam/

Copyright:  (c) 1996 The Ohio State University
Author:  The LAM Team <llamas@mpi.nd.edu>
Version:  6.1-3

pvm Parallel Virtual Machine

PVM is a software system that enables a collection of heterogeneous computers to be used as a coherent and flexible concurrent computational resource. The individual computers may be shared- or local-memory multiprocessors, vector supercomputers, specialized graphics engines, or scalar workstations, that may be interconnected by a variety of networks, such as ethernet, FDDI. User programs written in C, C++ or Fortran access PVM through library routines. After installation you find in /usr/doc/packages/pvm/ the documentation as PostScript file pvmbook.ps. Furthermore some examples are packed together in two tar archives. Those archives should be extracted into your HOME directory which leads to ˜ /pvm3/examples/ or ˜ /pvm3/gexamples/ in your HOME directory. The call ‘aimk all’ (see manual page aimk(1)) e.g. in ˜ /pvm3/examples/ compiles the examples for the PVM system. The PVM web home page is at http://www.epm.ornl.gov/pvm/pvm home.html .

Copyright:  (c) University of Tennessee, Oak Ridge National Laboratory, Emory University

Author:

J. J. Dongarra <pvm@msr.epm.ornl.gov>, G. E. Fagg <pvm@msr.epm.ornl.gov>, G. A. Geist <pvm@msr.epm.ornl.gov>, J. A. Kohl <pvm@msr.epm.ornl.gov>, R. J. Manchek <pvm@msr.epm.ornl.gov>, P. Mucci <pvm@msr.epm.ornl.gov>, P. M. Papadopoulos <pvm@msr.epm.ornl.gov>, S. L. Scott <pvm@msr.epm.ornl.gov>, V. S. Sunderam <pvm@msr.epm.ornl.gov>

Version:  3.4beta6-8

xpvm A graphical console and monitor for PVM

XPVM is a graphical console and monitor for PVM. It provides a graphical interface to the PVM console commands and information, along with several animated views to monitor the execution of PVM programs. These views provide information about the interactions among tasks in a parallel PVM program, to assist in debugging and performance tuning.

Copyright:  (c) (C) 1994 Oak Ridge National Laboratory, Oak Ridge TN.
Author:  James Arthur Kohl <kohl@msr.epm.ornl.gov>, G. A. Geist
Version:  1.2.5-0

0.1.4 D: Development (C, C++, Lisp, etc.)

autoconf An m4 macro package

This package is required for the configuration of many GNU source packages.

Contents: autoconf-2.12

Copyright:  (c) 1994 Free Software Foundation, Inc.
Version:  2.12-13

automake Tool for automatically generating GNU style Makefile.in’s

Automake is a tool for automatically generating ”Makefile.in”s from files called ”Makefile.am”. ”Makefile.am” is basically a series of ”make” macro definitions (with rules being thrown in occasionally). The generated ”Makefile.in”s are compatible to the GNU Makefile standards.

Contents: automake-1.2

Copyright:  (c) 1994, 1995, 1996 Free Software Foundation, Inc.
Author:  Tom Tromey <tromey@cygnus.com>
Version:  1.3-2

beg Back End Generator

This package contains only the BEG executable to generate Source code for the GMD MOCKA Compiler. The Sources are not free.

For further information: /usr/doc/beg-1.75/

Contents: beg-1.75

Copyright:  (c) 1993 GMD Karlruhe, Germany
Author:  Helmut Emmelmann <emmel@ipd.info.uni-karlsruhe.de>
Version:  1.75-10

bin86 8086 assembler and linker

Needed to compile the kernel.

Contents: bin86-0.4

Author:

Bruce Evans <bde@runx.oz.au> , Linus Torvalds <Linus.Torvalds@cs.helsinki.fi> , H.J. Lu <hjl@nynexst.com>

Version:  0.4-13

binutils GNU binutils

C compiler utilities: ar, gprof, ld, libconvert, nm, objdump, ranlib, size, strace, strip.

These utilities are needed whenever you want to compile a program or a kernel.

This version has been compiled by S.u.S.E. with the corresponding C library. They are only designed for i486-linux (ELF) and i486-linuxaout (a.out) but not for cross-compiling to other architectures such as Sparc, Alpha, and m68k.

Contents: binutils-2.8.1

D: Development (C, C++, Lisp, etc.)

Copyright:  (c) 1987-94 Free Software Foundation, Inc.

Author:

H.J. Lu <hjl@nynexst.com>, Ken Raeburn <raeburn@cygnus.com>, Ian Lance Taylor <ian@cygnus.com>, J.T. Conklin <jtc@cygnus.com>, Stan Shebs <shebs@cygnus.com>, Steve Chamberlain <sac@cygnus.com>, David J. Mackenzie <djm@cygnus.com>, Jeffrey A. Law <law@cygnus.com>, Sean Eric Fagan <sef@cygnus.com>, Jeffrey Osier <jeffrey@cygnus.com>, Jeffrey Wheat <cassidy@cygnus.com>, Mark Eichin <eichin@cygnus.com>, Roland H. Pesch <pesch@wrs.com>, Per Bothner <bothner@cygnus.com>, David D ”Zoo” Zuhn <zoo@cygnus.com>, Rob Savoye <rob@cygnus.com>, Jim Kingdon <kingdon@cygnus.com>, Mike Werner <mtw@cygnus.com>, Fred Fish <fnf@cygnus.com>, John Gilmore <gnu@cygnus.com>, Brendan Kehoe <brendan@cygnus.com>, D. V. Henkel-Wallace <gumby@cygnus.com>, K. Richard Pixley <rich@cygnus.com>, Bill Cox <bill@cygnus.com>, Peter Schauer <pes@regent.e-technik.tu-muenchen.de>, Eric Youngdale <eric@aib.com>, Richard Earnshaw <rwe@pegasus.esprit.ec.org>, Doug Evans <dje@cygnus.com>, Ian Dall <dall@hfrd.dsto.gov.au>, Jason Merrill <jason@cygnus.com>, Jim Wilson <wilson@cygnus.com>, Kung Hsu <kung@cygnus.com>, Michael Meissne <meissner@osf.org>, Michael Tiemann <tiemann@cygnus.com>, Richard Stallman <rms@gnu.ai.mit.edu>, Stu Grossman <grossman@cygnus.com>, Torbjorn Granlund <tege@cygnus.com>

Version:  2.9.1.0.4-2

bison GNU bison parser generator

Bison is a parser generator in the style of yacc(1).

Contents: bison

Copyright:  (c) 1984-93 Free Software Foundation, Inc.

Author:

David J. MacKenzie <djm@gnu.ai.mit.edu>, Noah Friedman <friedman@ai.mit.edu>, Richard Stallman <rms@ai.mit.edu>

Version:  1.25-13

bwbasic There is Basic under Linux, too

A Basic interpreter. Example programs are in /usr/local/lib/bwbtest .

Documentation: /usr/doc/misc/bwbasic-2.10.*

Contents: bwbasic-2.10

Copyright:  (c) 1993, Ted A. Campbell
Author:  Ted A. Campbell <tcamp@delphi.com>
Version:  2.10-13

clisp A Common-Lisp interpreter

Common Lisp is a high-level, all-purpose programming language. CLISP is an implementation of Common Lisp that closely follows the book ”Common Lisp - The Language” by Guy L. Steele Jr. This package includes an interactive programming environment with interpreter, compiler and debugger. This environment will be started by the command

clisp

The documentation of CLISP is placed in the directories

/usr/doc/packages/clisp/

/usr/doc/packages/clisp/doc/

As well as the conventional CLISP, the package also includes CLX, an extension of CLISP for the X Window System. Of course X11 has to be installed before the command

clx

is called. The description of this CLX version (new-clx) is placed in

/usr/doc/packages/clisp/clx/

with the file README. The subdirectory

/usr/doc/packages/clisp/clx/demos/

contains two nice applications.

Contents: clisp-1997-12-06

Copyright:  (c) 1992,93 Bruno Haible und Michael Stoll, 1994-97 Bruno Haible und Marcus Daniels

Author:

Bruno Haible <haible@ma2s2.mathematik.uni-karlsruhe.de>, Michael Stoll <michael@rhein.iam.uni-bonn.de>, Marcus Daniels <marcus@sysc.pdx.edu>, Gilbert Baumann <gilbert@ma2s2.mathematik.uni-karlsruhe.de>

Version:  97.12.06-9

cvs Concurrent Versions System

CVS is a frontend to the Revision Control System, which is included in the standard Linux distribution. pcl-cvs, an Emacs frontend to CVS, is also included. Contents: cvs-1.9.24

Copyright:  (c) 1989-92 Brian Berliner, Jeff Polk
Author:  Brian Berliner <berliner@sun.com>, Jeff Polk <polk@bsdi.com>
Version:  1.9.28-1

des Encryption routines

Version:  4.01-11

efence ElectricFence

Tool for debugging your C-programs (malloc debugger). Contents: ElectricFence-2.0.5

Version:  2.0.5-13

f2c A Fortran-77 to C translator

Contains libf2c.so.0.11, ported by Rob Hooft. This package uses an ’f77’ script that hides the C translation process from the user. Contents: f2c libf2c-0.11

Copyright:  (c) 1990-95 AT&T Bell Laboratories and Bellcore.

Author:

David Gay <dmg@bell-labs.com>, Stu Feldman , Mark Maimone , Norm Schryer , Rob Hooft (maintainer for Linux) <Hooft@EMBL-Heidelberg.DE>

Version:  0.11-15

f2caout a.out version of libf2c

Copyright:  (c) 1990-95 AT&T Bell Laboratories and Bellcore.

Author:

David Gay <dmg@bell-labs.com>, Stu Feldman, Mark Maimone , Norm Schryer, Rob Hooft (maintainer for Linux) <Hooft@EMBL-Heidelberg.De>

Version:  0.10-1

flex flex - fast lexical analyzer generator

flex is a tool for generating scanners: programs which recognize lexical patterns in text. Contents: flex-2.5.4

Copyright:  (c) 1990 The Regents of the University of California.
Author:  Vern Paxson <vern@ee.lbl.gov>
Version:  2.5.4-13

fpk Free Pascal Compiler

FPK is a 32-bit compiler for the i386 processor. It is designed to be, as much as possible, source compatible with Turbo Pascal 7.0 and Delphi II, but it also enhances these languages with elements like function overloading.

Copyright:  (c) 1997 Michael Van Canneyt
Author:  Michael Van Canneyt <michael@tfdec1.fys.kuleuven.ac.be>
Version:  0.99.0-2

g77 GNU Fortran Compiler

This version of the GNU Fortran compiler needs GNU gcc 2.7.2.1. The binaries of the C compiler must be under /usr/lib/gcc-lib/i486-linux/2.7.2.1.

Copyright:  (c) 1995 Free Software Foundation, Inc.
Author:  James Craig Burley <burley@gnu.ai.mit.edu>
Version:  0.5.19.1-1

gcc The GNU C compiler and support files

Compiled with 486 optimization, to maximize performance on 486 processors, but works fine with 386.

NOTE: Be sure to install at least the following packages besides this one, or you won’t be able to compile: binutils, include, libc, and the third part of the kernel source. (the include files) Original Version from FSF. Compiled with support for g77 and gnat.

Contents: g77-0.5.19.1 gcc-2.7.2.1 gnat-3.09-src gpc-2.0

Copyright:  (c) 1987-95 Free Software Foundation, Inc.

Author:

Adam Fedor <fedor@colorado.edu>, Alan Modra <alan@spri.levels.unisa.edu.au>, Allen Briggs <briggs@rrinc.com>, Andreas Schwab <schwab@issan.informatik.unidortmund.de>, Andrew Cagney <cagney@highland.com.au>, Andrew McCallum <mccallum@leopard.cs.rochester.edu>, Arne H. Juul <arnej@idt.unit.no>, Craig Burley <burley@gnu.ai.mit.edu>, DJ Delorie <dj@delorie.co)m, David Edelsohn <edelsohn@mhpcc.edu>, Doug Evans <dje@cygnus.com>, Douglas Rupp <drupp@cs.washington.edu>, Fergus Henderson <fjh@cs.mu.oz.au>, Gran Uddeborg <uddeborg@carmen.se>, Ian Lance Taylor <ian@cygnus.com>, J”orn Rennecke <amylaar@meolyon.hanse.de>, J.T. Conklin <jtc@cygnus.com>, Jason Merrill <jason@cygnus.com>, Jeffrey A. Law <law@cygnus.com>, Jim Meyering <meyering@na-net.ornl.gov>, Jim Wilson <wilson@cygnus.com>, John F. Carr <jfc@mit.edu>, Ken Raeburn <raeburn@wombat.gnu.ai.mit.edu>, Lee Iverson <leei@ai.sri.com>, Linus Torvalds <Linus.Torvalds@cs.Helsinki.FI>, Michael Gschwind <mike@lanai.vlsivie.tuwien.ac.at>, Michael Meissner <meissner@cygnus.com>, Mike Stump <mrs@cygnus.com>, Niklas Hallqvist <niklas@appli.se>, Oliver Kellogg <oliver.kellogg@space.otn.dasa.de>, Pat Rankin <rankin@eql.caltech.edu>, Paul Eggert <eggert@twinsun.com>, Per Bothner <bothner@cygnus.com>, Peter Flass <flass@lbdrscs.bitnet>, Philippe De Muyter <phdm@info.ucl.ac.be>, Richard Earnshaw <rearnsha@armltd.co.uk>, Richard Kenner <kenner@vlsi1.ultra.nyu.edu>, Richard Stallman <rms@mole.gnu.ai.mit.edu>, Roland McGrath <roland@churchy.gnu.ai.mit.edu>, Stan Cox <coxs@dgrtp.dg.com>, Stephen L Moshier <moshier@world.std.com>, Ted Lemon <mellon@toccata.fugue.com>, Tom Quiggle <quiggle@lovelace.engr.sgi.com>, Torbjorn Granlund <tege@cygnus.com>, Uwe Seimet <seimet@chemie.uni-kl.de>

Version:  2.7.2.1-6

gcl GNU Common Lisp

A Common Lisp interpreter with a Tcl/Tk interface.

Contents: gcl-2.2.2

Copyright:  (c) 1994 M. Hagiya, W. Schelter, T. Yuasa,
Author:  Bill Schelter <wfs@ma.utexas.edu>
Version:  2.2.2-18

gdb The GNU debugger

You can use GDB to debug programs written in C, C++, and Modula-2. GNU Fortran 77 is also partially included.

Contents: gdb-4.16

Copyright:  (c) 1986-95 Free Software Foundation, Inc.

Author:

Doug Evans <dje@cygnus.com>, Fred Fish <fnf@cygnus.com>, Ian Lance Taylor <ian@cygnus.com>, J.T. Conklin <jtc@cygnus.com>, Jason Molenda <crash@cygnus.com>, Jeffrey A. Law <law@cygnus.com>, Jim Kingdon <kingdon@cygnus.com>, Kung Hsu <kung@cygnus.com>, Michael Meissner <meissner@cygnus.com>, Per Bothner <bothner@cygnus.com>, Peter Schauer <pes@regent.e-technik.tu-muenchen.de>, Rob Savoye <rob@cygnus.com>, Stan Shebs <shebs@cygnus.com>, Stan Shebs <shebs@cygnus.com>, Steve Chamberlain <sac@cygnus.com>, Stu Grossman <grossman@cygnus.com>, Steve Chamberlain <sac@cygnus.com>

Version:  4.16-15

gdlib Gif-Manipulating Library

This is the gd gif-manipulating library, version 1.2. It was created to allow graphs, charts and the like to be generated on the fly for use on the World wide Web, but is useful for any application in which custom .GIFs are useful. It is not a paint program; it is a library. If you are looking for a paint program for the X Window System, check out David Koblas ”xpaint”, available by anonymous FTP from ftp.netcom.com in pub/ko/koblas. (I’m just a fan of Dave’s code.)

Author:  Tom Boutell <boutell@boutell.com>
Version:  1.2-7

gettext Tools for National Language Support (NLS)

This Library stands under the GPL License, not under LGPL. So please remember to read /usr/doc/packages/gettext/COPYING and remember that you have to publish the sources of a program linked against this library on request, if you distribute your program. Contents: gettext-0.10.32

Copyright:  (c) GPL
Version:  0.10.32-14

gforth GNU Forth

Copyright:  (c) 1997 Free Software Foundation, Inc.

Author:

Anton  Ertl  <anton@mips.complang.tuwien.ac.at>,  Bernd  Paysan

<paysan@informatik.tu-muenchen.de>, Jens Wilke <wilke@jwdt.com>

Version:  0.3.0-9

global common source code tag system

GLOBAL is a common source code tag system for C and Yacc. You can locate the specified function in C source files and move there easily. It is useful to hack a large project containing many subdirectories, many main() functions like MH, X or Linux kernel.

Copyright:  (c) 1998 Shigio Yamaguchi
Author:  Shigio Yamaguchi <shigio@wafu.netgate.net>
Version:  2.24-3

gmp GNU MP library

Library for calculating huge numbers (integer and floating point). Contents: gmp-2.0.2

Copyright:  (c) 1992-96 Free Software Foundation, Inc.
Author:  Torbjorn Granlund <tege@zevs.sics.se>
Version:  2.0.2-13

gnat GNU Ada Translator

The ”GNU Ada Translator”, an Ada83 and Ada9X front end to gcc-2.7.2. GCC 2.7.2 is required if you want to use this version of Ada.

Version:  3.09-2

gnats GNATS - Problem Report Management System

GNATS is a bug-tracking tool designed for use at a central support site. Software users who experience problems use electronic mail to communicate these problems to the the maintainers of that software; GNATS partially automates the tracking of these problems by: organizing problem reports into a database and notifying responsible parties of suspected bugs; allowing support personnel and their managers to edit, query and report on accumulated bugs; and providing a reliable archive of problems with a given program and a history of the life of the program by preserving its reported problems and their subsequent solutions. GNATS offers many of the same features offered by more generic databases. You can query and edit existing problem reports (PRs) as well as obtain reports on groups of PRs. The database itself is simply an ordered repository for problem reports; each PR receives a unique, incremental PRnumber, which identifies it throughout its lifetime. Contents: gnats-3.102-beta

Copyright:  (c) GPL

Author:

Brendan  Kehoe  <brendan@lisa.cygnus.com>,  Ian  Lance  Taylor

<ian@cygnus.com>, Jason Molenda <crash@godzilla.cygnus.co.jp>, Doug Evans <dje@canuck.cygnus.com>

Version:  3.102beta-15

gobjc GNU objective C Compiler

The source code is available in the gcc s package.

Version:  2.7.2.1-2

gpc GNU Pascal Compiler

A Pascal compiler based on the GCC 2.7.2.1 compiler.

Author:  Jukka Virtanen <jtv@hut.fi>, Peter Gerwinski <peter.gerwinski@uni-essen.de>
Version:  2.0-2

gpp GNU C++ Compiler

You can find the source code for gpp in the gcc s package.

Version:  2.7.2.1-3

guavac A free compiler for the Java language

Guavac is a free compiler for the Java language. Contents: guavac-0.3.1

Copyright:  (c) 1989, 1991 Free Software Foundation, Inc.
Author:  Effective Edge <info@eecs.com>
Version:  0.3.1-12

guile GNU’s Ubiquitous Intelligent Language for Extension

This is Guile, a portable, embeddable Scheme implementation written in C. Guile provides a machine independent execution platform that can be linked in as a library when building extensible programs. Contents: guile-1.2

Copyright:  (c) GLP

Author:

Jim Blandy <jimb@floss.red-bean.com>, George Carrette, Aubrey Jaffer, Gary Houston <ghouston@actrix.gen.nz>, Tom Lord, Anthony Green, Mikael Djurfeldt <mdj@mdj.nada.kth.se>, Mark Galassi <rosalia@papageno.lanl.gov>, Tim Pierce <twp@twp.tezcat.com>, Marius Vollmer <mvo@zagadka.ping.de>, Tom Tromey <tromey@cygnus.com>

Version:  1.2-14

icmake A Maker for C-Fans

A Maker using a C-like syntax. I like that :-) Documentation can be found in /usr/doc/misc/icmake-6.17.ps; examples are in /usr/local/lib/icmake-6.17-examples/ Contents: icmake-6.17

Copyright:  (c) ICCE 1992-1994
Author:  Frank B. Brokken <F.B.Brokken@icce.rug.nl>, Karel Kubat <K.Kubat@icce.rug.nl>
Version:  6.17-14

indent formats C source code

Version:  1.9.1-13

java Java Developers Kit v1.1.3

With this package you are able to create your own platform-independent JAVA applications. You can compile your it with the Java compiler and start it with the Java interpreter or the Appletviewer. The authors of this package stipulated that the sources should not cross hands more than two times. For this reason we are unable to provide a source package for this program. To get the sources, look at: http://www.blackdown.org/java-linux/Mirrors.cgi For more information please refer to /usr/doc/packages/java. Contents: linux-jdk.1.1.3-v2

Copyright:  (c) 1994, 1995, 1996, 1997 Sun Microsystems, Inc.
Author:  Steve Byrne <sbb@gnu.ai.mit.edu>
Version:  1.1.3v2-13

javarunt Java Runtime Environment v1.1.3-v2

Copyright:  (c) 1997 Sun Microsystems, Inc.
Author:  Steve Byrne <sbb@gnu.ai.mit.edu>
Version:  1.1.3v2-10

jdk115 Java Developers Kit 1.1.5

The Java Developers Kit version 1.1.5 Contents: jdk1.1.5

Copyright:  (c) 1994, 1995, 1996, 1997 Sun Microsystems, Inc.
Author:  Steve Byrne <sbb@gnu.ai.mit.edu>
Version:  1.1.5-5

jdk116 Java Developers Kit 1.1.6

The Java Developers Kit version 1.1.6

Copyright:  (c) 1994, 1995, 1996, 1997 Sun Microsystems, Inc.
Author:  Steve Byrne <sbb@gnu.ai.mit.edu>
Version:  1.1.6v2-3

jresep Java Runtime Environment 1.1.3v2

The separated version of JRE. There are two components which make up JRE: the basic VM runtime, and the internationalization support ( /usr/lib/rt and /usr/lib/i18n ). Please read /usr/doc/packages/jresep/* Contents: linux-rt.1.1.3-v2 rt

Copyright:  (c) 1997 Sun Microsystems, Inc.
Author:  Steve Byrne <sbb@gnu.ai.mit.edu>
Version:  1.1.3v2-10

kaffe A virtual machine to run Java(tm) code

Kaffe is a virtual machine design to execute Java byte code. Unlike other virtual machines available, this machine performs ”just-in-time” code conversion from the abstract code to the host machine’s native code. Please take a look at /usr/doc/packages/kaffe/* Kaffe is currently very beta!!!

Copyright:  (c) 1989, 1991 Free Software Foundation, Inc.
Author:  Tim Wilkinson <tim@transvirtual.com>
Version:  1.0.b1-5

lclint LCLint - a tool for statically checking C programs

With minimal effort, LCLint can be used as a better lint. If additional effort is invested adding annotations to programs, LCLint can perform stronger checks than can be done by any standard lint. Contents: lclint-2.2a

Copyright:  (c) Massachusetts Institute of Technology 1994, 1995, 1996.
Author:  David Evans <evs@larch.lcs.mit.edu>
Version:  2.2a-14

leda LEDA 3.5.2 a library of efficient data types and algorithms

LEDA is a library of the data types and algorithms of combinatorial computing. The main features are: LEDA provides a sizable collection of data types and algorithms in a form which allows them to be used by non-experts. In the current version, this collection includes most of the data types and algorithms described in the text books of the area. Any user of LEDA needs a license. This license is available free of charge for academic research and teaching. Any commercial use of this software requires a license which is distributed by LEDA Software GmbH. Any use by a firm, a commercial company or a government institution is considered to be commercial use. LEDA Software GmbH Postfach 151101 66041 Saarbruecken Germany email:leda@mpi-sb.mpg.de fax: +49 681 842502 For a research license please contact ftp://ftp.mpi-sb.mpg.de/pub/LEDA/www/ledaavailable.html Contents: leda-3.5.2 Any user of LEDA needs a license. This license is available free of charge for academic research and teaching. Any commercial use of this software requires a license which is distributed by LEDA Software GmbH. Any use by a firm, a commercial company or a government institution is considered to be commercial use.

LEDA Software GmbH Postfach 151101 66041 Saarbruecken Germany email:leda@mpi-sb.mpg.de fax: +49 681 842502

For a research license please contact ftp://ftp.mpi-sb.mpg.de/pub/LEDA/www/ledaavailable.html

Copyright:  (c) LEDA Software GmbH
Author:  LEDA Software GmbH <leda@mpi-sb.mpg.de>
Version:  3.5.2-12

libc Libraries for the C compiler

These libraries are needed to compile C code.

Contents: libc-5.4.44 libpthreads

Copyright:  (c) 1981-95 Regents of the University of California. , Free Software Foundation, Inc.

Author:

H.J. Lu <hjl@gnu.ai.mit.edu>, Alan Cox <Alan.Cox@linux.org>, Adam J. Richter <adam@yggdrasil.com>, Alan Modra <alan@spri.levels.unisa.edu.au>, Andreas Schwab <schwab@issan.informatik.uni-dortmund.de>, Andrew Walker <andy@keo.kvaerner.no>, Andries Brouwer <aeb@cwi.nl>, Charles Lopes <tjarls@infm.ulst.ac.uk>, Chip Salzenberg <chip@fin.uucp>, Christoph Kern <kern@informatik.uni-wuerzburg.de>, Corey Minyard <minyard@wf-rch.cirr.com>, Craig Smith <craig@rezrov.natinst.com>, Daniel Quinlan <quinlan@yggdrasil.com>, Delman Lee <delman@mipg.upenn.edu>, Dominik Kubla <dominik.kubla@unimainz.de>, Dsouza <m.dsouza@mrc-applied-psychology.cambridge.ac.uk>, Eric Youngdale <eric@aib.com>, Peter Eriksson <pen@signum.se>, Florian La Roche <florian@suse.de>, Fred N. Van Kempen <waltje@uwalt.nl.mugnet.org>, Gerhard Koerting <g.koerting@koerting.techem.ruhr-uni-bochum.de>, H. Peter Anvin <hpa@yggdrasil.com>, Harald Koenig <koenig@tat.physik.unituebingen.de>, Igor Sharfmesser <igor@irbis.alma-ata.su>, J. Alan Eldridge <alane@wozzle.linet.org>, Jaakko Hyvatti <jaakko.hyvatti@www.fi>, James H. Cloos Jr. <cloos@rahul.net>, Jeschke <i6210109@ws.rz.tu-bs.de>, Jim Van Zandt <jrv@vanzandt.mv.com>, Jochen Hein <Hein@Informatik.TU-Clausthal.de>, Jochen Kiemes <jochen kiemes@s2.maus.de> , John F. Haugh Ii <jfh@rpp386>, Jonathan Kamens <jik@cam.ov.com>, Karl Hakimian <hakimian@eecs.wsu.edu>, Kaz Sasayama <kaz@lilia.iijnet.or.jp>, Marc Gingold <gingold@amoco.saclay.cea.fr>, Martin Kuck <martin@unifix.de>, Martin Pirker <pirker@eiunix.tuwien.ac.at>, Mathew G. Monroe <mmex@shadowland.pc.cc.cmu.edu>, Matthias Urlichs <urlichs@smurf.noris.de>, Matthias Urlichs <urlichs@test.smurf.noris.de>, Michael Meskes <meskes@debian.org>, Michael Riepe <Michael.Riepe@stud.unihannover.de>, Mike Kennedy <kmb1@gate.net>, Mike Meissner <meissner@cygnus.com>, Miquel van Smoorenburg <miquels@drinkel.ow.nl>, Mitchum Dsouza <m.dsouza@mrc-applied-psychology.cambridge.ac.uk>, Nick Holloway <nick.holloway@alfie.demon.co.uk>, Nickolay Saukh <nms@ussr.eu.net>, Nicolai Langfeldt <janl@ifi.uio.no>, Niibe Yutaka <gniibe@mri.co.jp>, Olaf Flebbe <o.flebbe@science-computing.uni-tuebingen.de>, Patrick D’Cruze <pdcruze@orac.iinet.com.au>, Paul Eggert <eggert@twinsun.com>, Per Bothner <bothner@cygnus.com>, Peter Himmler <himmler@sparcy.mayn.sub.de>, Philip Gladstone <philip.gladstone@mail.citicorp.com>, Remy Card <remy.card@masi.ibp.fr>, Rick Sladkey <jrs@world.std.com>, Robert Andrew Ryan <rr2b+@andrew.cmu.edu>, Roland Mcgrath <roland@gnu.ai.mit.edu>, Rolf Rossius <rossius@hrz.tu-chemnitz.de>, Ross Biro <biro@yggdrasil.com>, Sean Mcneil <sean@mcneil.com>, Stephen L Moshier <moshier@world.std.com>, Stephen Tweedie <sct@dcs.ed.ac.uk>, Steve Losen <losen@delphi.com>, Steve M. Robbins <steve@nyongwa.montreal.qc.ca>, Swen Thuemmler <swen@uni-paderborn.de>, Theodore Ts’O <tytso@mit.edu>, Thomas Pundt <pundtt@goedel.uni-muenster.de>, Ulrich Drepper <drepper@gnu.ai.mit.edu>, Uwe Ohse <uwe@tirka.gun.de>, Vadim V. Vlasov <vvv@ms3.inr.ac.ru>, Wes Cowley <wcowley@cftnet.com>

Version:  5.4.46-1

libd Profiling and Debugging versions of libc.

Version:  5.4.46-1

libgpp C++-library

Version:  2.7.2.1-1

libobjc libobjects

The library for the GNU Objective C compiler.

Contents: libobjects-0.1.19

Version:  0.1.19-14

linclude Include files for the kernel

The kernel include files aren’t only needed for rebuilding the kernel but are also required to compile system programs.

So that you don’t have to re-install the complete kernel tree each time you build a program, we’ve placed the include files in an extra package. If you want to rebuild the kernel please install the ’linux’ package as well.

Author:  Linus Torvalds <torvalds@transmeta.com>
Version:  2.0.35-1

linux Rest of the kernel source code.

This package contains the rest of the Linux Kernel source files.

Contents: linux-2.0.35

Author:

Linus Torvalds <torvalds@transmeta.com>,  ,  Alain L. Knaff

<Alain.Knaff@poboxes.com>, Alan Cox <alan@lxorguk.ukuu.org.uk>, Alessandro Rubini <rubini@ipvvis.unipv.it>, Andrew Haylett <ajh@primag.co.uk>, Andrew Tridgell <Andrew.Tridgell@anu.edu.au>, Andries Brouwer <aeb@cwi.nl>, Angelo Haritsis <ah@doc.ic.ac.uk>, Arindam Banerji <axb@cse.nd.edu>, Avery Pennarun <apenwarr@bond.net>, Axel Boldt <boldt@math.ucsb.edu>, Bas Laarhoven <bas@vimec.nl>, Bernhard Kaindl <bkaindl@netway.at>, Bill Bogstad <bogstad@pobox.com>, Bjorn Ekwall <bj0rn@blox.se>, Bob Frey <bobf@advansys.com>, Bradley McLean <brad@bradpc.gaylord.com>, Bruno Haible <haible@ma2s2.mathematik.uni-karlsruhe.de>, Chih-Jen Chang <chihjenc@scf.usc.edu>, Chris Smith <csmith@convex.com>, Corey Minyard <minyard@wf-rch.cirr.com>, Craig Metz <cmetz@inner.net>, Craig Small <csmall@triode.apana.org.au>, Daniel Quinlan <quinlan@pathname.com>, Danny ter Haar <dth@cistron.nl>, Dario Ballabio <dario@milano.europe.dg.com>, Darren Senn <sinster@darkwater.com>, David C. Niemi <niemi@erols.com>, David Davies <davies@wanton.lkg.dec.com>, David Gentzel <gentzel@telerama.lm.com>, David Mosberger-Tang <davidm@azstarnet.com>, David S. Miller <davem@caip.rutgers.edu>, Derek Atkins <warlord@MIT.EDU>, Dirk Hohndel <hohndel@aib.com>, Dirk Melchers <dirk@merlin.nbg.sub.org>, Dirk Verworner, Dmitry S. Gorodchanin <begemot@bgm.rosprint.net>, Donald Becker <becker@cesdis.gsfc.nasa.gov>, Doug Evans <dje@cygnus.com>, Drew Eckhardt <drew@PoohSticks.ORG>, Drew Sullivan <drew@ss.org>, Eberhard Moenkeberg <emoenke@gwdg.de>, Ed Carp <ecarp@netcom.com>, Eddie C. Dost <ecd@skynet.be>, Eric S. Raymond <esr@thyrsus.com>, Eric Schenk <Eric.Schenk@dna.lth.se>, Eric Youngdale <eric@aib.com>, Florian La Roche <rzsfl@rz.uni-sb.de>, Frank Xia <qx@math.columbia.edu>, Fred Baumgarten <dc6iq@insu1.etec.uni-karlsruhe.de>, Fred N. van Kempen <waltje@uwalt.nl.mugnet.org>, Frederic Potter <Frederic.Potter@masi.ibp.fr >, Fritz Elfert <fritz@wuemaus.franken.de>, G”unter Windau <gunter@mbfys.kun.nl>, Gerard Roudier <groudier@iplus.fr>, Gero Kuhlmann <gero@gkminix.han.de>, Gordon Chaffee <chaffee@plateau.cs.berkeley.edu>, Grant Guenther <grant@torque.net>, Greg Hankins <gregh@cc.gatech.edu>, Greg Page <greg@caldera.com>, Greg Wettstein <greg@wind.rmcc.com>, H. Peter Anvin <hpa@zytor.com>, H.J. Lu <hjl@gnu.ai.mit.edu>, Hamish Coleman <hamish@zot.apana.org.au>, Hamish Macdonald <hamish@border.ocunix.on.ca>, Hannu Savolainen <hannu@voxware.pp.fi>, Hans Lermen <lermen@elserv.ffm.fgan.de>, Hans-Joachim Widmaier <hjw@zvw.de>, Harald Koenig <koenig@tat.physik.unituebingen.de>, Harald Milz <hm@seneca.linux.de>, Heiko Eissfeldt <heiko@colossus.escape.de heiko@unifix.de>, Hennus Bergman <hennus@cybercomm.nl>, Henrik Storner <storner@osiris.ping.dk>, Ian A. Murdock <imurdock@gnu.ai.mit.edu>, Ian Jackson <iwj10@cus.cam.ac.uk>, Ian Kluft <ikluft@thunder.sbay.org>, Jacques Gelinas <jacques@solucorp.qc.ca>, James B. MacLean <macleajb@ednet.ns.ca>, Jeff Tranter <Jeff Tranter@Mitel.COM>, Jeffrey A. Uphoff <juphoff@nrao.edu>, Jeremy Fitzhardinge <jeremy@sw.oz.au>, Jochen Hein <jochen.hein@delphi.central.de>, Joerg Reuter <jreuter@lykos.oche.de>, Johan Myreen <jem@vipunen.hut.fi>, John A. Martin <jam@acm.org>, John Aycock <aycock@cpsc.ucalgary.ca>, John Boyd <boyd@cis.ohio-state.edu>, John E. Gotts <jgotts@engin.umich.edu>, John Shifflett <john@geolog.com>, Jon Tombs <jon@gtex02.us.es>, Jonathan Naylor <g4klx@g4klx.demon.co.uk>, Jos Vos <jos@xos.nl>, Juan Jose Ciarlante <jjciarla@raiz.uncu.edu.ar>, Juergen Weigert <jnweiger@immd4.informatik.uni-erlangen.de>, J”urgen Fischer <fischer@etinf.fho-emden.de>, Kai Harrekilde-Petersen <khp@dolphinics.no>, Kai M”akisara <Kai.Makisara@metla.fi>, Kai Petzke <wpp@marie.physik.tu-berlin.de>, Karl Fogel <kfogel@cs.oberlin.edu>, Karl Keyte <kkeyte@koft.rhein-main.de>, Ken Pizzini <ken@halcyon.com>, Kenji Tsutomu Hollis <khollis@bitgate.com>, Kevin E. Martin <martin@cs.unc.edu>, Kevin Lentin <kevinl@cs.monash.edu.au>, Krishna Balasubramanian <balasub@cis.ohio-state.edu>, Lars Wirzenius <liw@iki.fi>, Laurence Culhane <loz@holmes.demon.co.uk>, Lawrence Foard <entropy@world.std.com>, Leo Spiekman <leo@netlabs.net>, Leonard N. Zubkoff <lnz@dandelion.com>, Marc Zyngier <maz@gloups.fdn.fr>, Marco van Wieringen <mvw@mercury.mcs.nl.mugnet.org>, Mark Lord <mlord@pobox.com>, Markus Kuhn <mskuhn@cip.informatik.uni-erlangen.de>, Martin Mares <mj@k332.feld.cvut.cz>, Martin Schulze <joey@linux.de>, Martin von Loewis <loewis@informatik.hu-berlin.de>, Matt Welsh <mdw@sunsite.unc.edu>, Matthias Urlichs <urlichs@noris.de>, Matti Aarnio <mea@utu.fi>, Michael A. Griffith <grif@cs.ucr.edu>, Michael Callahan <callahan@maths.ox.ac.uk>, Michael Hipp <mhipp@student.uni-tuebingen.de>, Michael K. Johnson <johnsonm@redhat.com>, Michael Meskes <meskes@topsystem.de>, Michael Neuffer <mike@i-Connect.Net>, Michael O’Reilly <michael@iinet.com.au>, Miguel de Icaza Amozurrutia <miguel@nuclecu.unam.mx>, Mike Jagdis <jaggy@purplet.demon.co.uk>, Mike McLagan <mike.mclagan@linux.org>, Mike Shaver <shaver@ingenia.com>, Miquel van Smoorenburg <miquels@cistron.nl>, Nick Holloway <Nick.Holloway@alfie.demon.co.uk>, Nigel Gamble <nigel@nrg.org>, Nigel Metheringham <Nigel.Metheringham@ThePLAnet.net>, Niibe Yutaka <gniibe@mri.co.jp>, Olaf Kirch <okir@monad.swb.de>, Orest Zborowski <orestz@eskimo.com>, Paal-Kristian Engstad <engstad@funcom.com>, Pat Mackinlay <pat@it.com.au>, Patrick Volkerding <volkerdi@ftp.cdrom.com>, Paul Gortmaker <gpg109@rsphy1.anu.edu.au>, Pauline Middelink <middelin@polyware.iaf.nl>, Peter Bauer <100136.3530@compuserve.com>, Peter De Schrijver <stud11@cc4.kuleuven.ac.be>, Peter MacDonald, Phil Lewis <beans@bucket.ualr.edu>, Philip Blundell <Philip.Blundell@pobox.com>, Philip Gladstone <philip@raptor.com>, Ralf Baechle <ralf@gnu.ai.mit.edu>, Ralf Flaxa <rfflaxa@immd4.informatik.uni-erlangen.de>, Randolph Bentson <bentson@grieg.seaslug.org>, Ray Burr <ryb@nightmare.com>, Ray Dassen <jdassen@wi.LeidenUniv.nl>, Raymond Chen <raymondc@microsoft.com>, Remy Card <Remy.Card@masi.ibp.fr>, Richard Henderson <richard@gnu.ai.mit.edu>, Rick Miller <rdmiller@execpc.com>, Rick Sladkey <jrs@world.std.com>, Rik Faith <faith@cs.unc.edu>, Rob W. W. Hooft <hooft@EMBL-Heidelberg.DE>, Robert Sanders <gt8134b@prism.gatech.edu>, Roger E. Wolff <wolff@dutecai.et.tudelft.nl>, Ron Holt <ron@novell.com>, Ross Biro <bir7@leland.Stanford.Edu>, Rudolf Koenig <rfkoenig@immd4.informatik.uni-erlangen.de>, Savio Lam <lam836@cs.cuhk.hk>, Scott Snyder <snyder@fnald0.fnal.gov>, Sebastian Hetze <she@lunetix.de>, Simmule Turner <sturner@tele-tv.com>, Simon Shapiro <shimon@i-Connect.Net>, Stefan Probst <snprobst@immd4.informatik.unierlangen.de>, Stefan Reinauer <stepan@home.culture.mipt.ru>, Stephen R. van den Berg (AKA BuGless) <berg@pool.informatik.rwth-aachen.de>, Stephen Rothwell <Stephen.Rothwell@canb.auug.org.au>, Stephen Tweedie <sct@dcs.ed.ac.uk>, Stuart Cheshire <cheshire@cs.stanford.edu>, Terry Dawson <terry@perf.no.itg.telecom.com.au>, Theodore Ts’o <tytso@mit.edu>, Thomas Bogendoerfer <tsbogend@alpha.franken.de>, Thomas Dunbar <tdunbar@vtaix.cc.vt.edu>, Thomas Sailer <sailer@ife.ee.ethz.ch>, Thomas Uhl <uhl@sun1.rz.fh-heilbronn.de>, Todd J. Derr <tjd@fore.com>, Tommy Thorn <Tommy.Thorn@irisa.fr>, Torsten Duwe <Torsten.Duwe@informatik.unierlangen.de>, Tsu-Sheng Tsao <tsusheng@scf.usc.edu>, Tuomas J. Lukka <Tuomas.Lukka@Helsinki.FI>, Volker Lendecke <lendecke@namu01.Num.Math.Uni-Goettingen.de>, Warner Losh <imp@village.org>, Wayne Davison <davison@borland.com>, Werner Almesberger <werner.almesberger@lrc.di.epfl.ch>, Werner Zimmermann <zimmerma@rz.fht-esslingen.de>, William (Bill) Metzenthen <billm@suburbia.net>, William E. Roadcap <roadcapw@cfw.com>, Willy Konynenberg <willy@xos.nl>, Winfried Tr”umper <winni@xpilot.org>, Yuri Per <yuri@pts.mipt.ru>

Version:  2.0.35-1

39 lthreads Linuxthreads - POSIX 1003.1c kernel threads for Linux

This is a release (BETA) of LinuxThreads, a BiCapitalized implementation of the Posix 1003.1c ”pthread” interface for Linux. Contents: linuxthreads-0.5

Copyright:  (c) 1996 Xavier Leroy
Author:  Xavier Leroy <Xavier.Leroy@inria.fr>
Version:  0.5-14

lx hack Source of the newest hacker’s kernel

The source tree of the newest development kernel. As these kernels contain the latest features and enhancements, you may run into problems when using this kernel. So we explicitly state that we do not give support for these kernels. Contents: linux-2.1.97 patches-until-2.1.97

Author:  Linus Torvalds <Linus.Torvalds@Helsinki.FI> et al.
Version:  2.1.115-0

lx suse The sources of the S.u.S.E. kernel

Kernel sources with the following modifications:

- New BSD-NCR driver

- New driver for ICP vortex RAID array controllers

- New driver for Tulip network cards

- Detection of more than 64 MB RAM

- New driver for Tekram DC390

- Bugfixes from pre-patch 2.0.36-1 Contents: linux-2.0.35.SuSE

Copyright:  (c) Linus Torvalds
Author:  Linus Torvalds <torvalds@transmeta.com>
Version:  2.0.35.SuSE-4

make GNU make command

Version:  3.76.1-11

mico Mico is CORBA

Free implementation of the basic CORBA 2.0 specification Documentation:

Copyright:  (c) 1997, Kay R”oder, Arno Puder
Author:  Kay R”oder <mico@cs.uni-frankfurt.de>, Arno Puder <mico@cs.uni-frankfurt.de>
Version:  2.0.5-1

minigcc The GCC Preprocessor without the C-Compiler

This Package contains just the Preprocessor which is used by the XFree86 packages.

Copyright:  (c) 1987-95 Free Software Foundation, Inc.
Version:  2.7.2.1-1

mocka Modula-2 compiler

mocka is a real modula2-compiler, now available as ELF version. The developers ask you to register to know how many of you use it since the money they get for developing depends on the amount of people using it. So please register! Documentation is to be found under: /usr/doc/packages/mocka Start with ’mocka’ (instead of ’mc’) Contents: mocka mocka9502main-linux mocka9502src

Copyright:  (c) 1993 by GMD Karlruhe, Germany
Version:  9502-13

mtc Modula-2 / C Converter

Unfortunately, there is no documentation available for this package. Contents: mtc-25sep92

Copyright:  (c) 1990 GMD - Universit”at Karlsruhe
Author:  Josef Grosch <grosch@karslruhe.gmd.de>
Version:  92.9.25-10

p2c A Pascal to C translator.

Translates PASCAL programs to C. These can then be compiled using gcc. Contents: p2c-1.20

Copyright:  (c) 1989-91 Free Software Foundation.
Author:  Dave Gillespie <daveg@csvax.cs.caltech.edu>
Version:  1.20-14

patch GNU patch program

Version:  2.1-13

pmake BSD make

This may be required if you’re going to port software with Makefiles designed for BSD. The binary is installed as usr/bin/pmake. Man pages are included. Contents: pmake-1.0

Copyright:  (c) 1988-90 Regents of the University of California
Author:  Karl London <karl@borg.demon.co.uk>
Version:  1.0-13

prcs PRCS -- Project Revision Control System

The Project Revision Control System, PRCS, is the front end to a set of tools that (like CVS) provide a way to deal with sets of files and directories as an entity, preserving coherent versions of the entire set. Contents: prcs-1.2.0

Copyright:  (c) 1997 Josh MacDonald
Author:  Josh MacDonald
Version:  1.2.0-11

python Python Interpreter

Python is an interpreted object-oriented programming language, and is often compared to Tcl, Perl, Scheme or Java. A survey of Python you could find in the documentation and tutorials included in the pyth doc (PostScript, HTML) package. If you have web access, point your browser to http://www.python.org. Contents: python-1.5

Copyright:  (c) 1991-1995 by Stichting Mathematisch Centrum, Amsterdam, The Netherlands.
Author:  Guido van Rossum <guido@python.org>
Version:  1.5.1-4

rcs Revision Control System

RCS, the Revision Control System, manages multiple revisions of files. RCS can store, retrieve, log, identify, and merge revisions. It is useful for files that are frequently revised e.g. programs, documentation, graphics, and papers. Contents: rcs-5.7

Copyright:  (c) 1982-89 Walter F. Tichy, 1990-95 Paul Eggert

Author:

Walter F. Tichy <tichy@ira.uka.de>, Paul Eggert <eggert@twinsun.com>, Adam Hammer <hammer@cs.purdue.edu>, Guy Harris <guy@sun.com>, Jay Lepreau <lepreau@cs.purdue.edu>

Version:  5.7-13

scheme UMB Scheme

UMB Scheme is an implementation of the language described in the IEEE Standard for the Scheme Programming Language (December, 1990). All syntax, variables and procedures are implemented. Integers are implemented as fixnums and bignums, rationals as pairs of integers, (inexact) reals as double-precision floats, and (inexact) complex numbers as pairs of double-precision floats. Contents: scheme-2.5

Copyright:  (c) Copyright: 1988
Author:  William R. Campbell <bill@cs.umb.edu>
Version:  2.5-15

slang Display control

Library with routines for display access. Similar to ncurses, but it has its own programming interface. Slang uses the entries in /usr/lib/terminfo.

Contents: slang

Copyright:  (c) 1993 John E. Davis
Author:  John E. Davis <davis@amy.tch.harvard.edu>
Version:  1.2.2-1

smaltalk GNU Smalltalk

GNU Smalltalk is an implementation which is almost Smalltalk-80 conform (^{TM}ParkPlace Systems) as described in Adele Goldberg and David Robson’s book ’Smalltalk-80: the Language and its Implementations’, which will be referred to as the ’Blue Book’. Smalltalk is an object oriented programming language, meaning that the programmer not only has to keep the data in his head but the operations which can be applied to this object. The ability of an object of keeping not only track of the data but also to those operations that are connected to itself are inseparable glued together. The amount of actions that can be performed by the object is precisely defined by the amount of operations of this special object (called methods in Smalltalk). You can’t ’see’ the contents of an object from outside the object. Seen from the outside an object is a black box which has some particular operations and states, but that’s about all you know. In Smalltalk everything is an object. This includes variables, executable procedures (methods) and stack frames (context of invoked methods or context of a block). Every object is an instance of a class. You can see a class as a datatype and its incorporated functions. An instance is a particular variable of this datatype. If you want to perform an operation on one of these objects you have to send it a message and the object reacts by performing the correspondent action. To print something, e.g., you could send the message like this: someObject print The message you send is indeed the name of a method (procedure) which should be carried out. If you send a message to an object Smalltalk tries to find out the method that has been defined for this kind of object. First it starts looking at the objects class. If there is no such class it bothers the parent class, in the grandfather class etc.. At the top of this class-construction there is a class called Object which has no ancestors. If Smalltalk couldn’t find a method until then it produces an error code. For a complete explanation of Smalltalk there is a tutorial which is provided some extra example programs. Additionally the X Smalltalk interfaces ’stix (gst stix) as well as the TCP interface (gst tcp) are included. Contents: smalltalk-1.2.alpha5

Copyright:  (c) 1990-1996 Free Software Foundation, Inc.
Author:  Steve Byrne <sbb@gnu.ai.mit.edu>
Version:  1.2.alpha5-13

smeiffel SmallEiffel

Implementation of the programming language Eiffel. If you want to use the programming environment, you should include the directory /usr/lib/SmallEiffel/bin into your PATH environment variable. Contents: SmallEiffel

Copyright:  (c) Dominique COLNET, Suzanne COLLIN
Author:  Dominique Colnet <colnet@loria.fr>
Version:  0.20-1

strace Trace system calls of a program.

With strace you can trace the activity of a program. Information as to any system calls the program made and the signals it has received/processed can be seen. Even child processes can be tracked. Contents: strace-3.1

Copyright:  (c) 1991-94 Paul Kranenburg, Branko Lankester, Rick Sladkey

Author:

Paul Kranenburg <pk@cs.few.eur.nl>, Branko Lankester <branko@hacktic.nl>, Rick Sladkey <rs@world.std.com>

Version:  3.1-18

svgaaout Svgalib Super-VGA Graphics Library (a.out)

This library is included for compatibility reasons. The ’real’ svgalib can be found in the package svgalib.

Copyright:  (c) 1993 Tommy Frandsen, 1993 Harm Hanemaayer

Author:

H. Hanemaayer <hhanemaa@cs.ruu.nl>, Michael Weller <eowmob@expmath.uni-essen.de>, Orest Zborowski <obz@Kodak.com>, David Wexelblat <dwex@XFree86.org>, Thomas Roell <roell@informatik.tu-muenchen.de>, Kevin E. Martin <martin@cs.unc.edu>, Rickard E. Faith <faith@cs.unc.edu>, Christopher Wiles <wileyc@fsr.net>

Version:  1.26-1

svgalib Svgalib Super-VGA Graphics Library

A video hardware graphics library, based on VGAlib. Supports most standard and non-standard VGA modes, and Super-VGA modes for Tseng ET4000, Cirrus 542x (including 15/16/24 bit modes), Mach32, Trident 8900c/9000, and other cards. Documentation is located in /usr/doc/packages/svgalib . Contents: svgalib-1.2.12

Copyright:  (c) 1993 Tommy Frandsen, 1993 Harm Hanemaayer

Author:

H. Hanemaayer <hhanemaa@cs.ruu.nl>, Michael Weller <eowmob@expmath.uni-essen.de>, Orest Zborowski <obz@Kodak.com>, David Wexelblat <dwex@XFree86.org>, Thomas Roell <roell@informatik.tu-muenchen.de>, Kevin E. Martin <martin@cs.unc.edu>, Rickard E. Faith <faith@cs.unc.edu>, Christopher Wiles <wileyc@fsr.net>

Version:  1.3.0-4

swi pl SWI-Prolog distribution

An extensive Prolog distribution. Please read the READMEs in /usr/doc/packages/swi-pl and the pl manpage. Contents: swi-pl

Copyright:  (c) 1990-1996 Jan Wielemaker
Author:  Jan Wielemaker <jan@swi.psy.uva.nl>
Version:  2.8.6-13

toolpack extensive collection of tools for Fortran programmers

An extensive collection of tools for Fortran programmers. Contents: toolpack-1.2

Copyright:  (c) Numerical Algorithms Group Ltd
Author:  Thomas Koenig (hacked for Linux) <ig25@rz.uni-karlsruhe.de>, Roman Augustyn
Version:  1.2-18

tvision tvision

Porting from Borland TurboVision for UNIX Contents: tvision-0.5

Copyright:  (c) Borland copyright and BSD-style copyright
Author:  Sergio Sigala <ssigala@globalnet.it>
Version:  0.5-4

tya TYA is a 100% inofficial JIT-compiler

TYA is a “100% inofficial” JIT-compiler. Documentation, test- and example-programs can be found at /usr/doc/packages/tya. Contents: tya07

Copyright:  (c) 1997 Albrecht Kleine
Author:  Albrecht Kleine <kleine@ak.sax.de>
Version:  1.0-0

xclass Linux Explorer + library

This package contains a library for uniform presentation of fvwm95 programs and (as a demo) the Linux Explorer, a file viewer, a Run utility and a Win95 clock. Best with fvwm95. Documentation in: /usr/doc/packages/xclass /usr/X11R6/lib/xclass/tutorial

Copyright:  (c) 1996, Hector Peraza
Author:  Hector Peraza <peraza@mitac11.uia.ac.be>
Version:  970707-11

xscheme An object-oriented scheme

XScheme is an implementation of the Scheme programming language with extensions to support object-oriented programming. There are currently implementations of XScheme running on the IBM-PC and clones under MS-DOS, on the Macintosh, the Atari-ST, and the Amiga. It is completely written in the programming language ’C’ and is easily extended with user written built-in functions and classes. It is available in source form to non-commercial users. XScheme follows the ”Revised 3 Report on the Algorithmic Language Scheme”. Contents: xscheme-0.16

Copyright:  (c) 1989 David Michael Betz
Author:  David Michael Betz
Version:  0.16-12

yacc BSD yacc

”bison -y” should react exactly the same. Contents: yacc

Version:  91.7.30-10

joystick joystick support (kernel module and binaries)

Kernel drivers for analog joysticks as a loadable module. Supports up to two joysticks. The kernel module is contained in package kernmod as /lib/modules/KERNEL-VERSION/misc/joystick.o Unless you are using the SuSE kernel sources, you might have to copy /usr/doc/packages/joystick/joystick.h to /usr/include/linux Contents: joystick-0.8.0

Copyright:  (c) Freely distributable
Author:  asmith@cbnewsd.att.com
Version:  0.8.0-18

0.1.5 DOC: Documentation

allman All man pages

This package contains duplicates of all man pages of every package on the CD (for those who don’t want to install every single package just to get the man pages online).

Version:  98.8.15-0

books Some books

This package contains some books, which are installed under /usr/doc/Books. The books are: The German ”Linux Anwender Handbuch” - Sebastian Hetze, Dirk Hohndel, Martin Mueller, Olaf Kirch a.o. ”Linux Installation and getting started” - Matt Welsh ”Linux Kernel Hackers’ Guide” - Michael K. Johnson ”Linux Programmers Guide” - Sven Goldt and Sven van der Meer ”Linux Network Administrators Guide” - Olaf Kirch ”Linux System Administrators Guide” - Lars Wirzenius ”Linux Users Guide” - Larry Greenfield ”DVIPS: A TeX Driver” ”The C Preprocessor” - Richard M. Stallman ”Using and porting GNU CC” - Richard M. Stallman ”The GNU C Library Reference Manual” - Sandra Loosemore, Roland McGrath, Andrew Oram and Richard M. Stallman The German ”Das kleine TeX Buch” - Fritz Cremer The Handbook for S.u.S.E. Linux can be found in package suse lxe.

DOC: Documentation

As well, there is an extensive list of Unix books (”A Concise Guide to UNIX Books”) by Samuel Ko, and an introduction to the theme ”World Wide Web server Software”.

Contents: install-guide-2.3 khg-0.6 lpg-0.4 nag-1.0.dvi nag-1.0.src sag-0.6 user-beta-1-src

Version:  98.8.4-0

booksfnt Pixmap fonts for the dvi files in the distribution

Installing this package will save you a lot of time waiting for the fonts to be computed for the dvi files in the distribution.

Since not all fonts used in the S.u.S.E. handbooks are included in the teTeX package, it is recommended that you install this package for them.

Contents: dvi-i386

Version:  98.7.8-0

cdb CDB - Components Database

With CDB (Component DataBase), you may get easily information about status of support for many hardware components. Start the Server with /sbin/init.d/cdb start. The frontend is done via WWW. Just start a browser on your computer with the URL: <your hostname>/cdb/english

Copyright:  (c) 1997 S.u.S.E. GmbH
Author:  Richard Jelinek <rj@suse.de>, Simon Pogarcic <sim@suse.de>
Version:  98.7.28-0

changes Change-Logs for the packages

Copyright:  (c) 1996 S.u.S.E. GmbH, F”urth, Germany
Version:  98.8.18-0

dns faq DNS-FAQ

The Domain-Name-Service (DNS) FAQ.

Copyright:  (c) 1997, Chris Peckham, cdp@pfmc.net
Author:  Chris Peckham <cdp@pfmc.net>
Version:  98.8.9-0

elispint Programming in Emacs Lisp: An Introduction

Programming in Emacs Lisp: An Introduction

Contents: emacs-lisp-intro-1.05

Copyright:  (c) FSF, GPL
Author:  Robert J. Chassell <bob@gnu.ai.mit.edu>
Version:  1.05-11

elispman GNU Emacs Lisp Reference Manual 19 (Version 2.4.2)

GNU Emacs Lisp Reference Manual edition 2.4.2, corresponding to GNU Emacs version 19.34.

Contents: elisp-manual-19-2.4.2

Copyright:  (c) Free Software Foundation, Inc., GPL
Author:  Bil Lewis, Dan LaLiberte, Richard Stallman <rms@gnu.ai.mit.edu>
Version:  2.4.2-14

fhs Filesystem Hierarchy Standard

This standard consists of a set of requirements and guidelines for file and directory placement under UNIX-like operating systems. The guidelines are intended to support interoperability of applications, system administration tools, development tools, and scripts as well as greater uniformity of documentation for these systems.

Contents: fhs-2.0

Copyright:  (c) 1994, 1995, 1996, 1997 Daniel Quinlan

Author:

Daniel  Quinlan  <quinlan@pathname.com>,  Brandon  S.  Allbery

<bsa@kf8nh.wariat.org>, Keith Bostic <bostic@cs.berkeley.edu>, Drew Eckhardt <drew@colorado.edu>, Rik Faith <faith@cs.unc.edu>, Stephen Harris <sweh@spuddy.mew.co.uk>, Ian Jackson <ijackson@cus.cam.ac.uk>, John A. Martin <jmartin@acm.org>, Ian McCloghrie <ian@ucsd.edu>, Chris Metcalf <metcalf@lcs.mit.edu>, Ian Murdock <imurdock@debian.org>, David C. Niemi <niemidc@clark.net>, Eric S. Raymond <esr@thyrsus.com>, Mike Sangrey <mike@sojurn.lns.pa.us>, David H. Silber <dhs@glowworm.firefly.com>, Theodore Ts’o <tytso@athena.mit.edu>, Stephen Tweedie <sct@dcs.ed.ac.uk>, Fred N. van Kempen <waltje@infomagic.com>

Version:  2.0-5

gawkdoc Documentation for gawk

Version:  3.0.3-21

gccinfo GNU info-pages for gcc

Version:  2.7.2.1-1

ge info Info files for GNU emacs

This package contains all info files for GNU emacs. All these files can be read online with GNU emacs and describe emacs and some of its modes.

Copyright:  (c) 1985-96 Free Software Foundation, Inc.

Author:

Richard Stallman <rms@gnu.ai.mit.edu>, Karl Heuer <kwzh@gnu.ai.mit.edu>, Geoff Voelker <voelker@cs.washington.edu>, David J. MacKenzie <djm@gnu.ai.mit.edu>, Morten Welinder <terra@gnu.ai.mit.edu>, Paul Reilly <pmr@gnu.ai.mit.edu>, Roland McGrath <roland@gnu.ai.mit.edu>, Brian Fox <bfox@gnu.ai.mit.edu>, Jim Blandy <jimb@gnu.ai.mit.edu>, Eric S. Raymond <esr@thyrsus.com>

Version:  20.2-5

gnuhtml HTML Documentation for GNU-Software

The GNU-info-files converted to html.

Version:  97.11.18-7

gppinfo Documentation for libg++

Version:  2.7-1

howto Collection of HOWTOs from the ’Linux Documentation Project’.

Documentation on many tasks relating to the setup and operation of Linux software. These documents will be installed in /usr/doc/howto. You can read it with ‘less <filename>‘.

Contents: howtos

Version:  98.8.14-0

howtodvi HOWTOs in DVI format

Version:  98.8.14-0

howtohtm HOWTOs in HTML format

Version:  98.8.12-0

howtoinf HOWTOs in info format

Version:  98.8.14-0

howtops HOWTOs as PostScript files

Version:  98.8.14-0

howtosgm SGML sources of the HOWTOs

Version:  98.8.14-0

i4ldoc Documentation for ISDN

More documentation about ISDN and Linux Includes:

- i4l-FAQ 1.1.1 (English version: pre 1.0.5)

- leafsite from Berhard Hailer (http://www.lrz-muenchen.de/˜ ui161ab/www/isdn) Version 2. Feb. 1997

Copyright:  (c) Berhard Hailer
Author:  Bernhard.Hailer@lrz.uni-muenchen.de
Version:  97.6.29-2

inf2htm info2html, converts INFO pages to HTML pages

This packages contains the CGI script info2html which creates html pages out of info documents. info2html generates by calling within a WWW browser a html page of the file /usr/info/dir You can easy click on the info references and nodes to get the on the fly generated html pages. Contents: arrows info2html1.1

Copyright:  (c) 1993-95 Karl Guggisberg
Author:  Karl Guggisberg <guggis@iam.unibe.ch>
Version:  1.1-13

isdnpop isdnpop

This package is for use in Germany only. It contains access software for German ISP. This package is for use in Germany only. It contains access software for a German ISP.

Version:  0.1-1

javadoc Java JDK1.1.3 documentation and demos

The documentation and demo apps and applets that belongs to the jdk1.1.3. ( includes the original sources of jdk1.1.3 at /usr/doc/packages/javadoc/src.zip ) All will be installed at /usr/doc/packages/javadoc. Contents: jdk113demo jdk113doc

Copyright:  (c) (c) 1994, 1995, 1996, 1997 Sun Microsystems, Inc.
Version:  1.1.3-8

ldp Linux Documentation Project HTML archive

A lot of Linux docu in HTML format. Contents: ldp

Copyright:  (c) GPL
Author:  A lot!, Greg Hankins <gregh@sunsite.unc.edu>
Version:  98.8.12-0

ldpman LDP Man-Pages

The man pages of the Linux Documentation Project. They are maintained by Andries Brouwer. Contents: man-pages-1.19

Author:

Alain L. Knaff <Alain.Knaff@imag.fr>, Alan Cox <Alan.Cox@linux.org>, Andries Brouwer <aeb@cwi.nl>, Arnt Gulbrandsen <agulbra@nvg.unit.no>, Bjorn Ekwall <bj0rn@blox.se>, Christopher Oliver <oliver@fritz.co.traverse.com>, Dan Miner <dminer@nyx.cs.du.edu>, Daniel Quinlan <quinlan@yggdrasil.com>, Darren Senn <sinster@darkwater.com>, David Metcalfe <david@prism.demon.co.uk>, Drew Eckhardt <drew@PoohSticks.ORG>, Giorgio Ciucci <giorgio@crcc.it>, Graeme W. Wilford <G.Wilford@ee.surrey.ac.uk>, Ian Jackson <iwj10@cus.cam.ac.uk>, Jens Schweikhardt <Jens.Schweikhardt@rus.uni-stuttgart.de>, Jim Van Zandt <jrv@vanzandt.mv.com>, Jochen Hein <Hein@Informatik.TU-Clausthal.de>, Jochen Karrer <cip307.cip.physik.uni-wuerzburg.de>, Luigi P. Bai <lpb@neosoft.com>, Mark D. Roth <roth@uiuc.edu>, Michael Elizabeth Chastain <mec@shell.portal.com>, Michael Haardt <michael@cantor.informatik.rwth-aachen.de>, Michael Shields <shields@tembel.org>, Mike Battersby <mike@starbug.apana.org.au>, Mitchum DSouza <Mitch.DSouza@NetComm.IE>, Nicolai Langfeldt <janl@ifi.uio.no>, Neil P. Soveran-Charley <athan@mordor.mersinet.co.uk>, Ralf Baechle <ralf@waldorf-gmbh.de>, Rik Faith <faith@cs.unc.edu>, Robert K. Nichols <Robert.K.Nichols@att.com>, Stephen Lee <sl14@cornell.edu>, Steven Greenland <stevegr@sugar.neosoft.com>, Thomas K. Dyas <tdyas@eden.rutgers.edu>, Thomas Koenig <ig25@rz.uni-karlsruhe.de>, Ulrich Drepper <drepper@gnu.ai.mit.edu>, Urs Thuermann <urs@isnogud.escape.de>

Version:  1.20-1

ldpmande LDP Man-Pages (german)

This are the german man pages of the Linux Documentation Project. Please note that they are normally older than the english versions. For reference you should use the english versions.

Copyright:  (c) 1993 Sebastian Hetze, diverse

Author:

Sebastian Hetze <she@soho.lunetix.de>, Andries E. Brouwer <aeb@cwi.nl>, Martin Schulze <joey@infodrom.north.de>, Rene Tschirley <gremlin@cs.tuberlin.de>, Norbert K”umin <norbert.kuemin@lugs.ch>, Elmar Jansen <ej@pumuckel.gun.de>, Aldo Valente <aldo@dagobar.rhein.de>, Lars J. Brandt <ljbrandt@jorma.ping.de>, Lutz Donnerhacke <lutz@as-node.jena.thur.de>, Michael Haardt <michael@cantor.informatik.rwth-aachen.de>, Christian Schmidt <c.schmidt@ius.gun.de>, Norbert Weuster <weuster@etecs0.uni-duisburg.de>, Ralf Baechle <ralf@waldorf-gmbh.de>, Patrick Rother <krd@burn.rhein-ruhr.de>, Jonas Rovan <jonas@blitz.de>, Wolfgang Jung <woju@keep.in-berlin.de>, Florian Brons <florian@merlin.pn.org>, Michael Burghardt <mb@suse.de>, Ralf Baumert <bau@heineken.chemie.uni-dortmund.de>, Heiko Schlittermann <heiko@lotte.sax.de>, Carsten Borchardt <carsten@bor.north.de>, Kai Dupke <dupke@koma.han.de>, Harald Wieland <dg9bcy@epc99.fbe.fh-wilhelmshaven.de>, Ralf W. Stephan <ralf@franken.de>, Jochen Hein <Hein@Informatik.TU-Clausthal.de>, Hanno Wagner <wagner@bidnix.bid.fh-hannover.de>

Version:  970513-0

ldpmanfr LDP Man-Pages (french)

This are the french man pages of the Linux Documentation Project. Please note that they are normally older than the english versions. For reference you should use the english versions.

Author:  Christophe Blaess <ccb@club-internet.fr>
Version:  0.5-0

ldpmanit LDP Man-Pages (italian)

This are the italian man pages of the Linux Documentation Project. Please note that they are normally older than the english versions. For reference you should use the english versions.

Author:  Michele Dalla Silvestra <dalla@psico.unipd.it>
Version:  0.2a-0

libcinfo Infofiles for the GNU C library and GDBM library

Version:  5.0.9-1

man9 Some man pages from section 9

This man pages are containing infos about kernel structures. Contents: man9

Author:

Cyrus Durgin <cider@speakeasy.org>, Keith Owens <kaos@ocs.com.au>, Kirk Petersen <kirk@speakeasy.org>, Stephen Williams <steve@icarus.com>

Version:  98.2.14-4

manyfaqs A collection of frequently asked questions/answers

Contents of the package: The full set of Linux HOWTOs. The comp.lang FAQs for Fortran, GCC, G++, TCL, Objective-C, Lisp, Scheme, and Perl. The PostScript FAQs. The Linux FAQ. The comp.unix.question FAQs. The comp.windows.x X FAQ and many other FAQs relating to X. manyfaqs: These FAQs will be installed under /usr/doc/faq. Contents: faq

Version:  98.2.3-5

odshtml HTML-Pages from Offenes Deutsches Schulnetz

Version:  97.11.18-7

perlref Perl 5 Reference Guide

Author:  Johan Vromans <jvromans@squirrel.nl>
Version:  5.001.2-1

ppp nt Linux connection to Windows NT

Documentation, but only in German: Dieses Dokument beschreibt, wie man eine PPP-Verbindung zwischen einem LINUX-Host und einem Windows NT Server unter Verwendung der PAP/CHAP Authentifizierung mit automatischem Rueckruf durch den Windows NT Server herstellt Contents: ppp nt-1.1

Copyright:  (c) Siegfried Schoen <sigi@f30x.erl.scn.de>
Author:  Siegfried Schoen <sigi@f30x.erl.scn.de>
Version:  1.1-5

pyth doc Python Documentation

The Python Reference, Tutorial, C-API Reference, Library Reference, Extending and Embedding Manuals in PostScript and HTML.

Copyright:  (c) 1991-1995 by Stichting Mathematisch Centrum, Amsterdam, The Nether lands.
Author:  Guido van Rossum <guido@python.org>
Version:  1.5.1-4

rfc The RFC’s (request for comments)

This is the reference concerning computer related questions. The discussed matter might be a little to theoretical for laymen but professionals do find all what they are looking for.

Copyright:  (c) many
Version:  98.8.14-0

sdb cgi Support Database search function

Search function in Support Database There will be installed two CGI-scripts for searching in Support Database via a local installed webserver (Apache). More info: /usr/doc/support-db/sdb e/sdb.html

Copyright:  (c) S.u.S.E. GmbH, Fuerth, Germany
Author:  Klaus Franken <kfr@suse.de>
Version:  0.80-0

sdb de S.u.S.E. Support DataBase (German)

A collection of frequently asked support questions. These are written in HTML and can be found under /usr/doc/support-db/sdb. The english sdb is in package sdb en.

Copyright:  (c) 1996-98 S.u.S.E. GmbH, Fuerth, Germany
Version:  98.8.15-0

sdb en S.u.S.E. Support DataBase (English)

A collection of frequently asked support questions. These are written in HTML and can be found under /usr/doc/support-db/sdb e. The English version is under construction.

Copyright:  (c) 1996-98 S.u.S.E. GmbH, Fuerth, Germany
Version:  98.8.15-0

selfhtml selfhtml

HTML4 Referenz and Tutorial Contents: selfhtml

Copyright:  (c) Stefan M”unz 1998
Author:  Stefan M”unz <muenz@compuserve.com>
Version:  7.0-2

suse lx Handbook S.u.S.E. Linux (german)

Copyright:  (c) S.u.S.E. GmbH, Fuerth
Version:  5.3.0.6-1

suse lxe Handbook S.u.S.E. Linux (english)

Copyright:  (c) S.u.S.E. GmbH, Fuerth
Version:  5.2.0.1-12

susebuch LaTeX macros for the S.u.S.E. Linux Handbook

Copyright:  (c) (c) S.u.S.E., GPL
Author:  Martin Scherbaum <maddin@suse.de>, Werner Fink <werner@suse.de>
Version:  1.3.2-0

susehilf S.u.S.E. Help-System (base)

This package contains the basics of the S.u.S.E. Help-System. To prevent unnecessary disk usage the man pages will be formatted ”on the fly”. Thus you can only view man pages that are already installed. Start ’susehelp’ by invoking help. susehelp --help shows a list of configuration possibilities. Contents: http-rman rman

Copyright:  (c) 1995-98 S.u.S.E. GmbH, F”urth, Germany
Author:  <suse@suse.de>
Version:  98.8.15-0

susepak S.u.S.E. Help-System (package descr. German)

This package contains German descriptions for all the packages in html format.

Copyright:  (c) 1995 S.u.S.E. GmbH, F”urth, Germany
Author:  <suse@suse.de>
Version:  98.5.4-0

susepake S.u.S.E. Help-System (package descr. English)

Copyright:  (c) 1995 S.u.S.E. GmbH, F”urth, Germany
Author:  <suse@suse.de>
Version:  98.5.4-0

toppp T-Online / PPP scripts

T-Online now offers PPP access. This package contains some PPP example scripts for T-Online. These scripts are copied to /usr/doc/packages/toppp. A description can be found under /usr/doc/packages/toppp/README.SuSE. Contents: toppp

Copyright:  (c) 1997 Klaus Franken, S.u.S.E. GmbH, Fuerth, Germany
Author:  Klaus Franken <kfr@suse.de>
Version:  98.2.13-4

uucpdoc Taylor UUCP Documentation

Copyright:  (c) 1991-93 Ian Lance Taylor
Author:  Ian Lance Taylor <ian@cygnus.com>
Version:  1.06.1-11

wie dvi German HOWTOs in DVI format

Version:  98.8.9-0

wie geht Collection of German HOWTOs

Author:

Brian Ward <bri@blah.math.tu-graz.ac.at>, Grant Taylor <gtaylor@picante.com>, Jens Vonderheide <Jens Vonderheide@wharfrat.fido.de>, Marco Budde <Budde@tuharburg.d400.de>, Matt Welsh <mdw@sunsite.unc.edu>, Peter S”utterlin <pit@unisw.gwdg.de>

Version:  98.8.9-0

wie html German HOWTOs in HTML format

Version:  98.8.9-0

wie info German HOWTOs in info format

Version:  98.8.9-0

wie pdf German HOWTOs in Acrobat (PDF) format

Version:  98.8.9-0

wie ps German HOWTOs as PostScript files

Version:  98.8.9-0

wie sgml SGML sources of the German HOWTOs

Version:  98.8.9-0

xe info Info files for X emacs

This package contains all info files for X emacs. All these files can be read online with X emacs and describe xemacs and some of its modes.

Copyright:

(c) 1994, 1995 Board of Trustees, University of Illinois, 1994 Free Software Foundation, Inc.

Author:

Chuck Thompson <cthomp@cs.uiuc.edu>, Ben Wing <wing@666.com>, and many other contributors

Version:  20.4-11

xf86html HTML Pages from XFree86

Version:  98.8.15-0

0.1.6 E: Emacs

elib Elib, a Lisp library for the GNU-Emacs

This package contains the completed GNU-Emacs Lisp Library Elib in Version 1.0. Elib was

\underline{E: Emacs}

developed as a library for Emacs-Lisp programs for the same reasons that libg++ was for C++ programs. Elib contains macros for:

- ‘container’ data structures (queues, stacks, AVL trees, etc.)

- string handling routines missing in standard emacs

- routines to handle so-called ‘cookies’ in a buffer

Contents: elib-1.0

Copyright:  (c) Inge Wallin, Per Cederqvist
Author:  Inge Wallin <inge@lysator.liu.se>, Per Cederqvist <ceder@lysator.liu.se>
Version:  1.0-22

emacs w3 Emacs/W3 - a WWW browser for Emacs

Implemented in Emacs Lisp.

Copyright:  (c) GPL

Author:

William M. Perry <wmperry@aventail.com>, Alf-Ivar Holm <alfh@ifi.uio.no>, Charles Levert <charles@comm.polymtl.ca>, Ching-Mo Chang <chang@wsu.edu>, Christian Limpach <chris@nice.ch>, Colin Rafferty <craffert@ml.com>, Cord Kielhorn <kielhorn@thphy.uni-duesseldorf.de>, Dave Love <d.love@dl.ac.uk>, Denys Duchier <Denys.Duchier@ps.uni-sb.de>, Dieter Maurer <dieter@hit.handshake.de>, Greg Stark <gsstark@mit.edu>, Hrvoje Niksic <hniksic@srce.hr>, Istvan Marko <istvan@cmdmail.amd.com>, Jens-Ulrik Holger Petersen <petersen@kurims.kyoto-u.ac.jp>, Joe Wells <jbw@cs.bu.edu>, John Palmieri <palmieri@math.mit.edu>, Jonathan Harris <jhar@tardis.ed.ac.uk>, Laurent Martelli <laurent@lolo.martelli.fr>, Michael Ernst <mernst@cs.washington.edu>, Paul Stodghill <stodghil@CS.Cornell.EDU>, Per Abrahamsen <abraham@dina.kvl.dk>, Roland McGrath <roland@ai.mit.edu>, Shuji Narazaki <narazaki@InetQ.or.jp>, SL Baur <steve@altair.xemacs.org>, Stefan Hornburg <racke@gundel.han.de>, Takahiro Hayata <hayata@sc511t.s.kobe-u.ac.jp>, T. V. Raman <raman@Adobe.COM>, Thierry Emery <Thierry.Emery@nmu.alcatel.fr>, Toby Speight <tms@ansa.co.uk>, Vladimir Alexiev <vladimir@cs.ualberta.ca>

Version:  4.0pre.23-5

ge auc AUC TeX: add-on for Emacs

These macros make it easier for you to write TeX documents with GNU Emacs. Documentation is to be found under: /usr/doc/packages/auctex/ and in the dvi-files math-ref.dvi tex-ref.dvi (reference cards) and in info files auctex in emacs infomode. AuC TeX is integrated in XEmacs 19.15, therefore you may not install this packages. Note that the binary formats of the byte-compiled lisp files of the both emacs editors are incompatible, therefore you can not use this package for XEmacs.

Contents: auctex-9.7p

Copyright:  (c) 1994 AUC TeX development

Author:

Kresten Krab Thorup <auc-tex mgr@iesd.auc.dk>, Per Abrahamsen <abraham@dina.kvl.dk>

Version:  9.7p-23

ge exec GNU Emacs

Eight Megabytes Always Continuously Swapping ;-) Love it or leave it.

Contents: emacs-20.2 leim-20.2

Copyright:  (c) 1985-98 Free Software Foundation, Inc.

Author:

Richard Stallman <rms@gnu.ai.mit.edu>, Karl Heuer <kwzh@gnu.ai.mit.edu>, Geoff Voelker <voelker@cs.washington.edu>, David J. MacKenzie <djm@gnu.ai.mit.edu>, Morten Welinder <terra@gnu.ai.mit.edu>, Paul Reilly <pmr@gnu.ai.mit.edu>, Roland McGrath <roland@gnu.ai.mit.edu>, Brian Fox, Jim Blandy <jimb@gnu.ai.mit.edu>, Eric S. Raymond <esr@thyrsus.com>

Version:  20.2-5

ge lisp Several lisp files for GNU emacs

Several lisp files not needed for running GNU emacs. Most of these files are pre byte compiled and therefore not necessary.

Copyright:  (c) 1985-96 Free Software Foundation, Inc.

Author:

Richard Stallman <rms@gnu.ai.mit.edu>, Karl Heuer <kwzh@gnu.ai.mit.edu>, Geoff Voelker <voelker@cs.washington.edu>, David J. MacKenzie <djm@gnu.ai.mit.edu>, Morten Welinder <terra@gnu.ai.mit.edu>, Paul Reilly <pmr@gnu.ai.mit.edu>, Roland McGrath <roland@gnu.ai.mit.edu>, Brian Fox <bfox@gnu.ai.mit.edu>, Jim Blandy <jimb@gnu.ai.mit.edu>, Eric S. Raymond <esr@thyrsus.com>

Version:  20.2-5

ge site Site-lisp extensions for GNU emacs

This package contains extensions for GNU emacs of the site-lisp path of emacs. It puts together the AUC-TeX package (version 9.7p) and a system wide startup file (site-start.el) which uses a few lisp macro files combined by us. For copyright have a look on these files. Note, that GNU-Emacs and XEmacs are incompatible. They use different binary code for compiled lisp files and they have different builtin lisp functions, not only names of such functions are different. Therfore you will find in /etc/skel/ an .emacs for your personal ˜ /.emacs which will switch between different init files. Fore GNU emacs there is a /etc/skel/.gnu-emacs for a personal ˜ /.gnu-emacs and therefore your ˜ /.emacs or ˜ /.gnu-emacs files should be checked against the /etc/skel versions. You can always make your preferred ˜ /.gnu-emacs by copying the /etc/skel/.emacs and /etc/skel/.gnuemacs file in your HOME directory and incorporating your extensions or changes.

Contents: site-lisp

Author:

Jonathan Stigelman <stig@hackvan.com>, Per Abrahamsen <abraham@iesd.auc.dk>, Werner Fink <werner@suse.de>, Alon Albert <alon@milcse.rtsg.ai.edu>, Wayne Mesard <wmesard@cs.stanford.edu>, Bruno Salvy <salvy@inria.inria.fr>, Marc Andreessen <marca@ncsa.uiuc.edu>, Inge Frick <inge@nada.kth.se>, Lawrence R. Dodd <dodd@roebling.poly.edu>, Mark B. Motl <motl@cssun.tamu.edu>

Version:  1.2-5

lacheck Syntax checker for LaTeX documents

Lacheck is a syntax checker for LaTeX documents. It is not any more a part of newer AuC-TeX packages. Therefore this packages is given as its own. You can use lacheck as standalone program too.

Contents: lacheck-1.23

Copyright:  (c) 1991,92 Kresten Krab Thorup, 199394 Per Abrahamsen

Author:

Kresten Krab Thorup <auc-tex mgr@iesd.auc.dk>, Per Abrahamsen <abraham@dina.kvl.dk>

Version:  1.23-14

psgml Emacs Add-On to edit SGML documents

‘psgml’ supports you while editing SGML documents. It respects the context of the used DTD (Document Type Definition) and offers the valid elements and attributes. Contents: psgml-1.0.1 Compiled for GNU Emacs (XEmacs has its own version!). Insert the content of the file /usr/doc/packages/psgml/dot.emacs in your own .emacs resp. .gnu-emacs!

Copyright:  (c) GPL

Author:

James Clark <jjc@clark.com>, Lennart Staflin <lenst@lysator.liu.se>, David Megginson <dmeggins@sprynet.com>, Matthias Clasen <mclasen@sun2.mathematik.unifreiburg.de>

Version:  1.0.1-21

xe exec X-Emacs

This is the current version of X-Emacs, former known as Lucid-Emacs. It’s got a nice GUI but is quite big and derived from GNU-emacs 19.13 (current is 19.31 or 19.34). Lisp macros are not necessarily interchangeable between GNU-emacs and Xemacs. This is mainly important for translated .elc files and the key macros.

EMU: Emulators

Contents: xemacs-20.4-elc xemacs-20.4-info xemacs-20.4-mule xemacs-20.4

Copyright:

(c) 1994, 1995 Board of Trustees, University of Illinois, 1994 Free Software Foundation, Inc.

Author:

Chuck Thompson <cthomp@cs.uiuc.edu>, Ben Wing <wing@666.com>, and many other contributors

Version:  20.4-11

xe lisp Several lisp files for X emacs

Several lisp files not needed for running X emacs. Most of these files are pre byte compiled and therefore not necessary. The true X-emacs addict will install them nevertheless.

Copyright:

(c) 1994, 1995 Board of Trustees, University of Illinois, 1994 Free Software Foundation, Inc.

Author:

Chuck Thompson <cthomp@cs.uiuc.edu>, Ben Wing <wing@666.com>, and many other contributors

Version:  20.4-11

0.1.7 EMU: Emulators

adamem A Coleco Vision and Coleco ADAM emulator

Adamem is an emulator for the Z80 based systems ColecoVision and Coleco ADAM. To get the emulator working you have to install the necessary ROMs. The package ei helps with this task as well. Contents: adamem

Copyright:  (c) 1996 Marcel de Kogel
Author:  Marcel de Kogel <m.dekogel@student.utwente.nl>
Version:  1.2-4

dosemu The DOS emulator

This package allows MS-DOS programs to be started in Linux. A virtual machine (the DOS box) provides the necessary BIOS functions. You will find the sources in the ZQ series. Documentation can be found in ’/usr/doc/packages/dosemu’ and in the man page, as well as in the sources. If you wish to allow non-root users to run the emulator, you must enter their names in the file ’/etc/dosemu.users’. Even so, some DOS programs will run only if DOSEMU was started by root. Please note as well that you should only allow ’trusted users’ to use the DOS emulator, as it inherently permits hardware access that would otherwise be forbidden to ’normal’ users. The parameter HogThreshold in ’/etc/dosemu.conf’ states when an ’idling’ DOSEMU should return the CPU to Linux. It is set to use the CPU as much as it wants (0). The lowest useful value is 10. If that isn’t fast enough, you (as UID root) can get maximum performance with nice -19 dos -D-a 2>/dev/null Don’t be surprised though if the other Linux processes then run very sluggishly. Some graphic cards with S3 chips can only be run in 16 color VESA modes, even if you have inserted video vga console graphics chipset s3 memsize 1024 in ’/etc/dosemu.conf’. On sensitive systems you should therefore use the ’secure on’ option in ’/etc/dosemu.conf’, which prevents the use of ’system.com’. This option also closes another security hole occurring with (mainly Dos4gw based) DPMI clients, which allows them access to the complete user space (including the DOSEMU code). The Aout DOSEMU has a second security hole (the possibility of replacing libdosemu) which has disappeared in ELF, since the ELF version no longer generates libdosemu. Contents: dosemu-0.66.7 In order to start dosemu, you will need a bootable hdimage. Please read the handbook, cd to /var/lib/dosemu and start ./setup-hdimage. ATTENTION: if you have already used dosemu: the configuration has changed slightly, so please read the dosemu-part of the handbook even if you think you know what you are doing :-)

Copyright:  (c) Matthias Lautner, Robert Sanders, Carnegie Mellon University

Author:

Michael E. Deisher <deisher@enws99.EAS.ASU.EDU>, James B. Maclean <macleajb@ednet.ns.ca>, Mike Batchelor <mikebat@netcom.com>, Lam Lai Yin <lam836@cs.cuhk.hk>, Uwe Bonnes <bon@elektron.ikp.physik.thdarmstadt.de>, Scott Buchholz <scottb@eecs.nwu.edu>, J. Lawrence Stephan <jlarry@ssnet.com>, John Davis <davis@space.mit.edu>, David Etherton <etherton@netcom.com>, Pasi Eronen <peronen@vipunen.hut.fi>, Jason E Gorden <gorden@jegnixa.hsc.missouri.edu>, Matthew Grant <grantma@ritz.equinox.gen.nz>, David Hansen <dhansen@metapath.com>, Jochen Hein <Hein@Informatik.TU-Clausthal.de>, Alan Hourihane <alanh@fairlite.demon.co.uk>, John Kohl <jtk@atria.com >, Rob Janssen <pe1chl@rabo.nl>, Harald Koenig <koenig@nova.tat.physik.uni-tuebingen.de>, Vinod G Kulkarni <vinod@cse.iitb.ernet.in>, Kang-Jin Lee <lee@tengu.inberlin.de>, Kevin P Lawton <bochs@world.std.com>, Marty Leisner <leisner@sdsp.mc.xerox.com>, Hans Lermen <lermen@suse.de>, Dong Liu <dliu@faraday-gw.njit.edu>, Tim Van der Linden <timvdl@innet.be>, Martin Ludwig <Martin.Ludwig@ruba.rz.ruhr-uni-bochum.de>, Alistair MacDonald <alistair@slitesys.demon.co.uk>, Amit Margalit <amitm@netvision.net.il>, Lawrence K Mao <lkmao@unf.edu>, Rod May <stoke@melbourne.DIALix.oz.au>, WAYNE P MEISSNER <n1046128@student.fit.qut.edu.au>, Lutz Molgedey <lutz@summa.physik.hu-berlin.de>, Erik Mouw <jakmouw@et.tudelft.nl>, Christoph Niemann <Christoph.Niemann@linux.org>, Max Parke <mhp@light.lightlink.com>, Mark Rejhon <mdrejhon@magi.com>, Ronnie <ronnie@epact.se>, Ed Sirett <ed@cityscape.co.uk>, Larry Stephan <jlarry@ssnet.com>, Corey Sweeney <corey@bbs.xnet.com>, Andrew Tridgell <Andrew.Tridgell@anu.edu.au>, Theodore Ts’o <tytso@mit.edu>, Robert Sanders <gt8134b@prism.gatech.edu>, Matthias Lautner, Tim Bird, Joel N. Weber II

Version:  0.66.7-22

ei Installation utility for various emulators.

Ei provides a comfortable and unified way to install ROM images of various emulators. This can be done from a MS-DOS formatted disk or via a given path in the Filesystem. You may also install software for these emulators via Internet.

Contents: ei

Copyright:  (c) (c) 1996,97 S.u.S.E GmbH, Fuerth Germany
Author:  Richard Jelinek <rj@suse.de>
Version:  0.4-9

ibcs2 The iBCS2 emulator

The iBCS2 emulator allows you to run binaries of other UNIX system such as SCO-Unix[tm] under Linux. The emulator is a loadable module that can be loaded into the kernel using the command ’insmod iBCS’. If you have activated the kernel daemon, the ibcs2 emulator module will be loaded automatically on demand. CAUTION: This package contains the libc.so.1 for SYSV 4.x systems! If you install this package you may encounter strange things! If you compile a program and specify the linker option ”-L/usr/lib”, the linker tries to link THIS library which of course doesn’t work. In such a case, either specify the option ”-L/lib -L/usr/lib” or do not specify any path. Then the program will be successfully linked. Contents: ibcs-2.0.970513 libnsl-960201 protlib-960201 sco-libs-940731 svr4-shlib-961222 CAUTION: This package contains the libc.so.1 for SYSV 4.x systems! If you install this package you may encounter strange things! If you compile a program and specify the linker option ”-L/usr/lib”, the linker tries to link THIS library which of course doesn’t work. In such a case, either specify the option ”-L/lib -L/usr/lib” or do not specify any path. Then the program will be successfully linked.

Copyright:  (c) 1993-95 Eric Youngdale, Joe Portman, Mike Jagdis, Al Longyear
Version:  970513-26

ines Nintendo Entertainment System Emulator

This is a Emulator for the Nintendo Entertainment System. It is fully operational. You just need the Games (*.NES-Images).

EMU: Emulators

Contents: ines

Copyright:  (c) 1996,97 Marat Fayzullin
Author:  Marat Fayzullin <fms@freeflight.com>
Version:  0.6-6

stella Stella - an Atari 2600VCS Emulator

The Atari 2600 Video Computer System (VCS), introduced in 1977, was the most popular home video game system of the early 1980’s. Now you can enjoy all of your favorite Atari 2600 games on your PC thanks to Stella!

Contents: stella

Copyright:  (c) 1996,1997 Bradford W. Mott
Author:  Bradford W. Mott <bwmott@eos.ncsu.edu>
Version:  0.6-13

stonx ST on X - an Atari emulator

This program allows you to emulate one (or more..;) ) ATARIs under Linux. You can run almost any GEM program as well as some ”hardware-close” games.

Contents: STonX-0.6.7-epsilon

Copyright:  (c) 1995-1997 Marinos Yannikos and Matrin D. Griffiths
Version:  0.6.7.epsilon-15

stonxsft Feed STonX with this ATARI software

A collection of selected software for the ATARI emulator. An overview of the enclosed programs you can find in suse.txt under /var/lib/stonx/software or on drive C: when you start stonx.

Contents: stonxsft

Version:  98.2.13-4

uae An Commodore Amiga emulator

UAE is an amiga emulator, that nearly completely simulates an A500. You’ll need a Kickstart ROM as well as a Workbench, to work with UAE. Some programs (demos/games) will work without these too.

Contents: uae-0.6.8

Copyright:  (c) Copyright 1995-1997 Bernd Schmidt & Co-Autoren
Author:  Berndt Schmidt <crux@Pool.Informatik.RWTH-Aachen.DE>
Version:  0.6.8-17

vgb A virtual Gameboy. This is real.

This emulator adds the whole world of Nintendo’s Gameboy to your PC. Call this depending on the color depth of your X-Server with vgb (8bit), vgb16 (16bit) and vgb32 (32bit).

Contents: vgb

Copyright:  (c) Marat Fayzullin 1994,1995,1996
Author:  Marat Fayzullin <fms@freeflight.com>
Version:  0.7-13

vice Commodore 8-bit Emulator (C64/C128/VIC20/PET)

This is a collection of Emulators for 8bit Commodore computers. The packet includes a C64, C128, VIC20 and a PET emulator. You’ll need the ROMS of these machines to get the emulators work. See the manual for this.

Contents: vice

Copyright:

(c) 1993-1996 Jarkko Sonninen, Teemu Rantanen, Jouko Valta, Ettore Perazzoli, Andre Fachat

Author:

Jarkko Sonninen <sonninen@lut.fi>, Teemu Rantanen <tvr@cs.hut.fi>, Jouko Valta <jopi@stekt.oulu.fi>, Ettore Perazzoli <ettore@comm2000.it>, Andre Fachat <a.fachat@physik.tu-chemnitz.de>

Version:  0.11.0-16

wine Wine - The MS Windows emulator (ALPHA!)

The Windows emulator. Please adapt /etc/wine.conf to your needs. Additionally you will need an installed MS-Windows (or at least certain parts of it) on your system.

Please have a look at /usr/doc/packages/wine/README.SuSE. There is more documentation available under this directory. Also consider using ’man wine’.

Please be aware that this software still is ALPHA and might damage system files!

After having adapted all necessary files (/etc/wine.conf) you can invoke wine by entering: ’wine program’ (e.g.)

Contents: wine-980104

Copyright:

(c) 1993-95 John Brezak, Erik Bos, Eric Youndale, Alexandre Julliard, David W. Metcalfe, Martin Ayotte, Martin von Loewis, Robert J. Amstadt, David Metcalfe, Johannes Ruscheinski, Miguel de Icaza, John Burton

Author:

Albrecht  Kleine  <kleine@ak.sax.de>,  Alex  Korobka

<alex@phm30.pharm.sunysb.edu>, Alexandre Julliard <julliard@di.epfl.ch>, Anand Kumria <akumria@ozemail.com.au>, Andreas Kirschbaum <ank@rbg.informatik.th-darmstadt.de>, Andrew Bulhak , Andrew Taylor <andrew@riscan.com>, Bernd Schmidt <crux@pool.informatik.rwth-aachen.de>, Bob Amstadt <bob@amscons.com>, Cameron Heide <heide@ee.ualberta.ca>, Carl Williams <crw@harris.mlb.semi.harris.com>, Charles M. Hannum <mycroft@mit.edu>, Constantine Sapuntzakis <csapuntz@mit.edu>, Dag Asheim <dash@ifi.uio.no>, Daniel Schepler <daniel@shep13.wustl.edu>, David Metcalfe <david@prism.demon.co.uk>, Duncan C Thomson <duncan@spd.eee.strath.ac.uk>, Eric Warnke <ew2193@csc.albany.edu>, Eric Youngdale <eric@aib.com>, Erik Bos <erik@hacktic.nl>, Erik Svendsen <z3esv@kmd-ac.dk>, Fons Botman <botman@rabo.nl>, Frans van Dorsselaer <dorssel@rulhm1.leidenuniv.nl>, Goran Thyni , Graham Menhennitt <gfm@werple.mira.net.au>, Greg Kreider <kreider@natlab.research.philips.com>, Gregory Trubetskoy <grisha@mira.com>, Hans de Graaff <J.J.deGraaff@twi.tudelft.nl>, Huw D. M. Davies <h.davies1@physics.oxford.ac.uk>, Itai Nahshon <nahshon@vnet.ibm.com>, James Youngman <mbcstjy@afs.man.ac.uk>, Jan Willamowius <jan@janhh.sh.sub.de>, Jeffrey Hsu <hsu@freefall.cdrom.com>, Jim Peterson <jspeter@birch.ee.vt.edu>, Jimmy Tirtawangsa <j0t2527@tam2000.tamu.edu>, Jochen Hoenicke <Jochen.Hoenicke@arbi.Informatik.Uni-Oldenburg.de>, Jochen Karrer <cip307@wpax01.Physik.Uni-Wuerzburg.DE>, Johannes Ruscheinski , John Brezak <brezak@apollo.hp.com>, John Harvey <john@division.co.uk>, John Richardson <jrichard@cs.uml.edu>, Jon Konrath , Jon Tombs <jon@gte.esi.us.es>, Juergen Marquardt <marqu@lunar.advantest.de>, Jukka Iivonen <iivonen@cc.helsinki.fi>, Karl Guenter Wuensch <hz225wu@unidui.uni-duisburg.de>, Keith Reynolds <keithr@sco.COM>, Kenneth MacDonald <K.MacDonald@ed.ac.uk>, Linus Torvalds <linus.torvalds@helsinki.fi>, Manfred Weichel <Manfred.Weichel@mch.sni.de>, Marcus Meissner <msmeissn@faui01.informatik.uni-erlangen.de>, Martin Ayotte <martin2@trgcorp.solucorp.qc.ca>, Martin von Loewis <loewis@informatik.hu-berlin.de>, Michael Patra <micky@marie.physik.tuberlin.de>, Michael Veksler <s1678223@t2.technion.ac.il>, Miguel de Icaza <miguel@roxanne.nuclecu.unam.mx>, Morten Welinder <terra@diku.dk>, Niels de Carpentier <niels@cindy.et.tudelft.nl>, Olaf Flebbe <o.flebbe@science-com-puting.uni-tuebingen.de>, Onno Hovers <onno@stack.urc.tue.nl>, Paul Falstad <pf@zoof.cts.com>, Peter Galbavy , Peter MacDonald , Philippe De Muyter <phdm@info.ucl.ac.be>, Ramon Garcia <ramon@ie3.clubs.etsit.upm.es>, Rick Sladkey <jrs@world.std.com>, Robert Pouliot <krynos@qbc.clic.net>, Roman Dolejsi <roman@sorry.vse.cz>, Ross Biro <biro@yggdrasil.com>, Scott A. Laird , Steffen Moeller <smoe0024@rz.uni-hildesheim.de>, Sven Verdoolaege <skimo@dns.ufsia.ac.be>, Thomas Sandford <t.d.g.sandford@prdsgrn.demon.co.uk>, Tristan Tarrant <tst@sthinc.demon.co.uk>, Ulrich Schmid <uschmid@mail.hh.provi.de>, Uwe Bonnes <bon@elektron.ikp.physik.thdarmstadt.de>, William Magro <wmagro@tc.cornell.edu>, William Smith <wos@dcs.warwick.ac.uk>, Yngvi Sigurjonsson <yngvi@hafro.is>

Version:  980614-5

xzx Sinclair ZX-Spectrum emulator (with ROM)

This program emulates a ZX-Spectrum with 48K, 128K and Interface1. Files for this computer can be loaded in sna or z80 format. Alternatively you may type your own programs in BASIC - of course. :-)

FUN: Games and more

Amstrad has kindly given permission to include the original ROM images. Amstrad further has the copyright on the ROM images. Contents: xzx-2.1.3

Copyright:  (c) 1993-97 Des Herriott, Erik Kunze
Author:  Des Herriott <dnh@mfltd.co.uk>, Erik Kunze <Erik.Kunze@fantasy.muc.de>
Version:  2.1.3-16

z81 Sinclair ZX81 (Timex TS1000) emulator (with ROM)

Z81 is a ZX81-Emulator. The ZX81 is also known as Timex TS1000 and a small predecessor of the ZX Spectrum. It used to have 1KB RAM of memory. Amstrad kindly gave permission to include the ROM-image containing the original OS. Amstrad retain it’s copyright. Contents: z81

Copyright:  (c) 1995-1997 Russell Marks, Amstrad (ROM)
Author:  Russel Marks <>
Version:  0.2-17

0.1.8 FUN: Games and more

3d chess 3 dimensional chess

Rules are to be found under /usr/doc/packages/3d-chess/3Dc-rules.html. Contents: 3Dc-0.8.1

Copyright:  (c) 1995 Paul Hicks
Author:  Paul Hicks <paulh@euristix.ie>
Version:  0.8.1-14

antipoli X11 Game

This game can be played with 2 - 10 players in a network. You will find the game and all available documentation under /usr/games/AntipoliX-2.1. Contents: AntipoliX-2.1

Copyright:  (c) 1994 Bernard Bross
Author:  Bernard Bross <bross@essi.fr>
Version:  2.1-12

asclock AfterStep’s clock

asclock was originally the AfterStep window manager’s clock. Documentation in: /usr/doc/packages/asclock man asclock Contents: asclock-1.0

Copyright:  (c) 1996, Beat Christen
Author:  Beat Christen <bchriste@iiic.ethz.ch>
Version:  1.0-12

astrolog An astrology program

This is an astrology program that, among many other things, will produce PostScript graphics and display charts in X. You can display the program parameters by typing astrolog -H. Documentation: /usr/doc/astrolog-4.1.README Contents: astrolog-4.10

Copyright:  (c) 1994 Walter D. Pullen
Author:  Walter D. Pullen <Astara@msn.com>
Version:  4.10-12

batalion A mesa-based 3D game

Battalion is a good example of how 3D Games can be implemented using Mesa and how existing OpenGL applications may be ported to Linux. Thanks to Mesa the work involved is quite low. Contents: batalion

Copyright:  (c) 1994-1996 by Andrew Johnson
Author:  Andrew Johnson <ajohnson@eecs.uic.edu>
Version:  97.6.29-14

battball Not so realistic military soccer

Play against the computer or other players (via network) soccer with helicopters or tanks.

Contents: battball-2.0

Copyright:  (c) 1996, 1997 Philip A. Hardin
Author:  Philip A. Hardin <pahardin@cs.utexas.edu>
Version:  2.0-5

bsdgames Several text mode games

This package copies these games into /usr/games: arithmetic, atc, backgammon, battlestar, bcd, bog, caesar, canfield, cfscores, cribbage, factor, fish, fortune, hangman, hunt, mille, monop, morse, number, paranoia, pom, ppt, primes, rain, robots, sail, snake, snscore, teachgammon, trek, wargames, worm, worms and wump. Contents: bsd-games-1.5

Copyright:  (c) 1989/1990 The Regents of the University of California.
Author:  Curt Olson <curt@sledge.mn.org>, Andy Tefft <teffta@erie.ge.com>
Version:  1.5-10

craft Craft - the vicious vikings

Craft is a strategy/war game similar to Warcraft and Dune II. Each player starts off with two workers and a knight, and tries to build a kingdom with which he can kill off all the other players. The workers can build buildings such as mills (which make archers), docks (for ships), and markets (to trade gold and lumber). The game takes a little while to figure out, but it’s worth it. Contents: craft-3.05

Copyright:  (c) Uwe Beyer
Author:  Uwe Beyer <uwe.beyer@gmd.de>
Version:  3.05-12

crosfire Graphical multi-player arcade game for X11

Adventure game, like nethack plus graphical interface. Can be played over a network, includes an editor. Documentation: man crossfire Contents: crossfire-0.93.5.client crossfire-0.93.5.maps crossfire-0.93.5

Copyright:  (c) 1992 Frank Tore Johansen, 1994 Mark Wedel
Author:  Frank Tore Johansen <frankj@ifi.uio.no>
Version:  0.93.5-15

daliclck Dali-Clock

An interesting clock, where the digits change smoothly into the next. Contents: xdaliclock-2.05

Copyright:  (c) (C) 1991, 1992, 1993, 1994, 1995 by Jamie Zawinski.
Author:  Jamie Zawinski <jwz@netscape.com>
Version:  2.05-12

emiclock One more Clock

Once again a clock..;). But this one is special. Nicely designed clock with little girls on the clock-face. You can choose one out of 20 sweetly designed dresses for the girl. Furthermore it has sound support (requires a new compilation) and an alarm clock included (who wants to miss his/her personal leisure time). This is the nicest ;-) Take a look at it. Docu: man emiclock Contents: EmiClock-1.0.2

Copyright:  (c) 1994 Masayuki Koba
Author:  Masayuki Koba <masa-k@sag.hitachi-sk.co.jp>
Version:  1.0.2-18

empire Well-known strategy game (text mode)

For more information, please read the man page or print /usr/doc/games/empire-5.0.dvi . Contents: empire-5.0

Copyright:  (c) 1987, 1988 Chuck Simmons
Author:  Chuck Simmons
Version:  5.0-13

figlet figlet

FIGlet can create characters in many different styles and can kern and ”smush” these characters together in various ways. FIGlet output is generally reminiscent of the sort of ”signatures” many people like to put at the end of e-mail and UseNet messages.

Contents: contributed figlet22 international ours

Copyright:  (c) Artistic License
Author:  Frank, Ian <figlet@uiuc.edu>, Glenn’s Letters <figlet@uiuc.edu>
Version:  2.2-5

fishtank An aquarium in the root window

A nice little aquarium with funny fish. This program only runs when the X server is running in 256 color mode.

Contents: xfishtank-2.1

Copyright:  (c) 1988-89 Patrick J. Naughton, Eric Bina
Author:  Patrick J. Naughton <naughton@wind.sun.com>, Eric Bina <ebina@ncsa.uiuc.edu>
Version:  2.1-13

flying Billiard Game

You can play a lot of billiard-like games: Snooker, Pool-Billiard, Karambolage, Carrom, Air-Hockey and Curling. Documentation: man flying

Contents: flying-6.20

Copyright:  (c) 1994 Helmut Hoenig
Author:  Helmut Hoenig <Helmut.Hoenig@hub.de>
Version:  6.20-14

fortune talk big...

Fortune displays a random text string from a set of files in a certain format. This occurs each time you start a login shell. To get this feature just uncomment the respective lines in the user’s .profile. Contents: fortune Together with a lot of funny sayings fortune also contains some sayings which may not be fully ”politically correct”. These sayings are nevertheless contains as S.u.S.E. Linux-Developers do not want to apply any censorship to fortune. Please note that not all users may appreciate all fortune sayings. Please do not install fortune in this case.

Copyright:  (c) 1986, 1993 The Regents of the University of California.
Author:  Ken Arnold
Version:  1.0-13

freeciv Free Civilization Clone

A clone of the well-known DOS-game Civilization by Microprose. Every player is the leader of an imaginary nation. The aim of the game can be to create a prospering civilization with commerce and knowledge exchange or (more often) the extinction of all other civilizations. To start a new game : First start the server ser, then start the client civ. Have fun !

Copyright:  (c) 1997, Allan Ove Kjeldbjerg, Claus Leth Gregersen, Peter Joachim Unold

Author:

Allan Ove Kjeldbjerg <allan@daimi.aau.dk>, Claus Leth Gregersen <leth@daimi.aau.dk>, Peter Joachim Unold <pjunold@daimi.aau.dk>

Version:  1.7.0-0

frisk X Window System Risk

A version of the Parker Brother’s classic as an X11 application. Can also be played over a network. The package includes: risk (stand-alone version), friskserver (server) & xfrisk (client). Documentation: /usr/doc/packages/frisk, as well as extensive online help.

Contents: Frisk-0.99b4

Copyright:  (c) 1993, 1994 Elan Feingold
Author:  Elan Feingold <feingold@zko.dec.com>
Version:  0.99b4-13

glchess 3D frontend for gnuchess

glchess is a fine 3D mesa frontend for gnuchess Contents: glchess-linux-bin

Copyright:  (c) GPL
Author:  Tomi.Sarvela@iki.fi
Version:  0.10b-6

gnuchess The GNU chess program

A worthy chess opponent. Runs in text mode. You can find an X11 interface in the xboard package. Contents: gnuchess-4.0.pl75

Copyright:  (c) 1988-90 John Stanback, 1992 Free Software Foundation

Author:

Hans Eric Sandstrom <hes@log-sv.se>, Stuart Cracraft <cracraft@ricechex.ai.mit.edu>, Mike McGann <mwm@hslrswi.hasler.ascom.ch>, Philippe Schnoebelen <phs@lifia.imag.fr>, Tim Mann <mann@src.dec.com>

Version:  4.0.pl75-13

gnugo Chinese game ”Go”

/usr/doc/gnugo/Documentation Contents: gnugo-1.2

Author:  Man L. Li <manli@cs.uh.edu>
Version:  1.2-13

gshogi A strategic game

Copyright:  (c) 1988-90 John Stanback, 1992 Free Software Foundation, Inc., 1993-95 Matthias Mutz
Author:  John Stanback, Matthias Mutz <mutz@fmi.uni-passau.de>
Version:  1.2p03-13

hextris Xhextris and CXhextris, hexagonal versions of Xtetris

This package provides two games under X Window System: Xhextris and the colored version CXhextris. Both are fast games involving dropping hexagonal blocks similar to xtetris. Both version are modified. They use the same font (but different letters) and they use the cursor keys. Contents: hextris

Copyright:  (c) Copyright 1990 David Markley, 1993 Khun Yee Fung (color version)
Version:  1.0-13

imaze X11/Xview - Multi-Player - 3D Game

a 3D TCP/IP-network-game under X11/XView. There are clients (imaze, ninja) working together with one server (imazesrv). To start the game, you’ve to fire up the server first: imazesrv /usr/games/imaze/labs/demolab.lab & Then you can start the clients with: ninja <Server> & imaze & Contents: imaze1.3

Copyright:  (c) 1993-1996 by Hans-Ulrich Kiel and Joerg Czeranski

Author:

Hans-Ulrich  Kiel  <Kiel@rz.tu-clausthal.de>,  Joerg  Czeranski

<Czeranski@informatik.tu-clausthal.de>

Version:  1.3-18

lincity lincity

Lincity is a well designed and easy to play SimCity clone. Besides the game itself it provides an editor for easily creating new levels and obstacles. Contents: lincity103

Copyright:  (c) I J Peters 1995,1997.
Author:  I J Peters <ijp@floot.demon.co.uk>
Version:  1.03-17

maelstr A high-quality Asteroids-Clone

You are on space patrol. Blast asteroids, fight UFOs, get extra-weapons. Beware of supernovae and black holes.

Contents: maelstr

Copyright:  (c) 1995 by Ambrosia Software, Inc.
Author:  Sam Lantinga <slouken@cs.ucdavis.edu>
Version:  1.4.3-17

manix Pacman Clone

X11 pacman clone. Contents: manix-19feb94

Copyright:  (c) 1994 Fernando Pereira
Author:  Fernando Pereira <fjp@minerva.inesc.pt>
Version:  94.2.19-9

meltflip Two nice Demos

Try it! Your display will turn upside down or it just melts away. Contents: xmelt-xflip-07feb91

Copyright:  (c) 1990 David Lemke
Author:  David Lemke <lemke@ncd.com>
Version:  07feb91-13

mirrma Arcade style game for color X11 with stereo sounds.

A game like ”Mindbender” (Amiga) or ”Deflektor” (Atari ST). This is a port of a program I wrote for the Amiga in 1989. Comes with full source code. Contents: mirrormagic-1.3

Copyright:  (c) 1995 by Holger Schemel.
Author:  Holger Schemel <aeglos@valinor.owl.de>
Version:  1.3-17

moontool The Moon in focus

One of the most important programs existing :-) Using this program, you can display all important information about the moon constantly. At last... Contents: xmoontool-22sep94

Copyright:  (c) 1993 John Walker
Author:  John Walker <kelvin@Autodesk.com>
Version:  94.9.22-11

net3d A networked, multi-player, 3d combat game

Copyright:  (c) Copyright (C) 1994 by Jamie Cameron
Author:  Jamie Cameron <jcam2@ccds.cc.monash.edu.au>
Version:  0.08-12

nethack A character based RPG

This RPG is somewhat cryptic with its character based output. But a true fan knows and appreciates its complexity and possibilities. Contents: nethack

Author:  Stephen L. Ericksen <stevee@cc.usu.edu>
Version:  3.2.2-13

oneko Oneko, a cat catches your mouse

A nice program which changes your cursor into a cat playing with your mouse cursor. The manual page will show more possibilities to change your cursor. Contents: oneko-1.1b

Copyright:  (c) 1990 Masayuki Koba

Author:

Tatsuya Kato <kato@ntts.co.jp>, John Lerchey <jl4l+@andrew.cmu.edu>, Eric Anderson <eanders+@cmu.edu>, Rick Niles <niles@axp745.gsfc.nasa.gov>

Version:  1.1b-5

oonsoo A Solitaire Card Game

Oonsoo is a solitaire card game. The goal of the game is to arrange twelve suits of cards, in order, onto twelve playing decks. When the game starts, one card face down and one card face up is dealt to each of the playing decks. Contents: oonsoo-1.0 oonsoo-1.1

Copyright:  (c) 1994 Bradford W. Mott
Author:  Bradford W. Mott <bwmott@eos.ncsu.edu>
Version:  1.1-9

pacman Pacman

Another pacman-game. Contents: pacman

Copyright:  (c) Roar Thronaes
Author:  Roar Thronaes
Version:  1.195.07.18-13

pipeman A game of skill

A well known game of skill. Docu: man xpipeman. Contents: xpipeman-18jul91

Copyright:  (c) 1991 Nigel Paver
Author:  Nigel Paver <pavern@uk.ac.man.cs>
Version:  91.7.18-9

playone Game - similar to monopoly

A game -similar to monopoly. Contents: play1-0.2

Copyright:  (c) 1994 TortiSoft, M”unchen
Author:  Torsten F”ahrmann <tfaehr@aol.com>
Version:  0.2-14

puzzle A puzzle

a puzzle..;) Contents: puzzle-21aug91

Copyright:  (c) 1988 Don Bennett
Author:  Don Bennett
Version:  91.8.21-8

rockdiam Arcade style game for color X11 with stereo sounds.

A game like ”Boulderdash” (C=64) or ”Emerald Mine” (AMIGA). Sound server takes advantage of some features of VoxWare(tm), but works on plain /dev/audio, too. Comes with full source code. Contents: rocks n diamonds-0.9b

Copyright:  (c) 1995 by Holger Schemel.
Author:  Holger Schemel <aeglos@valinor.owl.de>
Version:  0.9-17

sastroid Sasteroids

Sasteroids is a game similar to asteroids. The aim is (what else..;) to shoot every thing that appears on your monitor (asteroids, spaceships). You should avoid crashing into those asteroids. You collect points for every destroyed asteroid or spaceship. This started out as a direct port of xasteroids to run under svgalib (a wonderful super-vga graphics library by Harm Hanemaayer) Sasteroids was derived from Phil Goetz’s Xasteroids by Brad Pitzel. Contents: sast1.5-snd+gfx sast1.5-src

Copyright:  (c) 1994 Brad Pitzel
Author:  Brad Pitzel <pitzel@cs.sfu.ca>
Version:  1.5-16

space space

shooting game for console Contents: spaceboom-0.95.src

Copyright:  (c) FREEWARE
Author:  Joe Lee <hosxe@dds.nl>
Version:  0.95-8

thrust A gravity skill/action game

Nice action game. Thrust has some elements borrowed from xpilot. Contents: thrust-0.83c

Copyright:  (c) 1996,96 Peter Ekberg
Author:  Peter Ekberg <peda@lysator.liu.se>
Version:  0.83c-18

vgacard A small collection of card games (libvga)

The card games vga klondike, vga solitaire, vga ohhell & vga spider. X11 is not required to run these games, as they use libvga. However, especially S3 card tend to have problems using this library, so do not be disappointed if the programs do not run :-)

Contents: vga cardgames-1.3.1

Copyright:  (c) 1991-94 Evan Harris
Author:  Evan Harris <evan@cs.mu.oz.au>
Version:  1.3.1-14

x3d 3D animation

3D animation. The displayed objects can be turned using the mouse. Some example object files can be found in /usr/X11R6/lib/X11/x3d/ . The package includes the program hull, which can be used to create object files. See the file cube.poly for an example of the input format of the compiler.

Contents: x3d-2.2

Copyright:  (c) 1992 Mark Spychalla
Author:  Mark Spychalla <spy@castlab.engr.wisc.edu>
Version:  2.2-10

xabacus X-based abacus

There are books on how to use an abacus, but basically all it does is add and subtract, the rest you have to do in your head. Essentially, this is a proof by induction that a computer is more powerful than an abacus, since a computer program can contain an abacus. (But then again, you can simulate a computer within a computer, so what does that show). Actually, with a real abacus, one can move more than one row at a time with 10 fingers. But on the other hand, a real abacus does not have the current sum displayed as an integer.

Contents: xabacus-5.0

Copyright:  (c) 1994 David Albert Bagley
Author:  David Albert Bagley <bagleyd@bigfoot.com>
Version:  5.0-5

xancur animate cursor for x-window

animate cursor for x-window

Contents: xancur-1.4

Copyright:  (c) Christopher Alex.North-Keys Unix Systems Analyst Group Talisman
Author:  Christopher Alex <erlkonig@gnu.ai.mit.edu>
Version:  1.4-6

xaos Powerful fractal generator

XaoS is a fast portable real-time interactive fractal zoomer. It displays the Mandelbrot set (among other escape time fractals) and allows you zoom smoothly into the fractal. Various coloring modes are provided for both the points inside and outside the selected set. In addition, switching between Julia and Mandelbrot fractal types and displaying planes is provided.

Contents: XaoS-2.2

Copyright:  (c) 1997, Jan Hubicka
Author:  Jan Hubicka <hubicka@paru.cas.cz>, Thomas Marsh <tmarsh@austin.ibm.com>
Version:  2.2-13

xball Bouncing balls

This program simulates bouncing balls in a window. You create the balls in a variety of ways, and can set the gravity, elasticity, whether balls collide or not, etc.

Contents: xball-3.0

Copyright:  (c) 1993 David Nedde
Author:  David Nedde <daven@ivy.wpi.edu>
Version:  3.0-15

xbill Who is afraid of little Bill?

A small but mean hacker named Bill wants to infect some Workstations with his program. It is on you to save the world and to slap him for trying this.

Contents: xbill

Copyright:  (c) 1994-1996 Brian Wellington and Matias Duarte
Author:  Brian Wellington <bwelling@eng.umd.edu>, Matias Duarte <matias@eng.umd.edu>
Version:  2.0-13

xbl 3D Tetris

3D Tetris for X-Window

Copyright:  (c) GPL
Author:  Thierry EXCOFFIE <Thierry.EXCOFFIER@ligia.univ-lyon1.fr>
Version:  1.0h-4

xblast A multi-player arcade game for X11

XBlast is a multi-player arcade game for X11R5/R6. The game can be played with at least two players and up to four players. It was inspired by the video/computer game Bomberman(Dynablaster), which was to my knowledge first programmed for NEC’s PC Engine/Turbo Grafx. Other (commercial) versions of the original game exist for IBM-PC, Amiga, NES, GameBoy and Super NES. Contents: XBlast-2.0

Copyright:  (c) 1993/94 Oliver Vogel
Author:  Oliver Vogel <vogel@ikp.uni-koeln.de>
Version:  2.0-12

xboard X11 Interface for GNU Chess

Copyright:

(c) 1991 by Digital Equipment Corporation, Maynard, Massachusetts. 1991-1995 Free Software Foundation, Inc.

Author:

Tim Mann <mann@src.dec.com>, Chris Sears, Dan Sears, Elmar Bartel <bartel@@informatik.tu-muenchen.de>

Version:  3.4.pl0-14

xboing X11 breakout game

This version was compiled with rplay support. If you want to get the sounds, you should install the rplay daemon. Contents: xboing2.3

Copyright:  (c) 1993, 1994, Justin C. Kibell
Author:  Justin C. Kibell, <jck@citri.edu.au>
Version:  2.3-20

xbomb Another minesweeper clone

Copyright:  (c) 1994 1995 AMB Software
Author:  Andrew M. Bishop <amb@gedanken.demon.co.uk>
Version:  2.0-0

xbombs Minesweeper

Game of minesweeper. 30 on 16 fields :-) Contents: xmines

Copyright:  (c) 1993 Andrew Zimmerman
Author:  Andrew Zimmerman <zimmer@calvin.stanford.edu>
Version:  03dec93-13

xdemine Not just another version of the minesweeper game...

The object of the game is to find the location of the mines and to mark their location by placing a small flag. You have exactly as many flags as there are mines, no more, no less. The number of flags that remain to be placed is displayed in the top left corner and a timer is displayed in the top right corner. Contents: xdemineur-2.1

Copyright:  (c) 1993 Marc Baudoin
Author:  Marc Baudoin <babafou@agm-ita.ensta.fr>
Version:  2.1-5

xearth Our planet earth as animated background

Our planet earth as a background picture. Light and shades are time dependent. One can adjust lots of parameters..;) Contents: xearth-1.0

Copyright:  (c) 1994 Kirk Lauritz Johnson
Author:  Kirk Lauritz Johnson <tuna@cag.lcs.mit.edu>
Version:  1.0-13

xengine reciprocationg engine for X

Xengine displays a reciprocationg engine in a window and can be abused as a benchmark for X.

Contents: xengine.beta1

Copyright:  (c) Kazuhiko Shutoh, YAMAHA Corp.
Author:  Kazuhiko Shutoh <shutoh@isl.yamama.co.jp>
Version:  beta1-9

xfract Fractals under X11

Xfractint draws a wide variety of fractals (fast!). Xfractint is a port of the DOS program Fractint.

Contents: xfractint-3.02

Author:

Bert Tyler, Timothy Wegner, Mark Peterson , Pieter Branderhorst, Ken Shirriff <shirriff@eng.sun.com>

Version:  3.0-13

xgammon A backgammon program for Linux

xgammon contains an X11 interface, a simple file selector and saving dialog, a edit position menu option. a rollout option. a mail dump option for newsgroup posting of positions. and it is quite good; even for the experienced backgammon player.

Contents: xgammon-0.96 xgammon 0.98.orig

Copyright:  (c) 1994 Lambert Klasen & Detlef Steuer

Author:

Lambert Klasen <klasen@asterix.uni-muenster.de> , Detlef Steuer <steuer@amadeus.statistik.uni-dortmund.de>

Version:  0.98-9

xgas animated simulation of an ideal gas

xgas is a physical simulation of an ideal gas in a heated box. The temperature of each box can be independently controlled by a scrollbar.

Contents: xgas-29aug91

Copyright:  (c) 1989 Massachusetts Institute of Technology
Author:  Larry Medwin, Dave Sternlicht
Version:  91.8.29-9

xjewel X11 game - similar to tetris

X11 game - similar to tetris. Has a flavour of its own :-) Doku: /usr/doc/X11-misc/xjewel-1.6.README

Contents: xjewel-1.6

Copyright:  (c) 1990 Yoshihiro Satoh
Author:  Yoshihiro Satoh
Version:  1.6-12

xlife Conway’s game of life

Nice version of ”game of life”. Lots of interesting startpositions you can find under /usr/X11R6/lib/xlife/patterns. Documentation: /usr/doc/packages/xlife

Contents: xlife

Copyright:  (c) 1989 Jon Bennett

Author:

Jon Bennett <jcrb@cs.cmu.edu>, Daniel E Lovinger <dl2n@andrew.cmu.edu>, Charles Humphrey Silvers <Charles.Silvers@cs.cmu.edu>, Eric <esr@snark.thyrsus.com>

Version:  5.0-12

xlyap displays Lyapunov exponents

Copyright:  (c) 1991 Ronald Joe Record
Author:  Ronald Joe Record <rr@sco.com>
Version:  2.3-12

xmahjong mahjong

Mahjongg is a challenging chinese game similar to domino. It is usually played by four players. Xmahjongg is the solitair version designed for the X-Window System. More can be found in the appropriate manual page.

Contents: xmahjong

Copyright:  (c) Jeff S. Young
Author:  Jeff S. Young <jsy@cray.com>
Version:  97.9.21-8

xmemory xmemory - a memory game

This memory game is playable by many users on different X11 displays simultanously. Contents: xmemory-2.0

Copyright:  (c) 1993+1995, Helmut Hoenig, Heiligenhaus
Author:  Helmut Hoenig <Helmut.Hoenig@hub.de>
Version:  2.0-13

xmine Yet another minesweeper...;)

Version:  1.0.3-14

xmines Minesweeper clone

Copyright:  (c) 1993 Ashley Roll
Author:  Ashley Roll <ash@dstc.edu.au>
Version:  3.0-13

xmountns fractal landscape in the background

Xmountains is a fractal terrain generator written by Stephen Booth. It can be configured to calculate new ’mountains’ in a certain period of time. Contents: xmountains-2.4

Copyright:  (c) 1994 by Stephen Booth, the University of Edinburgh
Author:  Stephen Booth <S.Booth@ed.ac.uk>
Version:  2.4-13

xmris A nice X11 game

In this game, you are walking around, collecting fruits. The package includes an editor for new screens. This means that in case you get hooked you can create more difficult screens yourself. Contents: xmris

Copyright:  (c) 1993, 1992 Nathan Sidwell
Author:  Nathan Sidwell <nathan@pact.srf.ac.uk>
Version:  4.04-11

xpat2 Patience card game

Patience (card game); nice graphics. Contents: xpat2-1.04

Copyright:  (c) 1994 Heiko Eissfeldt, Michael Bischoff

Author:

Heiko  Eissfeldt  <heiko@colossus.escape.de>,  Michael  Bischoff

<mbi@mo.math.nat.tu-bs.de>

Version:  1.04-12

xphoon Shows moon phases..;)

Program which lets you see the actual moon phase. You can accelerate time to see how the moon will look like in a couple of days..;) Contents: xphoon

Copyright:  (c) 1988, 1991 Jef Poskanzer, Craig Leres
Author:  Craig Leres <>, Jef Poskanzer <>
Version:  91.9.18-9

xpilot A multiplayer arcade game

A shoot’em up which can be played by several players in a network. Please note that you have to start the server (xpilots) before starting the client (xpilot, plus the name of the computer which runs the server). Several maps can be found in /usr/local/games/lib/xpilot/maps/ Doku: /usr/doc/packages/xpilot Contents: xpilot-3.6.2

Copyright:  (c) 1991-95 Bjoern Stabell, Ken Ronny Schouten, Bert Gijsbers

Author:

Bjoern Stabell <bjoerns@staff.cs.uit.no> , Ken Ronny Schouten <kenrsc@stud.cs.uit.no>, Bert Gijsbers <bert@mc.bio.uva.nl>

Version:  3.6.2-16

xpinguin The ’Logo’ of Linux 2.0

Based on xteddy - really cute ;-) Contents: xpinguin

Copyright:  (c) 1994 Stefan Gustavson

Author:

Stefan Gustavson <stefang@isy.liu.se>, Jens P”onisch <J.Poenisch@wirtschaft.tuchemnitz.de>

Version:  1.0.2-13

xpuzzles Some game of skill under X11

Copyright:  (c) 1994 David Albert Bagley
Author:  David Albert Bagley <bagleyd@bigfoot.com>
Version:  5.0.1-12

xroach Some cockroaches on your root window

Xroach displays disgusting cockroaches on your root window. These creapy crawlies scamper around until they find a window to hide under. Whenever you move or iconify a window, the exposed orthopteras again scamper for cover. See manpage (man xroach). Contents: xroach

Copyright:  (c) 1991 by J.T. Anderson
Author:  J.T. Anderson <jta@locus.com>
Version:  12.6.97-12

xrobots A game of skill

Try to escape from a bunch of robots. Use your skill! Contents: xrobots-17nov89

Copyright:  (c) 1989 Brian Warkentine
Author:  Brian Warkentine <brianw@Sun.COM>
Version:  89.11.17-8

xshogi X11-interface for GNU Shogi

Copyright:  (c) 1992 Free Software Foundation, Inc., 1993-95 Matthias Mutz
Author:  Matthias Mutz <mutz@fmi.uni-passau.de>
Version:  1.2p03-13

xskat Skat

Skat is a popular game in Germany. Now with ”Ramsch” and more... WWW page: http://www.draeger.com/˜ gerhardt/xskat-latest.html Contents: xskat-2.2

Copyright:  (c) Frei
Author:  <Gunter Gerhardt> gerhardt@draeger.com
Version:  2.2-5

xsnow Happy X-Mas!

A really nice animation. Santa clause drives across the screen, pulled by his reindeer. It snows, and the snow accumulates on top of the windows :-) Documentation: man xsnow Contents: xsnow-1.40

Copyright:  (c) 1984 - 1994 Rick Jansen
Author:  Rick Jansen <rick@sara.nl>
Version:  1.40-12

xsok strategy game

Sokoban Contents: xsok-1.02

Author:  Michael <mbi@mo.math.nat.tu-bs.de>
Version:  1.02-12

xsol Solitair

The game Solitair. Contents: xsol-30nov92

Copyright:  (c) 1988 Sun Microsystems
Author:  Dave Lemke <lemke@sun.com>
Version:  92.11.30-9

xspringi a mass and spring simulation system

Xspringies is a simulator which allows you to interactively create and edit a system of masses and springs. The parameters of the masses and springs (such as mass elasticity and spring K) as well as those of the surrounding system (such as air viscosity and gravity) can be changed. Several configuration files can be found in /usr/X11R6/lib/X11/xspringies/ . Contents: xspringies-1.1

Copyright:  (c) 1991, 1992 Douglas M. DeCarlo
Author:  Douglas M. DeCarlo <dmd@gradient.cis.upenn.edu>
Version:  1.1-12

xtacy psychedelic animation program

This is a program similar to fractint. It may not contain as many mathematical functions, but it is command-line oriented and may be run as a root animation, in case your CPU load is too low :-) Contents: Xtacy

Copyright:  (c) 1994 Jer Johnson
Author:  Jer Johnson <mpython@gnu.ai.mit.edu>
Version:  1.13-12

xtartan xtartan

Xtartan offers you a wide variety of scottish tartans. They can be viewed or alternatively used as a background picture. One of the most exciting programs at all..;)) Contents: xtartan-2.1

Copyright:  (c) Jim McBeath Globetrotter Software, Saratoga, California
Author:  Jim McBeath <jimmc@hisoft.uucp>
Version:  2.1-12

xteddy A cuddly teddy bear for your X Windows desktop

Normally, xteddy just sits around doing nothing. After all, that’s what teddy bears are for. Look at him, talk to him, place heavy windows on top of him, zap him around until he becomes dizzy, do what you like; he will always be your true (albeit virtual) friend. You can move xteddy with the mouse by pointing at him and dragging him around. When clicked upon, he will pop up on top of all other windows. If you type ”q” on him, he will die (or, as I like to think of it, be tucked away in the file system until you need him next time). That’s it. But he’s cute. Contents: xteddy-1.0.3

Copyright:  (c) GPL
Author:  Stefan Gustavson <stefang@isy.liu.se>
Version:  1.0.3-11

xtetris X Tetris

Version:  2.6-11

xthing colorful images

This program creates psychadelic images. Sample files can be found in /usr/local/lib/xthing-sample . For more information see /usr/doc/X11-misc/xthing-15nov91.README . Contents: xthing-15nov91

Copyright:  (c) 1991 John L. Cwikla
Author:  John L. Cwikla <cwikla@uimrl7.mrl.uiuc.edu>
Version:  91.11.15-10

xtron A game of skill

Copyright:  (c) 1995, Rhett D. Jacobs
Author:  Rhett D. Jacobs <rhett@hotel.canberra.edu.au>
Version:  1.1a-12

xvier connect IV

xvier is a simple game where you and your opponent alternately throw stones into free columns. The stones pile up in the columns, and the goal is to get four stones in a row, in a column or diagonally. Contents: xvier-1.0

Author:  Norbert Jung <jung@dia.informatik.uni-stuttgart.de>
Version:  1.0-12

GNM: GNOME - GNU Network Object Model Environment

yahtzee Yahtzee in text mode

The well-known game of dice in text mode; up to 20 players (human or computer). Contents: colour-yahtzee yahtzee-1.0

Copyright:  (c) 1992 Orest Zborowski
Author:  Orest Zborowski
Version:  1.0-14

0.1.9 GNM: GNOME - GNU Network Object Model Environment

gnlibs GNOME Base package (Libs)

GNOME is a desktop environment. GNOME itself has no specific window manager. Many GNOME users like WindowMaker or Icewm as window manager for GNOME (see those packages). GNOME is started in S.u.S.E. Linux like a window manager: startx gnome or export WINDOWMANAGER=gnome This package contains the basic GNOME libraries. Documentation: GNOME online help /usr/doc/packages/gnome

Copyright:  (c) 1998, The Gnome Project

Author:

Carsten Schaar <nhadcasc@fs-maphy.uni-hannover.de>, Eckehard Berns <eb@berns.prima.de>, Federico Mena Quintero <federico@nuclecu.unam.mx>, George Lebl <jirka@5z.com>, Gregory McLean <gregm@comstar.net>, Havoc Pennington <hp@pobox.com>, Horacio J. Pea <horape@compendium.com.ar>, Jaka Mocnik <jaka.mocnik@kiss.uni-lj.si>, Maciej Stachowiak <mstachow@mit.edu>, Marc Ewing <marc@redhat.com>, Mark Galassi <rosalia@nis.lanl.gov>, Michael Fulbright <msf@redhat.com>, Miguel de Icaza <miguel@nuclecu.unam.mx>, Nathan Summers <rock@gimp.org>, Owen Taylor <owt1@cornell.edu>, Raja R Harinath <harinath@cs.umn.edu>, Szekeres Istvn <szekeres@cyberspace.mht.bme.hu>, Tom Tromey <tromey@cygnus.com>, Tristan Tarrant <ttarrant@etnoteam.it>

Version:  0.20-5

gncore GNOME Base package (Core)

This package contains some centrall GNOME applications like the panel or a taskbar. Documentation: /usr/doc/packages/gnome

Copyright:  (c) 1998, The Gnome Project

Author:

Carsten Schaar <nhadcasc@fs-maphy.uni-hannover.de>, Christopher Blizzard <blizzard@appliedtheory.com>, George Lebl <jirka@5z.com>, John Ellis <johne@bellatlantic.net>, Mark Galassi <rosalia@cygnus.com>, Martin Baulig <baulig@merkur.uni-trier.de>, Raja R Harinath <harinath@cs.umn.edu>, Tom Tromey <tromey@cygnus.com>

Version:  0.20-5

gnadmin GNOME Applications (Admin)

This package contains the GNOME logviewer. To be extended. Documentation: GNOME online help /usr/doc/packages/gnome

Copyright:  (c) 1997, The GNOME Project

Author:

Arturo Espinosa <arturo@nuclecu.unam.mx>, Carsten Schaar <nhadcasc@fs-ma-phy.uni-hannover.de>, Changwoo Ryu <cwryu@adam.kaist.ac.kr>, Gregory McLean <gregm@comstar.net>, Michael K. Johnson <johnsonm@redhat.com>, Raja R Harinath <harinath@cs.umn.edu>, Sebastian Wilhelmi <wilhelmi@ira.uka.de>, Tom Tromey <tromey@cygnus.com>

Version:  0.20-6

gngames GNOME Games

This package contains some games for GNOME. Documentation: GNOME online help /usr/doc/packages/gnome

Copyright:  (c) 1997, The GNOME Team

Author:

Carsten Schaar <nhadcasc@fs-maphy.uni-hannover.de>, Changwoo Ryu <cwryu@adam.kaist.ac.kr>, Gregory McLean <gregm@comstar.net>, Raja R Harinath <harinath@cs.umn.edu>, Tom Tromey <tromey@cygnus.com>

Version:  0.20-5

gngraph GNOME Graphics

This package contains some graphical applications for GNOME. Documentation: GNOME online help /usr/doc/packages/gnome

Copyright:  (c) 1998, The GNOME Team

Author:

Changwoo Ryu <cwryu@adam.kaist.ac.kr>, Raja R Harinath <harinath@cs.umn.edu>, Tom Tromey <tromey@cygnus.com>

Version:  0.20-5

gnguile GNOME Guile

A Scheme extension for GNOME programs. Documentation: GNOME online help /usr/doc/packages/gnome

Copyright:  (c) 1997, The GNOME Team

Author:

Carsten Schaar <nhadcasc@fs-maphy.uni-hannover.de>, Federico Mena Quintero <federico@nuclecu.unam.mx>, Lauri Alanko <la@iki.fi>, Marc Ewing <marc@redhat.com>, Marius Vollmer <mvo@zagadka.ping.de>, Martin Baulig <baulig@merkur.uni-trier.de>, Raja R Harinath <harinath@cs.umn.edu>, Tom Tromey <tromey@cygnus.com>

Version:  0.20-6

gnmedia GNOME Multimedia (Media)

This package contains some multimedia programs for GNOME. Documentation: GNOME online help /usr/doc/packages/gnome

Copyright:  (c) 1997, The GNOME Team

Author:

Changwoo Ryu <cwryu@adam.kaist.ac.kr>, Raja R Harinath <harinath@cs.umn.edu>, Tom Tromey <tromey@cygnus.com>

Version:  0.20-6

gnobjc GNOME Objective-C extension

With this extension package one is able to write GNOME programs in Objective-C. Documentation: GNOME online help /usr/doc/packages/gnome

Copyright:  (c) 1997, The GNOME Team

Author:

Carsten  Schaar  <nhadcasc@fs-maphy.uni-hannover.de>,  Elliot  Lee

<sopwith@cuc.edu>, Federico Mena Quintero <federico@nuclecu.unam.mx>, Kjartan Maraas <kmaraas@fib.hl.no>, Marc Ewing <marc@redhat.com>, Martin Baulig <baulig@merkur.uni-trier.de>, Miguel de Icaza <miguel@nuclecu.unam.mx>, Raja R Harinath <harinath@cs.umn.edu>, Timo Sirainen <a@sicom.fi>, Tom Tromey <tromey@cygnus.com>

Version:  0.20-5

0.1.10 GRA: All about graphics

bttv Video4Linux BT848 frame grabber driver

Driver for TV cards with a X11/Motifprogram for cards with the Bt848 chipset: Miro PCTV, Hauppauge Win/TV pci, STB PCI TV, Intel Smart Video III and the Diamond DTV2000. The driver may be allready stable, but it is a early version, so use it carefull, please! Have a lot fun with watching TV! ;-) You need a TV card for it. Read the README.SuSE and the original documentation in the directory /usr/doc/packages/bttv/, please!

Copyright:  (c) (C) 1996,97 Ralph Metzler & Marcus Metzler
Author:  Ralph Metzler <rjkm@thp.uni-koeln.de>, Marcus Metzler <mocm@thp.uni-koeln.de>
Version:  0.5.8-6

GRA: All about graphics

bttvgrab Video4Linux BTTV grab tool

BTTVGRAB is a ncurses grab tool for direct image recording from a frame grabber cards supported by the bttv driver. Can also be used for ”watching” TV via the WWW. Documentation could be find in /usr/doc/packages/bttvgrab/.

Copyright:

(c) (c) 1998 by Joerg Walter <trouble@moes.pmnet.uni-oldenburg.de>, (c) 1997,1998 by Adam Kopacz <Adam.K@idnet.de>

Author:

Joerg Walter <trouble@moes.pmnet.uni-oldenburg.de>, Adam Kopacz <Adam.K@idnet.de>

Version:  0.10.2-4

cgmdraw CGM Draw - library for generating CGM files

CGM Draw is a freely available library for generating CGM files from a C program. CGM (Computer Graphics Metafile) is a vector graphics format that can be read by many popular packages. With CGM Draw your code can quickly draw images complete with lines, arcs, rectangles, polygons, and text. CGM Draw is ideal for creating CGM files on the fly when you have a rapidly changing data set (such as in response to database queries.)

Copyright:  (c) Freely Redistributable
Author:  G. Edward Johnson <lorax@nist.gov>
Version:  1.3-7

faxedit A simple editor for pbm-files

Faxedit is a tool for editing pbm files like they are usually output or expected by fax programs. Its purpose is the easy answering of faxes. Parts of them can be cut out, rotated or resized. Input of text and the printing of the picture is also possible.

Contents: faxedit-1.14

Copyright:  (c) 1998 S.u.S.E. GmbH
Author:  Kalle Dalheimer <kalle@kde.org>
Version:  1.14-7

fbm Fuzzy Pixmap Manipulation Tools

Tools using the Fuzzy Pixmap Manipulation Library

Copyright:  (c) Rob Hooft, Michael Mauldin, (see netpbm)
Author:  Rob Hooft <hooft@EMBL-Heidelberg.DE>
Version:  1.2-4

fbm2fli Graphics Format Converter

Generation of fli-animations from fbm-files.

Copyright:  (c) 1993-1995 by Klaus Ehrenfried
Author:  Klaus Ehrenfried <klaus@es.go.dlr.de>
Version:  0.4b-4

films A few computer animation files

A small collection of FLI, GL, AVI, and MPEG files. Unfortunately, nobody knows where these films came from. In case you are one of the artists - please contact us; you will get the money back for the CD :-)

Contents: films

Version:  97.11.27-5

filmsbig Some more computer animation files

A few large computer animations. The interested viewer will no doubt discover a few well known themes. We would like to draw your attention to the film enrique.fli - one of the most beautiful films for us. More information on this film is to be found under /usr/doc/packages/filmsbig (in German). We thought it over, but we decided upon ”wasting” the space on CD for this package. We think it is worth it!..;)

Version:  97.11.27-5

freetype TrueType Font Engine

Library for working with TrueType Fonts. Documentation at /usr/doc/packages/freetype Contents: freetype-1.0

Copyright:  (c) 1996 David Turner / 1998 Robert Wilhelm / Werner Lemberg

Author:

David Turner <turner@enst.fr>, Robert Wilhelm <robert@physiol.med.tumuenchen.de>, Werner Lemberg <a7971428@unet.univie.ac.at>

Version:  1.1-3

gd an autoloadable interface module for libgd

This is a autoloadable interface module for libgd, a popular library for creating and manipulating GIF files. With this perl library you can create GIF images on the fly or modify existing files. Contents: GD.pm

Author:  Lincoln Stein <lstein@w3.org>
Version:  1.15-9

gif tools working with the gif library

convert gifs to various formats (and vice versa)

Copyright:  (c) Eric S. Raymond
Author:  Eric S. Raymond <esr@snark.thyrsus.com>
Version:  3.0-3

gimp Extendable graphic manipulation program

GIMP is the abbreviation of General Image Manipulation Program. It is a tool for creating, manipulating and saving graphical images. TIFF, JPEG, PNG, GIF and XPM and many others are supported. Contents: gimp-1.0.0, GimpUserManual-1.0.0

Copyright:  (c) 1995-98 Spencer Kimball and Peter Mattis
Author:  Spencer Kimball <spencer@xcf.berkeley.edu>, Peter Mattis <petm@xcf.berkeley.edu>
Version:  1.0.0-5

gimpman manual for gimp

manual for the gimp program as compressed postscript file.

Copyright:  (c) 1998 Olof & Karin Kylander
Author:  Olof Kylander, Karin Kylander
Version:  1.0.0-5

gltt Library for TrueType with OpenGL applications

gltt is a library that allows you to read and draw TrueType fonts in any OpenGL application. It supports bitmapped and anti-aliased font drawing as well as vectorized and polygonized drawing.

Copyright:  (c) GPL
Author:  Stephane Rehel <rehel@worldnet.fr>
Version:  2.2-3

gsview Ghostview

Ghostview offers you an X11 GUI for viewing PostScript files. This is an X11 interface to ghostscript. Contents: ghostview-1.5

Copyright:  (c) 1992 Timothy O. Theisen
Author:  Tim Theisen <tim@cs.wisc.edu>
Version:  1.5-12

gtk library for creation of graphical user interfaces

This a library that is needed for running the GIMP or SANE. Contents: gtk+-0.99.3

Copyright:  (c) 1995-98 Peter Mattis, Spencer Kimball and Josh MacDonald

Author:

Peter Mattis (petm@xcf.berkeley.edu), Spencer Kimball (spencer@xcf.berkeley.edu), Josh MacDonald (jmacd@xcf.berkeley.edu)

Version:  1.0.4-2

gv GV -- a ghostview derivate

GV offers you an X11 GUI for viewing PostScript files. This is an X11 interface to ghostscript. Contents: gv-3.5.8

Copyright:  (c) 1992 Timothy O. Theisen; 1995-97 Johannes Plass

Author:

Tim Theisen <tim@cs.wisc.edu>, Johannes Plass <plass@dipmza.physik.unimainz.de>

Version:  3.5.8-12

html2ps HTML to PostScript converter

HTML to PostScript converter written in Perl. html2ps know about many HTML 4.0 features. From the Homepage (http://www.tdb.uu.se/˜ jan/html2ps.html):

* Many possibilities to control the appearance; this is mostly done using configuration files. * Support for processing multiple documents, also automatically by recursively following links. * A table of contents can be generated, either from the links in a document, or automatically from document headings. * Configurable page headers/footers, that for example can contain document title, URL, page number, current heading, and date. * Automatic hyphenation and text justification can be selected. Contents: html2ps-1.0b1

Copyright:  (c) Jan Karrman and Jean-Philippe Argaud, GNU
Author:  Jan Karrman <jan@tdb.uu.se>, Jean-Philippe Argaud <no@e.mail>
Version:  1.0b1-11

imagemag Image / film viewer

Image Magick allows you to view and edit images. The tools read and write images in MIFF format. Other formats can be converted into MIFF. Contents: ImageMagick-4.0.1

Copyright:  (c) 1996 E. I. du Pont de Nemours and Company
Author:  John Cristy <cristy@dupont.com>
Version:  4.0.8-0

imfilm Example film for ImageMagick

An example film with a rotating DNA for demonstration of ImageMagick. The pictures are in the directory /usr/ImageMagick/scenes.

Version:  4.0.8-0

imlib Shared Library for loading and rendering 3D images

This is a general, all-purpose image loading and rendering (Image data to screen data) library. Imlib is capable of currently loading 25 different image formats. See the file /usr/include/X11/imlib.h or directory /usr/doc/packages/imlib for more details. This shared library is needed for the Enlightenment windowmanager and for GNOME. Contents: imlib DR-0.4

Copyright:  (c) 1997 Carsten Haitzler (Rasterman)
Author:  Carsten Haitzler <s2154962@cse.unsw.edu.au>
Version:  1.6-5

impict Example pictures for ImageMagick

Some example pictures for demonstration of ImageMagick. The pictures are in the directory /usr/ImageMagick/images.

Version:  4.0.8-0

jpeg Independent JPEG Group’s JPEG Software.

Software to implement JPEG image compression and decompression. JPEG pronounced ”jay-peg”) is a standardized compression method for full-color and gray-scale images. JPEG is intended for compressing ”real-world” scenes (most of the time there are pictures that have been scanned-in with a scanner). Cartoons and other non-realistic images are not its strong suit. You should note that the output of JPEG is not necessarily the same as its input. If this is a factor for you, then don’t use it. With typical real-world scenes, JPEG can achieve high compression rates, without noticeable differences. If you can accept pictures of lower quality, JPEG can achieve amazingly high compression rates. There are some library functions available for reading and writing JPEG files. The ’cjpeg’ and ’djpeg’ applications use the library to make conversions between JPEG and other popular graphics file formats possible. The library is meant to be used within other applications. Cjpeg compresses the input file, or standard input if no filename was given, and produces a JPEG/JFIF to standard output. Currently supported input file formats include: PPM (PBMPLUS color format), PGM (PBMPLUS Grayscale format), BMP, GIF, Targa and RLE (Utah Raster Toolkit Format). (RLE is only supported if the RLE library is available.) Djpeg decompresses a JPEG file in one of the above mentioned formats. Contents: jpeg-6

Copyright:  (c) 1991, 1992, 1993, 1994, Thomas G. Lane

Author:

Tom Lane <jpeg-info@uunet.uu.net>, Philip Gladstone <pjsg@ix.netcom.com>, Luis Ortiz <jpeg-info@uunet.uu.net>, Jim Boucher <jpeg-info@uunet.uu.net>, Lee Crocker <jpeg-info@uunet.uu.net>, George Phillips <phillips@cs.ubc.ca>, Davide Rossi <jpeg-info@uunet.uu.net>, Ge’ Weijers <jpeg-info@uunet.uu.net>

Version:  6a-15

libfbm Fuzzy Pixmap Manipulation Library

This is relese 1.0 of the Fuzzy Pixmap Manipulation (FBM) library. The sources are contained in the libgr source package

Copyright:  (c) Rob Hooft, Michael Mauldin, (see netpbm)
Author:  Rob Hooft <hooft@EMBL-Heidelberg.DE>, Michael Mauldin <mlm@cs.cmu.edu>
Version:  1.2-4

libgif the gif library

library to load, save and edit GIF-files

Copyright:  (c) Eric S. Raymond
Author:  Eric S. Raymond <esr@snark.thyrsus.com>
Version:  3.0-3

libgrout libgr - graphics libraries (a.out)

Shared graphics libraries with GIF, TIFF, JPEG support. These libraries are needed to run the XV image viewer when using the old a.out version.

Copyright:  (c) Rob Hooft, Michael Mauldin, (see netpbm)
Author:  Rob Hooft <hooft@EMBL-Heidelberg.DE>, Michael Mauldin <mlm@cs.cmu.edu>
Version:  1.3-1

libjpeg JPEG-Libraries

The libraries (static and dynamic) for the jpeg-graphics format. The sources are contained in the libgr source package

Copyright:  (c) Rob Hooft, Michael Mauldin, (see netpbm)
Author:  Rob Hooft <hooft@EMBL-Heidelberg.DE>, Michael Mauldin <mlm@cs.cmu.edu>
Version:  6.0.1-15

liblug multiple format graphics library

Supports (among others) GIF,PCX,TGA,RGB,TIFF,PIX,RLA,RLE,PBM,PGM,PNM,RAW,JPG

Copyright:  (c) Raul Rivero
Author:  Raul Rivero <rivero@pinon.ccu.uniovi.es>
Version:  1.0.8-6

libmpeg Mpeg library

Library with functions for access to mpeg files. Contents: mpeg lib-1.2.2

Copyright:  (c) 1992 The Regents of the University of California, 1994-95 Gregory P. Ward
Author:  Gregory P. Ward <greg@bic.mni.mcgill.ca>
Version:  1.2.2-10

libnetpb Libraries for the netpbm (NetPortableBitmap) gaphic formats

These are the libs for the netpbm graphic formats. Tools can be found in package netpbm. The sources are contained in the libgr source package

Copyright:  (c) Rob Hooft, Michael Mauldin, (see netpbm)
Author:  Rob Hooft <hooft@EMBL-Heidelberg.DE>, Michael Mauldin <mlm@cs.cmu.edu>
Version:  1.0.0-4

libpng Library for the PortableNetGraphics Format

This is the sixth (and hopefully last) beta release of libpng 1.0.

Copyright:  (c) Copyright (c) 1996, 1997 Andreas Dilger
Author:  Andreas Dilger <adilger@enel.ucalgary.ca>
Version:  1.0.1-2

librle rle-graphics library

The dynamic and static versions of librle. The sources are contained in the libgr source package

Copyright:  (c) Rob Hooft, Michael Mauldin, (see netpbm)
Author:  Rob Hooft <hooft@EMBL-Heidelberg.DE>, Michael Mauldin <mlm@cs.cmu.edu>
Version:  1.0.0-15

libtiff Tiff Library (with jpeg and compression support)

To link a program with this libtiff, you will have to add -ljpeg and -lz to include the necessary libjpeg and libz in the linking process.

Copyright:  (c) 1988-1997 Sam Leffler, 1991-1997 Silicon Graphics, Inc.
Author:  Sam Leffler <sam@engr.sgi.com>
Version:  3.4.37-17

mgp MagicPoint - a X11-based presentation tool

MagicPoint is a X11-based presentation tool. It is designed to make simple presentations easy while to make complicated presentations possible. Its presentation file (whose suffix is typically .mgp) is just text so that you can create presentation files quickly with your favorite editor (e.g. Emacs). The package also includes the tools mgp2html, mgp2ps and mgp2latex, which convert mgp-Presentations into other file formats.

Copyright:  (c) GPL
Author:  Yoshifumi Nishida <nishida@sfc.wide.ad.jp>
Version:  1.03a-0

mlcreat Moonlight Creator

Moonlight Creator is an OpenGL-based modeler and renderer (radiosity and raytracing). It is in early development stages, don’t expect too much! Documentation is below /usr/doc/packages/mlcreat. In order to test the example scenes, please copy /usr/doc/packages/mlcreat/moonlight to your home directory. Contents: ml-0.4.1 moonlight-db-0.3.1

Copyright:  (c) 1996-1998 By Stephane C.F. Rehel
Author:  Stephane C.F. Rehel <rehel@worldnet.fr>
Version:  0.4.1-9

mpegplay MPEG player

MPEG (Moving Pictures Experts Group) is a format for the compression of video and audio data. The compressed files can be played in real-time (hardware-decoder) or relatively slow (software solution) as a video sequence. This MPEG player plays video sequences using a 8, 16, or 32 bit display under X11. Contents: mpeg play-2.3-src

Copyright:  (c) 1992-93 University of California

Author:

Ketan Patel <tel@cs.berkeley.edu>, Brian Smith <bsmith@cs.berkeley.edu>, Henry Chi-To Ma <cma@cs.berkeley.edu>, Kim Man Liu <kliu@cs.berkeley.edu>

Version:  2.3-12

muscan Muscan, a scanner tool for Mustek scanners.

Muscan is a scanner program for Mustec scanners. Since there is a new project under development, muscan is not supported any more. Muscan contains Tkmuscan, a nice Tk tool which incorporates most of the features of a professional scanner tool. A list of supported scanners is to be found under: /usr/doc/packages/muscan Contents: muscan

Copyright:  (c) Copyright (C) 1995 Torsten Eichner
Author:  Thorsten Eichner <eichner@rhrk.uni-kl.de>
Version:  2.0.6-18

netpbm A powerful graphics conversion package.

The latest version of the Portable Bitmap Plus Utilities. This large package which is by now available for all hardware platforms provides tools for graphics conversion. Using these tools, images can be converted from virtually any format into any other format. A few of the supported formats include: GIF, PC-Paintbrush, IFF ILBM, Gould Scanner file, MTV ray tracer, Atari Degas .pi1 and .pi3, Macintosh PICT, HP Paintjet file, QRT raytracer, AUTOCAD slide, Atari Spectrum (compressed and uncompressed), Andrew Toolkit raster object, and many more. On top of that, man pages are included for all tools. The sources for this package are in the source-package of libgr. Contents: netpbm-1mar1994

Copyright:

(c) 1988-94 Alberto Accomazzi, Angus Duggan, Arthur David Olson, Bill Davidson, Bill Janssen, Bo Thide, Brian Moffet, Christos Zoulas, Dan Stromberg, David Beckemeyer, David Koblas, David K. Drum, David W. Sanderson, Doug Crabill, Douwe van der Schaaf, Frank Neumann, George Phillips, Ingo Wilken, James Darrell McCauley, Jamie Zawinski, Jef Poskanzer, John Tiller, Ken Yap, Mark Shand, Mark W. Snitily, Michael, Davidson, Michael Haberler, Mike Parker, Oliver Trepte, Paul Haeberli, Rainer Klute, ick Vinci, berg, Sam Leffler, Stephen Paul Lesniewski, Texas Agricultural Experiment Station, Tim Cook, Tony Hansen, Wilson Bent

Author:

Alberto  Accomazzi  <alberto@cfa.harvard.edu>,  Andre  Beck

<Andre Beck@IRS.Inf.TU-Dresden.de>, Angus Duggan <ajcd@dcs.ed.ac.uk>, Arnaud Le Hors <lehors@sophia.inria.fr>, Arthur David Olson <OLSONA@dc37a.nci.nih.gov>, Burkhard Neidecker-Lutz <neideck@nestvx.enet.dec.com>, Daniel Briggs <dbriggs@nrao.edu>, David Beckemeyer <bdt!david>, David Koblas <koblas@netcom.com>, Diomidis D. Spinellis <dds@cc.ic.ac.uk>, George Phillips <phillips@cs.ubc.ca>, Graeme W. Gill <graeme@labtam.oz.au>, Hadmut Danisch <danisch@ira.uka.de>, Ingo Wilken <Ingo.Wilken@informatik.uni-oldenburg.de>, James Darrell McCauley <mccauley@mcs.com>, John Tiller <tiller@galois.msfc.nasa.gov>, John Walker <kelvin@Autodesk.com>, Kanthan Pillay <svpillay@Princeton.EDU>, Ken Yap <ken@cs.rocester.edu>, Lyle Rains <lrains@netcom.com>, Marcel Wijkstra <wijkstra@fwi.uva.nl>, Marc Boucher <marc@PostImage.COM>, Oliver Trepte <oliver.trepte@dd.sigma.se>, Patrick J. Naughton <naughton@wind.sun.com>, Paul Haeberli <paul@manray.sgi.com>, Paul Breslaw <paul@mecazh.uu.ch>, S.Petra Zeidler <spz@specklec.mpifr-bonn.mpg.de>, Steve Belczyk <seb3@gte.com>, Warren Toomey <wkt@csadfa.cs.adfa.oz.au>, Wilson H. Bent. Jr. <whb@evtech.com>, Wolfgang Stuerzlinger <wrzl@gup.uni-linz.ac.at>

Version:  1mar94.p1-4

photocd PhotoCD image converter

This Programm includes an X11-Interface and programs for converting Photo CD images into TARGA or BITMAP (Windows 3.1) format. Documentation can be found in /usr/doc/misc/photocd.doc. Additionally included is the program ”hpcdtoppm” that converts Photo CD images int the PPM format. Contents: hpcdtoppm-0.6 photocd-1.12 xpcd-2.04

Copyright:  (c) 1994 Gerd Knorr
Author:  Gerd Knorr <kraxel@cs.tu-berlin.de>, Hadmut Danisch <danisch@ira.uka.de>
Version:  1.12-23

pictures Some pictures in GIF format.

So that you can try xv right away, we’ve put some pictures together. These pictures are installed under /usr/share/pictures. Please use xv’s special feature ”visual schnautzer” in the pictures’ directory to get an overview over the pictures. The example window manager menu contains a few items that load these pictures as desktop backgrounds.

Version:  98.3.13-3

pixmp Xpm pixel editor for X11

Pixmap is a programm which enables you to edit XPM-files (colour bitmaps). You can use them with every common used iconmanager and even incorporate them in your own desktop environment. Contents: pixmap2.6

Copyright:  (c) Copyright 1991,1992,1993,1994 - Lionel Mallet
Author:  Lionel Mallet <pixmap@sophia.inria.fr>, Tim Wise <pixmap@sophia.inria.fr>
Version:  2.6-11

plmagick perl-interface for ImageMagick

PerlMagick is an objected-oriented Perl interface to ImageMagick. Use the module to read, manipulate, or write an image or image sequence from within a Perl script. This makes it suitable for Web CGI scripts. You must have ImageMagick 4.0.8 above and Perl version 5.002 or greater installed on your system.

Copyright:  (c) 1998 E. I. du Pont de Nemours and Company
Author:  John Cristy <cristy@dupont.com>
Version:  1.42-0

povray Ray tracer

The Persistence of Vision Ray tracer creates three-dimensional, photo-realistic images using a rendering technique called ray tracing. It reads in a text file containing information describing the objects and lighting in a scene and generates an image of that scene from the view point of a camera also described in the text file. Ray tracing is not a fast process by any means, (the generation of a complex image can take several hours) but it produces very high quality images with realistic reflections, shading, perspective, and other effects.

Copyright:  (c) 1996 POV-Team(tm)

Author:

Steve Anger <sanger@hookup.net>, Dieter Bayer <100255.3074@compuserve.com>, Chris Cason <Chris.Cason@oaks.com.au>, Andreas Dilger <adilger@enel.ucalgary.ca>, Alexander Enzmann <xander@mitre.com>, Dan Farmer <74431.1075@compuserve.com>, Lutz Kretzschmar <100023.2006@compuserve.com>, Mike Miller <70353.100@compuserve.com>, Joel NewKirk <102627.1152@compuserve.com>, Anton Raves <100022.2603@compuserve.com>, Tim Rowley <trowley@geom.umn.edu>, Eduard Schwan <71513.2161@compuserve.com>, Zsolt Szalavari <zsolt@cg.tuwien.ac.at>, Timothy Wegner <twegner@phoenix.net>, Chris Young <76702.1655@compuserve.com>

Version:  3.02-20

pstoedit PostScript and PDF converter to vector grafik formats

pstoedit converts Postscript(TM) and PDF files to other vector graphic formats so that they can be edited graphically. pstoedit supports: tgif: Tgif .obj format (for tgif version >= 3) rpl: Real3D Programming Language Format lwo: LightWave 3D Object Format rib: RenderMan Interface Bytestream mif: (Frame)Maker Intermediate Format fig: .fig format for xfig xfig: .fig format for xfig pdf: Adobe’s Portable Document Format gnuplot: gnuplot format ps: Flattened PostScript debug: for test purposes dump: for test purposes (same as debug) dxf: CAD exchange format java: java applet source code idraw: Interviews draw format and: cgm: CGM Format (binary) cgmt: CGM Format (textual) if you will use the ”CGM Draw” library by G. Edward Johnson <lorax@nist.gov>. Cf. the document CGM projects at NIST http://speckle.ncsl.nist.gov/˜ lsr/cgm.htm. pstoedit needs an installed Ghostscript interpreter. Contents: cd1.3 pstoedit.2.60

Copyright:  (c) pstoedit: GPL, cd: Freely Redistributable

Author:

Wolfgang Glunz <Wolfgang.Glunz@mchp.siemens.de>, G. Edward Johnson <lorax@nist.gov>

Version:  2.60-7

sane SANE (Scanner Access Now Easy)

SANE (Scanner Access Now Easy) provides a graphical frontend to your Scanner. Various scanners are supported. Please notice that this project is still being developped. Contents: sane-0.70

Copyright:  (c) 1996, 1997, Free Software Foundation, Inc.

Author:

Andreas Beck <becka@sunserver1.rz.uni-duesseldorf.de>, David Etherton <etherton@netcom.com>, David Mosberger <David.Mosberger@acm.org>, Gordon Matzigkeit <gord@gnu.ai.mit.edu>, Michael K. Johnson <johnsonm@redhat.com>, Oliver Rauch <Oliver.Rauch@Wolfsburg.DE>, Patrick Reynolds <patrickr@virginia.edu>, Tristan Tarrant <ttarrant@etnoteam.it>

Version:  0.73-8

sced Scene editor and animator for ray tracers

Sced is a tool for three dimensional modeling. These scenes might as well be animated. Moreover it is able to export animated sequences to a renderer (e.g. PovRay, Rayshade, RenderMan, Radiance, Genray and Genscan). Support for PovRay is set by default. Documentation can be found under: /usr/doc/packages/sced. Contents: sced-0.94-

Copyright:  (c) 1995, 1996 Steve Chenney
Author:  Steve Chenney <schenney@cs.berkeley.edu>
Version:  0.94-20

smii Screen Machine II

A kernelmodule for the Screen Machine II. Contents: smii-2.0

Copyright:  (c) interpersonal-computing GmbH
Author:  interpersonal-computing GmbH <sm@ipc.de>
Version:  2.0-3

surf computer pictures of algebraic surfaces

This program was used to generate the pictures on the S.u.S.E. covers. Contents: surf-0.81

Copyright:  (c) 1996 Friedrich-Alexander-Universitaet Erlangen-Nuernberg
Author:  Stephan Endrass <endrass@mi.uni-erlangen.de>
Version:  0.81-20

tgif Vector-oriented drawing program

This program has nothing to do with the GIF format. It is an extremely powerful and easy to use vector-oriented drawing program which has every feature you would expect in a drawing utility. Contents: tgif-3.0-p12

Copyright:  (c) 1990-1996, William Chia-Wei Cheng
Author:  William Chia-Wei Cheng <william@cs.UCLA.edu>
Version:  3.0p12-12

tiff tools for converting from/to tiff

Library and support programs for the TIFF image format. The sources of this package are in the source-package of libgr. Contents: tiff-v3.4beta036

Copyright:  (c) 1988-1996 Sam Leffler, 1991-1996 Silicon Graphics Inc.
Author:  Sam Leffler <sam@engr.sgi.com>
Version:  3.4.37-17

tkdraw tkdraw

drawing tool save in eps format Contents: tkdraw

Copyright:  (c) (C) 1998 Sylvio Triebel (GPL)
Author:  Sylvio Triebel <silvio@inf-technik.tu-ilmenau.de>
Version:  1.1-4

xanim multiformat animation viewer for X

XAnim currently supports the following animation types: FLI, FLC, IFF, GIF87a, GIF89a, GIF89a animation extensions, DL, Amiga PFX, URT RLE, AVI (ULTI (16), CRAM (8 & 16), CVID (24), RGB (8), RLE8), Quicktime Animation (RAW (8), RLE (1,8,16 & 24), SMC (8), RPZA (16), CVID (16), YUV2 (24); and any combination of the above on the same command. After 11 pm it makes tea as well ;-) The program seems to be quite nice, anyway; there is even a small control panel included. Contents: xanim27064

Copyright:  (c) 1992-95 Mark Podlipec

Author:

Mark  Podlipec  <podlipec@shell.portal.com>,  Mark  Podlipec

<podlipec@wellfleet.com>

Version:  2.70.70-4

xbmbrosr xbmbrowser

view and manage X bitmap and X pixmap files Contents: xbmbrowser5.1

Author:  Ashley Roll <ash@dstc.edu.au>, Anthony Thyssen <anthony@cit.gu.edu.au>
Version:  5.1-12

xgrabsc A screengrabber for X.

You need screenshots easily? xgrab and xgrabsc give you a powerful tools to save screenshots in files (ps, xpm, etc.) or to print them. xgrabsc is invoked from a shell an sends the rectangle grabbed on the screen to standard output. xgrab is a menu driven frontend to xgrabsc. Please remember: this program will probably work with 8bpp only. Documentation may be obtained by man xgrab resp. man xgrabsc. The author’s original comments can be found in /usr/doc/packages/xgrabsc. :noitpircseD Notify: :Yfiton :Hsilgne Contents: xgrabsc-2.41

Copyright:

(c) 1990-95 Bruce Schuchardt, GemStone Systems Inc.; 1989 by the Massachusetts Institute of Technology

Author:

Bruce Schuchardt <bruce@slc.com>, Yves Arrouye, George Ferguson, David Lawrence, Alan Kent, Brian C. Smith, Johan Garpendal, Martin Birgmeier, Vasco Lopes Paulo, Hal R. Brand, Matt Freedman, Scott McCoy

Version:  2.41-13

xgrasp GRASP animation viewer

X11 GRASP animation viewer. Contents: xgrasp-1.7d

Copyright:  (c) 1991 Patrick J. Naughton
Author:  Patrick J. Naughton <naughton@sun.com>
Version:  1.7d-11

xli X11 Image Loading Utility

xli is a version of xloadimage. This utility will view several types of images under X11, or load images onto the X11 root window. xli can also be used on some (32 bit) MSDOS systems. The current version supports:

* CMU Window Manager raster files

* Faces Project images

* Fuzzy Bitmap (FBM) images

* GEM bit images

* GIF images (Including GIF89a compatibility)

* G3 FAX images

* JFIF style jpeg images

* McIDAS areafiles

* MacPaint images

* PC Paintbrush (PCX) images

* Portable Bitmap (PBM, PGM, PPM) images * Sun monochrome rasterfiles

* Sun color RGB rasterfiles

* Targa (TGA) files

* Utah Raster Toolkit (RLE) files

* X pixmap (XPM) files (Version 1, 2C and 3)

* X10 bitmap files

* X11 bitmap files

* X Window Dump (except TrueColor and DirectColor) Contents: xli-1.16

Copyright:  (c) 1992, 1993 David Koblas
Author:  David Koblas <koblas@netcom.com>
Version:  1.16-12

xmgr Plotting Tool

A plotting tool under X. Is you want to use the help function, you need Mosaic or any other html browser to read the help files. The name of the help viewer is set to the environmental variable GR HELPVIEWER. Due to legal reasons, we are not allowed to put an X11 html browser on the CD. Your friendly neighbour (the one with access to the internet) will surely be able to help you. Contents: xmgr-4.00beta.pl5

Copyright:  (c) 1995 Paul J Turner, Portland, OR
Author:  Paul J. Turner <pturner@teleport.com>
Version:  4.00pl5-13

xmorph A morphing program.

This morphing program reads two images in Targa format and computes the transitions from one image to the other (arbitrarily many steps). To avoid a simple fading effect, the program needs information about the shapes contained in the two images. This is done by using a grid consisting of connected checkpoints. The source grid is then transformed into the destination grid during the morphing process. Contents: xmorph-22jul94

Copyright:  (c) 1994 Michael J. Gourlay
Author:  Michael J. Gourlay <mike@penguin.gatech.edu>
Version:  94.7.22-10

xmrm Multi Resolution Morphing with GUI

This program provides a comfortable way to morph pictures. You’ll find a HTML-Manual in /usr/doc/packages/xmrm/manual. Just load the file mrm manual.html into your browser. Contents: xforms-0.81.ELF xmrm

Copyright:  (c) 1997 Gerhard Waldh”or, Andreas Artmann

Author:

Gerhard Waldh”or <xmrm@titus.cg.tuwien.ac.at>, Andreas Artmann <xmrm@titus.cg.tuwien.ac.at>

Version:  1.2-17

xpaint Xpaint

XPaint is a versatile bitmap/pixmap editing tool. The short features list: Usual paint operations: Brushes, Spray paint, Pencil, Lines, Arcs, Pattern Fill, Text, Boxes, Circles, Polygons. Works on multiple images simultaneously. Cut/Copy/Paste between all active images. Fatbits/Zoom on the image windows. Contents: xpaint-2.4.5

Copyright:  (c) 1992, 1993, David Koblas
Author:  David Koblas <koblas@netcom.com>, Greg Roelofs <newt@uchicago.edu>
Version:  2.4.5-17

xpdf a pdf-file viewer under X11

This tool serves for viewing files in PDF (Portable Document Format). It is used, e.g., by Adobe PDF Software ”Acrobat”. Contents: xpdf-0.7

Copyright:  (c) 1996 Derek B. Noonburg
Author:  Derek B. Noonburg <derekn@ece.cmu.edu>
Version:  0.7a-1

KDE: K Desktop Environment

xray XRay

XRay is a program in which you can visualize 3D bodies with various optical properties. It can handle objects such as polygons, spheres, ellipsoids and torii. This tool uses the following light sources: point, ray and surface. You can represent your data using the following algorithms: z-buffer, raytracing, enhanced radiosity or a two phase method. The package contains two main programs. One generates the scene, and the other is the user interface for the raytracer. If you have installed the linuxthread package, you can use the mulithreaded version. Before starting please read the documentation in /usr/doc/packages/xray/doku/ . The documentation in PostScript format is in the package xrayps.tgz. Contents: xray

Copyright:  (c) Helmut Fahrion
Author:  Helmut Fahrion <hf@suse.de>
Version:  98.5.7-3

xrayps XRay PostScript Documentation

This ist the documentation for XRay in PostScript format. This package is very large, 100Mb. Please install this package only if you want to print the documentation in book form. Contents: xray

Copyright:  (c) Helmut Fahrion
Author:  Helmut Fahrion <hf@suse.de>
Version:  98.5.7-3

xwarp Warp an image (GIF format)

Warp an image (GIF format) defining 2 grids (source and goal) under X11. The grid can be modified with the mouse. Example files can be found in /usr/local/lib/XWarp-1.00-example/ . Contents: XWarp-1.00

Author:  Luca Maranzano, Guido Cardino, Roberto Surlinelli, <calamaro@dist.dist.unige.it>
Version:  1.0-12

zgv Viewer (libvga)

Zgv is a GIF/JPEG/PNM/BMP graphic file viewer for VGA and SVGA displays. (It uses svgalib.) It provides a graphic-mode file selector to select file(s) to view. The program may be used as a command; in this case, the front end is not used. Contents: zgv2.8-src

Copyright:  (c) 1993, 1994 Russell Marks
Author:  Russell Marks <mr216@gre.ac.uk>
Version:  2.8-19

0.1.11 KDE: K Desktop Environment

kadmin KDE system administration

KDE is still in an early development stage (beta). Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). This package contains kdeadmin, some tools which are useful to administrate the system. Documentation in /usr/doc/packages/kde/kadmin Contents: kdeadmin-980407

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  1.0-6

kgames KDE games

Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). Some nice games for KDE.

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  1.0-5

kgoemon Patch set for kmidi

Using this patch-set will enlarge the functionality of the program kmidi.

Author:  goemon@anime.net
Version:  1.0-4

kgraph KDE graphics programs

Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). This package contains some graphics programs which sometimes can replace traditional programs: kdvi, kfax, kfract, kghostview, kpaint, kview

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  1.0-5

kicons Some usefull icons for KDE

With this package you have some very usefull icons for the KDE windowmanager Contents: kdeicons

Author:  Marcus Sch”afer <ms@suse.de>
Version:  1.0-2

klibs KDE base package (libs)

Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). This package contains kdelibs, one of the basic packages of the K Desktop Environment. It contains necessary libraries the KDE desktop. This package is absolutely necessary for using KDE! Documentation may be found in: /usr/doc/packages/kde/klibs

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  1.0-5

kmulti KDE-multimedia package

Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). This package contains some multimedia programs: kmix, kmedia, kscd

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  1.0-4

knet KDE networking package

Moreover, we cannot give technical support for KDE matters directly but we’d like to appoint you to the respective addresses in WWW (http://www.kde.org) or in case to our support database (http://www.suse.de/Support/sdb).

A mailer and a news reader for KDE.

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  1.0-5

ksupp KDE base package (support)

Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). This package contains kdesupport, one of the basic packages of the K Desktop Environment. It contains necessary libraries for managing graphics, mail and html processing. This package is absolutely necessary for using KDE! Documentation may be found in: /usr/doc/packages/kde/ksupp Contents: kdesupport-980404

Copyright:  (c) 1997,98 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  1.0-4

kutils KDE utility package

Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). Some utilities for KDE: karm, kcalc, kedit, khexdit, kjots, knotes, kzip

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  1.0-6

kbase KDE base package (base)

Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). This package contains kdebase, one of the basic packages of the K Desktop Environment. It contains among others KWM, the KDE window manager; KFM, the KDE file manager; KDisplay, the desktop configuration program. This package is absolutely necessary for using KDE! Documentation may be found in: /usr/doc/packages/kde/kbase

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  1.0-9

kdir KDE menu entries

Alpha stage! KDE is still in an early development stage (beta). Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is especially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). This package contains directories for the KDE Alpha Menu structure. Absolutely necessary! Alpha stage!

KDE is still in an early development stage (beta). Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is especially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb).

Copyright:  (c) 1998, S.u.S.E. GmbH
Author:  Marcus Sch”afer <ms@suse.de>, Martin Scherbaum <maddin@suse.de>
Version:  alpha-1

kadm KDE-Admin (Alpha!)

Alpha stage! Some KDE packages are still in an early development stage. Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is especially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). This package contains some additional admin tools. Documentation in: /usr/doc/packages/kde/kadm Contents: kadm klpc-0.1.3 Alpha stage! Some KDE packages are still in an early development stage. Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is especially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb).

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  alpha-19

kdevel KDE development tools (Alpha!)

Alpha stage! Some KDE packages are still in an early development stage. Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is especially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). Contains some tools to administrate software projects. Documentation in: /usr/doc/packages/kde/kdevel Contents: kdevel ktranslator-0.1.9 Alpha stage! Some KDE packages are still in an early development stage. Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is especially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb).

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  alpha-20

kmisc Some KDE tools (Alpha!)

Alpha stage! Some KDE packages are still in an early development stage. Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is especially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). Contains some very practical tools. Documentation in: /usr/doc/packages/kde/kmisc Contents: kmodplayer-0.3.6 konquest-0.0.3 kpov-0.1.5 krecord-0.4 ktamaga-0.2 Alpha stage! Some KDE packages are still in an early development stage. Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is especially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb).

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  alpha-22

knetutil KDE network utilities (Alpha!)

Alpha stage! Some KDE packages are still in an early development stage. Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is especially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). Contains some additional network tools. Documentation in: /usr/doc/packages/kde/knetutil Contents: kISDN-0.3.1 kfinger-0.6.1 khostchooser-0.2.2 kmodem0.0.3 knetutil ksamba-0.3 ktalkd-0.2.0 Alpha stage! Some KDE packages are still in an early development stage. Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is especially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb).

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  alpha-20

kscience KDE scientific tools (Alpha!)

Alpha stage!

Some KDE packages are still in an early development stage. Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is especially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). Some scientific tools. For example a simple fractal viewer. Documentation in: /usr/doc/packages/kde/kscience Contents: kplot-0.2 kscience Alpha stage! Some KDE packages are still in an early development stage. Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is especially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb).

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  alpha-19

ktools KDE tools (Alpha!)

Alpha stage! Some KDE packages are still in an early development stage. Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is especially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb). Some practical tools which you may need every day. For example: kdayplan Documentation in: /usr/doc/packages/kde/ktools Contents: kapm-0.0.2 kdbg-0.03 kgoodstuff-0.1.2 kless klyx-0.9.2 korganizer-0.9 ksendfax-0.1 ksnapshot-0.2.3 ktools ktop-0.9.7 kvmode-0.0.2 Alpha stage! Some KDE packages are still in an early development stage. Thus, KDE being relatively stable, it may nevertheless happen that any of the KDE applications crashes or produces strange effects. This package is especially concerned since it was built from not yet officially released programs. Newest KDE packages adepted to S.u.S.E. Linux may be found at ftp://ftp.suse.com/pub/suse update/KDE. General information for KDE can be found at the respective addresses in WWW (http://www.kde.org) or our support database (http://www.suse.de/Support/sdb).

Copyright:  (c) 1997,1998 The KDE-Team
Author:  The KDE Team <kde@kde.org>
Version:  alpha-20

0.1.12 N: Network-Support (TCP/IP, UUCP, Mail, News)

acmennrp acmennrp

ACME Labs Multithreaded NNTP Proxy Daemon written in Java

Copyright:  (c) Copyright (C)1996,1998 by Jef Poskanzer <jef@acme.com>. All rights reserved.
Author:  Jef Poskanzer <jef@acme.com>
Version:  19980618-1

N: Network-Support (TCP/IP, UUCP, Mail, News)

amanda Network Disk Archiver

Please note that this software is ALPHA.

Contents: amanda-2.4.0b6

Copyright:  (c) 1991, 1995 University of Maryland at College Park
Author:  James da Silva <jds@cs.umd.edu>
Version:  2.4.0b6-13

amd Automounter daemon

Automounter: mounts filesystems automatically. The daemon manages UFS and NFS mounts, and can mount and unmount filesystems based on demand. Contents: amd-upl102 This package is not maintained anymore, please use autofs instead.

Author:  Jan-Simon Pendry <jsp@doc.ic.ac.uk>
Version:  upl102-21

apache Apache httpd Web server

The Apache HTTP server. Very few programs have spread so quickly and attracted such a large following in such a short time as this HTTP server. According to Netcraft, Apache runs on about 49.05%% of all servers attached to the Internet. That represents about 1,182,142 systems. It has replaced within only a few months the NCSA server which, until then was the unchallenged favorite. This is undoubtedly because Apache was derived from the NCSA server sources and that its configuration is compatible to that of NCSA/1.3. The Apache project was established to provide a forum for the development of a freely available HTTP server. The goal is to develop a secure, efficient, and easily extendable server, which also closely parallels the development of the HTTP protocol. Thus Apache became a replacement to NCSA/1.3. Besides better performance, the removal of known errors, and better conformance to HTTP standards (HTTP/1.1), Apache offers the following useful features, which are often missing in other servers:

* password files, used to protect documents from unauthorized access, can be indexed in various ways. The search for an entry is as a result more efficient than in ordinary text files.

* client-request errors can be responded to in a flexible manner. CGI scripts can be used to analyze the error and generate more informative replies to both the client and the administrator than the HTTP status codes offer.

* there is no restriction on aliasing and redirection. The Alias and Redirect directives can be used in arbitrary number and depth.

* Apache can choose the best representation of a resource based on the browser-supplied preferences for media type, languages, character set and encoding (”Content Negotiation”).

* with Apache you can implement several virtual servers. Thus requests to different IP addresses (referring to the same computer) can be distinguished and, depending on the IP address, appropriately handled. Further advantages are the module concept and the well-documented API (Application Programmers Interface). Extensions are easily developed and new features are easily integrated into the server. For example, an implementation of the Secure Socket Layer (SSL) from Netscape exists for Apache. This permits secure communication between server and client. Apache is installed in the following directories: /etc/httpd -> configuration files /usr/sbin/httpd -> server binary /var/log/httpd.* -> protocol files /var/run/httpd.pid -> server PID /usr/local/httpd -> server root /sbin/init.d/apache -> Start/Stop script Contents: apache 1.3.0, mod perl

Copyright:  (c) 1995-98 The Apache Group.

Author:

Rob McCool <robm@netscape.com>, Robert S. Thau <rst@ai.mit.edu>, Roy T. Fielding <fielding@avron.ics.uci.edu>, Brian Behlendorf <brian@wired.com>, Harald Hanche-Olsen <hanche@imf.unit.no>, David Robinson <kluft@amdahl.com>, Wietse Venema <drtr@ast.cam.ac.uk>, Keith Shafer <Keith.Shafer@SanDiegoCA.ATTGIS.COM>, Kevin Steves <stevesk@mayfield.hp.com>

Version:  1.3.0-10

apassl apassl - Apache with SSL and PHP3 support

This is the apache WWW server enhanced with SSL from Ben Laurie and the DHTML language PHP3 for ADABAS D database access. Contents: apache, SSL, PHP3 To use apacheSSL a certificate has to be set up.

Copyright:  (c) 1995-1997 The Apache Group. / 1995-98 Ben Laurie

Author:

Rob McCool <robm@netscape.com>, Robert S. Thau <rst@ai.mit.edu>, Roy T. Fielding <fielding@avron.ics.uci.edu>, Brian Behlendorf <brian@wired.com>, Harald Hanche-Olsen <hanche@imf.unit.no>, David Robinson <kluft@amdahl.com>, Wietse Venema <drtr@ast.cam.ac.uk>, Keith Shafer <Keith.Shafer@SanDiegoCA.ATTGIS.COM>, Kevin Steves <stevesk@mayfield.hp.com> , Ben Laurie <ben@algroup.co.uk>

Version:  1.2.6-7

archie Where do I find what?

Using this program, you can query archie servers without having to login with telnet. This version uses the server archie.funet.fi as default. You may specify another host using the command line switch ”-h”. Contents: archie-1.4.1

Copyright:  (c) 1989-91 University of Washington

Author:

Alan Emtage <bajan@cs.mcgill.ca>, Peter Deutsch <peterd@cs.mcgill.ca>, Bill Heelan <wheelan@cs.mcgill.ca>, Clifford Neuman<bcn@isi.edu>, Brendan Kehoe <brendan@cygnus.com>,

Version:  1.4.1-13

argus Networkmonitoring Tool

Argus is a networkmonitoring tool start und parameters, look at /etc/rc.config information and documentation, look at /usr/doc/packages/argus Contents: libpcap-0.4a6.tar.Z tcp wrappers 7.4.tar.gz

Copyright:  (c) 1997 Carter Bullard
Author:  Carter Bullard <argus@sei.cmu.edu>
Version:  1.7.beta.1b-6

autofs autofs is a kernel-based automounter

autofs is a kernel-based automounter for Linux. It is still under development, although the features currently implemented seem to work well. There is not anywhere near enough documentation available yet; help in getting some written would be appreciated. This package replaces the amd package. Contents: autofs-3.1.0

Copyright:  (c) GPL, 1997 Transmeta Corporation
Author:  H. Peter Anvin <hpa@transmeta.com>
Version:  3.1.0-11

bind BIND - Name Server Utilities

The named daemon and support utilities including: dig, dnsquery, host, and nslookup. Documentation on setting up a name server can be found in /usr/doc/packages/bind . Contents: bind-4.9.7

Copyright:  (c) 1989 The Regents of the University of California.
Author:  Paul Vixie <vixie@vix.com>
Version:  4.9.7-13

bind8 BIND - Name Server Utilities (new version)

The named daemon and support utilities including: dig, dnsquery, host, and nslookup. Documentation on setting up a name server can be found in /usr/doc/packages/bind .

Copyright:  (c) 1989 The Regents of the University of California.
Author:  Paul Vixie <vixie@vix.com>
Version:  8.1.2-3

bing point-to-point bandwidth measurement tool

Bing determines the real (raw, as opposed to available or average) throughput on a link by measuring ICMP echo requests roundtrip times for different packet sizes for each end of the link. Contents: bing-1.0.4

Copyright:  (c) GPL
Author:  Pierre Beyssac <pb@fasterix.freenix.fr>
Version:  1.0.4-12

bulkmail bulk mailer

This is a C program to do ”bulk” mailing. For input, it takes a file of recipient addresses (one address per line) and a message (with headers already attached) to be sent to the recipients. It sorts the recipient list by reversed domain (so similar ones sort together), splits up the recipients into several groups containing no more than N domains each, creates an SMTP envelope for each group of recipients, and feeds that envelope to ”/usr/lib/sendmail -bs”. Splitting the envelopes up allows sendmail to perform delivery in parallel, so instead of having one large queue entry (for which sendmail might take awhile to get around to attempting delivery for some recipients), it has several smaller queue entries. Depending on your point-of-view, this can still be considered ”cluttering up your mail queue”, but it does seem to deliver messages more quickly to most recipients. The core of this program was extracted from a somewhat strange mailing list manager called na-net; it was designed to efficiently send out mail to 5000 people at a time. I have used this program to attempt delivery of a message to over 12000 recipients around the world, within a few hours. I’m currently using bulk mailer as a back-end for several mailing lists of modest size. However, the program is not extensively tested, and may not work well in all environments. (In particular, if your system has per-user process quotas, or a small number of process table entries, you will want to modify this program to recover gracefully.) There’s no warranty on this, but you’re welcome to use it if you want. Contents: bulk mailer-1.5

Author:  Keith Moore <moore@cs.utk.edu>
Version:  1.5-10

cipe Encrypted IPIP-Tunnel

CIPE (the name is shortened from *Crypto IP Encapsulation*) is a package for an encrypting IP tunnel device. This can be used to build encrypting routers for VPN (Virtual Private Networks) and similar applications. Contents: cipe-0.5.6

Copyright:  (c) 1996-1997 Olaf Titz
Author:  Olaf Titz <olaf@bigred.inka.de>
Version:  0.5.6-6

cnews Cleanup Release of C News.

Controls the spooling and transmission of Usenet news. Cnews is typically used with UUCP -- for NNTP connections INN is usually a better choice. To activate the cnews-system, you have to edit the crontab belonging to root or news. An example crontab can be found under /var/lib/news/crontab. Please read the file /var/lib/news/README.SuSE as well. Contents: c-news nntp-1.5.12.1 Please note, that this package is not supported by our support team. Please use inn if you need support for news. Feedback and bug fixes for cnews are of course welcome at feedback@suse.de.

Copyright:  (c) 1985-92 University of Toronto
Author:  Geoffrey Collyer
Version:  1.G-10

deliver A program for the delivery of email

A small and simple program that delivers electronic mail to local users. If you are attached to a network, you should install the program ”sendmail” as well.

Contents: deliver-2.1.13

Author:  Chip Salzenberg <chip@fin.uucp>
Version:  2.1.13-6

dhcp ISC DHCP server

Version 1of the ISC DHCP Distribution includes just a DHCP Server. Version 1 has been in feature freeze since late 1996, and is quite stable. This is the release that we would expect most sites to run in production. Please read the documentation in /usr/doc/packages/dhcp regarding configuration of the DHCP server.

Contents: dhcp-1.0pl2

Copyright:  (c) 1995, 1996, 1997 The Internet Software Consortium.
Author:  Ted Lemon <mellon@fugue.com>
Version:  1.0pl2-7

diald Automatic dialer

Demand dialing daemon for SLIP and PPP links. Supports demand dialed links with link timeouts and redialing. More info: /usr/doc/packages/diald Example configuration: /usr/doc/packages/diald/config

Contents: diald-0.14 diald-top-0.6

Copyright:  (c) 1994, 1995, 1996 Eric Schenk
Author:  Eric Schenk <schenk@cs.toronto.edu>
Version:  0.16-4

dochost Documentation server

With this package you are able to access a central documentation-server on which the S.u.S.E. onlinesupport system is stored. It should be installed on the server, too.

Contents: dochost-98.03.02

Copyright:  (c) Coypright (c) 1998 S.u.S.E. GmbH Fuerth, Germany
Version:  98.07.02-2

elm Menu-driven user mail program.

Elm is an interactive screen-oriented mailer program originally written by Dave Taylor. Elm is now maintained by the Elm Development Group.

Contents: elm2.4.ME+.37

Copyright:  (c) 1986,1987 Dave Taylor, 1988-1992 USENET Community Trust
Author:  Dave Taylor <taylor@intuitive.com>
Version:  2.4.37-15

emil Email filter

This program enables you to convert MIME, SUN mailtool and ’plain old style RFC822’. Could be interesting for ’elm addicted’ people..;-)

Contents: emil-2.1.0-beta9

Version:  2.1.0beta9-13

firewall Firewall and masquerading scripts

Two scripts to configure your machine as a firewall and/or masquerading router. See /usr/doc/packages/firewall for more details.

Copyright:  (c) 1997 S.u.S.E. GmbH
Author:  Bodo Bauer <bb@suse.de>, Volker Graf <volker.graf@student.uni-tuebingen.de> (fcii)
Version:  1.4.3-6

fortify Fortify extends your Netscape/Communicator to use 128-bit cryptography

This is Fortify for Netscape, a program that provides world-wide, unconditional, full strength 128-bit cryptography to users of Netscape Navigator (v3) and Communicator (v4).

Contents: Fortify-1.2.3-unix

Copyright:  (c) 1997 Farrell McKay
Author:  Farrell McKay <contact@fortify.net>
Version:  1.2.3-1

ftpdir A sample ”FTP configuration”

This package contains a ’change-root’-configuration for anonymous ftp.

Contents: fileutils-3.16 ftpdir gzip-1.2.4 ncompress-4.2.4 tar-1.11.8

Version:  98.8.4-0

gettyps Just another getty.

This getty is included because it supports ring-back. Documentation: man getty man gettydefs under /usr/doc/packages/gettyps you can find some examples.

Contents: getty ps-2.0.7i

Copyright:  (c) 1989,1990 by Paul Sutcliffe Jr.
Author:  Kris Gleason <gleasokr@boulder.colorado.edu>
Version:  2.0.7i-14

gup Group Update Program for INN and C-News

gup - A Group Update Program that accepts commands by mail to edit a newsgroup subscription file for subsequent use by news systems such as INN and C-News

Contents: gup

Copyright:  (c) 1993 Mark Delany & Andrew Herbert

Author:

Andrew Herbert <andrew@werple.apana.org.au>, Mark Delany <markd@bushwire.apana.org.au>

Version:  0.3-13

htdig WWW indexing and searching system

The ht://Dig system is a complete world wide web indexing and searching system for a small domain or intranet. This system is not meant to replace the need for powerful internet-wide search systems like Lycos, Infoseek, Webcrawler and AltaVista. Instead it is meant to cover the search needs for a single company, campus, or even a particular sub section of a web site. As opposed to some WAIS-based or web-server based search engines, ht://Dig can span several web servers at a site. The type of these different web servers doesn’t matter as long as they understand the HTTP 1.0 protocol.

Contents: htdig-3.0.8b2

Copyright:  (c) ht://Dig 1995, 1996, 1997 Andrew Scherpbier (GPL)
Author:  Andrew Scherpbier <andrew@contigo.com>
Version:  3.0.8b2-18

hylafax HylaFAX(tm) is a very complex and powerful fax server

HylaFAX(tm) is a very powerful fax server for UNIX systems. It includes the following features: Fax-Mail-Gateway (dispatching faxes to Email addresses via TSI) Mail-Fax-Gateway Adaptive Answer Support (dispatching of data or voice calls to external program’s like agetty, mgetty or vgetty) Outgoing documents can be any format Before executing the script ’faxsetup’ please DEFINITELY read /usr/doc/packages/hylafax/README.SuSE!!! Documentation can be found at /usr/doc/packages/hylafax/html.

Contents: hylafax-v4.0pl2 latex-cover-1.04 Before you can use this package you have to run ‘faxsetup‘ as root.

Copyright:  (c) 1988-1996, 1991-1996 Sam Leffler, Silicon Graphics, Inc.
Author:  Sam Leffler <sam@engr.sgi.com>
Version:  v4.0pl2-13

i4l ISDN for Linux

You need these tools to use an ISDN card under Linux. An explicite explanation will be found under /usr/doc/packages/i4l/README.SuSE and in /usr/doc/packages/i4l/README.Quick. Contents: isdn4k-utils-2.1b1 , isdnmon, isdnmonp, imon, xisdnload, isdnlog, imontty, isdnbar

Copyright:  (c) Fritz Elfert and more

Author:

Fritz Elfert <fritz@wuemaus.franken.de>, Karsten Keil <keil@temicech.spacenet.de>, Michael Hipp <Michael.Hipp@student.uni-tuebingen.de>, Andreas Kool <akool@Kool.f.EUnet.de>, Ullrich von Bassewitz <uz@musoftware.com>, Stefan Luethje <luethje@sl-gw.lake.de>, Carsten Paeth <calle@calle.in-berlin.de>

Version:  98.7.18-4

i4lfirm ISDN-firmware for active ISDN-cards

Firmware files for active ISDN-cards (ICN and AVM-B1)

Copyright:  (c) AVM, ICN
Version:  98.7.18-4

inetcfg Examples for SLIP and PPP-scripts

Some examples for DIP and PPP Scripts (Compuserve and German T-Online). The files live in /usr/doc/packages/inetcfg. Contents: inetcfg

Copyright:  (c) S.u.S.E. GmbH, Fuerth, Germany
Author:  Klaus Franken <kfr@suse.de>
Version:  98.2.3-5

inn Inter Net News

Rich Salz’s InterNetNews news transport system. Contents: inn-1.7

Copyright:  (c) 1983-1990 The Regents of the University of California.

Author:

Rich Salz <rsalz@osf.org>, Andreas Klemm <andreas@knobel.gun.de>, Vince Skahan <vince@atc.boeing.com>, Arjan de Vet <devet@adv.win.tue.nl>

Version:  1.7-12

ipfwadm Firewall and IP-accounting administration

Using this package you can activate, configure and maintain ’firewalling’ of the Linux kernel. Documentation: man ipfwadm Contents: ipfwadm-2.3.0

Copyright:  (c) 1995,1996 by X/OS Experts in Open Systems BV
Author:  Jos Vos <jos@xos.nl>
Version:  2.3.0-18

iptraf iptraf

TCP/IP Network Monitor

Copyright:  (c) GPL
Author:  Gerard Paul Java <riker@mozcom.com>
Version:  1.2.0-7

ipxrip IPX routing daemon

This is a RIP/SAP daemon for Linux. You can turn your Linux machine into an IPX router. Contents: ipxripd-0.7

Copyright:  (c) 1994, 1995 Ales Dryak; 1996, Volker Lendecke
Author:  Ales Dyrak <A.Dryak@sh.cvut.cz>, Volker Lendecke <lendecke@namu01.gwdg.de>
Version:  0.7-19

ircd Internet Relay Chat Server

Ircd is the server (daemon) program for the Internet Relay Chat Program.

Version:  2.9.5-6

ircii Internet Relay Chat

This is a version of IRC (Internet Relay Chat) for systems with direct TCP/IP access (including SLIP/PPP). It also includes voice support and other extra features. Contents: ircii-2.9

Copyright:  (c) 1990 Michael Sandrof

Author:

Michael  Sandrof  <ms5n+@andrew.cmu.edu>,  Matthew  Green

<mrg@mame.mu.oz.au>

Version:  2.9-14

irciihlp Help files for ircII

Version:  2.9-14

ldap A LDAP-Server

The Lightweight Directory Access Protocol (LDAP) is a protocol for accessing online directory services. It runs directly over TCP, and can be used to access a standalone LDAP directory service or to access a directory service that is back-ended by X.500

Copyright:  (c) 1992-1996 Regents of the University of Michigan
Author:  ldap support <ldap-support@umich.edu>
Version:  3.3-1

lprold BSD Print Spooling System

This is the Linux version of the BSD-compatible printer spooling system. The lpd binary included in this package expects to find TCP/IP running. Contents: lpr-tlr-971016

Copyright:  (c) 1983, 1991 The Regents of the University of California.
Author:  Thomas Roessler <roessler@guug.de>
Version:  3.0.1-14

lynx A text-based html browser

Lynx is an easy to use browser for html documents. It is purely text based and therefore makes it possible to use WWW resources on text terminals. We must mention that this program does not fall under the GPL license and that the Copyright-owner maintains the right to require the program to become licensed if used for commercial use. Contents: lynx2-7-2

Copyright:  (c) 1994 University of Kansas
Author:  Michael Grobe <grobe@kuhub.cc.ukans.edu>, Lou Montulli
Version:  2.8-2

mailx The BSD mailx mailer.

This program is a very good supplement to ”mail” and should be installed. Contents: mailx-8.1.1

Copyright:  (c) 1980-93 Regents of the University of California.
Version:  8.1.1-2

marsnwe Novell server emulation

Complete emulation of a Novell server. Supports file-service, bindery-service, print-service and routing-service. Your kernel has to be configured according to /usr/doc/packages/marsnwe/INSTALL. Contents: mars nwe-0.99.pl4

Copyright:  (c) Martin Stover
Author:  Martin Stover <mstover@freeway.de>
Version:  0.99.pl4-14

mdomo Majordomo

This is your majordomo. Provides an easy to configure means of maintaining mailing lists including user-administration. Contents: majordomo-1.94.4

Copyright:  (c) 1992-1995 D. Brent Chapman, John P. Rouillard

Author:

D. Brent Chapman <Brent@GreatCircle.COM>, Alan Millar <amillar@bolis.sfbay.org>, Paul Haas <paulh@hamjudo.mi.org>, Jon Luini <jon@luini.com>, John P. Rouillard <rouilj@dstar.iddis.com>

Version:  1.94.4-11

metamail MIME Mail-Handler

Metamail is required for reading multimedia mail messages (such as those using the Andrew toolkit) with elm. Contents: metamail-2.7-19

Copyright:  (c) 1991 Bell Communications Research, Inc.
Version:  2.7.19-14

mgetty mgetty

This package turns your computer into a fax machine. With a Zyxel modem, you can even use your computer as an answering machine. mgetty recognizes what kind of call it is receiving and does everything else automatically. You will find the documentation under /usr/doc/mgetty and TeX-Info files under /usr/info. The configuration files are in /usr/lib/mgetty+sendfax and /etc/mgetty+sendfax. Contents: mgetty-1.1.8

Copyright:  (c) 1993-1997 Gert D”oring
Author:  Gert D”oring <gert@greenie.muc.de>
Version:  1.1.8-20

mh Unix Mail Handler

MH is a popular mail handling system but includes only a command line interface. This disadvantage is on the other hand its most advantage over other graphical mail systems. Contents: mh-6.8.4

Author:

Michael Alan Dorman <mdorman@calder.med.miami.edu>, David Morris <bweaver@worf.netins.net>, Philip Hands <phil@hands.com>, Adam P. Harris <aph@debian.org>

Version:  6.8.4debian25-1

minicom minicom

Terminal program similar to Telix(tm) (program for calling other computers via modem) under MS-DOS. If you want to access your Modem with minicom, you have to be in the group uucp.

Copyright:  (c) 1991-1996 Miquel van Smoorenburg
Author:  Miquel van Smoorenburg <miquels@drinkel.ow.nl>
Version:  1.81.1-3

mininews Inews - post news from an NNTP client

Copyright:  (c) 1983-1990 The Regents of the University of California
Author:  Rich Salz <rsalz@osf.org>
Version:  1.25-1

mirror Perl scripts for mirroring ftp servers.

To use mirror, you must have done the following once: cd /usr/include; h2ph * sys/* linux/* Documentation: ”man mirror” and ”man mm”. Contents: mirror-2.8.f4

Copyright:  (c) 1993 Lee McLoughlin
Author:  Lee McLoughlin <lmjm@doc.ic.ac.uk>
Version:  2.8.f4-13

mrtg MRTG - The Mutli Router Traffic Grapher

MRTG - The Mutli Router Traffic Grapher A tool to visualise network traffic via a WebPage. Contents: gd1.2 mrtg-2.5.2

Copyright:  (c) 1996 Tobias Oetiker and all the Contributers to MRTG
Author:  Tobias Oetiker <oetiker@ee.ethz.ch> , Dave Rand <dlr@bungi.com>
Version:  2.5.2-6

mutt Mail program

Copyright:  (c) 1996 Michael R. Elkins
Author:  Michael Elkins <me@cs.hmc.edu>
Version:  0.93.1i-2

ncftp A comfortable ftp program

Author:  Mike Gleason <mgleason@cse.unl.edu>
Version:  2.4.2-13

ncpfs Tools for accessing Novell filesystems

You can mount Netware-server filesystems under Linux using this little tool. You can even print Linux documents on printers attached to the Novell server. Contents: ncpfs-2.0.10

Copyright:  (c) 1995 by Volker Lendecke
Author:  Volker Lendecke <lendecke@namu01.gwdg.de>
Version:  2.0.10-18

netacct netacct

This package logs network traffic. It provides a daemon (nacctd) that logs all traffic passing the machine it runs on (similiar to what tcpdump does).

Copyright:  (c) GPL
Author:  Ulrich Callmeier <uc@brian.lunetix.de>
Version:  0.4-8

netatalk Netatalk - Appletalk for Linux

Copyright:  (c) 1990,1996 Regents of The University of Michigan
Author:  Wesley Craig <netatalk@umich.edu>, Ann Arbor <netatalk@umich.edu>
Version:  1.4b2+asun2.0a18.2-4

netcat a simple but powerful network tool

Netcat is a simple Unix utility which reads and writes data across network connections, using TCP or UDP protocol. It is designed to be a reliable ”back-end” tool that can be used directly or easily driven by other programs and scripts. At the same time, it is a feature-rich network debugging and exploration tool, since it can create almost any kind of connection you would need and has several interesting built-in capabilities. You can find the documentation in /usr/doc/packages/netcat/README. Contents: netcat-1.10

Copyright:  (c) hobbit@avian.org
Author:  hobbit@avian.org
Version:  1.10-5

netdiag netdiag

Enables hardware level diagnosis without booting Dos Contents: 3c515 at1700-diag at1700 atp-diag e21 eepro100 eexpress el3 epic-diag hp+ mii-diag mii look ne2k-pci-diag ne2k pci-config rtl8139-diag tulip-diag vortex-diag yellowfin-diag

Copyright:  (c) Donald Becker
Author:  becker@cesdis.gsfc.nasa.gov
Version:  1.0-2

nn The nn news reader

This is the nn news reader (NetNews). It is configured for use as NNTP news reader. Contents: nn-6.5.1

Copyright:  (c) 1990 Kim F. Storm
Author:  Kim F. Storm <storm@texas.dk>
Version:  6.5.1-15

nn spool The nn news reader

This is the nn news reader (NetNews). It is configured to read the local /var/spool/mail directory.

Copyright:  (c) 1990 Kim F. Storm
Author:  Kim F. Storm <storm@texas.dk>
Version:  6.5.1-10

nntpd NNTP server daemon.

This NNTP server is necessary if you plan reading news with the cnews package. It has been configured so that it accesses the /var/spool/news directory to supply news to the news reader. If you want to compile the sources yourself please keep in mind to install and compile cnews first.

Please note, that this packages requires cnews. The packages cnews is not supported by our support team. Feedback and bug fixes for nntpd are of course welcome at feedback@suse.de.

Copyright:  (c) 1987 Regents of the University of California, 1989 Stan Barber, 1992 by Iain Lea

Author:

Brian Kantor <brian@ucsd.edu>, Stan Barber <nntp@academ.com>, Iain Lea <iain@ecrc.de>

Version:  1.5.12.1-10

pgp PGP encryption software

PGP is the abbreviation of ”Pretty Good Privacy”. This package contains a very good software to encrypt your private data.

Contents: pgp263is pgpdoc-html pgpdoc.html

Copyright:  (c) 1990-96 Philip Zimmermann
Author:  Philip Zimmermann <prz@acm.org>, Stale Schumacher <stale@hypnotech.com>
Version:  2.6.3i-13

pine The mail program Pine

A menu-driven user mail program. Originally based on Elm (Pine Is No-longer Elm), but with many additional features that make it easier to use. There is a completely revised Version 4, that has been released in July. Since we were not able to test this version very intensively, this older version will be installed as the default. However, we have made a package pine4, which enables you to have a look at the new version.

Contents: pine3.91 pine3.96

Copyright:  (c) 1986-94 University of Washington

Author:

Mike  Seibel  <mikes@cac.washington.edu>,  Steve  Hubert

<hubert@cac.washington.edu>, Laurence Lundblade <lgl@cac.washington.edu>

Version:  3.96-15

pine4 The mail program Pine (new version)

A menu-driven user mail program. Originally based on Elm (Pine Is No-longer Elm), but with many additional features that make it easier to use. This is version 4, whis has just been relased at the beginning of July. Since we were not able to test this package very intesively yet, we still install pine 3.96 as the default. This package also contains the text editor pico and the file manager pilot. Version 4 adds the following features:

- New screen navigation commands

- Hierarchical folder collections

- URL recognition and viewer dispatch

- Basic rendering of message bodies in HTML format

- Help screens with hyperlinks

- Various other improvements (Adressbook, MIME, IMAP4, LDAP) Contents: pine4 Please note, that this package ist rather new, therefore it has not been intensively tested yet!

Copyright:  (c) 1989-98 University of Washington

Author:

Mike  Seibel  <mikes@cac.washington.edu>,  Steve  Hubert

<hubert@cac.washington.edu>, Laurence Lundblade <lgl@cac.washington.edu>

Version:  4.02-3

plp Print spooler

A newer printer spooling system for Linux. Configuration is located in /etc/printcap, /etc/plp.conf, and /etc/printer perms. If problems occur shutdown the spooling system by entering ”/sbin/init.d/lpd stop”. Then you can switch to the BSD spooling system by removing package plp in YaST and installing package lprold. Restart the system with ”/sbin/init.d/lpd start”.

Contents: plp-4.0.3

Version:  4.1.2-1

pop POP Server/Client

This package contains (among others) fetchmail.

Contents: fetchmail-4.3.8 imap-4.1.BETA pop3d-1.00.4 popclient-3.0b6 popsrc qpopper2.2 upop

Copyright:

(c) 1991 Regents of the University of California, 1993-95 by Carl Harris, Jr., 1991-2, RSA Data Security, Inc., 1989 Mark Davies, 1990 Andy Linton, 1990 Victoria University of Wellington

Author:

Katie Stevens <dkstevens@ucdavis.edu>, Carl Harris <ceharris@mal.com>, Mark Erikson <mark@qualcomm.com>, <msharov@bev.net>

Version:  98.8.10-0

ppp PPP (”Point to Point Protocol”) for Linux

PPP kernel driver and the pppd daemon. An alternative to SLIP. Contents: ppp-2.2.0f

Author:

Paul  Mackerras  <paul.mackerras@cs.anu.edu.au>,  Michael  Callahan

<callahan@maths.ox.ac.uk>, Al Longyear <longyear@netcom.com>, Brad Parker <brad@faster.fcr.com>, Greg Christy <gmc@premises1.Qutoron.COM

Version:  2.2.0f-20

procmail Local email delivery

Sendmail calls procmail to deliver email into a local folder. Procmail can be configured to store email into different folders. Contents: procmail-3.11pre7

Author:  Stephen R. van den Berg <srb@cuci.nl>
Version:  3.11pre7-14

proftpd Full featured FTP daemon

Professional FTP Daemon offers features as configuration with a Apache like syntax in a single file and per directory with ”.ftpaccess”, support of multiple virtual FTP servers and anonymous FTP services, no SITE EXEC command to avoid security problems, hidden directories or files, wu.ftpd compatible logging (extended logging available) and utmp/wtmp support, Shadow password suite support, including support for expired accounts. Documentation in /usr/soc/packages/proftpd/ and http://www.proftpd.org/. Please install also the ftpdir-Package with FTP-Environment and activate the server in /etc/inetd.conf.

Copyright:  (c) (c) 1997, Public Flood Software
Author:  proftpd@proftpd.org
Version:  1.0.3-1

radiusc Radius client-software

Radiusclient provides some useful programs for doing authentication and accounting via a (remote) radius-server. Contents: radiusclient-0.3-alpha-4

Copyright:  (c) see /usr/doc/packages/radiusc/COPYRIGHT
Author:  Lars Fenneberg <lf@elemental.net>
Version:  0.3alpha4-6

radiusd Radius-Server

radiusd is the Livingston radius-server. It provides remote authentication an accounting services; it e.g. use bei Livingston Portmaster products (you guessed) for login-authentication and logging. Contents: radius-1.16.1

Copyright:  (c) Copyright 1992 Livingston Enterprises, Inc.
Author:  Livingston Enterprises, Inc.
Version:  1.16.1-7

rdist ”Remote file distribution” utility

This utility makes it possible to distribute files to several computers within a network. If possible, file attributes such as: owner, group, mode, and mtime are retained. It’s even possible to overwrite running binaries. Contents: rdist-6.1.3

Copyright:  (c) 1993 Michael A. Cooper
Author:  Michael A. Cooper <Michael.Cooper@Sun.Com>
Version:  6.1.3-13

rsync replacement for rcp/mirror that has many more features

rsync uses the ”rsync algorithm” which provides a very fast method for bringing remote files into sync.

It does this by sending just the differences in the files across the link, without requiring that both sets of files are present at one of the ends of the link beforehand. At first glance this may seem impossible because the calculation of diffs between two files normally requires local access to both files. A technical report describing the rsync algorithm is included with this package. Contents: rsync-1.6.3

Copyright:  (c) Andrew Tridgell and Paul Mackerras (GPL)
Author:  Andrew.Tridgell@anu.edu.au, paulus@cs.anu.edu.au
Version:  1.7.4-5

rzsz X-, Y- and Z-Modem data transfer protocols.

Allows you to Use ”sz filename” to send a file to your local system. Contents: lrzsz-0.12b

Version:  0.12b-13

samba Samba is a file server for Unix, similar to LanManager.

Samba is a suite of programs which work together to allow clients to access Unix filespace and printers via the SMB protocol (Seerver Message Block). CAUTION: The samba daemons are started by the init script /sbin/init.d/samba, not by inetd. The entries for /usr/sbin/smbd and /usr/sbin/nmbd must be commented out in /etc/inetd.conf. In practice, this means that you can redirect disks and printers to Unix disks and printers from LAN Manager clients, Windows for Workgroups 3.11 clients, Windows’95 clients, Windows NT clients and OS/2 clients. There is also a Unix client program supplied as part of the suite which allows Unix users to use an ftp-like interface to access filespace and printers on any other SMB server. Samba includes the following programs (in summary):

* smbd, the SMB server. This handles actual connections from clients.

* nmbd, the Netbios name server, which helps clients locate servers.

* smbclient, the Unix-hosted client program.

* smbrun, a little ’glue’ program to help the server run external programs.

* testprns, a program to test server access to printers.

* testparm, a program to test the Samba configuration file for correctness.

* smb.conf, the Samba configuration file.

* smbprint, a sample script to allow a Unix host to use smbclient to print to an SMB server. The suite is supplied with full source and is GPLed. This package expects its config file under /etc/smb.conf . Documentation: /usr/doc/packages/samba Contents: samba-1.9.18p4 smbfs-2.0.2

Copyright:  (c) 1992-95 Andrew Tridgell, Karl Auer, Jeremy Allison

Author:

Andrew  Tridgell  <Andrew.Tridgell@anu.edu.au>,  Karl  Auer

<Karl.Auer@anu.edu.au>, Jeremy Allison <jeremy@netcom.com>

Version:  1.9.18p8-5

sendfax sendfax part of mgetty

The sendfax part of mgetty. The sources are included in mgetty source package.

Copyright:  (c) 1993-1997 Gert D”oring
Author:  Gert D”oring <gert@greenie.muc.de>
Version:  1.1.8-20

sendmail BSD sendmail

The ”Unix System Administration Handbook” calls sendmail ”The most complex and complete mail delivery system in common use...” . Ready-made configuration files are included for systems connected by TCP/IP (with or without a name server) and for systems using UUCP. ’procmail’ is included as a local mail agent.

Contents: check sendmail-8.8.8

Copyright:  (c) 1983 Eric P. Allman
Author:  Eric Allman <sendmail@sendmail.ORG>
Version:  8.8.8-21

seyon telecommunications under X

The best telecommunications package for X available at present; it has separate windows for control, terminal, telephone register, etc. Seyon supports several protocols (Zmodem, Kermit, etc.) and auto download (only Zmodem). Using the built-in script language, automatic login or up/downloading is no longer a problem. Attention : For security reasons we had to remove the sgid bit for the uucp group. Please note, that all users that should be able to use seyon must be added to the group uucp (e.g. with YaST). If there are only only ”trusted users” in your system and you want to ensure, that seyon is usable for all of them, add the following line to /etc/permissions.local and call SuSEconfig /usr/X11R6/bin/seyon root.uucp 2755 The actual transmission is provided by the several programs like sz/rz/kermit. Since we are not allowed to supply kermit with our distribution you have to get it yourself from the Internet (ftp: kermit.columbia.edu). Please look at the various sites as sunsite.unc.edu or any server which mirrors sunsite (ftp.gwdg.de under /pub/Linux/X11/xapps/comm or /pub/Linux/apps/comm). You have to configure ˜ /.seyon/protocol to use it. Contents: seyon-2.14c seyon-kermit-1.0 Attention : For security reasons we had to remove the sgid bit for the uucp group. Please note, that all users that should be able to use seyon must be added to the group uucp (e.g. with YaST). If there are only only ”trusted users” in your system and you want to ensure, that seyon is usable for all of them, add the following line to /etc/permissions.local and call SuSEconfig /usr/X11R6/bin/seyon root.uucp 2755

Copyright:  (c) 1992 Muhammad M. Saggaf
Author:  Muhammad M. Saggaf <alsaggaf@mit.edu>
Version:  2.14c-17

sf An extensive firewall package

sf is an extensive firewall package. It includes its own description language, that allows you to create very complex configurations. filtering of all header fields in the IP,TCP,UDP,ICMP,IGMP packets intelligent RIP and FTP support dynamic rules, including counters and time-outs extensive logging, alerting, and counter intelligence prevention of packet and address spoofing Please read /usr/doc/packages/sf/README.SuSE first!!!

Copyright:  (c) 1996 Robert Muchsel and Roland Schmid
Author:  Bugreports <firewall-bugs@switch.ch>
Version:  0.2.9-5

slip dial out/in programs

This package include the two dial out/in programs dip and sliplogin. Users who want to use dip, have to be in group ”dialout”. This is also necessary for ”diplogins”.

Contents: dip-3.3.7p sliplogin-2.1.0

Copyright:

(c) 1988-1993 MicroWalt Corporation, 1991-1992 Ian Lance Taylor, 1990 The Regents of the University of California

Author:

Fred N. van Kempen <waltje@uwalt.nl.mugnet.org>, Sven Goldt <goldt@peagsus.inberlin.de>

Version:  97.10.27-14

slrn News reader.

Version:  0.9.5.2-1

smail The mail system ”Smail”.

Smail is a mail system which is easy to configure. A lot of examples are installed in addition to the program. Smail should NOT be used together with sendmail. Contents: smail-3.2 Please note, that this package is not supported by our support team. Please use sendmail if you need support for mail delivery. Feedback and bug fixes for smail are of course welcome at feedback@suse.de.

Author:  Ian Kluft (Portierung)
Version:  3.2-17

smtpd A SMTP store and forward proxy

This is a SMTP store and forward proxy. Please read /usr/doc/packages/README.SuSE!!! Please read /usr/doc/packages/README.SuSE!!!

Copyright:  (c) 1996, 1997 Obtuse Systems Corporation
Author:  Obtuse Systems Corporation <info@obtuse.com>
Version:  2.0-5

snmp CMU SNMP

SNMP is the Simple Network Management Protocol of the Internet defined as a standard in RFC 1155 and 1157 and 1901 up to 1908. The protocol allows mainly to retrieve and set variables, addressed as objects in the MIB (Management Information Base). The well-known MIB-2 is defined in RFC 1213 and contains variables related to the TCP/IP protocol suite.

Contents: cmu-snmp-linux-3.2

Copyright:  (c) Copyright 1988, 1989 by Carnegie Mellon University
Author:  Erik Schoenfelder <schoenfr@gaertner.de>
Version:  3.2-17

squid squid: proxy server for WWW

Copyright:  (c) GPL
Version:  1.NOVM.21-6

ssh secure shell client and server (remote login program)

Ssh (Secure Shell) a program for logging into a remote machine and for executing commands on a remote machine. It is intended to replace rlogin and rsh, and provides secure encrypted communication between two untrusted hosts over an insecure network. X11 connections and arbi- arbitrary TCP/IP ports can also be forwarded over the secure channel.

Contents: ssh-1.2.22

Copyright:  (c) 1995 Tatu Ylonen, Espoo, Finland
Author:  Tatu Ylonen <ylo@cs.hut.fi>
Version:  1.2.25-2

ssleay SSLeay - Secure Socket Library

The package includes libssl.a: My implementation of SSLv2, SSLv3 and the required code to support both SSLv2 and SSLv3 in the one server. libcrypto.a: General encryption and X509 stuff needed by SSL but not actually logically part of it. It includes routines for the following: Ciphers libdes - My libdes DES encryption package which has been floating around the net for a few years. It includes 15 ’modes/variations’ of DES (1, 2 and 3 key versions of ecb, cbc, cfb and ofb; pcbc and a more general form of cfb and ofb) including desx in cbc mode, a fast crypt(3), and routines to read passwords from the keyboard. RC4 encryption, RC2 encryption - 4 different modes, ecb, cbc, cfb and ofb. Blowfish encryption - 4 different modes, ecb, cbc, cfb and ofb. IDEA encryption - 4 different modes, ecb, cbc, cfb and ofb.

Digests MD5 and MD2 message digest algorithms, fast implementations, SHA (SHA-0) and SHA-1 message digest algorithms, MDC2 message digest. A DES based hash that is polular on smart cards. Public Key RSA encryption/decryption/generation. There is no limit on the number of bits. DSA encryption/decryption/generation. There is no limit on the number of bits. Diffie-Hellman keyexchange/key generation. There is no limit on the number of bits. X509v3 certificates X509 encoding/decoding into/from binary ASN1 and a PEM based ascii-binary encoding which supports encryption with a private key. Program to generate RSA and DSA certificate requests and to generate RSA and DSA certificates. Systems The normal digital envelope routines and base64 encoding. Higher level access to ciphers and digests by name. New ciphers can be loaded at run time. The BIO io system which is a simple non-blocking IO abstraction. Current methods supported are file descriptors, sockets, socket accept, socket connect, memory buffer, buffering, SSL client/server, file pointer, encryption, digest, nonblocking testing and null. Data structures A dynamically growing hashing system A simple stack. A Configuration loader that uses a format similar to MS .ini files. Programs in this package include enc - a general encryption program that can encrypt/decrypt using one of 17 different cipher/mode combinations. The input/output can also be converted to/from base64 ascii encoding. dgst - a generate message digesting program that will generate message digests for any of md2, md5, sha (sha-0 or sha-1) or mdc2. asn1parse - parse and display the structure of an asn1 encoded binary file. rsa - Manipulate RSA private keys. dsa - Manipulate DSA private keys. dh - Manipulate Diffie-Hellman parameter files. dsaparam- Manipulate and generate DSA parameter files. crl - Manipulate certificate revocation lists. crt2pkcs7- Generate a pkcs7 object containing a crl and a certificate. x509 - Manipulate x509 certificates, self-sign certificates. req - Manipulate PKCS#10 certificate requests and also generate certificate requests. genrsa - Generates an arbitrary sized RSA private key. gendh - Generates a set of Diffie-Hellman parameters, the prime will be a strong prime. ca - Create certificates from PKCS#10 certificate requests. This program also maintains a database of certificates issued. verify - Check x509 certificate signatures. speed - Benchmark SSLeay’s ciphers. s server- A test SSL server. s client- A test SSL client. s time - Benchmark SSL performance of SSL server programs. errstr - Convert from SSLeay hex error codes to a readable form. Documents avaliable are A Postscript and html reference manual (written by Tim Hudson tjh@cryptsoft.com). A list of text protocol references I used. An initial version of the library manual. Contents: SSLeay-0.8.1

Copyright:  (c) (C) 1997 Eric Young <eay@cryptsoft.com>
Author:  Eric Young <eay@cryptsoft.com>
Version:  0.8.1-12

strn News reader.

This version of strn is based on trn and allows scoring of a news article based on author and subject. Contents: strn

Version:  092q3-16

suck Reading News Offline

Author:

Tim Smith, Sven Goldt <goldt@math.tu-berlin.de>, Robert A. Yetman <boby@pixi.com>

Version:  3.9.3-1

susefax A fax client for the HylaFAX fax server

This is a fax-client for the HylaFax(tm) fax-server maintained and programmed by Sam Leffler (SGI). It is totally implemented in Java and requires the Java Developer Kit (JDK) at least version 1.1.1 (Package java, series d). It supports:

* integrated phone book

* sending faxes directly from the phone book * setting up job-parameters before sending a job and during processing of the job in the job-queue

* removing jobs from the job-queue

* direct viewing of received faxes (needs an external tiffg3 viewer or converter, for Linux/UNIX the converting is done by fax2ps from libtiff and viewing by ghostview or gv automatically by a script that comes with this package) * creating of serial-fax lists out of the phone book entries

* automatic cover page generation (not very fast :-( )

* auto-spooling (that’s important for WINDOWS users) YES!!! This program is also running on WINDOWS and hopefully all the other platforms, that support Java by the JDK from Sun!!! It has been tested on WindowsNT 4.0 and it seems to run!!!!!

* password prompting if the administrator of the fax server wants the users to access the server via password

Contents: SuSEFax

Copyright:  (c) 1997, 1998 S.u.S.E. GmbH, Fuerth
Author:  Carsten Hoeger <Carsten.Hoeger@suse.de>
Version:  0.84v1-2

suseppp SuSE PPP

Internet Provider Database including script to configure a PPP conection

Contents: suseppp

Author:  Reinhold Sojer <reinhold@suse.de>
Version:  1-10

tcpdump packet sniffer

With this program you can ”read” all or only certain packets going over the ethernet. It can be used to debug specific network problems.

Contents: tcpdump-3.4a6

Version:  3.4a6-12

tin The ”tin” News Reader

Tin is a full-screen, easy to use Netnews reader. It can read news locally (i.e. /var/spool/news) or remotely (rtin or tin -r option) via an NNTP (Network News Transport Protocol) server. Docu: man tin or man rtin.

Contents: tin-980226

Copyright:  (c) 1991-93 Iain Lea
Author:  Iain Lea <iain@ecrc.de>
Version:  980618-1

tleds Indicate network traffic via Keyboard LEDs

These programs help you monitor network traffic. They blink the Scroll-Lock LED when a network packet leaves the machine, and the Num-Lock LED when one is received.

Copyright:  (c) 1997 Jouni.Lohikoski@iki.fi
Author:  Jouni Lohikoski <jlohikos@cc.hut.fi>
Version:  1.04-6

tripwire A tool to observe the filesystem

By the use of tripwire it is possible to observe the filesystem. tripwire generates a database, controlled by a configuration file, of all files, their checksums, etc. and it reports changes. The location of the binary is /var/adm/tripwire/bin/tripwire. A configuration example can be found at /usr/doc/packages/tripwire. Tripwire wants its configurationfile to be found at /var/adm/tripwire/tw.config. The database has to be saved at /var/adm/tripwire/db.

Copyright:  (c) 1992, 1993, 1994 by the Purdue Research Foundation of Purdue University
Author:  Gene Kim <gkim@cs.purdue.edu>, Gene Spafford <spaf@cs.purdue.edu>
Version:  1.2-2

trn News reader (conf. for NNTP)

This trn is compiled so it needs the Network News Transfer Protocol Daemon (nntpd) to get its news.

Contents: trn-3.6

Copyright:  (c) 1991-92 Stan Barber

Author:

Wayne Davison <davison@borland.com>, Larry Wall <lwall@jpldevvax.jpl.nasa.gov>, Stan Barber <sob@bcm.tmc.edu>

Version:  3.6-16

trn spl Threaded News Reader (conf. for spool directory)

Compiled for reading news from the local spool-directory which resides under /var/spool/news. If you would like another configuration you have to recompile trn. This shouldn’t be a problem as trn is effortlessly compilable under Linux.

Copyright:  (c) 1991-92 Stan Barber

Author:

Wayne Davison <davison@borland.com>, Larry Wall <lwall@jpldevvax.jpl.nasa.gov>, Stan Barber <sob@bcm.tmc.edu>

Version:  3.6-12

url get Fetch data from their URL-addresses.

With url get you may get data from their URL-address. HTML, Gopher and FTP are supported. For further documentation see /usr/doc/packages/url get.

Contents: url get

Author:  Jack Lund <j.lund@cc.utexas.edu>
Version:  98.2.13-7

uucp Taylor UUCP

Ian Taylor’s Unix to Unix copy: mail and news over modem lines. This is the standard UUCP package of the Free Software Foundation. Configured to use HoneyDanBer and/or Taylor configuration files. Remark: We did not include an uucp guest account. So, if you want to create a guest account please make sure the directory /var/spool/uucppublic exists.

Contents: uucp-1.06.1

Copyright:  (c) 1991-93 Ian Lance Taylor
Author:  Ian Lance Taylor <ian@cygnus.com>
Version:  1.06.1-14

uucpcfg Taylor UUCP - sample configuration

Sample configuration files for Taylor UUCP.

Contents: uucpcfg

Version:  98.2.3-5

vacation Automatic reply to incoming mails

This program answers when you’re lying on the beach ;-) Documentation: man vacation

Contents: vacation-1.1.1

Copyright:  (c) 1983, 1987 University of California, 1993 Harald Milz
Author:  Eric P. Allman, Harald Milz <hm@seneca.muc.de>
Version:  1.1.1-13

vbox vbox; an ISDN answering machine for Linux

Package for realizing an answering machine under Linux using its ISDN capabilities. Please keep in mind that you need a kernel with CONFIG ISDN AUDIO support. Of course the ISDN package (i4l) has to be installed as well. Please have a look at /usr/doc/packages/vbox Contents: vbox-1.1 Please be aware that for running this package you need a kernel with ISDN AUDIO SUPPORT compiled in.

Copyright:

(c) Copyright: Michael Herold, Chris Benndorf, Jan Schoenepauck, Joachim Gassen, Marc Eberhard

Author:

Michael  Herold  <michael@abadonna.franken.de>,  Chris  Benndorf

<chrisb@ganzfix.duew.eunet.de>, Jan Schoenepauck <schoenep@wrcs3.urz.uniwuppertal.de>, Joachim Gassen <gassen@uni-muenster.de>, Marc Eberhard <Marc.Eberhard@Uni-Duesseldorf.DE>

Version:  1.1-22

vbox2b4 Vbox 2b4 An ISDN answering machine

Vbox is an ISDN answering machine. Besides recording incoming calls it offers a wide variety of features. For configuring vbox please have a look at: /usr/doc/packages/vbox2b4 and check an adapt the shipped configuration files under: /etc/isdn to your needs. Contents: vbox2b4 This package is still considered beta, that’s why this package is not yet included in the installation support.

Copyright:  (c) Michael Herold <michael@abadonna.mayn.de>

Author:

Michael  Herold  <michael@abadonna.mayn.de>,  Fritz  Elfert

<fritz@wuemaus.franken.de>, Carsten Keil <keil@temic-ech.spacenet.de>, Andreas Kool <akool@Kool.f.EUnet.de>, Stefan L”uthje <luethje@sl-gw.lake.de>, Marc Eberhard <Marc.Eberhard@Uni-Duesseldorf.DE>, Gert Doering <gert@greenie.muc.de>

Version:  2b4-10

w3mir HTTP copying and mirroring tool

The main focus of w3mir is to create and maintain a browsable copy of one, or several, remote WWW site(s). Used to the max w3mir can retrieve the contents of several related sites and leave the mirror browseable via a local web server, or from a filesystem, such as directly from a CDROM. w3mir’s goal is to be able to make useful mirrors of any reasonable WWW site. It specifically preserves link integrity within the mirrored documents as well as the integrity of links outside the mirror, if you want it to. w3mir has a powerful “multi scope” mechanism enabling the user to make mirrors of several related sites and have links between them refer to the mirrored documents rather than the original site. w3mir has several features directed at getting mirrors for CDROM burning and handling of some not too often seen problems when mirroring. Contents: MD5-1.7 MIME-Base64-2.05 libnet-1.0605 libwww-perl-5.34 w3mir-1.0.3

Author:

Lcolai Langfeldt <janl@ifi.uio.no>, Gorm Haug Eriksen <gorm@usit.uio.no>, Chris Szurgot <szurgot@itribe.net>, Ed Jordan <ed@olympus.itl.net>, Alex Knowles <aknowles@avs.com>

Version:  1.0.3-2

webalize tool for webserver logfile analysis

The Webalizer is a web server log file analysis program which produces usage statistics in HTML format for viewing with a browser. The results are presented in both columnar and graphical format, which facilitates interpretation. Yearly, monthly, daily and hourly usage statistics are presented, along with the ability to display usage by site, URL, referrer, user agent (browser) and country (user agent and referrer are only available if your web server procduces Combined log format files).

Author:  Bradford L. Barrett <brad@usagl.net>
Version:  0.98.16-12

weblint a syntax and minimal style checker for HTML

Weblint is a syntax and minimal style checker for HTML: a perl script which picks fluff off html pages, much in the same way traditional lint picks fluff off C programms. Contents: weblint-1.020

Author:  Neil Bowers <neilb@cre.canon.co.uk>
Version:  1.020-7

wget C tool for mirroring FTP and HTTP servers

Wget enables you to retrieve WWW documents or FTP files from a server. This might be done in script files or via command line.

Contents: wget-1.5.0

Copyright:  (c) GPL
Version:  1.5.2-4

whois whois client program (RIPE version)

This is intended for European users as it will query a RIPE whois server with extended whois capabilities rather than a standard whois protocol server. The whois program is fully compatible with the standard whois protocol and can thus replace your default whois program. Contents: ripe-whois-tools-2.2

Copyright:  (c) 1980 Regents of the University of California
Author:  <marten@ripe.net>
Version:  2.2-13

wuftpd wuftpd

This FTP daemon is a little different. It replaces the standard-in.ftpd included in series n. It is more flexible but more complicated to configure as well. The following manual pages are included too: ftpd (the programm) ftpshut (service-program) ftpcount (service-program) ftpwho (service-program) ftpaccess (explanation of /etc/ftpaccess) ftphosts (explanation of /etc/ftpaccess) xferlog (explanation of /var/log/xferlog) ftpconversions (explanation of /etc/ftpconversions) Furthermore there are some sample configurations which will be placed into /usr/doc/packages/wuftpd during installation. If you want to use it you have to replace the following line in /etc/services: ftp stream tcp nowait root /usr/sbin/tcpd in.ftpd with: ftp stream tcp nowait root /usr/sbin/tcpd wu.ftpd Then you have to send a HUP to inetd (kill(8)) or restart. Contents: wu-ftpd-2.4.2-beta-15

Copyright:  (c) 1994 Washington University in Saint Louis.
Author:  Anonymous (Internet, Washington University in Saint Louis)
Version:  2.4.2beta15-15

wwwoffle WWWOFFLE - World Wide Web Offline Explorer

A proxy HTTP server for computers with dial-up internet access.

- Caching of pages viewed while connected for review later.

- Browsing of cached pages while not connected, with the ability to follow links and mark other pages for download.

- Downloading of specified pages non-interactively.

- Multiple indices of pages stored in cache for easy selection.

- Interactive or command line option to select pages for fetching individually or recursively.

- All options controlled using a simple configuration file with a web page to edit it.

Copyright:  (c) Andrew M. Bishop 1996,97

Author:

Andreas Dietrich <quasi@baccus.franken.de>, Andrew M. Bishop <amb@gedanken.demon.co.uk>, Axel Rasmus Wienberg <2wienbe@informatik.unihamburg.de>, Tilman Bohn <tb@bohn.isdn.uni-heidelberg.de>, Ullrich von Bassewitz <uz@wuschel.ibb.schwaben.com>, Walter Pfannenmller <Walter.Pfannenmueller@tonline.de>, Yannick Versley <sa6z225@public.uni-hamburg.de>

Version:  2.2c-1

xibod xibod

This is the control panel for the ibod(1) ISDN Bandwidth On Demand daemon. It is a Motif application that communicates over a TCP socket port 6050 with the ibod daemon. This Package include the ibod daemon.

Copyright:  (c) GPL
Author:  Bjoern Smith <Smith@Compound.SE>
Version:  1.2beta-8

xinetd Another ’inetd’ with expanded functionality

xinetd takes the functionality of inetd and appends additional opportunities:

- Access Control

- Prevention of ’denial-of-access’ attacks

- Extensive logging abilities

- clear konfiguration file Contents: xinetd-2.2.1 Note that you have to set ’START INETD’ to ’no’ if you want to use xinetd.

Copyright:  (c) 1992 by Panagiotis Tsirigotis, 1995 by Charles Murcko

Author:

Panagiotis Tsirigotis <panos@cs.colorado.edu>, Chuck Murcko <chuck@topsail.org>, Jan Wedekind <jw@ping.de>, <jw@Germany.EU.net>

Version:  2.2.1-11

xntp Network Time Protocol daemon

xntpd is a daemon which sets and maintains a Unix system time-of-day in compliance with Internet standard time servers. xntpd is a complete implementation of the Net- work Time Protocol (NTP) version 3 standard, as defined by RFC 1305, but also retains compatibility with version 1 and 2 servers as defined by RFC 1059 and RFC 1119, respec- tively. Contents: xntp3-5.92

Author:  Dennis Ferguson, David Mills
Version:  3.5.93c-2

xrn X11 news reader

Version:  8.01-12

yaps Yet Another Pager Software

This software is primary designed to send message to so called pager devices including cellular phones which are able to receive textual messages (sometimes called SMS, short message system/service.)

Copyright:  (c) 1997 Ulrich Dessauer
Author:  Ulrich Dessauer <ud@cube.net>
Version:  0.96-0

ypclient YP - (NIS-) client

Use your Linux-PC as a NIS client (former YP-client) Included in this package are: ypbind, ypmatch, yppoll, yppasswd, ypset and ypwhich. You can configure a server or use a broadcast using ypbind and adapting /etc/yp.conf. Contents: ypbind-bsd-1.15-970316 ypclients-2.2-970318

Copyright:  (c) 1995 Theo de Raadt, Swen Th”ummler, Michael Rausch

Author:

Theo de Raadt <deraadt@fsa.ca>, Swen Th”ummler <swen@uni-paderborn.de>, Michael Rausch <rausch@julia.de>

Version:  2.2.970318-13

ypserv YP - (NIS)-server

Configure your Linux PC as YP-(NIS)-server using this package. Contents: ypserv-1.1.3

Copyright:

(c) 1989, 1991, Free Software Foundation, Inc., 1996, Florian La Roche, S.u.S.E. GmbH Fuerth, Peter Eriksson, Thorsten Kukuk, Swen Thuemmler, 1993,1994 Signum Support AB, Sweden, et.al.

Author:

Florian La Roche <florian@suse.de>, Peter Eriksson <pen@signum.se>, Thorsten Kukuk <kukuk@uni-paderborn.de>, Swen Thuemmler <swen@uni-paderborn.de>

Version:  1.2.4-13

faxprint Printer-filter for use with SuSEFax

This printer-filter creates a printer called ’fax’. If you use this one with any application, you can send faxes with the spooling mechanism of SuSEFax.

PAY: Commercial Software

Contents: faxprint-filter

Copyright:  (c) 1998, S.u.S.E. GmbH, Fuerth, Germany
Author:  Carsten Hoeger <choeger@suse.de>
Version:  0.1-8

0.1.13 PAY: Commercial Software

abuse Abuse - Arcade game

Abuse is an old style arcade game: find your way through the labyrinth, shoot every alien, and collect some items. To move the fighter, use the cursor keys. The mouse controls aim and shoot. This version contains only a demo level. However, a level editor is included. Abuse runs under X11 as well as on the console. Contents: abuse Please remember that this program is SHAREWARE.

Copyright:  (c) 1995 Crack dot Com
Author:  <abuse@crack.com>, <jc@crack.com>
Version:  1.10-6

ac3d dem Autocad look-alike graphic program

Ac3d is a graphic program which looks like autocad. Start the program with the command ac3d. Under /usr/doc/packages/ac3d/ you will find some demos in the directory acfiles/ and the documentation in html format in the directory manual/ . To read the documentation, start your WWW browser with the argument file://usr/doc/packages/ac3d/manual/ Please note the message in the dialog window at startup: this program is SHAREWARE and must be registered if you wish to use it. Contents: ac3d dem

Copyright:  (c) 1996 Andy Colebourne
Author:  Andy Colebourne <andy@comp.lancs.ac.uk>
Version:  1.56-6

acroread Acrobat Reader for PDF files

Copyright:  (c) 1997 Adobe Systems Inc.
Version:  3.01-7

adabas Adabas D Personal Edition

This package contains Adabas D Personal Edition. Log on as user ”adabas”, start the X Window System and run the program ”panel”. This personal edition has the following restrictions:

- max. 24 MB data - max. 3 concurrent users Contents: adabas-10.0.1

Copyright:  (c) 1997,98 by Software AG
Version:  10.0.1-6

adadoc Adabas D English documentation

This package contains the English documentation for Adabas D personal edition.

Copyright:  (c) 1997 by Software AG
Version:  10.0.1-6

adadok Adabas D German documentation

This package contains the German documentation for Adabas D personal edition.

Copyright:  (c) 1997 by Software AG
Version:  10.0.1-6

adbsmydb MYDB demo database

This package contains the MYDB demo database for Adabas D personal edition. Start the program ”panel” as user ”adabas”. Now log on as ”control” with the password ”control” and enter ”MYDB” as the database.

Copyright:  (c) 1997 by Software AG
Version:  10.0.1-6

antivir AntiVir - virus scanner for Linux

This is the AntiVir V virus scanner software package of H+BEDV Software. Further information can be found at http://www.antivir.de/

Copyright:  (c) 1994-98 by H+BEDV Datentechnik GmbH
Author:  Martin Ritter <mritter@antivir.de>
Version:  5.13.1.0-1

arcad ARCAD CAAD demo

Have a look at http://www.arcad.de Contents: arcad-0.074 This is a demo of the ARCAD CAAD System of ARCAD Systemhaus.

Copyright:  (c) ARCAD Systemhaus
Author:  Bernd Metzger <bernd@arcad.ic-marl.de>
Version:  0.076-0

arkeia ARKEIA network backup software

As root set the variable START ARKEIA in /etc/rc.config to yes and start the backup server: /sbin/init.d/arkeia start ARKEIA utilizes an exclusive multi-flow technology to deliver speeds that are 200 to 300% faster than rival software packages. Its unique transaction engine allows multiple backup and restore operations to be performed simutaneously with total reliabilty. This is especially valuable for multiple users backing up or restoring data on their own. (a multiple drive library is required). ARKEIA’s Java interface enables the system adminsitrator to manage multiple remote backup servers through internet as if they were local backups. ARKEIA is the most reliable software you can find. It has a powerful database capable of restoring sites to their original state, in the event of server disasters. ARKEIA integrates with all types of corporate environments and is available for UNIX or Windows NT servers. ARKEIA is the first software delivered with a JAVA interface, effectively neutralizing the problem of heterogeneous computer environments. Operators can use any Intranet station to access any remote network. Its universal interface makes it the first software able to backup via the Internet. ARKEIA is compatible with a wide range of devices and provides the best ”plug and play” solution for effective data security. Contents: arkeia-4.0 Please remember to start the backup server.

Copyright:  (c) Knox Software Comm
Version:  4.0-6

bartels The Bartels AutoEngineer

Bartels AutoEngineer (BAE) is a fully integrated CAE/CAD/CAM system for the design of printed circuit boards. The system is based on the famous Bartels Autorouter which has been incorporated in most leading PCB layout systems throughout the world, setting new industrial standards of autorouting success and dramatically reducing the requirement for manual routing. Contents: bartels-4.0.beta

Copyright:  (c) Oliver Bartels F+E
Version:  4.0.beta-5

blender 3D animation and rendering software

Being the in-house software of a high quality animation studio, Blender has proven to be an extremely fast and versatile design instrument. The software has a personal touch, offering a unique approach to the world of Three Dimensions. Use Blender to create TV commercials, to make technical visualizations, business graphics, to do some morphing, or design user interfaces. You can easy build and manage complex environments. The renderer is versatile and extremely fast. All basic animation principles (curves & keys) are well implemented. In order to work with blender, please remember to set the environment variable BLENDERDIR to /usr/X11R6/lib/blender This is a beta version (1.36) of Blender, the current stable version is 1.30. The Blender software is copyright by NeoGeo bv (www.neogeo.nl).

Copyright:  (c) NeoGeo bv (www.neogeo.nl)
Author:  blender@neogeo.nl
Version:  1.36-0

c16 demo CONZEPT 16 demo

CONZEPT 16 - Demo version CONZEPT 16 is a program development tool based on a relational database, which can be used under a variety of operating systems, on many different computers, and in every PC-networking environment. CONZEPT 16 is completely developed in Germany and is oriented to the latest generation of microcomputers. To exploit the capabilities of each PC generation, there are special versions adapted to the corresponding CPU (8086/88, 80286, 80386, 80486, and Pentium). CONZEPT 16 is not only an ideal development tool for programmers wanting to develop applications on a relational database platform, but also for those users who want to configure easily their personal environment. You can easily generate structures, masks,and lists using the menu choices, which means that you don’t have to learn a new programming language. CONZEPT 16 is presently available for MS-DOS, Novell Netware, UNIX System V/386, Solaris, OSF/1 and all PC networks which support DOS version 3.1 or higher (LAN-server, Netware-Lite, Network OS etc.). CONZEPT 16 has multiuser and multi-platform capability. Since data security is most important in this context, CONZEPT 16 uses automatic as well as user-driven transactions. For maximum process parallelization, CONZEPT 16 uses ”two-phase-locking”, not only on Novell but on all other networks, including OS/2 and UNIX. Give it a try and start the demo as root from text console: usr/lib/c16 demo/startdemo Contents: c16 demo

Copyright:  (c) vectorsoft GmbH
Version:  4.0.14-6

cactus Cactus

Cactus is a tool for manipulating and designing chemical formula and structures. It is capable of loading and exporting various file formats.

Copyright:  (c) Dr. Wolf-D. Ihlenfeldt
Author:  Dr. Wolf-D. Ihlenfeldt <Wolf-Dietrich.Ihlenfeldt@CCC.Chemie.Uni-Erlangen.de>
Version:  3.11-3

cint A C/C++ interpreter

Cint is a C/C++ interpreter. It is solid enough to interpret its own over 60,000 lines of source code and let the interpreted interpreter interprete a (small) program. ”makecint” makes it possible to encapsulate arbitrary C/C++ object as a precompiled library. Commercial use of CINT requires registration to Hewlett-Packard Japan. Send e-mail to the author (gotom@jpn.hp.com) describing your name, company or organization, address and purpose of using CINT. After the installation further information can be found in the file /usr/doc/packages/cint/README

Copyright:  (c) 1991˜ 1996 Hewlett-Packard Japan
Author:  Masaharu Goto <gotom@jpn.hp.com>
Version:  5.13-2

crisp A programmable editor

CRISP is an editor for programmers which has many improved features.

Wherever possible, CRISP has no internal restrictions regarding the number of files that can be edited simultaneously, the maximum line length, etc. CRISP is only constrained by the main memory and the hard disk capacity of your system. As a result, CRISP combines some of the best features of other available editors such as vi, emacs, BRIEF. Please note that the program is SHAREWARE. Further information can be found in /usr/doc/packages/crisp . Contents: crisp

Copyright:  (c) 1996 Foxtrot Systems Ltd; Paul David Fox
Author:  Paul Fox fox@crisp.demon.co.uk
Version:  5.1.1b-7

cysched Synchronize server software and Cyberscheduler software

Edit /etc/rc.config and set START SYNCHRONIZE to ”yes”, then run ‘/sbin/init.d/synchro start‘ as root Go to the URL http://localhost/synchro/En US/login.html and log-on as ”demo”. Product information: Industrial strength calendaring and scheduling for your changing enterprise. Synchronize is the scheduling solution that fits both your end-user and your IT requirements! Since 1990, enterprises of all sizes have relied on Synchronize for scalable, enterprise-wide, efficient real-time calendaring and scheduling. This proven client/server software supports enterprise-wide coordination and collaboration among team members, whether they’re located on different floors, multiple buildings, across the country, or around the world, bringing the enterprise closer together. The coordination and collaboration features designed into the Synchronize solution work together to enhance decision-making and communication. This helps your users increase their productivity, enhance teamwork, simplify collaborative efforts, and minimize project cycle times. Get the Synchronize advantage Designed specifically for enterprise scheduling, only Synchronize gives you all these important advantages: No hidden costs. Synchronize is framework independent, which means it does not require a proprietary messaging infrastructure and operates in mixed application or mixed framework environments. The broadest cross-platform support. No other calendaring and scheduling software supports as many different desktop environments and servers. Synchronize includes Web and native support for all these platformsWindows, Macintosh, UNIX/X and Character, in addition to the most complete UNIX and NT server line-up. The ability to support mixed Web and native desktops enables you to roll out your traditional native client and/or Intranet solution on the platforms you want and in the time frame that works best for your organization. Excellent technical support. Our Technical Support Department is staffed by a team of knowledgeable engineers trained in supporting cross-platform environments. CrossWind’s customers enjoy a rapid response with personal attention to each and every support question. Seamless UNIX & NT integration. Our servers are offered on both UNIX and NT platforms supporting seamless coexistence, or future integration, or migration plans. Scalability. A distributed database architecture lets you quickly and transparently add as many new users and user groups as you need. You can administer calendaring centrally or at multiple locations. No-hassle upgrades. Synchronize is the only enterprise calendaring and scheduling product that offers true backward compatibility. Our commitment to cross-release compatibility makes the upgrade process straightforward and discretionarynot an enterprise-wide crisis. Maximizing the productivity of every resource A powerful suite of real-time group calendaring and scheduling applications, Synchronize contains important coordination tools that save your organization time in scheduling meetings and company resources.

Users can distribute agendas and memos, schedule meetings, allocate resources, assign and track tasks, and send remindersno matter what platform they’re running. Work teams in multi-vendor computing environments can achieve consensus on both short-term and ongoing issues through the use of shared notes in private or public forums … easily and transparently. Synchronize enables group and personal task management as well as flexible resource management capabilities. Users have access to individual and group calendars, shared notes, and updated to-do lists. Support for the Internet, printers and portable computers makes it easy to access or take your schedule and project notes with you wherever you go. Built for performance and simple administration Synchronize is based on a robust client/server architecture using multi-threaded servers. All schedules, meetings, deadlines, reserved resources, project notes, messages, and work assignments are maintained in a common database stored in the Synchronize server, enabling quick and easy sharing of vital data. Administrators can easily distribute Synchronize’s common database into many networked databases that take advantage of existing servers and networks. The ability to communicate directly across TCP/IP provides real-time access to critical information without the delays typically associated with file-based access methods, e-mail based scheduling, and slow transport protocols. Advanced architecture makes Synchronize highly scalable: It can be used for enterprise collaboration on LANs, WANs and corporate INTRANETs in organizations ranging from 5 to 50,000 users (or more). Synchronize’s distributed database architecture allows administrators to add new users easily and transparently, with no interruption in service. CrossWind is dedicated to building solutions based on industry standards. As a result Synchronize’s interoperability with Web servers and applications is simple and straightforward. As a member of the Internet Engineering Task Force working group, responsible for establishing Internet standards for calendaring and scheduling, CrossWind continues its commitment for interoperability through standards support. Technical Information Designed for cross-platform deployments, Synchronize runs on most all commercial UNIX and NT servers, with desktop access through Web browsers as well as native software clients for Microsoft Windows, Macintosh, X.11/Motif and ASCII desktops. Affordable with No Hidden Costs Although some competitive products appear to have a very low per-user pricea good look at the total cost of ownership will reveal that a substantial, continuing investment will be required to assure smooth operations for these products. Synchronize is easy to deploy and requires no additional expenses for integration into an enterprise’s existing environment. Ready ”off the shelf,” Synchronize does not require programming, extra database servers, or email/messaging platforms. All Synchronize clients and servers are available (at no charge) for unrestricted download from our ftp site 24 hours a day. You can reconfigure, migrate and ”webify” your system at anytime at no charge. Since upgrades are backward compatible, they can be done (or not) on a workgroup or departmental basis. Your organization will not be compelled to buy into new releases under the duress of lost support or lack of compatibility. Contents: cysched-2.1.0

Copyright:  (c) Crosswind Technologies
Version:  2.1.0-5

eloquen HP Eloquence (B1368B)

HP ELOQUENCE Product Description: HP Eloquence is a integrated environment for development and execution of commercial applications. It is easy to use and ideally suited for small to midsize companies. Designed to operate in a distributed network environment it supports fast, flexible application development, implementation and maintenance. HP Eloquence has a broad installed base on HP-UX systems worldwide with a large number of ready-to-use commercial applications. HP Eloquence is available for HP-UX, Windows NT and LINUX platforms. The product includes: - State-Of-The-Art Interactive Development Environment on Windows 95 and Windows NT platforms. - Standard, character-oriented Development Environment on HP-UX and LINUX platforms - Powerful Programming Language, a superset of HP Business Basic - Integrated Database Management System (IMAGE compatible) - Runtime Environment for Graphical User Interfaces HP Eloquence is a product of Hewlett-Packard GmbH, Germany. -------------------------------------------------------------------- Licence Information: The HP Eloquence ”Personal Edition” is available for free. It provides a free license for personal or evaluation usage. It is limited to two users and a maximum database size of 50 MB. Be sure to read and agree to the ”Personal Edition License Agreement” before you use the software. -------------------------------------------------------------------- Support Information: We provide 60 days free support via E-mail after registration on our Web-Page http://www.hpeloquence.com/support/register.html Information about commercial support for HP Eloquence are available on: http://www.hp-eloquence.com/info/ -------------------------------------------------------------------- Product and Price Information of the commercial version of HP Eloquence are available on : http://www.hp-eloquence.com/info/order.html If you need further information please contact: Marxmeier Software GmbH Phone: +49 202 243 1440 Fax: +49 202 243 1420 Email: info@hpeloquence.com Contents: eloquen-A.06.01 The license of this demo expires in 30 days!

Copyright:  (c) 1997 Marxmeier Software Entwicklung GmbH
Version:  A.06.01-4

executor A demo of ARDIs Mac emulator.

A demo version of the Executor. It is limited to 15 minutes runtime per invocation and fully operational after installation (start with executor-demo). There is some Mac shareware included, e.g. a Lemmings demo. The emulator also supports sound. Contents: executor This is a demo version.

Copyright:  (c) 1986-1996 ARDI
Author:  ARDI <questions@ardi.com>
Version:  2.0-7

flagship Clipper compiler (DEMO) FlagShip for Linux

FlagShip is a database development and porting system for porting all applications based on the xBase language to UNIX. Completely source and file compatible to Clipper 5.2 (and previous versions) including macros, code blocks, objects, etc. The motto is ”the same sources and data for both DOS and UNIX”. (quote) You will find the documentation under /usr/doc/packages/flagship. To install, change to the directory /usr/lib/FlagShip and start the program FSinstCD. After installation the directory /usr/lib/FlagShip can be deleted. The full version is also included, you only need to order a key from S.u.S.E.

Copyright:  (c) 1996 Multisoft Datentechnik GmbH
Author:  multisoft Datentechnik GmbH <info@fship.com>
Version:  4.42-1

flagstry FlagShip test-version for Linux

FlagShip is a database development and porting system for porting all applications based on the xBase language to UNIX. Completely source and file compatible to Clipper 5.2 (and previous versions) including macros, code blocks, objects, etc. The slogan is ”the same sources and data for both DOS and UNIX”. (quote) As this program may only be used for 10 days, we do not offer an automatic installation. The pre-installation will be done in /usr/FlagShip/freedemo. Please read /usr/FlagShip/readme.engl for more information. Note: a full version of Flagship can be found in the package ”flagship”.

Copyright:  (c) 1995 Multisoft Datentechnik GmbH
Author:  <74431.552@compuserve.com>
Version:  4.42-1

graffy graffyDOC DEMO - a package for technical design / CAD

Info: /usr/doc/packages/graffy Start: ’/usr/X11R6/bin/graffy’

Copyright:  (c) DURST CAD/CONSULTING GmbH
Author:  Guenther <gl@durstcad.s.eunet.de>
Version:  L.02.41-0

heitml HeiTML

heitml (pronounced ”Hi”-TML) significantly extends and enhances the functionality of HTML by definable tags and full programming features. This makes dynamic content and database applications possible simply within the HTML world, without CGI and without external scripting or programming languages. This means you, as an HTML author, can embed applications in your web pages, simply by using some new tags without CGI and without programming. As an advanced user or programmer on the other hand you can create and program powerful tag libraries. This approach makes heitml suitable for HTML newcomers and professional programmers alike. Contents: heitml-1.2 HEITML (CGI) demo version, needs ADABAS D

Copyright:  (c) 1996-1997 H.E.I.
Version:  1.2-12

insure Insure++ detects run-time errors in C/C++ programs

Insure++ is an automatic runtime error detection tool for C/C++. Insure++ detects errors that your compiler and debugger may find trivial or even overlook entirely. You can find further information at www.fastcase.de Contents: Insure, Codewizard, Inuse, TCA This is commercial software. Please read /usr/doc/packages/insure/README.SuSE on how to obtain a license key.

Copyright:  (c) 1997,98 fastcase Technology
Author:  Patrick Walter <walter@fastcase.de>
Version:  4.1-1

ivs demo I.V.S demo of a time registration and calculation program

This is a ready to be used demo version of I.V.S. You can calculate and organize projects during a certain period of time. We used to have the problem that by the end of a month we did not have an overview of the ”performance” of our employees. That made it difficult for us to calculate. I.V.S is thought for companies, independent professionals and all others. Input of hours-of-work is achieved as follows: Customer <-> Project <-> Hours-of-work The licensed version offers multiple customers, projects (customer-specific) and tasks. With the supplied demo version you do not have the builtin user-administration; therefore everybody is able to see the data of the others! This isn’t true with the full licensed version.

Copyright:  (c) K + R Informationssysteme Konstanz
Author:  K + R Informationssysteme Konstanz <info@iridium.de>
Version:  9604-1

ixware IXware - Enterprise Solution Server

IXware is a german enterprise solution software. English documentation is not available. Contents: ixware-2.2.2.1 This software is in German only!

Copyright:  (c) Steiner, Sch”afer, R”uhl GbR
Author:  Winfried Sch”afer <wschaefer@s2r-ixware.de>
Version:  2.2.2.1-5

mapledmo Demo Maple V R3

Maple V is a comprehensive problem-solving system that supports a wide variety of powerful mathematical operations such as symbolic analysis, numerical analysis, and graphics. Maple V also features a complete programming language that uses syntax similar to that of BASIC, FORTRAN, Pascal or C. This language is indispensable for those wishing to develop solutions to larger problems. This is a demo version of the well-known mathematical package. Maple V can be found in /usr/local/maple . The binary was compiled using X11R5 and therefore requires the old XFree86[tm] 2.x libraries. If you are interested in a licensed version, please contact S.u.S.E. GmbH. Contents: mapledmo

Copyright:  (c) 1993 Waterloo Maple Software
Author:  <info@maplesoft.on.ca>
Version:  3-6

mars e Nice Editor for console and for X.

This nice Editor works both on the console and under X. It is completely in German language. Please read the file /usr/lib/mars e/README for licensing terms! Start the Editor simply with ”e”!

Copyright:  (c) Martin Stover
Author:  Martin Stover <mstover@Kool.f.EUnet.de>
Version:  4.11b-2

mhitdemo WYSIWYG document editor

Magic HIT is a modern document editor. Each document is based on a document class which determines the layout and structure of its contents. Additional pluses include: Easy to use, no interruptions in having to deal with formal or layout-related parameters. A modern, graphical user interface concept also makes it possible to integrate Magic HIT into a workflow installation. Magic HIT will be installed in the directory /usr/lib/magicHIT. Technical information is in the file /usr/lib/magicHIT/README. Magic HIT is integrated as an editor tool in the object oriented card and databank system, Xcard. Magic HIT runs in demo mode (no printing or saving). You can activate the full version with a key that you can get for free from Interface Connection GmbH. Documentation and a detailed description of the included document classes is available through Interface Connection GmbH. Also available: MagicHIT for Windows. InterFace Connection GmbH, Leipziger Strasse 16, 82008 Unterhaching, Telephone: +49 89 610 49-60, Telefax: +49 89 610 49-80

Copyright:  (c) 1996 InterFace Connection GmbH
Author:  InterFace Connection GmbH <service-center@ifconnection.de>
Version:  1.1-1

mind Motif Interface Designer

Mind is a GUI designer which generates C code for use with the Motif libraries. The graphical elements can be positioned directly on the screen and properties like look and behavior can be specified. The compiler can be started directly from Mind. For documentation/installation please see file /usr/doc/mind/INSTALL.SuSE. Contents: mind-3.0.r2 mind30 r2d

Copyright:  (c) flux GbR, Berlin
Author:  flux GbR, Berlin <flux.service@tonline.de>
Version:  3.0.r2-4

netscape Netscape (including script for starting netscape)

This package contains a start script for netscape. It makes it possible to set some variables an libraries for netscape. In this packages a netscape commuicator[tm] 4.05 is integrated. Note: the the program netscape is installed as netscape but the script will be called by its name Netscape. Attention: Please read the license conditions in /usr/doc/packages/netscape/ Contents: communicator-v405-export.x86-unknown-linux2.0 libXpm.4.6

Copyright:  (c) (C) 1994-1997 Netscape Communications
Version:  4.05-7

not demo NOTARIUS development tools with mask generator and DB connectivity

NOTARIUS is an development tool for UNIX systems with database access such as YARD, Informix and more. This is a demo version for Linux and offers the possibility to easily test the capabilities of NOTARIUS. More infos are to be found under: /usr/lib/not demo/README There is a PostScript document containing more than 100 pages.

Copyright:  (c) K + R Informationssysteme Konstanz
Author:  K + R Informationssysteme Konstanz <info@iridium.de>
Version:  1.2-1

rk512 Simatic S5 - Linux interface

rk512 has been designed to enable a cheap interface between Simatic S5 and Linux computers using a serial interface and the rk512 protocol. This trial version runs only 4 hours.

Copyright:  (c) Heisch Automatisierungstechnik
Version:  1.01.b-1

simcity Demo version of the well known game.

The demo runs 15 minutes, or if you try and save the game. If you would like to purchase a key to unlock the game follow the directions when the demo ends. Start the game to get more information. Contents: simcity

Copyright:  (c) 1993 DUX Stoftware Corporation
Author:  Mark Mattes <RonnieW252@aol.com>
Version:  95.4.10-6

snavig Cygnus Source-Navigator

Source-Navigator, the source-code comprehension tool from Cygnus Solutions, helps you analyze, reuse, and develop software in C/C++, Java, Tcl, and FORTRAN. Source-Navigator dramatically reduces the time required to understand existing code and assists with the re-structuring and reengineering of existing software. Source-Navigator, with its intuitive graphical user interface, directs you through analysis, change, build, and debug cycles. Its built-in database enables true project scalability, even in projects exceeding one million lines of code. Source-Navigator Lite Try a version of Source-Navigator. Source-Navigator Lite is available for use free of charge and provides full browsing, navigation, and project editing capability. The Lite configuration does have some limitations to its scalability:

- Supports projects smaller than 50,000 lines of code - Does not expose the API for integration or customization - Allows limited project printing Source-Navigator Lite can be used free of charge, and provides an easy way to evaluate the functionality. If you decide to upgrade to a complete version of Source-Navigator, you simply need a license code from Cygnus, and your application and projects will be automatically updated. Contents: SN-4.0-linux The Lite version is resticted to 50,000 lines of source code

Copyright:  (c) 1997,98 Cygnus Solutions
Version:  4.0.3-1

sniff SNiFF+ Integrated Cross-Platform Development Environment

SNiFF+ is an open, extensible and scalable cross-platform programming environment for C, C++, FORTRAN, JavaTM, CORBA IDL and other languages. The main goal in developing SNiFF+ was to create an efficient and portable environment with a comfortable user interface. SNiFF+ is tailored for large multi-developers projects. Contents: sniff-3.0 SNiFF+ comes with a preinstalled evaluation license which is valid until 31 Oct. 1998.

Copyright:  (c) TakeFive Software GmbH
Version:  3.0-0

solid Free SOLID Desktop 2.2 database

Free SOLID Desktop 2.2 for Linux is a special campaign version of SOLID Desktop 2.2 for Linux, which is restricted to SOLID Server version 2.2 and allows 2 local concurrent connections from the same computer where the server is running. cd /opt/solid ./setup The database engine still needs to be configured.

Copyright:  (c) Solid Information Technology Ltd
Version:  2.2-1

terroid Zap ’em arcade game.

TERROID is a classic-arcade game made to bring the quality of R-Type, SWIV, Wings of Death or Lethal Excess to your workstation. This is a DEMO of a shareware product.

Copyright:  (c) 1997 DIGITAL Nightmares
Author:  Michael Thayer <M.J.Thayer@ed.ac.uk>
Version:  1.3-1

uc Unix Cockpit

”Unix Cockpit” (UC) is a new Unix file manager. UC integrates file browsers, a directory tree, and your standard shell into a useful, mouse driven tool. Using a source and target browser, file operations like copy, move, delete, or edit can be executed. Additional operations (compress, gzip, find, etc.) can be chosen via a menu which can also hold user defined commands. In a separated shell even more commands can be entered. The current directory tree can be displayed in an additional window, providing the possibilities to navigate through the file system, select the working directory, and move directories. Using a small control window, all used windows can be pushed down, popped up, or be hidden. UC also provides an online help system. Summarizing, UC is a good X11 shell. Please notice that this program is shareware. Contents: uc Please notice that this program is shareware.

Copyright:  (c) 1989 - 1995 Henrik A. Klagges
Author:  Henrik A. Klagges <henrik@UniX11.com>
Version:  2.03-8

vdx VDX (libc5) - A Motif Interface Designer (Demo)

The View Designer/X is a motif interface builder, which enables application developers to interactively design user interfaces and generate C and C++ source code. The VDX provides following features: WYSIWYG and Tree View with drag & drop, Resource Editor, reusable components, templates, adaptable code generation. This package is a demo version compiled for libc5. Contents: vdx-1.0.4

Copyright:  (c) Copyright 1996-1998, Dirk Laessig
Author:  Dirk Laessig <lazy@de.ibm.com>
Version:  1.0.4-4

vishop VisualShop light - german version

Edit the file /etc/vshop.conf and make local configurations, then start der database: /sbin/init.d/vshopdb start Contents: vishop-1.1.d VisualShop light - german version

Copyright:  (c) 1997,98 intraDAT Netzwerktechnik und Software GmbH
Author:  Ralf Schwoebel <rschwoebel@intradat.com>
Version:  1.1.d-1

vse Visual SlickEdit - Demo

Demo version of the well-known ’Visual SlickEdit’ code editor. Invoke the demo by entering ’vsedemo’. Further information can be found under: /usr/doc/packages/vse Contents: vse-1.73

Copyright:  (c) 1993-1996 MicroEdge Inc.
Author:  Siegfried Hirsch <shirsch@infos.de>
Version:  1.73-5

wingz A easy to use spreadsheet program for X (SHAREWARE)

Wingz is a spreadsheet similar to Excel or Xess3, but it is shareware. Extensive graphics are possible. Wingz can totally be controlled by scripts. The registration fee is about $50,- per user. Wingz is available for several other UNIX systems and Mac/Windows machines and is extensible (real time versions, hyperscript). The manufacturer offers support for this product. Contents: wingz This program is SHAREWARE! Please read /usr/Wingz/SHAREWARE.

Copyright:  (c) Investment Intelligent Systems Corporation, USA
Author:  Investment Intelligent Systems Corporation <linux-bug@wingz.com>
Version:  1.4.1-7

xcept4 X11 btx program with a Tcl/Tk GUI

XCept4.0 is a demo version which only runs for 15 minutes and expires on 01.10.98. For installation see the file ”/usr/doc/packages/xcept4/INSTALL.SuSE”. For a Full-Version, please contact: Config Informationstechnik e.G. Dr. C.-U. Linster Lederhosenstra”se 30 91341 R”ottenbach Telefon: +49 9131 85 7280 Fax: +49 9131 85 8732 E-Mail: linster@informatik.uni-erlangen.de Contents: xcept4 XCept4.0 is a demo version which only runs for 15 minutes and expires on 31.03.98.

Copyright:  (c) Copyright (c) 1992-1995 Arno Augustin, Frank H”oring,

Author:

Arno Augustin <aoaugust@cip.informatik.uni-erlangen.de>, Frank H”oring <fkhoerin@cip.informatik.uni-erlangen.de>

Version:  4.0.2-3

xess3 Xess3 Spreadsheet - Demo

A spreadsheet program which can certainly be compared to the programs it was modelled after. Sample files can be found in /usr/local/lib/xess/examples. Xess also offers an API interface for C programs. Attention: This program will install into /usr/local.

Copyright:  (c) 1992-1996 Applied Information Systems, Inc. and others
Author:  <support@ais.com>
Version:  3.1.3b-7

xgal A Galaga-clone

This is a clone of the arcade hit Galaga. Contents: xgal Please remember that this program is shareware.

Copyright:  (c) 1995 Joe Rumsey
Author:  Joe Rumsey <ogre@atomic.com>
Version:  1.6c-16

xv GIF/TIFF/JPEG/PostScript image viewer

XV is an interactive image viewer for the X Window System. The XV program displays images in the GIF, JPEG, TIFF, PBM, PGM, PPM, X11 bitmap, Utah Raster Toolkit RLE, PDS/VICAR, Sun Rasterfile, BMP, PCX, IRIS RGB, XPM, Targa, XWD, possibly PostScript, and PM formats on workstations and terminals running the X Window System, Version 11. The XV program is also capable to manipulate pictures. It can zoom in and out of picture, and rotate and flip them. It also provides a wide variety of operations to modify colors. Surley, one of the most famous programs under Linux. You will find the documentation in /usr/doc/packages/xv/. Please remember that this program has shareware status for commercial use. Contents: xv-3.10a

Copyright:  (c) 1989, 1994 John Bradley (SHAREWARE)
Author:  John Bradley <bradley@dccs.upenn.edu>
Version:  3.10a-22

yardsql YARD-SQL Private Edition

This ”private edition” of YARD-SQL differs from the full version only in the amount of data which can be managed (in the private edition it is limited to 20 MB). YARD-SQL means: - standard SQL - dynamical SQL - CLient-Server architecture - Network capabilitiy - Referential integrity - Transaction security - embedded SQL for C - German dokumentation - Constraints - Interactive SQL YARD-SQL database server: Relational Client-Server architecture based database system for Linux providing ANSI/X/Open and SQL interface. Remote access via YARD-NET, extended SQL language and SELECT possibilities, 8 bit keymaps for national language support, Query Optimizer, support of different datatypes, data-integrity, power tools for database maintainement etc. YARD-ESQLC: SQL language embedding for C (without restrictions) YARD-NET: Distribution of Client-Server processes within one network via TCP/IP; including embedding of MS-Windows PCs via YARD-ODBC (without restrictions). YARD-X: Graphical user interface under OSF/Motif. YARD is also available for other operating systems; please contact us for questions or more informations.

SGM: Components for a SGML system

Please read the file /usr/lib/YARD/README. Contents: yardsql-4.04.0

Copyright:  (c) 1996-98 YARD Software GmbH
Author:  YARD Software GmbH <yard@yard.de>
Version:  4.04.0-5

0.1.14 SGM: Components for a SGML system

docbk30 DocBook: DTD and DSSSL-files

Copyright:

(c) Copyright 1992, 1993, 1994, 1995, 1996 HaL Computer Systems, Inc., O’Reilly & Associates, Inc., ArborText, Inc., and Fujitsu Software Corporation.

Author:

Eve L. Maler <elm@arbortext.com>, Terry Allen <tallen@sonic.net>, Norman Walsh <norm@berkshire.net>

Version:  3.0-17

gf A ”general formatter” for SGML documents

‘gf’ from Gary Houston is short for ”general formatter”, i.e., it can work on documents which use the ISO ”general” document type definition (DTD). It can convert SGML documents conforming to a small number of DTDs into various output formats: LaTeX, ASCII, RTF and Texinfo. However not every output format can be generated for every DTD. Apart from the general DTD, gf supports the HTML DTD used in the WWW project and Gary’s Snafu DTD. ‘gf’ is not intended as a flexible system for hacking up a formatter for a random DTD, but as a usable document production system for a few DTDs. Contents: gf-0.46

Copyright:  (c) (c) Gary Houston, GPL
Author:  Gary Houston <ghouston@actrix.gen.nz>
Version:  0.46-16

html dtd HTML DTDs (Document Type Definitions) + Docs

Document Type Definitions (DTDs) for HTML 2.0 (standard), HTML 3.2 (recommendation of the W3 Consortium), HTML 4.0 (Specification of the W3 Consortium, Draft) and ISO-HTML (draft). The package also contains the documentations (v. /usr/doc/packages/html dtd). Contents: html-3.2 html-spec-19950922 iso-html

Version:  0.4-1

hyperref Resolve LaTeX cross-references for the HyperTeX project

Copyright:  (c) Sebastian Rahtz
Author:  s.rahtz@elsevier.co.uk
Version:  6.14-7

iso ent Character Entity Sets f”ur ISO 1986:8879 (SGML)

Copyright:  (c) ISO
Version:  1.0-5

jade dsl DSSSL-Engine for SGML documents

Jade is an implementation of DSSSL (Document Style, Semantics and Specification Language); pronounce it as ”dissl” -- it rimes with whistle. The parser ”nsgmls” and helper tools like ”sgmlnorm”, ”spam”, ”spent”, and ”sx” are now included in the separate package ”sp”. You’ll find the documentation at /usr/doc/packages/jade dsl/.

Copyright:  (c) (c) James Clark, GPL
Author:  James Clark <jjc@jclark.com>, David Megginson <dmeggins@microstar.com>
Version:  1.1.1-2

jadetex JadeTeX - LaTeX macros to process TeX output from Jade (jade dsl)

With Sebastian Rahtz’ macro package ‘jadetex’ is is possible to process the output of the TeX backend of Jade (jade dsl). Resulting DVI files are viewable e.g., with ‘xdvi’ or printable like any other DVI file.

Contents: jadetex

Copyright:  (c) 1996,1997 Sebastian Rahtz <s.rahtz@elsevier.co.uk>
Author:  Sebastian Rahtz <s.rahtz@elsevier.co.uk>
Version:  0.57-7

linuxdoc linuxdoc-sgml

The file doc/guide.sgml is SGML source for the User’s Guide. As you probably don’t know how to format it yet, there is formatted plain ASCII in doc/guide.txt. There’s also guide.dvi and guide.ps, if you wish to print them instead. The first thing you should read is doc/guide.txt, which explains how to install the software, and how to start writing documents with it. The file doc/example.sgml is a simple example document which you can use as a tutorial. For real-life examples of uses of Linuxdoc-SGML, look at the Linux HOWTOs on sunsite.unc.edu:/pub/Linux/docs/HOWTO/other-formats/sgml. These HOWTOs are written in SGML and automatically translated into PostScript, dvi, and ASCII using this system. You can also look at the URL http://sunsite.unc.edu/mdw/HOWTO/ which has HTML formats of many Linux HOWTOs, produced using Linuxdoc-SGML. Contents: linuxdoc-sgml-1.5

Author:  Matt Welsh <mdw@sunsite.unc.edu>, Greg Hankins <gregh@cc.gatech.edu>
Version:  1.5-16

sgmltool SGML-Tools - a text-formatting package

SGML-Tools is a text-formatting package based on SGML (Standard Generalized Markup Language), which allows you to produce LaTeX, HTML, GNU info, LyX, RTF, and plain ASCII (via groff) from a single source. This system is tailored for writing technical software documentation, an example of which are the Linux HOWTO documents. It should be useful for all kinds of printed and online documentation. SGML-Tools is not able to process arbitrary SGML documents; in such a case, give jade dsl a try and write your own DSSSL scripts (take the docbk30 package as an example). Contents: sgml-tools-1.0.5 This package is considered as the successor of the Linuxdoc package. Production systems should probably prefere Linuxdoc for the moment.

Copyright:  (c) GPL

Author:

Cees A. de Groot <cg@pobox.com>, Arun Sharma <asharma@sco.com>, Andre Schoorl <aschoorl@engr.UVic.CA>, Daniele Giacomini <daniele@tv.shineline.it>, Dave Love <d.love@dl.ac.uk>, Dave Platt <dplatt@iq.TVSoft.com>, Eric S. Raymond <esr@thyrsus.com>, Francisco J. Montilla <pacopepe@nova.es>, Fabrizio Polacco <fpolacco@megabaud.fi>, Frank Pavageau <pavageau@imaginet.fr>, Jan Echternach <echter@informatik.uni-rostock.de>, Henry S. Thompson <ht@cogsci.ed.ac.uk>, Karsten Tinnefeld <tinnefeld@irb.informatik.uni-dortmund.de>, Kees de Bruin <kees de bruin@tasking.nl>, Ken MacLeod <kmacleod@midamerican.com>, Lee Marzke <marzke lee@si.com>, Marco Budde <Marco Budde@hqsys.fido.de>, Mitchell Perilstein <mnp@compass-da.com>, Sven Rudolph <sr1@os.inf.tudresden.de>, Tommy Marcus McGuire <mcguire@cs.utexas.edu>, Ulrich Windl <ulrich.windl@rz.uni-regensburg.de>

Version:  1.0.7-1

sgrep search a file for a structured pattern

Copyright:  (c) GPL
Author:  Jani Jaakkola <jjaakkol@cs.helsinki.fi>, Pekka Kilpelainen <kilpelai@cs.helsinki.fi>
Version:  0.99-14

sp SGML parser tools

The tools of this package provide the possibility to manage SGML and XML documents. It contains the parser ‘nsgmls’ and the supporting programs ‘sgmlnorm’, ‘spam’, ‘spent’, and ‘sx’. ‘sx’ is useful as a converting tool from SGML to XML, the comming WWW standard. You’ll find the documentation for all the programs under /usr/doc/packages/sp/.

Copyright:  (c) (c) James Clark, GPL
Author:  James Clark <jjc@jclark.com>
Version:  1.3-2

SND: Sound related stuff

0.1.15 SND: Sound related stuff

aumix A curses and command line based mixer

With aumix you may set up the mixer parameters via script. Contents: aumix-1.8

Copyright:  (c) 1993, 1996, 1997 the authors

Author:

Trevor Johnson <trevor@jpj.net>, Alessandro Rubini <alessandro.rubini@linux.it>, Kenn Humborg <kenn@wombat.ie>, Dan Fandrich <dan@fch.wimsey.bc.ca>, Jeff Spirko <spirko@topquark.physics.lehigh.edu>, Harley Silver <harley@widow.aracnet.net>, Rainer Canavan <canavan@cs.bonn.edu>

Version:  1.8-6

dap Digital Audio Processor

DAP is a comprehensive audio sample editing and processing suite. DAP currently supports AIFF and AIFF-C audio files, 8 or 16 bit resolution and 1, 2 or 4 channels of audio data. Note however that on Linux and Solaris, compressed AIFF-C files are not currently supported, only non-compressed AIFF and AIFF-C files. Call as /usr/bin/DAP

Copyright:  (c) 1997 Richard Kent
Author:  Richard Kent <richardk@cee.hw.ac.uk>
Version:  1.3-3

gmod Linux music module player

A module player which - similar to xmp - supports the Gravis Ultrasound and AWE soundcards. xgmod has a Qt-based frontend. Contents: gmod-3.1

Copyright:  (c) 1996 by Andrew Robinson
Author:  Andrew J. Robinson <robinson@cnj.digex.net>
Version:  3.1-15

jazz Midi sequencer

Complete midi sequencer which supports recording and playing as a bundle of editing functions. There are two main windows. One displays the complete track the other further details. It requires either a midi card supported by /dev/sequencer or a complete MPU 401 compatible device. Contents: jazz-2.1d

Copyright:  (c) 1994 Andreas Voss

Author:

Andreas  Voss  <andreas@avix.rhein-neckar.de>,  Per  Sigmond

<Per.Sigmond@hiagder.no>

Version:  2.1d-19

mxv WAV-player, modify and analyze samples, sound-synthesis too

With mxv you may play and modify sound files of different formats (wav, au, etc.). Mxv also provides the function of recording via /dev/dsp.

Copyright:  (c) Douglas Scott
Author:  Douglas Scott <MixViews@ccmrc.ucsb.edu>
Version:  1.10-1

rosegard Rosegarden Editor and Sequencer suite

Rosegarden is a free integrated musical notation editor and MIDI sequencer for Unix/X platforms, with specific support for Linux PCs and SGI IRIX workstations.

Copyright:  (c) Chris Cannam, Andrew Green, Richard Bown, 1997
Author:  Chris Cannam <cannam@zands.com>
Version:  2.1-0

rplay Sound daemon

rplay is a sound package which handles local as well as network sound output. A lot of programs use this daemon for sound output. Contents: rplay-3.3.0

Copyright:  (c) 1993, 1994 Mark Boyns
Author:  Mark Boyns <boyns@sdsu.edu>
Version:  3.3.0-6

sapphire An Audio-Tracer

Sapphire is an audio-tracer. You could imagine this as a ”ray tracer for sounds”. Sapphire has the ability to compute whole music pieces given by a description language. There is a File andrew.sp under /usr/lib/sapphire. Let your PPro working its little heart out. :-) Contents: sapphire

Copyright:  (c) 1993-1995 J. Finnis
Author:  Jim Finnis <>
Version:  2.1-13

slab slab

SLab Recording Studio Software Contents: slab-2.1

Copyright:  (c) Free
Version:  3.0-0

snd au Sound files in AU-format. (Sun Audio)

These are sound files in the AU format under /usr/share/sounds/au. So if you have a sound card and a kernel with sound support you can listen to those files with a simple ’cat <file> > /dev/audio’.

Version:  98.2.13-4

snd mod Some MOD-files (music on disc)

These are files for modplayers. They are containing sampled instruments and sequencer data which are interpreted by the players. One such player can be found in the package ”tracker”.

Copyright:  (c) Diverse
Version:  98.2.13-4

snd wav Soundfiles in WAV-format.

These are sound files in the WAV format under /usr/share/sounds/wav. So if you have a sound card and a kernel with sound support you can listen to those files with a simple ’play <file>’. Contents: sounds

Version:  98.2.13-4

sox Sound Conversion Tools and Library

SOX is intended as the Swiss Army knife of sound processing tools. It doesn’t do anything very well, but sooner or later it comes in very handy. SOX is really only useable day-to-day if you hide the wacky options with one-line shell scripts.

Copyright:  (c) 1991 Lance Norskog And Sundry Contributors
Author:  Chris Bagwell <cbagwell@sprynet.com>
Version:  12.14-6

speakf Voice Communication Over Data Networks

Voice Communication Over Data TCP/IP Networks. Supported compression algorithms are: GSM, ADPCM, and others. Sound can be encrypted either with a key supplied in a file, DES and/or IDEA with given key(s). Contents: speak freely-6.1c Please install a mixer (i.e. xmix) - a sound card, you need also.

Copyright:  (c) 1996 John Walker, Release 6.1b

Author:

John Walker <kelvin@fourmilab.ch>, ”Unmoderated list for discussion” <speak-freely@fourmilab.ch>, ”Periodic digest of messages” <speak-freely-digest@fourmilab.ch>, ”Announcements of general interest” <speak-freely-announce@fourmilab.ch>

Version:  6.1c-11

tcd Compact curses based CD player

Compact curses based CD player, with easier key control and more features than other CD players, such as track repeat, continuous play, track database, etc. Contents: tcd-1.0

Copyright:  (c) GPL
Author:  Tim P. Gerla <timg@means.net>
Version:  1.0-15

SND: Sound related stuff

tclmidi language designed for creating standard midi files

Tclmidi is language designed for creating and editing standard MIDI files. With the proper device interface (a kernel module in package kernmod or in this package as /usr/doc/packages/tclmidi/midi.o) it will also play and record MIDI files.

Copyright:  (c) 1993-96 Michael B. Durian
Author:  Michael B. Durian <durian@boogie.com>
Version:  3.1p3-6

tracker Tracker is a portable MOD-player for Linux.

Tracker allows to play standard-MOD files (MOD = Music On Disc) independent of the used sound card (/dev/audio support has to be in kernel). You’ll find examples of MOD-files in /usr/share/sounds/mod. Contents: tracker

Version:  4.3-19

wavplay Player / recorder for WAV files

From the man-page of wavplay : wavplay plays a WAV format file to the audio device, wavrec samples the audio device and writes a WAV file. xltwavplay provides a Motif-based graphical interface. Contents: wavplay-1.0

Copyright:  (c) GPL
Author:  Warren W. Gay <bx249@freenet.toronto.on.ca>
Version:  1.0-5

workbone text-mode program for playing audio CD Roms

WorkBone is an interactive control panel in text-mode to play audio CDs under Linux. It is a text-modus interface to the ”hardware.c” module of the X11 CD-program WorkMan. WorkBone is designed to easily play audio CDs in text-mode under Linux. WorkBone by purpose doesn’t have all those features WorkMan has. It is thought for those people who want to listen to their music with as minimal expense as possible. WorkBone uses the number block as a control panel to enable jumping from track to track, play CDs, stop etc. Furthermore the actual playing time is permanently actualized and displayed. WorkBone works with most CD-ROM drives. But there still are some NEC CD-ROMs that don’t work because they us not standardized own audio commands which results in a malfunction of WorkBone. WorkBone expects to find the device: /dev/cdrom . If you don’t have this device, change to /dev and soft link your current cd rom device to ’cdrom’. Contents: WorkBone-2.3

Copyright:  (c) 1994 Thomas McWilliams
Author:  Thomas McWilliams <thomas.mcwilliams@f615.n109.z1.fidonet.org>
Version:  2.3-18

workman Workman

Audio CD player based on OpenLook. Contents: WorkMan-1.3

Author:  Steven Grimm <koreth@hyperion.com>
Version:  1.3-19

xmcd Motif(tm) CD Audio Player

xmcd is an X11/Motif-based CD player utility; and cda is a command-line driven, non-graphical CD audio player. After installation you still have to configure your CD-player. Go to the directory /usr/X11R6/lib/X11/xmcd/config and run configure.sh as root. Documentation: man xmcd Contents: xmcd-2.2

Copyright:  (c) 1995-1997 Ti Kan
Author:  Ti Kan <ti@amb.org>
Version:  2.2-20

xmitsumi CD-Player panel for Mitsumi CD-ROM drives

A simple CD-Player for Mitsumi CD-ROM drives.

Contents: xmitsumi-0.75

Copyright:  (c) 1994 Liem Bahneman
Author:  Liem Bahneman <roland@cac.washington.edu>
Version:  0.75-19

xmix X11 interface for the Linux sound driver

If your sound card is supported by Linux, you can use this program to adjust the sound.

Copyright:  (c) 1992 Steve Haehnichen
Author:  Steve Haehnichen, Hal Brand
Version:  2.1-18

xmp eXtended Module Player

With xmp you may play several music formats: MOD, S3M, XM, 669, STM, MTM, PTM, OKT, FAR. Furthermore xmp uses the special hardware found in the GUS and AWE soundcards. Regular soundcards are only supported with the commercial OSS sound driver. Contents: xmp-0.99c

Copyright:  (c) 1996-1998 C Matsuoka, H Carraro Jr

Author:

Claudio Matsuoka <claudio@cpgei.cefetpr.br>, Takashi Iwai <iwai@dragon.mm.t.utokyo.ac.jp>

Version:  1.1.5-9

xqmixer Sound Mixer

a mixer for X-Windows using qt-libraries Contents: XqMixer-1.4

Copyright:  (c) GPL
Author:  Rajko Albrecht <rajko.albrecht@informatik.tu-chemnitz.de>
Version:  1.4-9

0.1.16 SPL: Spell checking utilities and databases

dicts Ispell dictionary files

This package contains sources of eight ispell dictionaries: for German (ready compiled for ispell in package igerman) for Danish (ready compiled for ispell in package idanish) for Spanish (ready compiled for ispell in package ispanish) for French (ready compiled for ispell in package ifrench) for Italian (ready compiled for ispell in package iitalian) for Dutch (ready compiled for ispell in package idutch) for Swedish (ready compiled for ispell in package iswedish) for Norwegian (ready compiled for ispell in package inorsk) Please read the README’s under /usr/src/dicts/i*/ . Contents: dutch96 espa˜ nol-1.5 francais-IREQ-1.4 hk2-deutsch idanish-1.2 iswedish-1.2 italian norsk prepare-dicts

Author:

Heinz  Knutzen  <hk@informatik.uni-kiel.d400.de>,  Martin  Boyer

<mboyer@robot.ireq.ca>, Marco Roveri <marco@mrg.dist.unige.it>, G”oran Andersson <goran@debian.org>, Santiago Rodriguez <srodri@fi.upm.es>, Jesus Carretero <jesus@fi.upm.es>, Piet Tutelaers <rcpt@urc.tue.nl>, Anders Andersen <aandersen@acm.org>

Version:  1.2-8

iamerica American and british ispell dictionaries

This packages includes ready american and british dictionaries for ispell. After installation of this package english will be american english (in oposite of the package iamerica). A short usage description is given in /usr/doc/packages/ispell/README of the packages ispell. The sources for this dictionary are included in the source package of ispell.

Copyright:  (c) 1983, by Pace Willisson, 1987 - 1993, 1995 by Geoff Kuenning, Granada Hills, CA.
Version:  3.1.20-15

SPL: Spell checking utilities and databases

ibritish British and american ispell dictionaries

This packages includes ready british and american dictionaries for ispell. After installation of this package english will be british english (in oposite of the package iamerica). A short usage description is given in /usr/doc/packages/ispell/README of the packages ispell. The sources for this dictionary are included in the source package of ispell.

Copyright:  (c) 1983, by Pace Willisson, 1987 - 1993, 1995 by Geoff Kuenning, Granada Hills, CA.
Version:  3.1.20-15

idanish Danish ispell dictionary

This packages includes a ready danish dictionary for ispell. A short usage description for ispell is given in /usr/doc/packages/ispell/README of the packages ispell. The sources for this dictionary are included in the package dicts.

Copyright:  (c) 1997 G”oran Andersson
Author:  G”oran Andersson <goran@debian.org>
Version:  1.2-8

idutch Dutch ispell dictionary

This packages includes a ready dutch dictionary for ispell. A short usage description for ispell is given in /usr/doc/packages/ispell/README of the packages ispell. The sources for this dictionary are included in the package dicts.

Copyright:  (c) 1997 Piet Tutelaers
Author:  Piet Tutelaers <rcpt@urc.tue.nl>
Version:  1.2-8

ifrench French ispell dictionary

/usr/doc/packages/ispell/README of the packages ispell. The sources for this dictionary are included in the package dicts.

Copyright:  (c) 1993-96 Martin Boyer
Author:  Martin Boyer <mboyer@ireq-robot.hydro.qc.ca>
Version:  1.2-8

iitalian Italian ispell dictionary

This packages includes a ready italian dictionary for ispell. A short usage description for ispell is given in /usr/doc/packages/ispell/README of the packages ispell. The sources for this dictionary are included in the package dicts.

Copyright:  (c) 1997 Marco Roveri
Author:  Marco Roveri <marco@mrg.dist.unige.it>
Version:  1.2-8

inorsk Norwegian ispell dictionary

This packages includes a ready norwegian dictionary for ispell. A short usage description for ispell is given in /usr/doc/packages/ispell/README of the packages ispell. The sources for this dictionary are included in the package dicts.

Copyright:  (c) Anders Andersen, Ivar Aavatsmark(?)
Author:  Anders Andersen <aandersen@acm.org>, Ivar Aavatsmark
Version:  1.2-8

ispanish Spanish ispell dictionary

This packages includes a ready spanish dictionary for ispell. A short usage description for ispell is given in /usr/doc/packages/ispell/README of the packages ispell. The sources for this dictionary are included in the package dicts.

Copyright:  (c) 1994-96 Santiago Rodriguez, Jesus Carretero
Author:  Santiago Rodriguez <srodri@fi.upm.es>, Jesus Carretero <jesus@fi.upm.es>
Version:  1.2-8

ispell ispell

Ispell is a fast screen-oriented spelling checker that shows you your errors in the context of the original file, and suggests possible corrections when it can figure them out. Compared to UNIX spell, it is faster and much easier to use. Ispell can also handle languages other than English. Ispell has a long history, and many people have contributed to the current version - some of the major contributors include R. E. Gorin, Pace Willisson, Walt Buehring, and Geoff Kuenning. You find a short description in the directory /usr/doc/packages/ispell/.

Contents: ispell-3.1.20

Copyright:  (c) 1983, by Pace Willisson, 1987 - 1993, 1995 by Geoff Kuenning, Granada Hills, CA.

Author:

R. E. Gorin, Pace Willisson <pace@ai.mit.edu>, Walt Buehring, Goeff Kuenning <geoff@ITcorp.com>

Version:  3.1.20-15

iswedish Swedish ispell dictionary

This packages includes a ready swedish dictionary for ispell. A short usage description for ispell is given in /usr/doc/packages/ispell/README of the packages ispell. The sources for this dictionary are included in the package dicts.

Copyright:  (c) 1997 G”oran Andersson
Author:  G”oran Andersson <goran@debian.org>
Version:  1.2-8

words English dictionary

This packages contains an English dictionary which will be installed as /usr/share/dict/words This dictionary is maybe used by look(1) and ispell(1).

Contents: linux.words.2

Author:  Rik Faith <faith@cs.unc.edu>
Version:  words.2-14

igerman German ispell dictionary

This packages includes a ready german dictionary for ispell. A short usage description is given in /usr/doc/packages/ispell/README of the packages ispell. The sources for this dictionary are included in the package dicts.

Copyright:  (c) 1993-95 Heinz Knutzen
Author:  Heinz Knutzen <hk@informatik.uni-kiel.d400.de>
Version:  1.2-8

0.1.17 TCL: Tcl/Tk/TclX, Tcl-Language and Tk-Toolkit for X

blt Tk extension

An add-on library for Tk. Does not require patching of the Tk and Tcl libraries. Provides some additional Widgets and Functions. These extensions only work with the tk old and tcl old libraries.

Contents: blt-1.9

Author:  George Howlett <george.howlett@att.com>
Version:  1.9-17

convert convert

As you see when you ran Convert this is a unit conversion software. (over 700 conversions)

Contents: convert-1.3

Copyright:  (c) (c) 1995 Costas Vlassi
Author:  Costas Vlassis <lonewolf@compulink.gr>
Version:  1.3-5

dejagnu dejagnu

DejaGnu contains an extension to tcl, called ’expect’. Dejagnu is a library of expect scripts that are used by GNU for running test suites on GNU tools. Expect is great for making e.g. login scripts for telnet, ftp or anything else. Or for any other automated programs that have to react to the output of an arbitrary program started by expect.

Contents: dejagnu-970221

Copyright:  (c) Copyright (C) 1992-1995 Free Software Foundation, Inc.
Author:  Cygnus
Version:  970221-15

TCL: Tcl/Tk/TclX, Tcl-Language and Tk-Toolkit for X

eddi eddi

X editor based on the TiX shell, with text highlighting and several other useful features. Contents: eddi-0.9b12

Copyright:  (c) Holger Schanz
Author:  Holger Schanz <holger.schanz@student.uni-ulm.de>
Version:  0.9b12-6

expect expect

Expect is a tool primarily for automating interactive applications such as telnet, ftp, passwd, fsck, rlogin, tip, etc. Expect really makes this stuff trivial. Expect is also useful for testing these same applications. Expect is described in many books, articles, papers, and FAQs. There is an entire book on it available from O’Reilly. Expect is necessary for dejagnu.

Copyright:  (c) public domain
Author:  libes@nist.gov
Version:  5.22-16

filerun FileRunner

FileRunner is an X11 file manager for Unix. It is simple and efficient and has a built-in FTP client.

Copyright:  (c) Henrik Harmsen
Author:  Henrik Harmsen <hch@cd.chalmers.se>
Version:  2.4.2-2

itcl object oriented extension for Tcl

incr Tcl adds object-oriented programming facilities to Tcl. It was NOT designed as yet another whiz-bang object-oriented programming language; indeed, it is patterned somewhat after C++. It was designed to support more structured programming in Tcl. Scripts that grow beyond a few thousand lines become extremely difficult to maintain. [incr Tcl] attacks this problem in the same way that any object- oriented programming language would, by providing mechanisms for data encapsulation behind well-defined interfaces. Please read the file /usr/doc/packages/itcl/README.SuSE. Contents: itcl2.2

Copyright:  (c) 1993-1995 AT&T Bell Laboratories,

Author:

Michael J. McLennan <michael.mclennan@att.com>, Dr. John Ousterhout <john.ousterhout@eng.sun.com>

Version:  2.2p2-3

jstools jstools

This is version 4.2/4.5 of the jstools distribution. jstools is a suite of applications, and some libraries they share. The most interesting application is multi-mode editor with support for WYSIWYG generation of HTML and PostScript, among other goodies. There are also a number of small graphical utilities intended to be used from within shell scripts or window-manager configuration files. Contents: jstools-tk4.2v4.5 jstools src

Copyright:  (c) Jay Sekora
Author:  Jay Sekora <js@aq.org>
Version:  4.5-16

listexec List browser

With this little program one can list all files in a specified directory and e.g. hear them with a specified sound player or view them as pictures with a picture viewer. The syntax may be obtained by listexec -h Contents: listexec-0.4.3

Copyright:  (c) 1997, S.u.S.E. GmbH
Author:  Marcus Sch”afer <ms@suse.de>, Martin Scherbaum <maddin@suse.de>
Version:  0.5-1

omnimoni System Monitor

A Tcl/Tk based fully configurable system-monitor. See /usr/X11R6/lib/omnimoni for details.

Contents: OmniMoni-1.0

Copyright:  (c) 1995 Rainer Mager
Author:  Rainer Mager <rvm@fore.com>
Version:  1.0-15

samt samt

SAmT is a tool for the System Administrator. It helps keeping a list of important system files without having to look for them each time you need them. It provides an easy means of rebooting the system or drive it into a certain runlevel respectively. Furthermore it can create a backup of your given files, thus giving you the opportunity of switching back to a ”running system”. Contents: samt Please be aware that this software is still considered beta!

Copyright:  (c) Michael Burghart S.u.S.E. GmbH 1997
Author:  Michael Burghart <mb@suse.de>
Version:  2.8-9

scotty Scotty - Tcl Extensions for Network Management Applications

Scotty is the name of a software package which allows to implement site specific network management software using high-level, string-based APIs. The software is based on the Tool Command Language which simplifies the development of portable network management scripts. The scotty source distribution includes two major components. The first one is the Tnm Tcl Extension which provides access to network management information sources. The second component is the Tkined network editor which provides a framework for an extensible network management system. Contents: scotty-2.1.7

Copyright:  (c) Juergen Schoenwaelder
Author:  J”urgen Sch”onw”alder <schoenw@cs.utwente.nl>
Version:  2.1.7-19

spectcl SpecTcl

A Interface builder for Tcl/Tk and Java directly from Sun Microsystems. Documentation: /usr/doc/packages/spectcl

Copyright:  (c) 1997 Sun Microsystems
Author:  Raymond Johnson <spectcl@tcl.eng.sun.com>, Ioi Lam, Allan Pratt
Version:  1.1-2

tcl The Tcl script language

Tcl is a simple to use text-based scripting language with many built-in features that make it especially nice for writing interactive scripts similar to perl. The TCL shell is /usr/bin/tclsh7.6 This package is required if you want to install Tk. Contents: tcl7.6p2

Copyright:  (c) The Regents of the University of California, Sun Microsystems, Inc,
Author:  Dr. John Ousterhout <john.ousterhout@eng.sun.com>
Version:  7.6p2-16

tcl new The Tcl script language

Tcl is a simple to use text-based scripting language with many built-in features that make it especially nice for writing interactive scripts similar to perl. The TCL shell is /usr/bin/tclsh8.0 This package is required if you want to install Tk. The bytecode compiler is new. The core of the Tcl interpreter has been replaced with an on-the-fly compiler that translates Tcl scripts to byte codes. Contents: tcl8.0b2

Copyright:  (c) The Regents of the University of California, Sun Microsystems, Inc,
Author:  Dr. John Ousterhout <john.ousterhout@eng.sun.com>
Version:  8.0p1-11

tcl old The Tcl script language version 7.4

Tcl is a simple to use text-based scripting language with many built-in features that make it especially nice for writing interactive scripts similar to perl. The TCL shell is /usr/bin/tclsh7.4 This package is required if you want to install Tk old.

Contents: tcl7.4p3

Copyright:  (c) The Regents of the University of California, Sun Microsystems, Inc,
Author:  Dr. John Ousterhout <john.ousterhout@eng.sun.com>
Version:  7.4p3-15

tclx TclX - Extended Tcl: Extended command set for Tcl

Extended Tcl is a superset of standard Tcl and is built alongside the standard Tcl sources. Extended Tcl has three basic functional areas: A set of new commands, a Tcl shell (i.e. a Unix shell-style command line and interactive environment), and a user-extensible library of useful Tcl procedures, any of which can be automatically loaded on the first attempt to execute it. In addition, a detailed help system is available for TCL/TK: tclhelp.

Contents: tclX7.6.0

Copyright:  (c) Copyright 1992-1995 Karl Lehenbauer and Mark Diekhans.
Author:  Karl Lehenbauer and Mark Diekhan <info@NeoSoft.com>
Version:  7.6.0-16

tix Tools for tk

The Tix library has by far the greatest collection of widgets for programming with Tcl/Tk. Highlights include: Hierarchical List box, Directory List/Tree View, Spreadsheet, Tabular List box, Combo Box, Motif style FileSelectBox, MS Windows style FileSelectBox, PanedWindow, NoteBook, Spin Control widget .... and many more. With these new widgets, your applications will look great and interact with your users in intuitive ways.

Contents: Tix4.1.0

Copyright:  (c) 1996, Expert Interface Technologies
Version:  4.1.0-16

tix new tix new

The Tix library has by far the greatest collection of widgets for programming with Tcl/Tk8.0. Highlights include: Hierarchical List box, Directory List/Tree View, Spreadsheet, Tabular List box, Combo Box, Motif style FileSelectBox, MS Windows style FileSelectBox, PanedWindow, NoteBook, Spin Control widget .... and many more. With these new widgets, your applications will look great and interact with your users in intuitive ways.

Copyright:  (c) Copyright (c) 1996, Expert Interface Technologies
Version:  4.1.0.006-7

tixlpq Tix-lpq

tix frontend for lpq

Contents: tixlpq-0.1

Copyright:  (c) GPL
Version:  0.1-7

tk TK toolkit for TCL

Tk is an extension to Tcl that allows you to quickly and easily build X11 applications that have the look and feel of Motif apps. The TK-Shell is /usr/X11R6/bin/wish4.2 These applications might as well be run on Windows 95/NT as well as on apple. Demo: /usr/X11R6/lib/tk4.2/demo/widget

Contents: tk4.2p2

Copyright:  (c) The Regents of the University of California. 1994-95 Sun Microsystems, Inc.
Author:  John Ousterhout <john.ousterhout@eng.sun.com>
Version:  4.2p2-14

tk new TK toolkit for TCL

Tk is an extension to Tcl that allows you to quickly and easily build X11 applications that have the look and feel of Motif apps. The TK-Shell is /usr/X11R6/bin/wish8.0 Native look and feel. The widgets have been rewritten to provide (nearly?) native look and feel on the Macintosh and PC. Demo: /usr/X11R6/lib/tk8.0/demo/widget Contents: tk8.0b2

Copyright:

(c) Copyright: The Regents of the University of California. 1994-95 Sun Microsystems Inc,

Author:  John Ousterhout <john.ousterhout@eng.sun.com>
Version:  8.0p1-11

tk old TK toolkit for TCL

Tk is an extension to Tcl that allows you to quickly and easily build X11 applications that have the look and feel of Motif apps. The TK-Shell is /usr/X11R6/bin/wish4.0 These applications might as well be run on Windows 95/NT as well as on apple. Demo: /usr/X11R6/lib/tk4.0/demo/widget Contents: tcl7.4p3 tk4.0p3

Copyright:  (c) The Regents of the University of California. 1994-95 Sun Microsystems, Inc.
Author:  John Ousterhout <john.ousterhout@eng.sun.com>
Version:  4.0p3-16

tkdesk TkDesk - an ultra comfortable graphical file/desktop manager

TkDesk is a graphical desktop and file manager for UNIX (especially Linux) and the X Window System. Compared with other file managers available, it offers the most complete set of file operations and services, plus gives the user the ability to configure most aspects of TkDesk in a powerful way. The reason for this is the use of Tcl/Tk as the configuration and (for the greatest part of TkDesk) implementation language. TkDesk has been influenced by various other systems and file managers: NeXT, for laying out the file browser windows, Apple Finder, for the idea of file annotations and, (shock horror), Windows 95, for some other (of course minor and unimportant ;-)) inspirations. This is a brief overview of the most prominent features of TkDesk: o Arbitrary number of automatically refreshed file browsers and file list windows, o Configurable file-specific pop-up-menus, o Drag and drop, o Files and directories may also be dropped onto the root window, o Configurable application bar, with several displays and cascadable pop-up menus for each button, files can also be dropped here, o History of visited directories, opened files, executed commands, and others, which is automatically saved to disk, o Find files through their annotation, name, contents, size or age, o Trash can for safe deletion of files and directories, o Calculation of disk usage for directory hierarchies, o All file operations (find, copy, disk usage, etc.) are carried out in the background, o Traversal of directory hierarchies through recursive cascaded menus, o Bookmarks, create menu entries for often used files/directories, o Comprehensive hypertextish online help, o Built-in multi-buffer and undo-capable editor, o Close coupling with Netscape and XEmacs, o Sound support, o Powerful configuration of nearly all aspect of TkDesk through Tcl/Tk, this also allows the Tcl-literate to extend TkDesk in arbitrary ways, o Free of charge! But see the file COPYING, or menu entry Help/License for information on usage and redistribution of TkDesk.

Copyright:  (c) Christian Bolik
Author:  Christian Bolik <zzhibol@rrzn-user.uni-hannover.de>
Version:  1.0-0

tkfont tkfont

It is a program similar to xfontsel. Contents: tkfont-1.1

Copyright:  (c) Neil Grant
Author:  Neil Grant <grantj@sfu.ca>
Version:  1.1-13

tkinfo tkinfo

TkInfo is a tk script to read GNU ”info” files and display them. TkInfo can be used stand alone (via WISH), or embedded within an application to provide integrated, on-line help. Contents: tkinfo-1.4

Copyright:  (c) Axel Boldt
Author:  Axel Boldt <boldt@math.ucsb.edu>
Version:  0.8-13

tkinspec Browser for Tk programs

Tkinspect serves for supervising running Tk applications. All variables, resources and procedures can be viewed and changed. Contents: tkinspect-5.1.6 tkinspect src

Author:  Sam Shen <SLShen@lbl.gov>
Version:  5.1.6-17

tkirc A pretty IRC client with GUI

Tkirc is a frontend for IRCII, written in Tcl/Tk. It can be customized with a file named .tkircrc in the users home-directory. An example .tkircrc can be found in /usr/doc/packages/tkirc/tkircrc-example Contents: tkirc1.202

Copyright:  (c) 1997 Andreas Gelhausen
Author:  Andreas Gelhausen <atte@gecko.North.DE>
Version:  1.202-5

tkmail TkMail

TkMail is an X windows interface to mail built using Tcl/Tk. Reading, sending, and managing mail messages can almost all be done using only the mouse (except for the body of the message, of course). TkMail writes its folders in the format of the standard Unix ’Mail’ command. It also uses ’sendmail’ commands to do the mail delivery work. Contents: tkmail-4.0beta9

Copyright:  (c) Christophe Le Bars
Author:  Christophe Le Bars <clebars@debian.org>
Version:  4.0beta8-23

tkman Manual-browser for X

A manual-browser for X with hyperlinks, history and more. Contents: tkman-1.8

Copyright:  (c) 1993-1996 T.A. Phelps
Author:  T.A. Phelps <phelps@ACM.org>
Version:  1.8-15

tkman8 Manual-browser for X

A manual-browser for X with hyperlinks, history and more. Contents: tkman-2.0b4

Copyright:  (c) T.A. Phelps
Author:  T.A. Phelps <phelps@ACM.org>
Version:  2.0b4-16

tkppp tkppp

tkppp is a graphical front end to manually control a PPP connection. It presents in a small window a button for starting/quitting PPP, a small status message area, informational readouts and option buttons. Contents: tkppp-1.4-alpha

Copyright:  (c) Eric Jeschke
Author:  jeschke@cs.indiana.edu
Version:  1.4-17

tkrat TkRat

TkRat is a graphical Mail User Agent (MUA) which handles MIME. It is mainly written in C but the user interface is done in tcl/tk.

Contents: tkrat-1.0.5

Copyright:  (c) Martin Forssen
Author:  maf@dtek.chalmers.se
Version:  1.0.5-10

tkstep TkSTEP

This is Tk 4.1p1 with some modifications to make it look like N*XTSTEP’s interface, plus some new widgets inspirated on N*XTSTEP. It also incorporates the DND drag and drop patches. Contents: tkstep4.2

Copyright:  (c) Alfredo K. Kojima
Author:  Alfredo K. Kojima <kojima@inf.ufrgs.br>
Version:  0.5b1-15

tkworld tkWorld

tkGrep: tkGrep is a Graphical User Interface to the Unix Grep command. It provides point and click access to Command Line Options, as well as a Toolbar for rapid execution of your grep command. tkSort: tkSort is just like tkGrep, except that it is built for the Unix sort command. tkREM: tkREM provides a Graphical User Interface for building Regular Expressions. The major modes (grep, Perl, Tcl, Python, Emacs, Sed & Awk) have buttons with meta-characters as their label. Contents: tkWorld-0.04 tkWorld src

Copyright:  (c) Wes Bailey
Author:  Wes Bailey <bailey@phys.ksu.edu>
Version:  0.04-14

tkxcd tkxcd

This is a diff front end with a look and feel based on Atria Clearcase xcleardiff. Both files are displayed in a window each and the differences are marked in different colors. Contents: tkxcd 1.1.0

Copyright:  (c) John C. Quillan
Author:  John C. Quillan <quillan@doitnow.com>
Version:  1.1.0-12

tuba tuba

debugger for Tcl/Tk Contents: tuba-2.1.p1

Copyright:  (c) GPL
Author:  John Stump <iliad@doitnow.com>
Version:  2.1.p1-3

vtcl Visual Tcl

Visual Tcl is a freely-available, high-quality application development environment for UNIX, Windows and Macintosh platforms. Written entirely in Tcl and generating pure Tcl should make porting either unnecessary or trivial. Contents: vt-1.10 vtcl

Copyright:  (c) Stewart Allen
Author:  Stewart Allen <stewart@neuron.com>
Version:  1.10b-13

xcdroast X-CD-Roast

CD-Writer-Program for X. Frontend for CD-related programs like ”cdwrite” and ”mkisofs”. Successor of the cdwtools-0.93. Tcl/Tk/Tix-based. Contents: xcdroast-0.96d-beta2 Please handle with extreme care! If a wrong SCSI device is entered this partition might be overwritten irreversibly (including partition tables and partitions belonging to other OS)

Copyright:  (c) Thomas Niederreiter
Author:  Thomas Niederreiter <p7003ad@sunmail.lrz-muenchen.de>
Version:  0.96b2-12

xdmcfg XDMConfig

XDMConfig enables you to easily setup your XDM environment. It provides a means of setting e.g. fonts, colours and more. Even the layout.

TEX: TeX/LaTeX and applications

Please note that this is still a Beta version!

Copyright:

(c) Dr.Stephan Reck <woody@rockaway.franken.de> Michael Burghart <mb@suse.de>

Author:  Dr.Stephan Reck <woody@rockaway.franken.de>, Michael Burghart <mb@suse.de>
Version:  0.1-1

xf Interface builder for TCL/TK

XF allows you to interactively create and modify an user-interface based on the Tcl/Tk package. Tcl/tk (version 4.0 or later) is needed to run XF. XF creates a Tcl/Tk script as its output, which can then be interpreted using wish or any other Tcl/Tk intepreter. See: /usr/doc/packages/xf. A HTML-Documentation for XF and TCL/Tk is also included. The new version xf4.0 is configured for the use with Tcl/Tk 7.5/4.1. Contents: xf

Copyright:  (c) 1992-1996 Sven Delmas
Author:  Sven Delmas <sven@cimetrix.com>
Version:  4.0-17

xmclock a tcl based screen-locker

a menu based tcl program for creating various pictures which you will see after the screen locks. Contents: xmclock-1.1.2

Version:  1.1.2-6

zircon zircon

Zircon is an Tcl/Tk based IRC II client. This application is well designed and offers various features and possibilities. Contents: zircon-1.17p4b12

Copyright:  (c) Lindsay F. Marshall 1993
Author:  Lindsay F. Marshall <zircon@catless.newcastle.ac.uk>
Version:  1.17-16

0.1.18 TEX: TeX/LaTeX and applications

bibview X11 front-end for BibTeX databases

Using bibview, entries in several BiB databases can be moved, manipulated, created and searched. If you prefer German to English menus, all you have to do is copy the file var/X11R6/lib/appdefaults/BibView.ger to var/X11R6/lib/app-defaults/BibView . Contents: bibview-2.1

Copyright:  (c) 1992 TU MUENCHEN
Author:  Holger Martin, Peter Urban, Armin Liebl <liebla@informatik.tu-muenchen.de>
Version:  2.1-12

chess chess

This package contains (La)TeX macros (LaTeX2.09) and Fonts for typesetting chess figures within a (La)TeX document. Please have a look at the examples and documentation under /usr/lib/teTeX/texmf/doc/latex/chess/. Contents: chess12

Copyright:  (c) Febr. 1992, Piet Tutelaers
Author:  Piet Tutelaers <rcpt@urc.tue.nl>
Version:  1.2-18

colortbl colortbl - Add color to LaTeX tables

Contents: colortbl

Author:  David Carlisle
Version:  97.6.29-17

detex TeX -> ASCII Converter

This filter removes (La-)TeX macros. Contents: deTeX-2.6

Copyright:  (c) 1986 Daniel Trinkle
Author:  Daniel Trinkle <trinkle@cs.purdue.edu>
Version:  2.6-14

dvi2tty A TeX to ASCII converter

dvi2tty converts a TeX-dvi file into a readable ASCII text file. You can then do full text searches without finding the TeX macros as well. For documentation see the man page (man dvi2tty) and /usr/doc/packages/dvi2tty.

Contents: dvi2tty-5.1

Copyright:  (c) 1989, 1995 Svante Lindahl / Marcel Mol
Author:  Svante Lindahl <zap@cs.kth.sunet>, Marcel Mol <marcel@duteca.et.tudelft.nl>
Version:  5.1-13

dviutils Tools for working on DVI files

This package contains some utilities for manipulating DVI documents. Page selection and rearrangement are supported, including arrangement into signatures for booklet printing, and page merging for n-up printing. dvibook rearranges pages into book signatures, adds blank pages as needed dviselect selects pages and page ranges dvitodvi performs general page rearrangement of selected pages dviconcat concatenation of DVI files Included you will find three simple scripts for scaling DIN A4 pages and rearranging two by two onto a DIN A4 page (a4toa5). But the number of pages must be even (which can be done by odd2even or dvibook).

Contents: Dviutils

Copyright:  (c) 1987, 1989 University of Maryland
Author:  Chris Torek (University of Maryland), Angus Duggan
Version:  94.3.19-9

l2h LaTeX2HTML converter

LaTeX2HTML lets you easily convert not too complex LaTeX-documents into the WWW format HTML. This enables one to have both a written and online version of even older LaTeX texts. Please look out for the corresponding variables in /etc/rc.config!!! Documentation may be found in: /usr/doc/packages/l2h

Contents: latex2html-96.1.h

Copyright:  (c) 1993 Nikos Drakos
Author:  Nikos Drakos <nikos@cbl.leeds.ac.uk>
Version:  96.1.h-14

l2hhtml Documentation for the l2h package

The LaTeX2html manual, in HTML, readable by any HTML browser. Call it e.g. by saying: arena file:/usr/doc/packages/l2h/manual/manual.html

Contents: l2hhtml-96.1

Copyright:  (c) 1993, Nikos Drakos
Author:  Nikos Drakos <nikos@cbl.leeds.ac.uk>
Version:  96.1-13

lyx LyX - LaTeX based WYSIWYG editor

LyX is a frontend for LaTeX under X11. It combines the comfortable use of a word processor with the high quality of LaTeX type setting. Documents are almost WYSIWYG-like displayed. Low level attributes (”large italic”) are not directly chosen, but high level layouts for each paragraph. Of course, low level formatting is still possible. Using the LaTeX compiler, the LyX document is translated into a .dvi file. The reverse process, to process an already existing LaTeX document with the LyX editor, is not possible (except as a plain ASCII file).

TEX: TeX/LaTeX and applications

Contents: lyx-0.12.0

Copyright:  (c) 1995-97 Matthias Ettrich
Author:  Matthias Ettrich <ettrich@informatik.uni-tuebingen.de>
Version:  0.12.0pre8-5

musictex MusiXTeX

This packages is only usable for TeX-experts with musical ambitions. It contains a music Tex format file musixtex for direct usage onto TeX music files. LaTeX users should use the corresponding style file. For a description you should look under /usr/lib/teTeX/texmf/doc/musixtex/. Contents: musixexa-T86 musixtex-T86

Copyright:  (c) 1995 Daniel Taupin

Author:

Daniel Taupin <taupin@lps.u-psud.fr>, Ross Mitchell <rmm@dar.csiro.au>, Andreas Egler <Andreas Egler@wob.maus.de>

Version:  5.15-13

psutils Tools for manipulating PostScript files

This archive contains some utilities for manipulating PostScript documents. Page selection and rearrangement are supported, including arrangement into signatures for booklet printing, and page merging for n-up printing. psbook rearranges pages into signatures psselect selects pages and page ranges pstops performs general page rearrangement and selection psnup put multiple pages per physical sheet of paper psresize alter document paper size epsffit fits an EPSF file to a given bounding box You will find a README under /usr/doc/packages/psutils/ which also describes several perl-scripts for importing PostScript files. A manual page for each ps utility is also included. Contents: psutils-p17

Copyright:  (c) 1991-1995 Angus J. C. Duggan
Author:  Angus J. C. Duggan <angus@harlequin.co.uk>
Version:  p17-9

tetex TeX, LaTeX, BibTeX, makeindex, MetaFont, MetaPost, dvips, xdvi

After installing teTeX you will find a large selection of documentation for TeX, LaTeX, and various layout styles under /usr/lib/teTeX/texmf/doc. TeX (pronounced tech) is an interpreter for text formating and was developed by Donald E. Knuth. It works with control and macro commands on a text file. Working with TeX is similar to good old typesetting methods. LaTeX or better LaTeX2e is a complex macro package which removes the cryptical TeX interface, and does most of the work for the user. TeX uses special fonts produced by the MetaFont program. Various printer drivers and an X11 viewer is also included in this package. The teTeX package is based on the standard TeX package of Karl Berry which makes configuration much easier. It’s also possible to use PostScript fonts; a real PostScript printer is required however. If the ghostscript (gs) package is installed all drivers for printing and viewing can use these fonts. Note, however that the fonts included in the ghostscript package are not identical to Adobe’s PostScript fonts. The copyright prohibits us from including them on the CD. Beside these features there are the programs MakeIndex (for producing indexes), BibTeX (for literature data-processing), and the script xtexsh (requires tcl/tk). Note the differences between LaTeX2.09 and LaTeX2e. LaTeX2e can process original LaTeX2.09 text but style files that require LaTeX2.09 internal macros may not work correctly. Under /usr/lib/teTeX/texmf/doc/latex/base/ you will find the documentation for LaTeX2e. New users should review the documentation stored there. The teTeX package from Thomas Esser includes a full texmf tree, many programs (tex, dvips, ...), shell scripts configuration, and a big collection of documentations. This package is simple configured by the script texconfig and has thereby multi-language options. Do NOT install this package together with ntex! teTeX is NOT an upgrade or update for the ntex package!! Contents: ec-ready-mf-tfm teTeX-lib-0.4pl8 teTeX-src-0.4pl7 texmf-src-0.4pl8

Author:

Karl Berry , Tim Morgan, Tomas Rokicki, Donald Ervin Knuth, Thomas Esser <te@informatik.uni-hannover.de>

Version:  0.4pl8-26

texcad Drawing program for LaTeX pictures

A really nice drawing program which is especially useful in cooperation with LaTeX. Please note the documentation in /usr/doc/packages/texcad and the file COPYRIGHT there in. Contents: xtexcad-2.4

Copyright:  (c) Klaus Zitzmann, 1993-1996 Johannes Sixt

Author:

Klaus  Zitzmann  <zitzmann@infko.uni-koblenz.de>,  J.K.Wight

<J.K.Wight@newcastle.ac.uk>, Byron Rakitzis <byron@netapp.com>, Johannes Sixt <Johannes.Sixt@eudaptics.co.at>, Notker Amann <amann@isr.uni-stuttgart.de>

Version:  2.4-12

tktex tktex

TkTeX is a Tk-based interface to the UN*X tools related to TeX, such as editors, previewers, printing programs, etc. Contents: tktex-2.2

Copyright:  (c) Norman Danner
Author:  Norman Danner <ndanner@indiana.edu>
Version:  2.2b1-15

transfig Graphic converter

TransFig is a set of tools for creating TeX documents with graphics which are portable, in the sense that they can be printed in a wide variety of environments. The transfig directory contains the source for the transfig command which generates a Makefile which translates Fig code to various graphics description languages using the fig2dev program. In previous releases, this command was implemented as a shell script. Documentation: man transfig Contents: transfig.3.2

Copyright:

(c) 1985-1994 Supoj Sutantavibul, Frank Schmuck, Konrad Kwock, Herbert Bauer, Micah Beck

Author:

Anthony Starks <ajs@merck.com>, George Ferguson <ferguson@cs.rochester.edu>, Herbert Bauer <heb@regent.e-technik.tu-muenchen.de>, Micah Beck <supoj@sally.utexas.edu>, Supoj Sutantavibul <beck@cs.utk.ecu>

Version:  3.2-15

ts Tcl/Tk based shell for LaTeX

My aim was writing a TeX environment similar to the DOS-TeXShell by Schlegelmilch. That is, you can use their short keys for editing (F3), composing (F6), or showing (F8). The complete list of keys is shown in the online help (press F1). Menus, Keys, and Warnings are in English, but you can use German or English online help. Keys, warnings end menus are in English (since they are shorter) but help is available in German and English. Will be installed in /usr/lib/ts and some READMEs under /usr/doc/packages/ts If you want to use the English help-system, change the symlink under /usr/lib/ts to the right path (there is no parameter): cd /usr/lib/ts ln -sf eng help Contents: ts-9710

Copyright:  (c) 1996 Jens P”onisch
Author:  Jens P”onisch <J.Poenisch@wirtschaft.tu-chemnitz.de>
Version:  9804-2

xfig Facility for Interactive Generation of figures under X11

Xfig is a menu-driven tool that allows the user to draw and manipulate objects interactively in an X window. The resulting pictures can be saved, printed on PostScript printers or converted to a variety of other formats (e.g. to allow inclusion in LaTeX documents).

X: Base X Window System - XFree86^{TM}3.3

Documentation: /usr/doc/packages/xfig & man xfig Examples: /usr/doc/packages/xfig/Examples Contents: xfig.3.2

Copyright:  (c) 1985 Supoj Sutanthavibul, 1991-94 Paul King, Brian V. Smith

Author:

Anthony Dekker <dekker@ACM.org>, Brian V. Smith <bvsmith@lbl.gov>, Jim Daley <jdaley@cix.compulink.co.uk>, Ross Martin <martin@trcsun3.eas.asu.edu>, Uwe Bonnes <bon@lte.e-technik.uni-erlangen.de>

Version:  3.2-17

xtem a TeX menu system for X-Window

xtem is a TeX environment under X. This package is rather powerful and offers a huge help and reference system which should help the TeX novices. TeX administrators estimate this environment. The authors of this package ask for feedbacl for continuing their work! For more information lookup /usr/doc/packages/xtem. Contents: xtem src xtem texmenu.6.18

Copyright:  (c) 1996 G. Lamprecht, W. Lotz, R. Weibezahn; IWD, Bremen University
Author:  G. Lamprecht, W. Lotz, R. Weibezahn <weibezahn@iwd.uni-bremen.de>
Version:  6.18-7

0.1.19 X: Base X Window System - XFree86^{TM}3.3

sax S.u.S.E. advanced X Configuration

SaX stands for ( SuSE advanced XF86 configurator ) and is an alternativ possibility to configure the X window system rather easily.

Copyright:  (c) 1998,S.u.S.E. GmbH,F”urth
Author:  Marcus Sch”afer <ms@suse.de>
Version:  1.2-3

xaw3d 3D Athena Widgets

This is a library that can be used instead of the standard Athena-Widget-Library. It was tried to keep the standard of the libXaw library. There are also programs which explicitely use this library (this is the reason why the library was included). NOTE: Do NOT replace /usr/X11R6/lib/libXaw.so.6.1! Contents: Xaw3d-1.4

Copyright:  (c) 1989, 1994 X Consortium

Author:

Kaleb S. Keithley <kaleb@x.org>, Achille Hui <eillihca@drizzle.stanford.edu>, Robert Withrow <witr@rwwa.com>, Malcolm Strickland <chuckstrickland@orl.mmc.com>, David Flanagan <david@ora.com>, Jamie Zawinski <jwz@netscape.com>, Frank Terhaar-Yonkers <fty@bizarre.trpnc.epa.gov>, Mike Schulze <mike@cs.curtin.edu.au>, Gustaf Neumann <neumann@mohegan.wiinf.uni-essen.de>, Jerry Whelan <guru@stasi.bradley.edu>, Mark Rawling <mwr@mel.dit.csiro.au>, Tim Theisen <tim@cs.wisc.edu>, Mitch Trachtenberg <mitch@mta.com>, Keith Packard <keithp@ncd.com>, Dimitri P. Bouras <dimitri@ee.ubc.ca>, Uri Blumenthal <uri@watson.ibm.com>, Heiko Schroeder <heiko@pool.informatik.rwth-aachen.de>

Version:  1.4-5

xaw3dd 3D Athena Widgets as replacement for Athena Widgets

This version of libXaw3d will be installed in /usr/X11R6/lib/Xaw3d/ as libXaw.so.6.1. Note that this version is linked with the option ‘-soname libXaw.so.6’ and therefore it will be able to replace the regular libXaw.so.6.1 in /usr/X11R6/lib/. NOTE: Do NOT replace /usr/X11R6/lib/libXaw.so.6.1. All necessary changes in /etc/ld.so.conf will be done by S.u.S.E configuration done by YaST. On any problem using or starting X11 programs please remove this package and delete any remaining symbolic link under /usr/X11R6/lib/Xaw3d/. On any problem starting or using X11 programs, you should remove this package. Please read the package description.

Author:

Kaleb S. Keithley <kaleb@x.org>, Achille Hui <eillihca@drizzle.stanford.edu>, Robert Withrow <witr@rwwa.com>, Malcolm Strickland <chuckstrickland@orl.mmc.com>, David Flanagan <david@ora.com>, Jamie Zawinski <jwz@netscape.com>, Frank Terhaar-Yonkers <fty@bizarre.trpnc.epa.gov>, Mike Schulze <mike@cs.curtin.edu.au>, Gustaf Neumann <neumann@mohegan.wiinf.uni-essen.de>, Jerry Whelan <guru@stasi.bradley.edu>, Mark Rawling <mwr@mel.dit.csiro.au>, Tim Theisen <tim@cs.wisc.edu>, Mitch Trachtenberg <mitch@mta.com>, Keith Packard <keithp@ncd.com>, Dimitri P. Bouras <dimitri@ee.ubc.ca>, Uri Blumenthal <uri@watson.ibm.com>, Heiko Schroeder <heiko@pool.informatik.rwth-aachen.de>

Version:  1.4-5

xaw95 3D Athena Widgets with W95 look and feel

This is a library that can be used instead of the standard Athena-Widget-Library. It was tried to keep the standard of the libXaw library. NOTE: Do NOT replace /usr/X11R6/lib/libXaw.so.6.1, use packages xaw95d instead. Contents: Xaw95-1.1

Copyright:  (c) 1989, 1994 X Consortium

Author:

Kaleb S. Keithley <kaleb@x.org>, Achille Hui <eillihca@drizzle.stanford.edu>, Robert Withrow <witr@rwwa.com>, Malcolm Strickland <chuckstrickland@orl.mmc.com>, David Flanagan <david@ora.com>, Jamie Zawinski <jwz@netscape.com>, Frank Terhaar-Yonkers <fty@bizarre.trpnc.epa.gov>, Mike Schulze <mike@cs.curtin.edu.au>, Gustaf Neumann <neumann@mohegan.wiinf.uni-essen.de>, Jerry Whelan <guru@stasi.bradley.edu>, Mark Rawling <mwr@mel.dit.csiro.au>, Tim Theisen <tim@cs.wisc.edu>, Mitch Trachtenberg <mitch@mta.com>, Keith Packard <keithp@ncd.com>, Dimitri P. Bouras <dimitri@ee.ubc.ca>, Uri Blumenthal <uri@watson.ibm.com>, Heiko Schroeder <heiko@pool.informatik.rwth-aachen.de>, Eddie Hiu-Fung, Lau <eddielau@cuhk.edu.hk>

Version:  1.1-13

xaw95d Replacement for Athena Widgets with W95 look and feel

This version of libXaw95 will be installed in /usr/X11R6/lib/Xaw95/ as libXaw.so.6.1 and libXaw3d.so.6.1. Note that this versions are linked with the option ‘-soname libXaw.so.6’ respectively ‘-soname libXaw3d.so.6’ and therefore they will be able to replace the regular libXaw.so.6.1 or libXaw3d.so.6.1 in /usr/X11R6/lib/. NOTE: Do NOT replace /usr/X11R6/lib/libXaw.so.6.1 or /usr/X11R6/lib/libXaw3d.so.6.1. All necessary changes in /etc/ld.so.conf will be done by S.u.S.E configuration done by YaST. On any problem using or starting X11 programs please remove this package and delete any remaining symbolic link under /usr/X11R6/lib/Xaw95/.

Copyright:  (c) 1989, 1994 X Consortium

Author:

Kaleb S. Keithley <kaleb@x.org>, Achille Hui <eillihca@drizzle.stanford.edu>, Robert Withrow <witr@rwwa.com>, Malcolm Strickland <chuckstrickland@orl.mmc.com>, David Flanagan <david@ora.com>, Jamie Zawinski <jwz@netscape.com>, Frank Terhaar-Yonkers <fty@bizarre.trpnc.epa.gov>, Mike Schulze <mike@cs.curtin.edu.au>, Gustaf Neumann <neumann@mohegan.wiinf.uni-essen.de>, Jerry Whelan <guru@stasi.bradley.edu>, Mark Rawling <mwr@mel.dit.csiro.au>, Tim Theisen <tim@cs.wisc.edu>, Mitch Trachtenberg <mitch@mta.com>, Keith Packard <keithp@ncd.com>, Dimitri P. Bouras <dimitri@ee.ubc.ca>, Uri Blumenthal <uri@watson.ibm.com>, Heiko Schroeder <heiko@pool.informatik.rwth-aachen.de>, Eddie Hiu-Fung, Lau <eddielau@cuhk.edu.hk>

Version:  1.1-13

xdevel Developing under XFree86^{TM}

This package contains the include files, libraries and the imake templates for the X-Window system. It is needed for the development and the compilation of X11 programs.

Version:  3.3.2.3-5

xdmbgrd S.u.S.E. Linux background

The S.u.S.E. Linux background for your xdm workstation.

X: Base X Window System - XFree86^{TM}3.3

Contents: xdmbgrd

Copyright:  (c) Copyright (c) 1997 S.u.S.E. GmbH Fuerth
Author:  Werner Fink <werner@suse.de>
Version:  0.2-8

xdmsc Xterminal -- Usage S.u.S.E. Linux as X terminal

Some useful scripts for using S.u.S.E. Linux as X terminal. You will find the documentation in the following directory /usr/doc/packages/xdmsc/

Contents: Xterminal-0.2

Copyright:  (c) 1997 S.u.S.E GmbH
Author:  Werner Fink <werner@suse.de>
Version:  0.1-13

xdoc Japanese documentation of XFree86^{TM}

The English documentation is part of xf86.

Version:  3.3.2.3-5

xextra Two hardware independent X-servers, Xvfb and Xnest

Xvfb - a virtual framebuffer X server. Xvfb is an X server that can run on machines with no display hardware and no physical input devices. It emulates a dumb framebuffer using virtual memory. Xnest -- a nested X server. Xnest is a client of the real server which manages windows and graphics requests on its behalf. Xnest is a server to its own clients. Xnest manages windows and graphics requests on their behalf.

Version:  3.3.2.3-5

xf86 Basic X11 package

Contains XFree86^{TM}and some other programs from the contrib-directory.

Contents: XFree86-3.3.2.3 latin2-fonts

Copyright:  (c) 1994-1998 The XFree86 Project, Inc.

Author:

Alan  Hourihane  <alanh@fairlite.demon.co.uk>,  Bernhard  Bender

<br@elsa.mhs.compuserve.com>, Craig Groeschel <craig@metrolink.com> , David Dawes <dawes@XFree86.Org>, David Wexelblat <dwex@XFree86.Org>, Dirk Hohndel <hohndel@suse.de>, Erik Nygren <nygren@mit.edu>, Gertjan Akkerman <akkerman@dutiba.twi.tudelft.nl>, Glenn Lai <glennlai@netcom.com>, Hans Nasten <nasten@everyware.se>, Harald Koenig <koenig@tat.physik.unituebingen.de>, Harm Hanemaayer <hhanemaa@cs.ruu.nl>, Harry Langenbacher <harry@brain.jpl.nasa.gov>, Henrik Harmsen <harmsen@eritel.se>, Henry Worth <henry.worth@amail.amdahl.com>, Jim Tsillas <jtsilla@ccs.neu.edu>, Joe Moss <joe@morton.rain.com>, Jon Block <block@frc.com>, Jon Tombs <tombs@XFree86.org>, Kevin Martin <martin@cs.unc.edu>, Leonard N. Zubkoff <lnz@dandelion.com>, Marc Evans <marc@XFree86.org>, Marc La France <Marc.La-France@ualberta.ca> , Michael Rohleder <michael.rohleder@stadtfrankfurt.de>, Mike Hollick <hollick@graphics.cis.upenn.edu>, Norbert Distler <Norbert.Distler@physik.tu-muenchen.de> , Orest Zborowski <orestz@eskimo.com>, Paolo Severini <lendl@dist.dist.unige.it>, Pascal Haible <haible@izfm.uni-stuttgart.de>, Peter Trattler <peter@sbox.tu-graz.ac.at>, Rich Murphey <rich@XFree86.org>, Rik Faith <faith@cs.unc.edu>, Robin Cutsha, <robin@XFree86.org>, Simon Cooper <sc@sgi.com>, Stuart Anderson <anderson@metrolink.com>

Version:  3.3.2.3-5

xfnt100 100dpi screen fonts for the X-server

These fonts are compressed to save disk space.

Version:  3.3.2.3-5

xfntbig Big fonts for the X-server

These fonts are compressed to save disk space. These are the asiatic fonts (including Kanji, Jiskan) which the ’average’ european user won’t need..;-)

Version:  3.3.2.3-5

xfntcyr Cyrillic fonts

These fonts are compressed to save disk space.

Version:  3.3.2.3-5

xfntscl Scalable fonts for X11

Speedo and Type 1 fonts for XFree86^{TM}.

Version:  3.3.2.3-5

xfsetup The XFree86 configuration tool ”XF86Setup”.

Please also install the X server for your graphics board and invoke the configuration tool by entering XF86Setup...

Copyright:  (c) 1994, 1995, 1996 The XFree86 Project, Inc.
Version:  3.3.2.3-5

xlkit XFree86^{TM}Linkkit

If you want or need to rebuild an XFree86^{TM}server but don’t want to install the full source for XFree86-3.3.1, you need this package in addition to the base configuration and the programmers kit.

Version:  3.3.2.3-5

xman on-line manuals for XFree86^{TM}

Manpages for programs that come with XFree86^{TM}.

Version:  3.3.2.3-5

xpm The X-Pixmap libraries

The shared libraries are needed to run programs compiled with Xpm, such as xfilemanager and fvwm. This package contains libXpm.so.4. Contents: xpm-3.4j

Author:  Arnaud Le Hors <lehors@sophia.inria.fr>
Version:  3.4j-10

xshared X11 shared libs

Since some programs were linked with the X11 libraries but do work on a text console as well (e.g. emacs), we have put the X11 dynamic libraries into an extra package so you can use such programs without installing the entire X11 system.

Version:  3.3.2.3-5

xxprt X Print Server

Contains the X print server.

Copyright:  (c) 1994 - 1997 The XFree86 Project, Inc.
Version:  3.3.2.3-5

oldlibs5 Libraries from XFree86^{TM}2.1.1 (X11R5)

This package installs libX11.so.3.1.0, libXaw.so.3.1.0 and libXt.so.3.1.0 libraries in the /usr/X11R6/lib directory. Almost all programs are linked with the new libraries, however there are a few binaries that require the old libraries. Contents: oldlibs5

Version:  97.6.29-6

oldlibs6 a.out (DLL) shared libraries for X11 binaries

These libraries are needed by X programs in the old a.out binary format. They will be installed in /usr/i486-linuxaout/lib. This package also contains the ”.sa” libraries which allow the generation of a.out binaries using ’gcc -b i486linuxaout’.

Version:  97.7.23-3

modegen Roll your own X video mode

Two ’sc’ spreadsheets to create X mode lines. Remember that this package is ’for experts only’ - normally XF86Setup (in package xfsetup) does everything you need.

X3D: 3D stuff for X11 and console

Contents: modegen

Copyright:  (c) 1994 Mike Jagdis
Author:  Mike Jagdis <jaggy@purplet.demon.co.uk>
Version:  98.2.3-5

0.1.20 X3D: 3D stuff for X11 and console

dev3dfx Device driver for 3Dfx Voodoo boards

This package installs the 3Dfx device driver to allow access to 3Dfx Voodoo boards without the user having root privledges.

Copyright:  (c) Daryll Strauss <daryll@harlot.rb.ca.us>
Author:  Daryll Strauss <daryll@harlot.rb.ca.us>
Version:  1-0

glide v2 Glide Library for 3Dfx Voodoo II boards

This package installs the Glide Library for 3Dfx Voodoo II boards.

Copyright:  (c) 3Dfx Interactive, Inc.
Author:  Daryll Strauss <daryll@harlot.rb.ca.us>, 3Dfx Interactive, Inc. <info@3dfx.com>
Version:  2.51-2

glide vg Glide Library for 3Dfx Voodoo Graphics boards

This package installs the Glide Library for 3Dfx Voodoo Graphics boards.

Copyright:  (c) 3Dfx Interactive, Inc.
Author:  Daryll Strauss <daryll@harlot.rb.ca.us>, 3Dfx Interactive, Inc. <info@3dfx.com>
Version:  2.43-2

glide vr Glide Library for 3Dfx Voodoo Rush boards

This package installs the Glide Library for 3Dfx Voodoo Rush boards.

Copyright:  (c) 3Dfx Interactive, Inc.
Author:  Daryll Strauss <daryll@harlot.rb.ca.us>, 3Dfx Interactive, Inc. <info@3fx.com>
Version:  2.43-2

glidesdk Glide2x SDK for 3Dfx Voodoo boards

This package installs the Glide2x SDK for 3Dfx Voodoo boards.

Copyright:  (c) 3Dfx Interactive, Inc.
Author:  Daryll Strauss <daryll@harlot.rb.ca.us>, 3Dfx Interactive, Inc. <info@3dfx.com>
Version:  2.1-2

mesa Mesa

Mesa is a 3-D graphics library with an API which is very similar to that of OpenGL*. To the extent that Mesa utilizes the OpenGL command syntax or state machine, it is being used with authorization from Silicon Graphics, Inc.

Contents: MesaDemos-2.6 MesaLib-2.6

Copyright:  (c) Brian Paul
Author:  Brian Paul <brianp@elastic.avid.com>
Version:  2.6-6

mesa3dfx Mesa w/ 3Dfx Glide driver

Mesa is a 3-D graphics library with an API which is very similar to that of OpenGL*. To the extent that Mesa utilizes the OpenGL command syntax or state machine, it is being used with authorization from Silicon Graphics, Inc. This version of Mesa is hardware accelerated for 3Dfx Voodoo Chips.

Copyright:  (c) Brian Paul
Author:  Brian Paul <brianp@elastic.avid.com>
Version:  2.6-0

rushlibs Additional X11 Libraries for 3Dfx Voodoo Rush boards

This package installs additional X11 Libraries for 3Dfx Voodoo Rush boards.

Copyright:  (c) Daryll Strauss <daryll@harlot.rb.ca.us>
Author:  Daryll Strauss <daryll@harlot.rb.ca.us>
Version:  3.2.2-2

rushsvga Hardware accelerated 3D X Server for 3Dfx Voodoo Rush boards

This package installs the Hardware-accelerated 3D X Server for 3Dfx Voodoo Rush boards.

Copyright:  (c) Daryll Strauss <daryll@harlot.rb.ca.us>
Author:  Daryll Strauss <daryll@harlot.rb.ca.us>
Version:  3.2.2-2

0.1.21 XAP: X Applications

addrbo Address book

The program ”addressbook” lets you manage your addresses, phone numbers, e-mail addresses, and fax numbers.

Contents: addrbo src addressbook-0.6.1

Copyright:  (c) Clemens Durka
Author:  Clemens Durka <durka@informatik.tu-muenchen.de>
Version:  0.6-16

amaya WebBrowser by W3Consortium

A real nice browser with good capabilities but still a beta release. This is the current w3-testbed-client.

Contents: amaya-src-1.2a

Copyright:  (c) Copyright 1995 (MIT) (INRIA)
Author:  Irene Vatton <irene.vatton@w3.org>
Version:  1.3-0

arena WWW-Browser

This WWW browser is one of the few that do not fall under the restrictive copyright conditions of the HTML library. We greatly regret that we are not allowed to distribute popular browsers such as Mosaic and Chimera on our CDs, although these are freely available for private use in the Internet. The copyright conditions set by Mosaic forbid us from doing so.

Contents: arena

Copyright:  (c) 1994 CERN

Author:

Hakon W Lie <arena@info.cern.ch>, Dave Raggett <arena@info.cern.ch> , Henrik <arena@info.cern.ch>, Phill <arena@info.cern.ch>

Version:  beta.1e-13

argoups UPS management software

With the programs of this package you can controll the actual condition of an USV. If the power fails the system makes a shutdown before you get crazy :-)

Contents: argoups

Copyright:  (c) COPYRIGHT (c) CORES INFORMATICA
Version:  1.0-1

arrow Arrow

Arrow was designed to be a full-featured, yet easy to use mail program. Arrow still has a great deal of functionality, including secure mail via PGP and a spell checker. One unique feature of Arrow is its drag-and-drop functionality. Messages can be dragged from one mail file and moved to another mail file, with each mail file viewed in a separate window.

Copyright:  (c) 1998 by Glenn Bach
Author:  Glenn Bach <glenn@cco.wonderland.caltech.edu>
Version:  0.5-5

asedit An Editor

A simple ASCII text editor for X11/Motif. Mouse support, dialog boxes, hypertext online help, undo/redo. The help box for the online help has a little bug: it is too small, you must enlarge the box with the mouse.

Contents: asedit-1.3.2

Copyright:  (c) 1994 Andrzej Stochniol
Author:  Andrzej Stochniol <A.Stochniol@ic.ac.uk>
Version:  1.3.2-18

XAP: X Applications

auis ful Andrew User Interface System - Full System

AUIS is a collection of tools which enable you to design, view and mail multimedia documents. This is a complete AUIS Distribution. It is designed for users who want to create their own AUIS applications. If this isn’t what you want, you should consider installing the smaller packages. Please read also the README under /usr/andrew. Contents: auis1 auis2 auis3 auis4 contrib doc This package also includes the other AUIS packages. So please install either this package, or the other packages!

Copyright:  (c) 1992 Carnegie Mellon University & IBM
Author:  Andrew Consortium <info-andrew-request@andrew.cmu.edu>
Version:  635L1-0

auis ml Andrew User Interface System - Mail System

AUIS is a collection of tools which enable you to design, view and mail multimedia documents. This package is aimed at all those who want a ”better mail system” and are interested in a multi-media mail-interface (MIME). This package requires that the package ”auis-wp” be installed on your system. The X Window System, a running mail system and the ”csh” shell are required as well. Please read also the README files under /usr/andrew.

Copyright:  (c) 1992 Carnegie Mellon University & IBM
Author:  Andrew Consortium <info-andrew-request@andrew.cmu.edu>
Version:  635L1-0

auis src Andrew User Interface System - programming environment

AUIS is a collection of tools which enable you to design, view and mail multimedia documents. This package is designed for application programmers who want to design their own applications using AUIS. This is **NOT** the source code of AUIS! Please read also the README files under /usr/andrew.

Copyright:  (c) 1992 Carnegie Mellon University & IBM
Author:  Andrew Consortium <info-andrew-request@andrew.cmu.edu>
Version:  635L1-0

auis wp Andrew User Interface System - word processing

AUIS is a collection of tools which enable you to design, view and mail multimedia documents. This package contains the main word processor on which all the other packages of AUIS are based. Please read also the README files under /usr/andrew.

Copyright:  (c) 1992 Carnegie Mellon University & IBM
Author:  Andrew Consortium <info-andrew-request@andrew.cmu.edu>
Version:  635L1-0

axe A simple to use X11 editor

This simple to use editor features: multiple windows, multiple buffers, configurable menus and buttons, access to external filters, keyboard macros, comprehensive on-line help, ... This version of aXe-6.1.2 is usable only with the real shared Xaw library. For Xaw3d please use aXe from the axe3d package.

Contents: aXe3D-6.1.2

Copyright:  (c) 1992, 1994 The University of Newcastle upon Tyne
Author:  Jim Wight <j.k.wight@newcastle.ac.uk>
Version:  6.1.2-13

axe3d A simple to use X11 editor

This simple to use editor features: multiple windows, multiple buffers, configurable menus and buttons, access to external filters, keyboard macros, comprehensive on-line help, ... This version of aXe-6.1.2 is usable only with the real shared Xaw library. For Xaw3d please use aXe from the axe3d package.

Contents: aXe3D-6.1.2

Copyright:  (c) 1992, 1994 The University of Newcastle upon Tyne
Author:  Jim Wight <j.k.wight@newcastle.ac.uk>
Version:  6.1.2-14

babaya Tool to exit X11 gently

A tool to exit X11 gently. Have you always been annoyed by the error messages (... broken pipe) X11 produces when exiting ? Click the right mouse button in the babaya window for a ”clean” shutdown of X11. Documentation: /usr/doc/babaya

Contents: babaya-1.2b

Copyright:  (c) 1992-1993 - Muhammad M. Saggaf
Author:  Muhammad M. Saggaf <alsaggaf@mit.edu>
Version:  1.2b-12

bitchx BitchX an IRC client.

BitchX is one of the well-known IRC clients. Very well designed and offers lots of possibilities and settings.

Copyright:  (c) Colten Edwards
Author:  Colten Edwards <edwards@sk.sympatico.ca>
Version:  75p1-3

calctool A calculator

A simple scientific calculator, which can also represent the values also in binary, hexadecimal or decimal format.

Contents: calctool-2.4.12

Copyright:  (c) Rich Burridge
Author:  Rich Burridge <richb@sunaus.oz.au>
Version:  2.4.12-13

cellular Cellular automate simulation program

automates named cellang and a compiler that makes programms for the X-Window-System from it. You can find the Documentation for cellang and a tutorial in /usr/doc/packages/cellular. Examples are in /usr/lib/cellular/examples. Documentation: man cellc, avcam, tpipe

Contents: cellular-07apr94

Copyright:  (c) 1992, 1993, 1994 J Dana Eckart
Author:  J Dana Eckart <dana@rucs.faculty.cs.runet.edu>
Version:  94.4.7-10

conv Conversion utility (units of measurement)

Copyright:  (c) GPL
Author:  Stefan Heimers <sheimers@ezinfo.vmsmail.ethz.ch>
Version:  1997.08.11-14

coolmail Cool XBiff clone

Shows if new mail has arrived. The main advantage of this program is that you can click into the window to start the mailer. Documentation: man coolmail

Contents: coolmail-1.3

Copyright:  (c) 1994 Byron C. Darrah
Author:  Byron C. Darrah <darrah@kaiwan.com>, Randall K. Sharpe <rsharpe@ncsa.uiuc.edu>
Version:  1.3-20

elem Table of chemical elements

Copyright:  (c) GPL
Author:  Stefan Heimers <sheimers@ezinfo.vmsmail.ethz.ch>
Version:  1998.01.20-8

fader Screen brightness adjuster

Adjusts the screen brightness. There is a version for X11 and one for the console.

Contents: fader-1.01

Copyright:  (c) 1994 Klaus Thiele
Author:  Klaus Thiele <100520.634@compuserve.com>
Version:  1.01-13

fct FCT

Firweall Cobfiguration Toolkit Please note, that there is no support for this package by S.u.S.E. Use at your own risc!

Copyright:  (c) GPL
Author:  Jens Hellmerichs-Friedrich <ft114@fen.baynet.de>
Version:  1.0.9.p1-2

felix Computer algebra system FELIX

More information about this package can be found in /usr/doc/felix-2.1. Unfortunately, the sources are not available. Contents: felix-2.1

Copyright:  (c) 1990-93 Joachim Apel & Uwe Klaus

Author:

Joachim Apel <japel@informatik.uni-leipzig.de>, Uwe Klaus <uklaus@informatik.unileipzig.de>

Version:  2.1-13

gnuplot gnuplot

Gnuplot is a command-line driven interactive function plotting utility. Gnuplot supports many different types of terminals, plotters, and printers (including many color devices, and pseudo-devices like LaTeX) and can easily be extended to include new devices. Contents: gnuplot-beta336

Copyright:  (c) 1986 - 1993 Thomas Williams, Colin Kelley

Author:

Mailingliste  <bug-gnuplot@dartmouth.edu>,  Mailingliste  <info-

gnuplot@dartmouth.edu>, Mailingliste <info-gnuplot-beta@dartmouth.edu>, Mailingliste <info-gnuplot-request@dartmouth.edu>

Version:  3.5beta6.347-5

gvim vi with GUI

README.SuSE for package ’gvim’. This package includes only a new binary with GUI (3D-Athena-Widgets) and a resource file (/etc/gvimrc). The normal package ’vim’ is needed! Start: /usr/X11R6/bin/gvim Copy and modify /etc/gvimrc to ˜ /.gvimrc if needed.

Author:  Bram Moolenaar <mool@oce.nl>
Version:  4.5-15

hfsutils Tools used for Hierarchical File System used on Mac

HFS is the ”Hierarchical File System” used on modern Macintosh computers. With this package, you can read and write Macintosh-formatted media such as floppy disks, CD-ROMs, and SCSI hard disks on most UNIX platforms. You can also format raw media into an HFS volume.

Copyright:  (c) 1996-98 Robert Leslie
Author:  Robert Leslie <rob@mars.org>
Version:  3.2-7

knews Threaded news reader under X11

Knews is a threaded news reader with an X Window interface that uses NNTP to get news. It displays the threads in a graphical tree. For more information read the man page (”man knews”). Contents: knews-0.9.8

Copyright:  (c) Copyright 1995 by Karl-Johan Johnsson
Author:  Karl-Johan Johnsson <mathematik.su.se>
Version:  0.9.8-13

kradio Video4Linux frontend for radio on bttv cards.

Video4Linux frontend for radio on bttv cards, i.e. WinTV/radio.

Copyright:  (c) (C) 1997,98 Gerd Knorr <kraxel@cs.tu-berlin.de>
Author:  Gerd Knorr <kraxel@cs.tu-berlin.de>
Version:  2.14-5

maxwell maxwell

Maxwell is a fully featured Word Processor application for Linux.

Contents: maxwell-0.5.2

Copyright:  (c) Andrew Haisley, Dave Miller, Tom Newton

Author:

Andrew Haisley <maxwell@eeyore-mule.demon.co.uk>, Dave Miller <maxwell@eeyore-mule.demon.co.uk>, Tom Newton <maxwell@eeyoremule.demon.co.uk>

Version:  0.5.2-3

mfm a file manager for X11

Copyright:  (c) 1995, Andrew T. Renalds
Author:  Andrew T. Renalds <renalds@abn.com>
Version:  1.3-18

mfsm mfsm

X-Windows Motif utility that monitors free space and user quotas of filesystems.

Copyright:  (c) GPL Shane D. Hill
Author:  Shane D. Hill <Shane.Hill@dsto.defence.gov.au>
Version:  1.3-0

mgdiff X-diff

Mgdiff is a graphical front end to the Unix diff command based upon X11 and the Motif widget set. It allows the user to select two files for comparison, runs the diff command, parses the output and presents the results graphically.

Contents: mgdiff-1.0

Copyright:  (c) Copyright (c) 1994 Daniel Williams
Author:  Daniel Williams <dan@sass.com>
Version:  1.0-19

mupad MuPAD - Multi Processing Algebra Data Tool

MuPad is a program system for doing computer algebra. It is being developed at the University of Paderborn, Germany. Computer algebra systems providei symbolic handling of mathematical formulae and, supplied with expert knowledge, make it easier to handle. They allow interactive manipulation of mathematical objects, effortless functional programming of particular mathematical problems. Their immense effect on mathematical lessons at universities and high schools has already been noted. Even schools are starting to use it as part of their lessons. MuPAD is one of the first universal computer algebra systems. It is already well known for its object oriented language as well as for its steady growing implementations of mathematical algorithms. An extraordinary comfortable GUI, an interactive on-line help system, the included debugger and a powerful graphics tool are the main advantages of MuPAD. In 1993 MuPAD has been awarded with the ”German-Austrian software price” followed by the First European Software Award in 1994. Before you start using MuPAD you should read the README file which is located under /usr/MuPAD. This file explains how to accustom your environment to fully use MuPAD.

Contents: mupad-1.3.1

Copyright:  (c) B. Fuchssteiner, Automath, University of Paderborn, Germany

Author:

Prof. Dr. Benno Fuchssteiner <benno@uni-paderborn.de> , Ralf Hillebrand <tonner@uni-paderborn.de>

Version:  1.3.1-5

nedit Full featured plain text ”Nirvana Editor”

NEdit is a GUI style plain text editor for workstations with X Windows and Motif. NEdit provides all of the standard menu, dialog, editing, mouse support, as well as macro extension language, Syntax highlighting, and a lot other nice features (and extensions for programmers).

Contents: nedit-5.0.2

Copyright:  (c) (c) 1991 Universities Research Association, Inc.
Author:  Mark Edel <edel@fnal.gov>
Version:  5.0.2-2

offix some programs with drag and drop features

This package contains a clipboard, editor, executor, file manager and a printer frontend with drag and drop features.

Contents: offix-2.1.orig

Copyright:  (c) 1996 Csar Crusius, Randolf Werner, Andr Hentz
Author:  Csar Crusius <c.crusius@ieee.org>, Randolf Werner , Andr Hentz <a.hentz@ieee.org>
Version:  2.1.2debian-15

pan note manager

Post a Note. A quite useful note manager.

Contents: pan-2.04

Copyright:  (c) 1992 Jeffrey W. Bailey
Author:  Jeffrey W. Bailey <jeff@rd1.interlan.com>
Version:  2.04-15

papyrus Word processor

A nice word processor. Documentation can be found: /usr/doc/packages/papyrus

Copyright:  (c) 1996-1997 Carlos Nunes
Author:  Carlos Nunes <loscar@mime.univ-paris8.fr>
Version:  1.1.1-2

pcb Interactive printed circuit board design

Pcb is a handy tool for the X Window System built to design printed circuit boards. All coordinate units are 1/1000 inch. For details see the manual which can be found under /usr/doc/packages/pcb. There is also a man page available.

Contents: pcb-1.5.3.ALPHA

Copyright:  (c) 1994-1996 Thomas Nau
Author:  Thomas Nau <Thomas.Nau@rz.uni-ulm.de>
Version:  1.5.3.ALPHA-11

pixedit Pixel Color Editor

A tool to modify colors using several scrollbars. The program to find the colors you ”really” like :-)

Contents: pixedit-1994

Copyright:  (c) 1994 Mike Yang
Author:  Mike Yang <mikey@sgi.com>
Version:  1994-14

plan Clear and easy to use schedule planner

This is a really useful program. It is statically linked with the Motif library. If you have Motif installed on your system, you can recompile the program and save a lot of main memory.

Contents: plan-1.7.1-src-stat

Copyright:  (c) 1993-97 Thomas Driemeyer
Author:  Thomas Driemeyer <thomas@bitrot.in-berlin.de>
Version:  1.7.1-6

plotmtv A multipurpose X11 plotting program

PLOTMTV’s capabilities include 2D line and scatter plots (x-vs-y), contour plots, 3D surface, line and scatter plots as well as vector plots. The program has an rough but functional Graphical User Interface, through which it is possible to zoom in, zoom out, pan, toggle between 2D and 3D plots, and rotate 3D plots. Both color and gray-scale PostScript output are supported.

Contents: Plotmtv1.4.1

Copyright:  (c) 1995 Kenny K.H. Toh
Author:  Kenny K.H. Toh <ktoh@td2cad.intel.com>
Version:  1.4.1-12

procmtr Display actual system parameters

With procmeter one can display various system parameters as e.g. processor load, network load, etc. Documentation may be found under /usr/doc/packages/procmtr Contents: procmeter-2.2

Copyright:  (c) 1994,95,96 Andrew M. Bishop
Author:  Andrew M. Bishop <amb@gedanken.demon.co.uk>
Version:  2.2-13

qps visual process status monitor

qps is an X11 version of top or ps, with motif- or windows-like interface. Contents: qps-1.2

Author:  Mattias Engdegrd <f91-men@nada.kth.se>
Version:  1.2-17

rasmol Molecular Graphics Visualization tool

RasMol is an X Window System tool intended for the visualization of proteins and nucleic acids. It reads Brookhaven Protein Database (PDB) files and interactively renders them in a variety of formats on either an 8bit or 24/32bit color display. Documentation is in /usr/doc/rasmol & man rasmol Examples are in /usr/lib/rasmol Contents: RasMol-2.5.1

Copyright:  (c) 1992-1994 by Roger Sayle
Author:  Roger Sayle <ras32425@ggr.co.uk>
Version:  2.5.1-12

roottail print text directly to X11 root window

Displays a given file anywhere on your X11 root window with a transparent background.

Copyright:  (c) 1998 Mike Baker
Author:  Mike Baker <mjbaker@mtu.edu>
Version:  0.0.4-2

rxvt The ’other’ X-Terminal.

rxvt is smaller as xterm. That’s why many people like it. This package also includes rclock - it is now included in the rxvt source package. Contents: rxvt-2.4.5

Author:

Mark  Olesen  <olesen@me.QueensU.CA>,  Robert  Nation

<nation@rocket.sanders.lockheed.com>

Version:  2.4.5-7

scilab Scilab - large mathematical package

A large mathematical package containing a lot of demos. Please take a look at the license information in /usr/lib/scilab/notice.* . This program may not have a very appealing look, but it is as powerful as some commercial products. Scilab even includes a converter for Maple files. Contents: scilab-2.3.1-src

Copyright:  (c) 1994-97 INRIA, France
Author:  <Scilab@inria.fr>
Version:  2.3.1-13

snns The ”Stuttgart Neural Network Simulator”

Nice simulator of neural networks which has been developed at the university of Stuttgart. All parameters are presented in individual windows. To really use this program a fundamental background of neural networks is required. SNNS (Stuttgart Neural Network Simulator) is a software simulator for neural networks on Unix workstations developed on the university of Stuttgart. The aim of the project is the development of an flexible simulation interface for research and usage of neural networks. SNNS consists mainly of two parts: o simulator kernel written in C o GUI under X11 The simulator kernel works with the internal network-data structures of the neural network and performs all operations by learning or doing. It can also be used separately as a C program embedded in own applications. It supports multiple network topologies and almost like RCS supports the sites concept. SNNS can be extended by the individual user with user-specific activation functions, output functions, site-functions and learning procedures which can easily be connected as simple C programs with the simulation kernel. At the moment of this writing the following algorithms are supported: o back propagation o vanilla (online) BP o BP with Momentum Term and Flat Spot Elimination o Batch BP o Time Delay BP o Counter propagation o Quick prop o Back percolation 1 o Generalized radial-base-functions (RBF) o RProp o ART 1 o ART 2 o ARTMAP o Cascade Correlation o Return Cascade Correlation o Dynamic LVQ o Back propagation through time (for recursive networks) o Batch Back propagation through time (for recursive networks) o Quick propagation through time ” ” ” o Kohonen Feature Maps The GUI XGUI based on the kernel presents you 1/2/3-dimensional views of the neural network and controls the kernel during the runtime of the simulation. Additionally the 2D user-interface provides an integrated network-editor, which can be used to design, manipulate and present neural networks. A complete user manual can be found in the package snnsdoc. The current version of SNNS is to be obtained from ftp.informatik.uni-stuttgart.de /pub/SNNS Contents: SNNSinfo-1.03 SNNSv4.1.Manual.src SNNSv4.1

Copyright:

(c) 1990-96 SNNS-Gruppe, Institut f”ur Parallel und Verteilte H”ochstleistungsrechner (IPVR), Universit”at Stuttgart, Breitwiesenstr. 20-22, D-70565 Stuttgart

Author:

Andreas Zell <zell@informatik.uni-stuttgart.de>, G”unter Mamier, Michael Vogt, Niels Mache, Ralf H”ubner, Kai-Uwe Herrmann, Tobias Soyez, Michael Schmalzl, Tilman Sommer, Artemis Hatzigeorgiou, Sven D”oring, Dietmar Posselt, Tobias Schreiner

Version:  4.1-19

snnsdoc User Manual for SNNS

SNNS (Stuttgart Neural Network Simulator) User Manual, Version 3.3, Universitaet Stuttgart, Fakultaet Informatik, Report No. 3/94 (revised).

Copyright:

(c) 1990-95 SNNS-Gruppe, Institut f”ur Parallel und Verteilte H”ochstleistungsrechner (IPVR), Universit”at Stuttgart, Breitwiesenstr. 20-22, D-70565 Stuttgart

Author:

Andreas Zell <zell@informatik.uni-stuttgart.de>, G”unter Mamier , Michael Vogt , Niels Mache , Ralf H”ubner , Kai-Uwe Herrmann , Tobias Soyez , Michael Schmalzl , Tilman Sommer , Artemis Hatzigeorgiou , Sven D”oring , Dietmar Posselt , Tobias Schreiner

Version:  4.1-19

spice An analog circuit simulator

Spice is an analog circuit simulator Pieces of advice and examples at /usr/doc/packages/spice Contents: spice3f4

Copyright:  (c) 1990 Regents of the University of California
Version:  3f4-15

spicecad SPICECAD is a schematic entry for SPICE

SPICECAD is a graphical frontend to Spice. Spice does NOT have to be installed since it is statically linked to the SpiceCAD binary. Please read also /usr/doc/packages/spicecad/README and README.SuSE!!! Contents: spicecad-1.5

Copyright:  (c) Martin Maschmann

Author:

Martin  Maschmann  <martin.maschmann@t-online.de>,  Martin
Maschmann  <martin.maschmann@vlsi.com>,  Martin  Maschmann

<martin.maschmann@bigfoot.com>

Version:  1.5-12

susecurs S.u.S.E. logo as cursor in X11

With this package a script do chamaeleon is supplied which transform the boring X-cursor into the S.u.S.E. chamaeleon. Contents: do chameleon0.3

Author:  Werner Fink <werner@suse.de>
Version:  0.3-1

tn5250 5250 Emulator

The 5250 is most commonly used for connected to IBM’s AS/400. While one can connect to an AS/400 with a VT100 emulator, its not ideal. The problem is that the 5250 is a screen at a time terminal, whereas the VT100 is a character at a time device. The emulator uses the binary mode of telnet to transfer the 5250 data stream.

Copyright:  (c) Michael Madore
Author:  Michael Madore <mmadore@blarg.net>
Version:  0.12.5-2

tosh ut Toshiba Linux Utilities

Toshiba Linux Utilities:

1. fan: a tool to turn the internal fan on/off.

2. hotkey: to give grafical feedback (under X) to the user while using the ”Fn” key to switch:

* Battery Save mode

* Resume/Boot mode

3. svpw: an utility to change the user and the supervisor password. Here you’ll find further information: http://www2.prestel.co.uk/hex/toshiba.html Please, read the documentation under /usr/doc/packages/tosh ut If you don’t use these tools with care, it’s possible to lock the system completely!

Copyright:  (c) GPL
Author:  Jonathan Buzzard <jab@hex.prestel.co.uk>
Version:  1.0-0

vidtext Viewer for Videotext

A program that allows you to view videotext pages under X11. The program is accompanied with a kernel loadable module that enables access to the hardware of the videotext decoder. Contents: VideoteXt-0.6 CAUTION: This precompiled module supports videotext cards (kits) as described in c’t 11/91. So, if you have another card installed you need to install the sources and to generate the module vtx.o. Information may be found under /usr/doc/packages/vidtext

Copyright:  (c) 1994-1996 Martin Buck
Author:  Martin Buck <martin-2.buck@student.uni-ulm.de>
Version:  0.6-21

vnc A virtual X-Server

With vnc you can start a virtual X-Server and get access to it from different machines with the vncviewer. You can change the machine without terminating your session. Viewer and server are also available for Windows - they allow access to vnc under windows. You can find it under /dosutils/vnc on the first CD. Contents: vnc

Copyright:  (c) GPL
Author:  <vnc@orl.co.uk>
Version:  3.3.2-1

x3270 IBM host access tool.

x3270 opens a telnet connection to an IBM host in an X window. Contents: x3270-3.1.1.6

Author:  Paul Mattes <pdm@cnt.com>
Version:  3.1.1.6-14

xarchie X11 tool to search archie servers

This program allows you to search the various archie servers automatically for a given file. You need an internet connection, though. Contents: xarchie-2.0.10

Copyright:  (c) 1989-91 University of Washington

Author:

Alan Emtage <bajan@cs.mcgill.com>, Bill Heelan <wheelan@cs.mcgill.ca>, Brendan Kehoe <brendan@cygnus.com>, George Ferguson <ferguson@cs.rochester.edu>, Peter Deutsch <peterd@bunyip.com>

Version:  2.0.10-12

xautolck lock the screen after a certain time automagically

This program uses another locker program to lock the screen automagically after an idle period. It may e.g. be called as follows: xautolock -locker ”xlock -mode blank” using xlock as locker program. Contents: xautolock

Copyright:  (c) Stefan De Troch, Michel Eyckmans

Author:

Stefan De Troch <stefan@specs.philips.com>, Michel Eyckmans <eyckmans@imec.be>

Version:  pl10-12

xawtv Video4Linux TV Program using the bttv driver

xawtv is a ”simple” program for watching TV with your linux box. It uses the Athena widgets. xawtv supports the bttv driver and uses the video4linux interface. Its supports also radio on the card. Contents: xawtv v4l-conf set-tv radio kradio

Copyright:  (c) (C) 1997,98 Gerd Knorr <kraxel@cs.tu-berlin.de>
Author:  Gerd Knorr <kraxel@cs.tu-berlin.de>
Version:  2.14-5

xbanner X-Background writings and images

How about your company’s name in the background of the login screen? Or your company’s logo? It will surely work with xbanner which is designed for the use with e.g. XDM login screens. Many example configurations may be found in /usr/X11R6/lib/X11/xbanner and its subdirectories. Documentation lies in /usr/doc/packages/xbanner, in the subdir html you find a HTML version of the documents. Contents: xbanner-1.2

Copyright:  (c) 1996 Amit Margalit
Author:  Amit Margalit <amitm@netvision.net.il>
Version:  1.2-12

xbarcode program to create barcodes

Barcodes of the following types can be created, viewed, and saved in PPM, GIF, PCX, PBM, LaTeX or PostScript format: EAN13, EAN8, UPC, CodeABar, Code 128, Code 3/9, Code 2/5, Code 2/5 Matrix, Code 2/5 Interleaved. By using xbarcode, barcodes can be inserted into documents without any problems (and without doing real ”cut and paste”). A version which runs without X11 is also included (bar).

Contents: xbarcode-2.11

Copyright:  (c) 1994 Martin Bauer, Jan Laitenberger
Author:  Martin Bauer <Martin Bauer@S2.MAUS.DE>, Jan Laitenberger
Version:  2.11-13

xbench benchmark for X11

A benchmark for X11.

Contents: xbench-0.2

Copyright:  (c) 1989 Siemens
Author:  Claus Gittinger <unido!sinix!claus@uunet.uu.net>
Version:  0.2-10

xcb X11 cut&paste utility

Xcb provides access to the cut buffers built into every X server. It allows the buffers to be manipulated either via the command line, or with the mouse in a point and click manner. The buffers can be used as holding pens to store and retrieve arbitrary data fragments, so any number of different pieces of data can be saved and recalled later. The program is designed primarily for use with textual data.

Contents: xcb-2.3

Copyright:  (c) 1992,1993,1994 by Farrell McKay
Author:  Farrell McKay <Farrell.McKay@mpx.com.au>
Version:  2.3-12

xcnf Make settings in X

With xcnf you may take some settings under X: - mouse - system beep - builtin X screen saver The settings may be stored in ˜ /.Xcnf and can be recalled every time starting X by calling xcnfset e.g. in ˜ /.xinitrc Documentation: xcnf --help xcnfset --help /usr/doc/packages/xcnf/README

Contents: xcnf-0.1

Copyright:  (c) 1997, Martin Scherbaum, S.u.S.E. GmbH
Author:  Martin Scherbaum <maddin@suse.de>
Version:  0.1-13

xcolors displays colors defined in rgb.txt

Xcolorsel displays colors defined in rgb.txt. You can create an RGB file by redirecting the output of showrgb to a file.

Contents: xcolors-04oct91

Copyright:  (c) 1989,1991 - Stephen Gildea
Author:  Stephen Gildea <gildea@intouchsys.com>
Version:  91.10.4-10

xcoral A multiwindow mouse-based text editor for the X Window System

Xcoral is a multiwindow mouse-based text editor for the X Window System. A built-in browser enables you to navigate through C functions, C++ classes, methods and files. A SMall Ansi C Interpreter (Smac) is also built-in to dynamically extend the editor’s possibilities (user functions, key bindings, modes etc). Xcoral provides variable width fonts, menus, scrollbars, buttons, search, regions, killbuffers, macros and undo. An on-line manual box, with a table of contents and an index, helps you to use and customize the editor. Commands are accessible from menus or key bindings. Xcoral is a direct Xlib client and runs on color/bw X Display.

Contents: xcoral-3.1

Copyright:  (c) 1989-94 Lionel Fournigault, Bruno Pages and Dominique Leveque.
Author:  Lionel Fournigault <lf@lri.lri.fr>
Version:  3.1-12

xdf disk usage under X11

a graphical df (Disk free).

Contents: xdf-1.3

Copyright:  (c) 1996 Georg Greve
Author:  Georg Greve <development@fusebox.hanse.de>
Version:  1.3-7

xdtm The X Desktop Manager

The X Desktop Manager is a graphical shell for the X Window System, it provides mechanisms for all the usual shell activities and is intended to reduce the training necessary for UNIX novice to become productive on your system.

Contents: xdtm-2.5.8

Copyright:  (c) 1991 Edward Groenendaal

Author:

Edward Groenendaal, Ramon J. Santiago <rjs@pencom.com> , Lionel Mallet <Lionel.Mallet@sophia.inria.fr>

Version:  2.5.8-13

xephem Space tool

A program which displays all sorts of information about our planetary system. Lots of nice, colourful buttons.

Contents: xephem-3.0

Copyright:  (c) 1990-93 Elwood Charles Downey
Author:  Elwood Charles Downey <ecdowney@noao.edu>
Version:  3.0-12

xfileman xfilemanager

Xfilemanager is an X11 based Unix file manager with full color displays.

Contents: xfilemanager-0.4.1

Author:  Ove Kalkan <kalkan@ramz.ing.tu-bs.de>
Version:  0.4.1-13

xflame GUI for cdrecord und mkisofs

A little tool to create CD-Roms

Copyright:  (c) GPL
Author:  Dr. Frank Rieg <dr.frank.rieg@t-online.de>
Version:  1.0-6

xfm xfm, a file manager for X.

This is a powerful file and applications manager program written using the X toolkit. It contains virtually all of the features that you would expect in a file manager - move around your directory tree in multiple windows, and move, copy or delete files with simple mouse operations. The integrated application manager allows you to load files into your favorite applications from the file manager by dragging icons. You should invoke xfm.install before launching xfm for the first time.

Contents: xfm-1.3.2

Copyright:  (c) 1990 - 1995 Simon Marlow, Albert Graef

Author:

Albert Graef <ag@MuwiInfb.Geschichte.Uni-Mainz.DE>, Simon Marlow <simonm@dcs.gla.ac.uk>

Version:  1.3.2-13

xfmail XForms based Email reader system

XFMail is an X-Windows application for sending and receiving electronic mail. Features include: MIME and PGP support, address books with mailing list functions, message filtering rules, and multilingual support. It retrieves mail both from POP servers and from a local mail spool. Mail can be sent using sendmail, an SMTP gateway, or a POP server. It also supports IMAP4 remote folders.

Contents: xfmail-1.2

Copyright:  (c) 1995, 1996, 1997 Gennady B. Sorokopud, 1995 Ugen J. S. Antsilevich
Author:  Gennady B. Sorokopud <gena@NetVision.net.il>
Version:  1.3-8

xformat xformat

Little X frontend to fdformat, mke2fs and mformat using the Xforms library.

Copyright:  (c) GPL
Author:  Christoph Neerfeld <Christoph.Neerfeld@mail.bonn.netsurf.de>
Version:  1.0.1-8

xftp X11 frontend to ’ftp’

Makes ftp much more easy! Contents: llnlxftp-2.0.4

Copyright:  (c) 1993-1994 University of California
Author:  Neale G. Smith <llnlxftp@llnl.gov>
Version:  2.0.4-18

xgopher A gopher client for the X Window System

A gopher client for the X Window System. This program offers you a comfortable GUI for accessing gopher clients. You must be connected to the internet, though. Documentation can be found under /usr/doc/X11-misc/xgopher-1.3 & man xgopher Contents: xgopher.1.3.3

Copyright:  (c) 1992, 1993 University of Illinois
Author:  Allan Tuchman <a-tuchman@uiuc.edu>
Version:  1.3.3-12

xgrok Graphical Resource Organizer Kit

Grok is a simple database manager and UI builder that can -

* keep phone and address lists (like a rolodex)

* store phone call logs

* keep todo lists

* manage any other database after simple GUI-driven customization More precisely, grok is a program for displaying and editing strings arranged in a grid of rows and columns. Each row is presented as a ”card” consisting of multiple columns, or ”fields”, that allow data entry. The presentation of the data is programmable; a user interface builder that allows the user to arrange fields on a card graphically is part of grok. Grok also supports a simple language that allows sophisticated queries and data retrieval. Grok comes with the above examples and a few others as pre-built applications. Contents: xmbase-grok-1.4.1

Copyright:  (c) 1995 Thomas Driemeyer
Author:  Thomas Driemeyer <thomas@bitrot.in-berlin.de>
Version:  1.4-20

xgroups file manager - similar to the Windows program manager

This file or desktop manager is similar to the Windows program manager. Its configuration is somehow unconventional, but the program is useful nevertheless. The one or other will sureley like this program. Contents: xgroups-1.3

Copyright:  (c) 1993 Giuseppe Zanetti
Author:  Giuseppe Zanetti <beppe@maya.dei.unipd.it>
Version:  1.3-12

xhtalk a talk program under X11

You should copy a sample configuration file as .xhtalk into your home directory before using xhtalk the first time. You find one in /usr/doc/packages/xhtalk/SAMPLE.xhtalk Contents: xhtalk-2.9p1-linux

Copyright:  (c) 1992 Klaus Hartenstein
Author:  Klaus Hartenstein <kdh@ira.uka.de>, Angelo Haritsis <ah@doc.ic.ac.uk>
Version:  2.9p1-13

xinfo X-Browser for GNU Info Pages

Makes reading of info files a lot easier. Contents: xinfo-1.01.01

Copyright:  (c) 1989, 1990 Jordan K. Hubbard
Author:  Jordan K. Hubbard <jkh@meepmeep.pcs.com>
Version:  1.01.01-12

xirc xirc -a chat programm under X11

An X-based irc-client

Copyright:  (c) 1996-1997 Joe Croft
Author:  Joe Croft <jcroft@unicomp.net>
Version:  2.0-17

xkeycaps A frontend to xmodmap

xkeycaps is a graphical frontend for xmodmap. A virtual keyboard is shown and you can redefine any key on runtime change your ˜ /.Xmodmap or use other applications (provided that you use Click-to-Focus). For layout you should choose ”PC 102 key keyboard #1 - Linux/XFree86 German”. Contents: xkeycaps-2.29

Copyright:  (c) 1991, 1992, 1993, 1994, 1995 Jamie Zawinski
Author:  Jamie Zawinski <jwz@netscape.com>
Version:  2.29-13

xled xled - a LED for vbox the ISDN answering machine for vbox.

xled is your interface to your answering machine. Clicking onto the icon launches your personal vbox. Contents: xled-1.3.1

Copyright:  (c) Jan Schoenepauck und Joachim Gassen
Author:  Jan Schoenepauck <schoenep@uni-wuppertal.de>
Version:  1.3.1-15

xless text browser

xless 1.7 - the well-known text browser. Necessary for displaying the various HOWTOs. Contents: xless-1.7

Copyright:  (c) 1994 Dave Glowacki
Author:  Dave Glowacki <dglo@cs.berkeley.edu>
Version:  1.7-14

xlock A screen saver

Patrick J. Naughton’s xlock screensaver/locker for X, with many more modes added by David A. Bagley. For security reasons we now have xlock sgid shadow instead of suid root. If you use a yp-server which only gives passwords to programs running as user ”root”, you have to change xlock back to suid-root. Contents: xlockmore-4.04

Copyright:  (c) 1988-1995 Patrick J. Naughton, David A. Bagley

Author:

David Albert Bagley <bagleyd@bigfoot.com>, Patrick J. Naughton <naughton@eng.sun.com>

Version:  4.04-12

xlogin xlogin, xtelnet

This package contains two scripts, which called with the name of a remote host, will open an xterm window on the local X display with a remote session or login shell. The sript xlogin starts by a rsh call (remote shell) an xterm on the remote host. If necessary the access will be allowd by sending the magic key (cookie) of the display to the remote host. The script xtelnet starts a local xterm with a telnet session on the remote host. Contents: xlogin-0.1

Copyright:  (c) 1997 S.u.S.E. GmbH
Author:  Werner Fink <werner@suse.de>
Version:  0.1-12

xlogmas watch various logfiles

Displays logfiles under X, configurable via a popup menu to select which files to watch and to choose tail/cat mode.

Copyright:  (c) 1998 Georg Greve
Author:  Georg Greve <greve@fusebox.hanse.de>
Version:  1.0b-1

xlpq frontend to manage printer queues

xforms based frontend to manage printer queues. I invokes lpq, lprm and lpc. It shows what is currently in the printer queue, lets you select different queues and lets you delete printjobs. You may even disable queueing and printing if you are root.

Copyright:  (c) GPL
Author:  Dimitrios P. Bouras <dbouras@hol.gr>
Version:  1.0.2-1

xmail X Window System interface to the mail program

xmail is an X-based interface to the Berkeley mail program. This version of xmail includes support for decompressing and displaying the content of an X-Face mail header, if such a header exists in the message being read. This feature is a compile time option, and requires the existence of the compface library routines, which are NOT supplied with the xmail sources. X-Face headers are compressed bitmap images, typically of the face of the person owning such a header. The compressed header contains only printable characters, which allows it to be included in a mail message. The USENIX FaceSaver project is typically the source for these bitmaps, which are compressed into X-Face header format using the compface software tools written and distributed by James Ashton <jaa@cs.su.oz.au>. The compface library is also provided with the distri- bution of the faces software, written by Rich Burridge, and is available by email from his archive server <rbarchive-server@Aus.Sun.COM>. Contents: xmail-1.6

Copyright:  (c) 1990-94 by National Semiconductor Corporation
Author:  Michael C. Wagnitz <michael@berlioz.nsc.com>
Version:  1.6-12

xmftp xmftp

X/Motif FTP client, include static and dynamic version Contents: xmftp-1.0.4

Copyright:  (c) GPL
Author:  Viraj Alankar <valankar@bigfoot.com>
Version:  1.0.4-5

xnetload This program displays network traffic and uptime for the network

This program displays network traffic and uptime for the network connection named on the command line. Contents: xnetload-1.2

Author:  Roland Smith <rsmit06@ibm.net>
Version:  1.5.1-1

xosview info on the system load

A small program - similar to xsysinfo which is mostly configurable using resources via ˜ /.Xresources. Shows actual CPU, swap, memory, active interrupts and -if wanted- netpacket statistics. If you want to get informations about ’network traffic’ you still have to call ipfwadm (included in the same package) to activate this as root. This might look like: /sbin/ipfwadm -A -a -P all -S <own IP-address> -D 0/0 /sbin/ipfwadm -A -a -P all -S 0/0 -D <own IP-address> Please have a look at /usr/doc/packages/xosview/README.linux. For showing shared pages, the module memstat.o have to loaded with insmod. The sources of the module is located in /usr/doc/packages/xosview/memstat/. The installation is only after an upgrade of the kernel necessary. Contents: xosview-1.6.0 If you use the ipfwadm package from our distribution the ipfwadm will be invoked at boot time.

Copyright:  (c) 1994-98 Mike Romberg
Author:  Mike Romberg <romberg@md.fsl.noaa.gov>
Version:  1.6.0-8

xplot An interactive 3D plotting program

Xplot is an interactive 3D plotting program for X windows. It is capable of displaying simple objects that have easily described geometries. Currently, it only runs on 8bit pseudo color visuals. The way of using xplot is as follows. First one defines the object by specifying its geometry. Then plot it. Once an object is displayed on the screen, its appearance can be modified interactively by mouse input. Docu: /usr/doc/X11-misc/xplot-beta-09feb95 & man xplot

Contents: xplot-beta-09feb95

Copyright:  (c) 1986-89 Zou Maorong, Thomas Williams, Colin Kelley
Author:  Zou Maorong, Thomas Williams, Colin Kelley
Version:  95.2.9-11

xplz German zip codes

This program displays the zip codes of German towns. For more information see /usr/doc/X11misc/xplz-3.4.README .

Contents: xplz-3.4 xplz-data

Copyright:  (c) 1993 Rainer Klute
Author:  Rainer Klute <klute@irb.informatik.uni-dortmund.de>
Version:  3.4-13

xprompt small tool for prompting users

Small tool to ask the user for one or more responses (e.g. from batch files). Documentation: /usr/doc/X11-misc/xprompt-28sep91.README

Contents: xprompt-28sep91

Copyright:  (c) 1989-91 - Barry Brachman & University of British Columbia
Author:  Barry Brachman <brachman@cs.ubc.ca>
Version:  91.9.28-9

xrolodex rolodex-like application

A ”virtual business card” manager. For more information see the xrolodex manpage and /usr/doc/packages/xrolodex. This is a demo version of a commercial product.

Copyright:  (c) 1991, 1992 Iris Computing Laboratories
Author:  Jerry Smith <jsmith@spectro.com>
Version:  1.5-12

xrpm X frontend for RPM

Author:  George A. Farris <george@gmsys.com>
Version:  2.2-4

xscrns a screensaver applikation

The xscreensaver program waits until the keyboard and mouse have been idle for a period, and then runs a graphics demo chosen at random. It turns off as soon as there is any mouse or keyboard activity.

Contents: xscreensaver-2.14

Author:  Jamie Zawinski
Version:  2.14-7

xselect Manipulate the XSelection

With this little tool one can pipe the currently selected text under X into a file or vice versa. Help can be found by reading the manpage with man xselection

Contents: xselection-1.6.1

Copyright:  (c) 1990 Richard Hesketh, Computing Lab. University of Kent at Canterbury, UK
Author:  Richard Hesketh <rlh2@ukc.ac.uk
Version:  1.6.1-12

xsession A session manager

The xsession program is a session manager. It is normally executed by your ˜ /.xinitrc (or ˜ /.xsession) script and controls your X Window session. As soon as it is started, xsession launches a window manager and some applications of your choice. At anytime during your session, you may switch to another window manager or execute some other applications from the xsession menus. Examples may be found under /usr/doc/packages/xsession/examples. Contents: xsession-1.1

Copyright:  (c) 1993, 1994, Alain Nissen & Raphael Quinet

Author:

Alain  Nissen  <nissen@montefiore.ulg.ac.be>,  Raphael  Quinet

<quinet@stud.montefiore.ulg.ac.be>

Version:  1.1-11

xsky A Star Atlas

You find it under /usr/X11R6/lib/X11/xsky. Docu: man xsky Contents: xsky-2.0.1

Copyright:  (c) 1992-93 Terry R. Friedrichsen
Author:  Terry R. Friedrichsen <terry@venus.sunquest.com>
Version:  2.0.1-11

xspline draws colorful spline functions in the root window

draws colorful spline functions in the root window Contents: xspline-02may92

Copyright:  (c) 1992 Jef Poskanzer
Author:  Jef Poskanzer <jef@netcom.com>
Version:  92.5.2-9

xspread An X-Windows spreadsheet

A spreadsheet for the X Windows System. Contents: xspread-2.3

Copyright:  (c) 1992 Board of Regents of the University of Wisconsin

Author:

Robert Parbs, James Cornelius, Michael Frey, Dan Gruber, Fang Wang, <softeng@cs.uwm.edu>

Version:  2.3-12

xsysinfo Displays the system load

This program displays information about the load of your system. CPU Load, CPU idle, Memory and Swap are displayed constantly. Contents: xsysinfo-1.2 xsysinfo-1.5

Copyright:  (c) 1992, 1993 - Gabor Herr
Author:  Gabor Herr <herr@iti.informatik.th-darmstadt.de>
Version:  1.5-10

xtail X11 Version of tail

This tool (Motif, statically linked) may especially used as file monitor. Nice to view files. Contents: xtail-1.03

Copyright:  (c) 1994 Klaus Thiele
Author:  Klaus Thiele <100520.634@compuserve.com>
Version:  1.03-19

xtalk X11 client for talk

XTalk is a talkclient and it was written for use under the graphical user interface X11. It is full useable with mouse or the keyboard via shortcuts. The important functions are also accessible over the buttonstrip with easy understand symbols. The functionality compared with the ”normal” talkcommand was expand considerable. It is possible to save the written and received text with an any name. Moreover you can scroll in the text. No informations are lost. The telephonbookfunction is very useful for frequent used talkpartner. The connection parameter will show permanent on the statusline under the buttonstrip.Besides you can open a second or third connection without great difficulty, by using the menupoint ”File->New XTalk”.

Very useful is the build in FTP function. She is based on a own protocol and therefore only enabled, when the partner runs also a XTalk client. By construct a connection, XTalk looks, if XTalk is running on the other end. In this case, the transfer menu will enabled automatically.

Copyright:  (c) 1997 Stephan Pilz
Author:  Stephan Pilz <spilz@htwm.de>
Version:  1.2-11

xtar View contents of a tar file

XTar is a Motif program for viewing and manipulating tar archives (which may be compressed). The contents of the archive are displayed in the main window and various options are available from pulldown and popup menus. Docu: /usr/X11R6/lib/X11/xtar/xtar.doc & man xtar Contents: xtar-1.4

Copyright:  (c) 1993 University of Liverpool
Author:  Rik Turnbull <rik@csc.liv.ac.uk>
Version:  1.4-19

xtexit Prompt on exiting X

‘xtexit‘ sends an request to all clients to shut down. If the application still needs an user interaction (e.g. if a file should be saved) this is possible. If you answer by the affirmative all applications will be closed. This method is not fully water-proof but better than killing each and every client without being able to interfere. xterm applications anyway are killed immediately! If this package is installed, it will be automatically integrated into the sample user fvwm menu. Contents: xtexit

Copyright:  (c) 1990 Front End Oy

Author:

Klaus Franken <kfr@suse.de>, Teemu Torma, Front End Oy <tot@frend.fi>, Paul Raines <raines@slac.stanford.edu>

Version:  0.42-13

xvhelp A help system based on Openlook

This package includes a help browser (xvhelp - similar to the MS Windows Help system) and a compiler (hcc) to create your own help files. This version of xvhelp uses /usr/openwin/lib/help as the directory for help files. Documentation: /usr/doc/packages/xvhelp Contents: XVhelp-1.00

Copyright:  (c) Public Domain
Author:  Roberto Surlinelli, Luca Maranzano, Guido Cardino <calamaro@dist.dist.unige.it>
Version:  1.0-12

xvi Multi-window vi

Xvi is a portable multi-window version of vi. Different files can be displayed in separate windows (instead of displaying them in one window, separated by horizontal status lines). It is also possible to display different parts of one file in different windows. Contents: xvi-3.1

Copyright:  (c) 1990-92 Chris and John Downey
Author:  Chris Downey, John Downey
Version:  3.1-13

xvnews A News reader based on Openlook

If you find trn too obscure, this news reader might save you a lot of trouble when reading News. Requirements: access to an NNTP server. Contents: xvnews-2.3.2

Copyright:  (c) 1988-91 Ellen M. Sentovich, Rick L. Spickelmier
Author:  Ellen M. Sentovich , Rick L. Spickelmier, J.J. deGraaff <J.J.deGraaff@twi.tudelft.nl>
Version:  2.3-14

xvtdl A todo list manager

XVTDL has the following features:

* You can manage multiple todo lists; lists can have sublists.

* Each todo list item is prioritized; each list can be displayed in a combination of priority, chronological, and alphabetical order.

* List items, if you do not ”check them off”, will propagate from one day to the next.

* List items can be recurring items, i.e., items that will automatically pop up weekly, biweekly, monthly, etc. These items will propagate like ”normal” list items.

* List items can have deadlines associated with them, with several types of actions that execute on or after a deadline.

* You can set list items on any day, with easy calendar management.

* Lists can be printed on a ”plain” or PostScript printer.

* Printing/display supports ISO 8859.1 character set.

* List items can be logged when checked on/off with optional comments added to the log file.

* List items can be retained or deleted when checked off.

* Multiple lists can be combined for viewing or printing.

* Multiple lists from multiple files can be manipulated.

* Changes to the todo list can be made and are detected by the running programs.

Contents: xvtdl-5.2

Copyright:  (c) 1991-94 Mike Jipping, Hope College
Author:  Mike Jipping <jipping@cs.hope.edu>
Version:  5.2-11

xwpick X Windows Image Pick up Utility

Xwpick lets you pick an image from an arbitrary window or rectangular area of an X Window (DEC Window) server and write it to a file in a variety of formats.

Contents: xwpick-2.20

Copyright:  (c) 1993, 1994 by Evgeni Chernyaev
Author:  Evgeni Chernyaev <chernaev@mx.ihep.su>
Version:  2.20-12

xzap tool to view X11 programs

This very useful tool displays the running applications. Using this tool, applications which are not responding or can not be found on the desktop any more can be killed easily.

Contents: xzap-1.1

Copyright:  (c) 1993 The University of Newcastle upon Tyne
Author:  Gerry Tomlinson <gerry.tomlinson@newcastle.ac.uk>
Version:  1.1-13

xzet Graphical System Activity Monitor

XZET is an X windows application which is Zmans’ vesion of Xetat and Top, hence the name XZET. It provides 3 different types of real time system resource monitoring windows. The first is a global system activity window which displays stripcharts of global idle, user, system and nice percentages. It also displays stripcharts for free swap space and free physical memory. Finally, it has displays for global process counts and states, and free space monitoring of up to three file systems. A second window lists information on every process running in the system. The information varies based on whether a short or long format is selected. The long format is quite complete. The user can toggle the state of this list between ”frozen” and ”auto-updating”. If an entry in the list is selected with the mouse, a third and final window type will appear. This is an individual process window which features an updating tachometer widget for both the processes percent cpu load and the percent physical memory load. Information similar to that provided in the short format of the global process list is also enumerated.

Copyright:  (c) 1992-94 Loganville Technical Resources, Inc

Author:

Branko Lankester, Roger Binns, Michael K. Johnson <johnsonm@stolaf.edu>, Robert J. Nation <nation@rocket.sanders.lockheed.com>

Version:  1.0-16

XDEV: Development under X11

0.1.22 XDEV: Development under X11

ddd Debugger with graphical user interface

The DDD debugger (Data Display Debugger) is a comfortable GUI to the well-known debuggers GDB and DBX. Data structures can be represented as graphs and interactively shown. You can debug programs in C, C++, Pascal, MODULA-2, FORTRAN and ADA, and even at the assembler code level! Contents: ddd-2.2.3

Copyright:  (c) 1995-98 Technische Universitaet Braunschweig, Germany.

Author:

Andreas Zeller <zeller@ips.cs.tu-bs.de>, Dorothea Luetkehaus <luetke@ips.cs.tubs.de>

Version:  3.0-1

dlgedit QT Dialog Editor

A nice small tool to design interactively input dialogs for the QT Toolkit. Documentation: /usr/doc/packages/dlgedit Contents: dlgedit-1.2

Copyright:  (c) 1997, Jeff Harris, Klaus Ebner
Author:  Jeff Harris <jharris@cis.ohio-state.edu>, Klaus Ebner <klaus@gaspode.ndh.com>
Version:  1.2-17

ezwgl ezwgl

EZWGL stands for ”The EZ Widget and Graphics Library”. It is a library of about 400 C functions written on top of Xlib. It is designed for developing graphics user interface under the X window system programming environment.

Copyright:  (c) GPL
Author:  <mzou@math.utexas.edu>
Version:  1.38-6

iv InterViews - a GUI.

Copyright:  (c) 1987-1991 Stanford University; 1991 Silicon Graphics, Inc.

Author:

Michael N. Lipp <mnl@dtro.e-technik.th-darmstadt.de>, Herbert Thielen <thielen@lpr.e-technik.tu-muenchen.de>

Version:  3.1g-1

jx jx, jcc, C/C++ Development Environment

Code Crusader is a complete code development environment, inspired by MetroWerks CodeWarrior. It can run on any UNIX machine that uses the X Window System. JX is a full-featured application framework and widget library for use with the X Window System. It provides support for all facets of application development, including distributed applications, and aims to combine the best of MacOS and NeXTSTEP. It is built directly on top of Xlib and has been carefully optimized for performance.

Copyright:  (c) John Lindal
Author:  John Lindal <jafl@alice.wonderland.caltech.edu>
Version:  1.1.16-5

qtdevel QT development environment

This package contains the QT runtime libraries. The former QT package was split into the following ones: qtlib runtime libraries qtdevel developpment environment qtext extensions If you just want to run KDE, it is sufficient to install the package qtlib. Please read in /usr/doc/packages/qt/LICENSE for licensing issues! Please read in /usr/doc/packages/qt/LICENSE for licensing issues!

Copyright:  (c) 1992-1997, Troll Tech AS
Author:  Troll Tech AS <info@troll.no>
Version:  1.33-8

qtext QT extensions

This package contains some QT extensions (Netscape plugin, ImageIO, OpenGL). The former QT package was split into the following ones: qtlib runtime libraries qtdevel developpment environment qtext extensions If you want to develop programs with QT or simply e.g. recompile KDE, then you have to install the qtdevel package. Please read in /usr/doc/packages/qt/LICENSE for licensing issues! Please read in /usr/doc/packages/qt/LICENSE for licensing issues!

Copyright:  (c) 1992-1997, Troll Tech AS
Author:  Troll Tech AS <info@troll.no>
Version:  1.33-8

qtlib QT runtime libraries

This package contains the QT runtime libraries. The former QT package was split into the following ones: qtlib runtime libraries qtdevel developpment environment qtext extensions If you want to develop programs with QT or simply e.g. recompile KDE, then you have to install the qtdevel package. Please read in /usr/doc/packages/qt/LICENSE for licensing issues! Contents: qtlib-1.33 Please read in /usr/doc/packages/qt/LICENSE for licensing issues!

Copyright:  (c) 1992-1997, Troll Tech AS
Author:  Troll Tech AS <info@troll.no>
Version:  1.33-8

vtk Visualization ToolKit

The Visualization ToolKit (vtk) is a software system for 3D Computer Graphics and Visualization. Vtk includes a textbook published by Prentice-Hall (”The Visualization Toolkit, An Object-Oriented Approach To 3D Graphics, 2nd edition” ISBN 0-13-954694-4), a C++ class library, and several interpreted interface layers including Tcl/Tk, Java, and Python. Vtk has been implemented on nearly every Unix-based platform and PC’s (Windows NT and Windows95). The design and implementation of the library has been strongly influenced by object-oriented principles. Contents: vtk20p1src

Copyright:  (c) Ken Martin, Will Schroeder, Bill Lorensen
Version:  2.0-10

vtkdata Visualization ToolKit - Datafiles

Datafiles for the Visualization ToolKit.

XDEV: Development under X11

Contents: vtk20data.part1 vtk20data.part2 vtk20data.part3 vtk20data.part4 vtk20data.readme vtk20manhtml

Copyright:  (c) Ken Martin, Will Schroeder, Bill Lorensen
Version:  2.0-4

wxwdemo Demo programs for wxwin

Some examples to show the possibilities of wxwindows. These programs have been compiled using OpenLook.

Copyright:  (c) 1993-97, AIAI, University of Edinburgh
Author:  Julian Smart <J.Smart@ed.ac.uk>
Version:  1.67-16

wxwdocht Documentation for wxwin in HTML

Documentation for wxwin in HTML.

Copyright:  (c) 1993-97, AIAI, University of Edinburgh
Author:  Julian Smart <J.Smart@ed.ac.uk>
Version:  1.67-16

wxwdocps Documentation of wxwin in PostScript

Documentation of wxwin in PostScript

Copyright:  (c) 1993-97, AIAI, University of Edinburgh
Author:  Julian Smart <J.Smart@ed.ac.uk>
Version:  1.67-16

wxwin GUI Library for Motif, Openlook and MS-Windows

wxWindows is a Class Library for C++, which provides a GUI for more than one platform. The current version supports OpenLook (XView), Motif and Windows 3.1. WxWindows has originally been developed at the Artificial Intelligence Applications Institute at the University of Edinburgh. This version has been compiled for the use with OpenLook but there is a MS-Windows version as well. The whole package can be found under /usr/lib/wxwin.

Copyright:  (c) 1993-97, AIAI, University of Edinburgh
Author:  Julian Smart <J.Smart@ed.ac.uk>
Version:  1.67-16

wxwstat Static libraries for package wxwin

This package contains the static libraries for package wxwin.

Copyright:  (c) 1993-97, AIAI, University of Edinburgh
Author:  Julian Smart <J.Smart@ed.ac.uk>
Version:  1.67-16

wxxt Port of wxwin with motif look-and-feel

The package wxxt is a port of wxwin that imitates the look and functionality of a motif interface but does not need the motif libraries to be compiled or run.

Copyright:  (c) 1995, AIAI, University of Edinburgh (Julian), 1995, GNU (Markus)
Author:  Markus Holzem <mho@comnets.rwth-aachen.de>
Version:  1.66d-10

xforms Forms Library

XForms is a GUI toolkit based on Xlib for X Window Systems. It features a rich set of objects, such as buttons, sliders, and menus etc. integrated into an easy and efficient object/event callback execution model that allows fast and easy construction of X-applications. In addition, the library is extensible and new objects can easily be created and added to the library. Contents: xforms-0.88 Please, note the Copyright under /usr/doc/packages/xforms/Copyright; extract: If you intend to use xforms commercially, this includes in-house use and consulting, you must contact the authors at xforms@world.std.com for a license arrangement. Running a publicly available freeware that requires xforms is not considered commercial use, even in a commercial setting.

Copyright:  (c) 1995 T.C. Zhao and Mark Overmars
Author:  T.C. Zhao <zhao@bragg.phys.uwm.edu>, Mark Overmars
Version:  0.88.1-3

xtpanel Interface Builder using scripts

This program allows you to build (complex) programs using a simple scripting language. Included are a tutor program and some sample scripts (xt-colors, xt-lister, xt-load, xt-lpq, xt-page). Documentation: /usr/doc/packages/xtpanel Contents: xtpanel-3.01

Copyright:  (c) 1992, 1993 - Board of Trustees - Leland Stanford Junior University
Author:  Steve Cole <steve@sep.stanford.edu>, Dave Nichols <dave@sep.stanford.edu>
Version:  3.01-9

xwb software development environment for openwin

Graphical user interface using the Openwin look. Supports editing, compiling and launching of C programs with a mouse-click. Contents: xwb v3.0

Copyright:  (c) 1996 by Michael Sohmen
Author:  Michael Sohmen <SOMI0011@RZ03.FH-Karlsruhe.de>
Version:  3.0-14

xwpe A X-window programming environment

xwpe is a X-window programming environment designed to use on UNIX-systems. It is similar to ’Borland C++ or Turbo Pascal’ environment. The difference between the programming environments from Borland an xwpe is that many compilers and linkers may be started. From a menu three different debuggers can be chosen. And of course not only key strokes can select these options but you can also use a mouse. Contents: xwpe-1.4.2

Copyright:  (c) 1994 Fred Kruse
Author:  Fred Kruse <kruse@rrzn-nis.uni-hannover.de>
Version:  1.4.2-14

xxgdb xxgdb, user interface to the GNU debugger under X

Xxgdb is a graphical user interface to the gdb debugger under the X Window System. Contents: xxgdb-1.12

Copyright:

(c) 1989 The University of Texas at Austin, 1990 Microelectronics and Computer Technology Corporation

Author:  Po Cheung <cheung@sw.mcc.com>
Version:  1.12-12

ygl a GL graphics library with sample file

Besides the Ygl library the glwidget library is included. You can check what to do with this library by trying one of the examples that come along: coltest, lines, lissa, rgbtest, simple and smile Contents: Ygl-3.1

Copyright:  (c) 1994 Fred Hucht, Michael Staats

Author:

Fred  Hucht  <fred@thp.Uni-Duisburg.DE>  ,  Michael  Staats

<michael@hal6000.thp.Uni-Duisburg.DE>

Version:  3.1-14

0.1.23 XSRV: Several X Servers (XFree86^{TM}3.3 and other)

xvga16 server for vga cards (16 colors)

This server uses standard video modes but only supports 16 colors. It should work with almost any video-hardware.

Version:  3.3.2.3-5

x8514 server for 8514-based cards

Please only install if you have a fully 8514 compatible graphics card.

Version:  3.3.2.3-5

xagx server for agx-cards

X-Server for graphic cards based on the AGX chipset. If there is no card installed that uses this chipset you must not install this server.

Version:  3.3.2.3-5

XSRV: Several X Servers (XFree86^{TM}3.3 and other)

xi128 Server for Number Nine Imagine 128 graphic cards

Version:  3.3.2.3-5

xmach32 Mach32 Server

X-Server for graphic cards based on the Mach32 chipset. If this card is not installed on your system you should not install this server.

Version:  3.3.2.3-5

xmach64 Mach64 Server

If you don’t have this special card installed in you computer you should not install this server.

Version:  3.3.2.3-5

xmach8 Mach8 server

If you don’t have this special card installed in you computer you should not install this server.

Version:  3.3.2.3-5

xmono X-monochrome server

If you don’t have a monochrome card installed in your computer you should not install this server.

Version:  3.3.2.3-5

xp9k Accelerated server for P9000-based cards

This server is designed for the following cards: Diamond Viper VLB, Diamond Viper PCI, Orchid P9000 and some clones (STAR 2000). Cards like Viper Pro and other P9100 and P9130 cards are not yet supported.

Version:  3.3.2.3-5

xs3 server for S3-based cards (excluding ViRGE and ViRGE/VX)

This server supports the following chipsets: 911, 924, 801/805, 928, 864, 964, Trio32, Trio64, Trio64V+, Trio64UV+, Aurora64V+, Trio64V2/DX and /GX, but not ViRGE and ViRGE/VX. It does not necessarily work with every card that is using one of the supported chipsets. For more information please refer to the chapter about X11 in our handbook.

Version:  3.3.2.3-5

xs3v server for S3 (ViRGE and ViRGE/VX)-based cards

This server is exclusively for cards with S3 ViRGE and ViRGE/VX. Only install it if you have such a card. Staring with XFree86-3.3 the SVGA server has S3V support as well and is to be prefered over the standalone S3V server.

Version:  3.3.2.3-5

xsvga server for svga cards

If you don’t have a SVGA card installed you shouldn’t install this server. Please check the handbook on details about the chipsets supported. This server contains experimental suport for the Matrox Millennium II.

Version:  3.3.2.3-5

xw32 server for W32 cards

This is the server for ET4000-W32 and ET6000 cards. If you don’t have such a card you shouldn’t install it. The SVGA server has better and faster support for most of these cards and offers high and true color on many W32 and all ET6000 cards.

Version:  3.3.2.3-5

xcyrix alpha quality server for Cyrix MediaGX

This is a alpha quality server for the Cyrix MediaGX CPU with graphics included. Only install it if you have this CPU installed in your system. This server can only be configured with xf86config. This server can only be configured using xf86config

Copyright:  (c) 1997-1998 S.u.S.E. GmbH ; 1992-1998 The XFree86 Project, Inc.
Version:  1.0-2

xglint Accelerated server for GLINT/PERMEDIA/PERMEDIA-2 based cards

This server supports cards based on the 3DLabs chipsets GLINT 500TX/MX, Permedia, Permedia2 and Permedia 2v. Cards that use a Ti Ramdac together with the 500TX or Permedia are currently not supported. Only install it if you have one of the supported cards installed in your system. Please note that this X server can ONLY be configured with ”xf86config”. Unfortunately the configuration with ”XF86Setup” is not possible.

Copyright:  (c) 1997-1998 S.u.S.E. GmbH / 1992-1998 The XFree86 Project, Inc.

Author:

Alan Hourihane <alanh@fairlite.demon.co.uk>, Dirk Hohndel <hohndel@suse.de>, Stefan Dirsch <sndirsch@suse.de>, Helmut Fahrion <hf@suse.de>,

Version:  4.11-0

xp9x00 non accelerated server for Weitek P9100 based cards

This server supports Weitek P9100 based cards. Only install it if you have one of these motherboards installed in your system. This server can only be configured with xf86config. This server can only be configured using xf86config

Copyright:  (c) 1997-1998 S.u.S.E. GmbH ; 1992-1998 The XFree86 Project, Inc.
Version:  1.0-0

xrenditn non accelerated server for Rendition Verite based cards

This server supports card based on Rendition Verite V1000, V2100, V2200. Only install it if you have one of these cards installed onto your system. Please note that this X server can ONLY be configured with ”xf86config”. Unfortunately the configuration with ”XF86Setup” is not possible.

Copyright:  (c) 1997-1998 S.u.S.E. GmbH / 1992-1998 The XFree86 Project, Inc.
Version:  1.0-1

xsis alpha quality server for SiS 5597

This is a alpha quality server for the SiS onboard UMA graphics chipset 5597. Only install it if you have one of these motherboards installed in your system. This server can only be configured with xf86config. This server can only be configured using xf86config

Copyright:  (c) 1997-1998 S.u.S.E. GmbH ; 1992-1998 The XFree86 Project, Inc.
Version:  2.2-0

xneomag Server for graphics cards based on NeoMagic chipsets

For exact information about supported cards, color depths and the license of this binary only server please consult the files in /usr/doc/packages/xneomag.

Copyright:  (c) 1998 Red Hat Software

Author:

Jens  Owen  <jens@precisioninsight.com>,  Kevin  E.  Martin

<kevin@precisioninsight.com>

Version:  1.0.0-1

0.1.24 XV: XView (OpenLook, Applications)

ssbase Slingshot extensions

These are extensions for XView programming; it contains static and dynamic libraries in elf-format. Also included are the include files and the documentation. Contents: SlingShot2.1

Copyright:  (c) 1991 Sun Microsystems, Inc.
Version:  2.1-17

ssexmp Slingshot examples

A little collection of Slingshot examples. You can see the use and the advantage of the Slingshot extension package.

Copyright:  (c) 1991 Sun Microsystems, Inc.
Version:  2.1-17

XWM: Window Manager and Desktop

xtoolpl xtoolplaces and xtoolwait

This program offers you information about every window existing on your current desktop; providing a means of starting it at exactly the same position. Furthermore there is the program xtoolwait that just waits after an X application has been launched and presents its window.

Contents: xtoolplaces-1.3.1 xtoolwait-0.3

Copyright:  (c) 1993 Xerox Corporation.

Author:

Adam  Stein  <adam@iset.scan.mc.xerox.com>,  Richard  Huveneers

<richard@hekkihek.hacom.nl>

Version:  1.3.1-12

xv32aout Shared XView Libraries (a.out)

These are the a.out libraries of the XView system.

Version:  3.2-1

xv32base XView (OpenLook) Base system

Programs, dynamic libraries, configuration files and documentation for XView 3.2 for Linux. It contains: xgettext, msgfmt, olwm, svenv, toolwait, winsysck, owplaces, openwin, xcenter, clock, cmdtool, olwmslave, props, textedit, meminfo

Contents: xview-3.2p1.4.src

Copyright:  (c) 1990-1992 Sun Microsystems, Inc.
Version:  3.2p1.4-12

xv32dev XView (OpenLook) development kit

Imake-configuration and include files for developing under XView 3.2.

Copyright:  (c) 1990-1992 Sun Microsystems, Inc.
Version:  3.2p1.4-12

xv32exmp Sample programs for Xview

Sample programs for Xview which demonstrate the Slingshot and UIT extensions, libraries which greatly simplify programming a user interface under X.

Copyright:  (c) 1990-1992 Sun Microsystems, Inc.
Version:  3.2p1.4-12

xv32slib XView version 3.2 static libraries

Static libraries for XView 3.2

Copyright:  (c) 1990-1992 Sun Microsystems, Inc.
Version:  3.2p1.4-12

0.1.25 XWM: Window Manager and Desktop

3dpixm 3d pixmap icons for fvwm

Very nice 3d pixmaps for the fvwm window manager, in two sizes, for menus and button bars. You should install this package.

Contents: 3dpixm

Author:  Joerg Mertin <joerg@pc50.zrz.tu-berlin.de>
Version:  98.2.4-5

3dpixms Small 3d pixmap icons for fvwm

Very nice 3d pixmaps for the fvwm window manager. These pixmaps are used for the pop-up menus.

Author:  Joerg Mertin <joerg@pc50.zrz.tu-berlin.de>
Version:  98.2.4-5

afterstp NeXTStep-like window manager

This window manager follows the tradition of NeXTStep. The functionality of the bowman window manager was enhanced with some configuration options. As well, there are shadowed menus, the button-bar Wharf, animated windows, a proper sound support, and many more features. It should be started with a color depth of at least 16 bits! Documentation: /usr/doc/packages/afterstp Contents: AfterStep-1.0

Copyright:

(c) 1988 Evans & Sutherland Computer Corporation, 1989 Massachusetts Institute of Technology,1994 Robert Nation, 1996 by Beat Christen

Author:

Frank Fejes <frank@ssax.com> , Dan Weeks <dan@mango.sfasu.edu>, Jonathan B. Leffert <jonl@geom.umn.edu>, Bo Yang <eric@coeus.ucsd.edu>, Alfredo Kengi Kojima, Rob Nation, Dave Adkins, Beat Christen, Robert Depenbrock, Amadeus Foudray, Martin Scherbaum <maddin@suse.de>

Version:  1.0-11

amiwm AMIGA-like window manager

For those who still want their AMIGA desktop :-) Documentation: /usr/doc/packages/amiwm Contents: amiwm-0.20pl28

Copyright:  (c) 1996, Marcus Comstedt
Author:  Marcus Comstedt <marcus@lysator.liu.se>
Version:  0.20pl28-12

bowman Window manager with the NeXT-Step ”look and feel”

Based on the fvwm window manager, imitating the NeXTSTEP GUI. Can be installed without problem with fvwm and fvwm95. Documentation: /usr/doc/packages/bowman Contents: bowman-0.93e

Copyright:  (c) Bo Yang
Author:  Bo Yang <eric@coeus.ucsd.edu>
Version:  0.93e-14

cdesim CDE simulator

CDEsim simulates a window manager. With the help of fvwm (version 2), it simulates the ”look and feel” of the popular CDE user interface. More the look than the feel. It is fully integrated into the susewm window manager configuration tool, and will be automatically preconfigured. Documentation and original configuration files under: /usr/doc/packages/cdesim Please read /usr/doc/packages/cdesim/README.cdesim (also available from the cdesim menu!).

Copyright:  (c) October 9, 1996, Michael A. Haas
Author:  Michael A. Haas <gnosis@infinet.com>, Martin Scherbaum <maddin@suse.de>
Version:  0.42-17

ctwm Window Manager

A very popular window manager among users with fast video cards. ctwm is an improvement of twm (Tab Window Manager). If you launch ctwm from the fvwm2 menu, it will return to fvwm2 when you click on the restart menu button in the ctwm root menu. Documentation: man ctwm or in /usr/doc/packages/ctwm Contents: ctwm-3.4pl2

Copyright:  (c) 1992 Claude Lecommandeur

Author:

Claude Lecommandeur <lecom@sic.epfl.ch>, Dave Payne, Dave Sternlicht, Steve Pitschke, Tom LaStrange

Version:  3.4pl2-12

fvwm THE window manager for Linux

The fvwm window manager (version 2) is the standard window manager for the X window system under Linux. fvwm is highly configurable and has a similar look-and-feel to the Motif standard. The personal configuration is stored in the file ˜ /.fvwm2rc in the user’s home directory. The system wide configuration is generated at installation time by SuSEconfig and may be found in /usr/X11R6/lib/X11/fvwm2 . Documentation: man fvwm2 /usr/doc/packages/fvwm

XWM: Window Manager and Desktop

Contents: cdesim-0.42 fvwm-2.0.46

Copyright:  (c) 1994 Robert Nation, Mark Boyns, Mark Scott

Author:

Charles  Hines  <chuck hines@vnet.ibm.com>,  Robert  Nation

<fvwm@wonderland.org>, Andrew Davison <davison@cs.monash.edu.au>, Drew Moseley <ddm@cis.ufl.edu>, Jason L Tibbitts <tibbs@tcamc.uh.edu>, John Heidemann <johnh@ficus.CS.UCLA.EDU>, Mark Boyns <boyns@sdsu.edu>, Mark Scott <mscott@mcd.mot.com>, Mike Finger <mfinger@mermaid.micro.umn.edu>, Mr. Per Persson <pp@solace.mh.se>, Nobutaka Suzuki <nobuta-s@is.aist-nara.ac.jp>

Version:  2.0.46-21

fvwm1 THE window manager for Linux (old version)

Since versions 1.x and 2.x are not compatible, we have kept the old version in the distribution. This version is configured with the system-wide configuration file /usr/X11R6/lib/X11/fvwm/system.fvwmrc, as well as the user’s personal configuration file ˜ /.fvwmrc in the home directory. The program susewm is able to configure this window manager,too. Important! Please read the file /usr/doc/packages/fvwm1/README.SuSE! Documentation: man fvwm fvwmman <manpage> /usr/doc/packages/fvwm1 Contents: fvwm-1.24r

Copyright:  (c) 1994 Robert Nation, Mark Boyns, Mark Scott

Author:

Andrew  Davison  <davison@cs.monash.edu.au>,  Drew  Moseley

<ddm@cis.ufl.edu>, Jason L Tibbitts <tibbs@tcamc.uh.edu>, John Heidemann <johnh@ficus.CS.UCLA.EDU>, Mark Boyns <boyns@sdsu.edu>, Mark Scott <mscott@mcd.mot.com>, Mike Finger <mfinger@mermaid.micro.umn.edu>, Mr. Per Persson <pp@solace.mh.se>, Nobutaka Suzuki <nobuta-s@is.aist-nara.ac.jp>, Robert Nation <fvwm@wonderland.org>

Version:  1.24r-16

fvwm95 Virtual window manager, look’n’feel of Win95

The look’n’feel of Win95 without Win95! With icons in the menus, frames and handling like win95, a task-bar with start button a.m.o.m. By installing the package susewm one can configure fvwm95 similar to fvwm, fvwm2 or bowman. Important1! The config file’s name was changed to .fvwm95rc ! Important2! Please read the file /usr/doc/packages/fvwm95/README.SuSE! Documentation: /usr/doc/packages/fvwm95 man fvwm95 fvwm95man <man page> Contents: fvwm95-2.0.43a-Imake

Copyright:  (c) 1993, Robert Nation
Author:  Rob Nation <fvwm@wonderland.org>, Hector Peraza <peraza@mitac11.uia.ac.be>
Version:  2.0.43a-22

icewm Windowmanager with Taskbar

Icewm is a windowmanager often used together with GNOME. Simple and fast. Documentation: /usr/doc/packages/icewm

Copyright:  (c) 1997, Marko Macek
Author:  Marko Macek <Marko.Macek@snet.fri.uni-lj.si>
Version:  0.9.8-1

icons Lots of icons

An extensive collection of more than 1800 icons for any purpose. Among them black/white and color ones, small and huge ones, simple and artistic ones. Documentation in: /usr/doc/packages/icons Contents: icons

Version:  98.2.4-5

lxuser Configuration for the sample user

This package contains the main dependencies for the sample user.

Copyright:  (c) 1995,1996,1997 S.u.S.E. GmbH
Version:  97.6.29-2

mlvwm MacIntosh like window manager

Mac users will feel right at home! This is a very early version. Documentation can be found: /usr/doc/packages/mlvwm man mlvwm Contents: mlvwm-current

Copyright:  (c) 1996, Takashi HASEGAWA / Uchikawa-Lab.
Author:  Takashi HASEGAWA <tak@bioele.nuee.nagoya-u.ac.jp>
Version:  0.7.0-11

olvwm olvwm - OpenLook Virtual Window Manager

This is a virtual window manager based on the older OpenLook window manager olwm with exactly the same ”look & feel”.

Copyright:  (c) (c) Copyright 1989 Sun Microsystems, Inc.
Author:  Scott Oaks <scott.oaks@East.Sun.COM>
Version:  3.2p1.4-12

pixmaps Icons and pixmaps for window managers

In this package you find all icons of the window managers of the S.u.S.E. Linux CD. The window managers concerned are:

- fvwm (package fvwm1) - fvwm2 (package fvwm) - fvwm95 (package fvwm95) - bowman (package bowman) - AfterStep (package afterstp) - CDEsim (package cdesim) - ctwm (package ctwm) - qvwm (package qvwm) - mlvwm (package mlvwm) - lxuser (former package lxuser) Also you find in this package the program xpmroot, formerly contained in the fvwm* window manager family. The reason why we put the icons together in this package is to avoid conflicts and redundancy whilst installing or removing packages. If there were icons in the former package with identical name but different contents these icons were renamed (icon1,icon2,iconN). Nevertheless, you may find the original pixmaps in the source packages (package s) of the above mentioned window managers. Documentation: /usr/doc/packages/pixmaps/README.SuSE man xpmroot Contents: pixmaps-0.1

Copyright:  (c) 1997, S.u.S.E. GmbH, Fuerth
Author:  Martin Scherbaum <maddin@suse.de>
Version:  0.1-16

qvwm Win95 like window manager

Q V stands for 9 5 in Japanese. It is a window manager that resembles the Win95 GUI Documentation in: /usr/doc/packages/qvwm Contents: qvwm-1.0b6g

Copyright:  (c) 1996, KOURAI Kenichi
Author:  KOURAI Kenichi <kourai@is.s.u-tokyo.ac.jp>
Version:  1.0b6g-12

susewm Window manager configuration tool

This package allows you to create in one step the configuration files for the fvwm, fvwm2, fvwm95, bowman, afterstep, cdesim, mwm, and ctwm window managers. In the menus you’ll only find programs which are really installed in the system. The desktop of the seven window managers will look the same (considering the particularities of the window managers). In version 1.5 for mwm and ctwm there is only support for keeping the menus actual since these two are very different from the five other window managers.

0.2. Disk usage of the packages

The icons used in the fvwm2 / fvwm95 menus are taken from the package 3dpixms which themselves are generated from the 3dpixm package. You can find some information in /usr/doc/packages/susewm. The documentation is also accessible by the desktop menu. Contents: susewm-1.6

Copyright:  (c) 1996-1997 S.u.S.E. GmbH, Fuerth, Germany
Author:  Martin Scherbaum <maddin@suse.de>
Version:  1.61-3

wm2 A minimalist window manager

For all who hate popup menus and icons. Just set WINDOWMANAGER in ˜ /.bashrc to wm2. Dokumentation in: /usr/doc/packages/wm2 Contents: wm2-4

Copyright:  (c) 1996, Chris Cannam
Author:  Chris Cannam <cannam@zands.demon.co.uk>
Version:  4-13

wmaker WindowMaker

WindowMaker is a colourful and felxible window manager and the designated successor of AfterStep. Especially the so called ”themes” make it very interesting. Those themes may be found in package wmthemes. To use it just place the following line in your ˜ /.bashrc: export WINDOWMANAGER=windowmaker Documentation in: /usr/doc/packages/wmaker

Copyright:  (c) 1997, Alfredo K. Kojima
Author:  Alfredo K. Kojima <kojima@windowmaker.org>
Version:  0.15-3

wmthemes Themes for WindowMaker

This packages contains more than 70 themes for the WindowMaker window manager. (package wmaker). Documentation: /usr/doc/packages/wmthemes

Copyright:  (c) many authors
Version:  0.1-1

enlight The Enlightenment windowmanager (beta)

The Enlightenment windowmanager by Carsten ”Rasterman” Haitzler. This is version 0.14-SNAP (Beta!).

Copyright:  (c) 1997 Carsten Haitzler (Rasterman)
Version:  0.14snap-4

esound The sound daemon for the Enlightenment window manager

For enabling Enlightenment to produce some sounds..

Copyright:  (c) Carsten Haitzler
Author:  Carsten Haitzler <raster@redhat.com>
Version:  0.2-4

fnlib The fnlib (font lib)

The fnlib is necessary for the Enlightenment window manager.

Copyright:  (c) Carsten Haitzler
Author:  Carsten Haitzler <raster@redhat.com>
Version:  0.2-4

0.2 Disk usage of the packages

Package  /  /usr/X11  /usr  /var  Σ

Series a1

aaa base  426  16  99  22  564
aaa dir  29  17  63  27  136
at  0  0  55  1  57
base  96  0  288  0  385
bash  403  0  748  0  1152
bash1  329  0  40  0  369
bdflush  6  0  1  0  8
compress  0  0  17  0  17
cpio  0  0  75  0  75
cron  0  0  45  2  48
db  0  0  539  0  539
devs  51  0  0  0  51
diff  0  0  118  0  118
eazy  24  0  1  0  25
elvis  0  0  745  0  745
ext2fs  352  0  109  0  461
file  0  0  173  0  173
fileutil  167  0  187  0  355
find  0  0  110  0  110
gawk  0  0  420  0  420
gdbm  0  0  470  0  470
gppshare  0  0  1760  0  1760
groff  0  0  2480  0  2480
gzip  44  0  22  0  66
kbd  41  0  1462  1  1504
kernmod  5262  0  0  0  5262
ldso  336  0  106  0  442
less  0  0  101  0  102
libaout  620  0  1507  0  2128
libz  0  128  214  0  343
lilo  63  0  530  0  593
loadlin  152  0  0  0  152
localedb  0  0  3680  0  3680
man  2  0  209  109  321
modules  183  0  10  0  193
ncurses  1312  0  4356  0  5668
net tool  99  0  35  0  135
netcfg  15  0  0  1  16
nkita  12  0  584  0  597
nkitb  12  0  967  0  980
nvi  0  0  373  0  373
pcmcia  143  15  217  0  376
perl  2  0  7129  2  7133
ps  43  17  233  0  294
rpm  527  0  615  1  1144
sh utils  111  0  335  0  447

Package  /  /usr/X11  /usr  /var  Σ
shadow  300  0  519  0  819
shlibs  1851  0  463  0  2314
shlibs6  301  0 31484  0  31785
syslogd  1  0  53  0  54
sysvinit  221  0  21  0  242
termcap  0  0  186  0  186
terminfo  0  0  884  0  884
texinfo  0  0  489  0  489
textutil  8  0  320  0  328
timezone  0  0  364  0  364
umsprogs  35  0  3  0  38
util  387  0  413  0  800
vim  1  0  1162  0  1164
yast  2278  0  433  1  2712
Σ  16245  193 68022  167  84646

Series a2

compat  0  0  522  0  522
oldshlbs  1820  0  9  0  1829
Σ  1820  0  531  0  2351

Series ap1

a2ps  0  0  30  0  30
alien  0  0  100  1  101
apmd  0  60  107  0  167
aps  0  0  33  135  169
arc  0  0  122  0  122
ash  60  0  33  0  93
autolog  0  0  19  0  20
bc  0  0  214  0  214
buffer  0  0  13  0  13
bzip  0  0  121  0  121
cdcons  0  0  21  0  21
cdparano  0  0  120  0  120
cdrecord  0  0  242  0  242
cdwrite  0  0  20  0  20
cdwtools  0  0  458  0  458
cfengine  0  0  764  0  764
chipmunk  0  0  266  0  266
ctags  0  0  29  0  29
deco  0  0  216  0  216
dialog  0  0  34  0  34
dosfsck  91  0  1  0  93
duconv  0  0  4  0  4
dump  119  0  11  0  131
ed  0  0  57  0  57
edy  0  0  70  0  70
enscript  4  0  1745  0  1749
gcal  0  0  1535  0  1535

Package  /  /usr/X11  /usr  /var  Σ
gdtmon  299  0  0  0  299
git  0  0  406  0  406
glimpse  0  0  1086  0  1086
gpm  2  0  177  1  181
gs both  0  0  1252  0  1252
gs fonto  0  0  1488  0  1488
gs fonts  0  0  2156  0  2156
gs serv  0  0  1211  0  1211
gs vga  0  0  1217  0  1217
gs x11  0  0  1246  0  1246
hello  0  0  34  0  34
id utils  0  0  291  2  294
idled  0  0  43  0  43
ingres  0  0  2376  0  2376
ingrtool  0  0  312  0  312
inlinem  0  0  44  0  44
isapnp  36  0  60  0  96
jed  0  270  1088  0  1358
joe  0  0  303  0  303
jove  0  0  411  3  415
lha  0  0  46  0  46
logsurf  0  0  162  0  162
lsof  0  0  442  0  442
makewhat  0  0  2  0  2
mbedit  0  0  353  0  353
mc  0  0  716  0  716
mdutils  151  0  33  0  184
mkdosfs  15  0  5  0  21
mkisofs  0  0  622  0  622
mktemp  5  0  2  0  8
mmv  0  0  115  0  115
mpegutil  0  0  93  0  93
mt st  0  0  14  0  14
mtools  1  0  854  0  856
nc  0  0  853  0  853
obst  0  0  9935  0  9935
octave  0  0 47590  0  47590
onyx  0  0  557  0  557
pdksh  0  0  307  0  307
pg datab  0  0  0  876  876
pg ifa  0  158  479  0  637
pg iface  0  1541  388  0  1930
pico  0  0  123  0  123
postgres  0  0  4409  1  4411
ppa  0  0  47  0  47
quota  16  0  65  0  82
recode  0  0  199  0  199
rman  0  0  132  0  132

Package  /  /usr/X11  /usr  /var  Σ
rpm2html  0  0  375  1  376
rpmfind  0  0  1130  0  1130
sc  0  0  213  0  213
screen  0  0  255  0  255
scsiinfo  0  0  163  0  163
seejpeg  0  0  486  0  486
sharutil  0  0  571  0  571
splitvt  0  0  83  0  83
su1  0  0  32  0  33
sudo  0  0  338  1  340
svgatext  44  0  584  1  630
taper  294  0  23  0  317
tcsh  245  0  65  0  311
teapot  0  0  657  0  657
unarj  0  0  35  0  35
units  0  0  208  0  208
unzip  0  0  170  0  170
wdiff  0  0  156  0  156
xdelta  0  0  527  0  527
yamm  0  0  47  0  47
ytree  0  0  112  0  112
z88  0  4708  1131  0  5840
zip  0  0  124  0  124
ziptool  0  0  8  0  8
zoo  0  0  104  0  104
zsh  0  0  646  0  646
Σ  1382  6737 98372  1022  107531

Series ap2

gs lib  0  0  2370  0  2370
Σ  0  0  2370  0  2370

Series beo1

bwatch  0  48  6  0  55
lam  0  0 13463  0  13463
pvm  0  0  6238  0  6238
xpvm  0  686  839  0  1526
Σ  0  734 20546  0  21282

Series d1

autoconf  0  0  558  0  558
automake  0  0  561  0  561
beg  0  0  1709  0  1709
bin86  0  0  71  0  71
binutils  0  0  2219  0  2219
bison  0  0  227  0  227
bwbasic  0  0  149  0  149
clisp  0  0 12954  0  12954
cvs  0  0  751  0  751
des  0  0  113  0  113

Package  /  /usr/X11  /usr  /var  Σ
efence  0  0  41  0  41
f2c  0  0  806  0  806
f2caout  0  0  693  0  693
flex  0  0  183  0  183
fpk  0  0  2675  0  2675
g77  0  0  2387  0  2387
gcc  0  0  2986  0  2986
gcl  0  0  3480  0  3480
gdb  0  0  1154  0  1154
gdlib  0  0  182  0  182
gettext  0  0  1091  0  1091
gforth  0  0  788  0  788
global  0  0  133  0  133
gmp  0  0  300  0  300
gnat  0  0 15576  0  15576
gnats  0  0  1130  8  1138
gobjc  0  0  2574  0  2574
gpc  0  0  2604  0  2604
gpp  0  0  2897  0  2897
guavac  0  0  2719  0  2719
guile  0  0  1608  0  1608
icmake  0  0  221  0  221
indent  0  0  148  0  148
java  0  0 26099  0  26099
javarunt  0  0  9477  0  9477
jdk115  0  0 33295  0  33295
jdk116  0  0 30461  0  30461
jresep  0  0  6066  0  6066
kaffe  0  0  3462  0  3462
lclint  0  0  2394  0  2394
leda  0  0 14936  0  14936
libc  0  0  2325  0  2325
libd  0  0  5751  0  5751
libgpp  0  0  3744  0  3744
libobjc  0  0  2852  0  2852
linclude  0  0  2474  0  2474
linux  0  0 24351  0  24351
lthreads  55  0  113  0  169
lx hack  0  0 68905  0  68905
lx suse  0  0 27074  0  27074
make  0  0  223  0  223
mico  0  0 11774  0  11774
minigcc  0  0  78  0  78
mocka  0  0  1620  0  1620
mtc  0  0  586  0  586
p2c  0  0  593  0  593
patch  0  0  134  0  134
pmake  0  0  124  0  124

Package  /  /usr/X11  /usr  /var  Σ
prcs  0  0  490  0  490
python  0  0  7592  0  7592
rcs  0  0  452  0  452
scheme  0  0  162  0  162
slang  0  0  1159  0  1159
smaltalk  0  0  1393  0  1393
smeiffel  0  0  4068  0  4068
strace  0  0  89  0  89
svgaaout  0  0  166  0  166
svgalib  21  0  1469  0  1490
swi pl  0  0  1855  0  1855
toolpack  0  0  5585  0  5585
tvision  0  0  4708  0  4708
tya  0  0  153  0  153
xclass  0  5461  40  0  5501
xscheme  0  0  218  0  218
yacc  0  0  72  0  72
Σ  76  5461 374270  8  379816

Series d2

joystick  0  0  44  0  44
Σ  0  0  44  0  44

Series doc1

allman  0  0 12607  0  12607
books  0  0 23873  0  23873
booksfnt  0  0  0 11938  11938
cdb  0  0  1680  0  1681
changes  0  0  1783  0  1783
dns faq  0  0  617  0  617
elispint  0  0  1386  0  1386
elispman  0  0  5247  0  5247
fhs  0  0  363  0  363
gawkdoc  0  0  1611  0  1611
gccinfo  0  0  399  0  399
ge info  0  0  980  0  980
gnuhtml  0  0  9958  0  9958
gppinfo  0  0  101  0  101
howto  0  0  4062  0  4062
howtodvi  0  0  9503  0  9503
howtohtm  0  0 11221  0  11221
howtoinf  0  0  359  0  359
howtops  0  0 19930  0  19930
howtosgm  0  0  8047  0  8047
i4ldoc  0  0  1185  0  1185
inf2htm  0  0  26  0  26
isdnpop  0  0  7  0  7
javadoc  0  0 11030  0  11030
ldp  0  0 14861  0  14861

Package  /  /usr/X11  /usr  /var  Σ
ldpman  0  0  763  0  763
ldpmande  0  0  565  0  565
ldpmanfr  0  0  993  0  993
ldpmanit  0  0  293  0  293
libcinfo  0  0  430  0  430
man9  0  0  52  0  52
manyfaqs  0  0  4887  0  4887
odshtml  0  0  307  0  307
perlref  0  0  875  0  875
ppp nt  0  0  63  0  63
pyth doc  0  0  5280  0  5280
rfc  0  0 25900  0  25900
sdb cgi  0  0  11  0  11
sdb de  0  0  8772  0  8772
sdb en  0  0  4438  0  4438
selfhtml  0  0  6733  0  6733
suse lx  0  0 24601  0  24601
suse lxe  0  0 15397  0  15397
susebuch  0  0  829  0  829
susehilf  5  0  1415  1  1422
susepak  0  0 20715  0  20715
susepake  0  0 20662  0  20662
toppp  0  0  8  0  8
uucpdoc  0  0  1536  0  1536
wie dvi  0  0  2400  0  2400
wie geht  0  0  552  0  552
wie html  0  0  2150  0  2150
wie info  0  0  88  0  88
wie pdf  0  0 11798  0  11798
wie ps  0  0  3517  0  3517
wie sgml  0  0  1728  0  1728
xe info  0  2442  0  0  2442
xf86html  0  0  7175  0  7175
Σ  5  2442 315769 11939  330157

Series e1

elib  0  0  189  0  189
emacs w3  0  0  1854  0  1854
ge auc  0  0  1418  0  1418
ge exec  0  0 20220  0  20220
ge lisp  0  0 19283  0  19283
ge site  14  0  975  0  990
lacheck  0  0  51  0  51
psgml  0  0  831  0  831
xe exec  0  34030  0  2  34033
xe lisp  0  26059  0  0  26059
Σ  14  60089 44821  2  104928

Series emu1

Package  /  /usr/X11  /usr  /var  Σ
adamem  0  143  22  0  166
dosemu  8  7  2121  626  2763
ei  0  0  33  0  33
ibcs2  200  0  581  0  781
ines  0  96  84  0  181
stella  0  113  177  0  290
stonx  0  911  931  2  1845
stonxsft  0  0  2 10001  10004
uae  0  0  687  0  687
vgb  0  151  245  0  396
vice  0  0  1598  0  1598
wine  1  0 13597  0  13598
xzx  0  345  121  0  466
z81  0  56  180  0  236
Σ  209  1822 20379 10629  33044

Series fun1

3d chess  0  41  5  0  47
antipoli  0  0  103  0  103
asclock  0  24  5  0  29
astrolog  0  256  234  0  490
batalion  0  521  1607  0  2129
battball  0  286  11  0  297
bsdgames  0  0  1487  14  1502
craft  0  20488  489  0  20977
crosfire  0  24431  3179  0  27611
daliclck  0  63  0  0  63
emiclock  0  245  0  0  245
empire  0  0  105  0  105
figlet  0  0  7148  0  7148
fishtank  0  387  9  0  396
flying  0  154  0  0  154
fortune  0  0  3574  0  3574
freeciv  0  93  1662  0  1756
frisk  0  87  67  0  155
glchess  0  22  116  0  139
gnuchess  0  0  1812  0  1812
gnugo  0  0  38  0  38
gshogi  0  0  558  0  558
hextris  0  75  0  1  76
imaze  0  1135  517  0  1652
lincity  0  1567  1613  0  3181
maelstr  0  253  2246  0  2499
manix  0  17  12  1  31
meltflip  0  18  0  0  18
mirrma  0  74  800  21  896
moontool  0  786  1  0  788
net3d  0  201  0  0  201

Package  /  /usr/X11  /usr  /var  Σ
nethack  0  0  1633  0  1633
oneko  0  35  7  0  43
oonsoo  0  474  4  0  479
pacman  0  183  0  0  183
pipeman  0  38  0  0  38
playone  0  1011  157  0  1168
puzzle  0  21  0  0  21
rockdiam  0  123  1241  248  1613
sastroid  0  0  311  2  313
space  0  0  1867  0  1867
thrust  0  0  256  0  256
vgacard  0  0  120  0  120
x3d  0  1060  6  0  1067
xabacus  0  13  0  0  13
xancur  0  270  14  0  285
xaos  0  0  279  0  279
xball  0  1133  16  0  1149
xbill  0  33  0  152  186
xbl  0  502  13  0  515
xblast  0  340  0  0  340
xboard  0  293  30  0  324
xboing  0  599  378  0  978
xbomb  0  22  0  0  22
xbombs  0  18  0  1  19
xdemine  0  27  0  0  27
xearth  0  173  0  0  173
xengine  0  11  0  0  11
xfract  0  1216  4  0  1220
xgammon  0  3274  0  0  3274
xgas  0  34  0  0  34
xjewel  0  45  0  0  45
xlife  0  467  65  0  533
xlyap  0  29  0  0  29
xmahjong  0  47  7  0  54
xmemory  0  27  2  0  29
xmine  0  24  0  0  24
xmines  0  20  0  0  20
xmountns  0  35  5  0  41
xmris  0  185  0  0  185
xpat2  0  475  0  0  475
xphoon  0  88  0  0  88
xpilot  0  580  822  0  1402
xpinguin  0  181  0  0  181
xpuzzles  0  614  0  0  614
xroach  0  18  0  0  18
xrobots  0  20  0  0  20
xshogi  0  110  0  0  110
xskat  0  180  2  0  182

Package  /  /usr/X11  /usr  /var  Σ
xsnow  0  24  5  0  30
xsok  0  56  59  5  121
xsol  0  29  0  0  29
xspringi  0  59  412  0  471
xtacy  0  67  6  0  74
xtartan  0  59  0  0  59
xteddy  0  68  8  0  77
xtetris  0  40  0  0  40
xthing  0  3891  3  0  3894
xtron  0  117  0  0  117
xvier  0  37  0  0  37
yahtzee  0  0  19  2  21
Σ  0  69719 35149  447  105340

Series gnm1

gnlibs  15019  0  69  0  15088
Σ  15019  0  69  0  15088

Series gnm2

gncore  11745  1  42  0  11789
Σ  11745  1  42  0  11789

Series gnm3

gnadmin  767  0  38  0  805
gngames  2673  0  37  0  2710
gngraph  1091  0  37  0  1129
gnguile  2474  0  38  0  2513
gnmedia  648  0  37  0  685
gnobjc  9521  0  64  0  9586
Σ  17174  0  251  0  17428

Series gra1

bttv  0  1124  113  0  1238
bttvgrab  0  0  69  0  69
cgmdraw  0  0  132  0  132
faxedit  0  489  66  0  555
fbm  0  277  0  0  277
fbm2fli  0  51  4  0  56
films  0  0 18452  0  18452
filmsbig  0  0 151384  0  151384
freetype  0  0  869  0  869
gd  0  0  230  0  231
gif  0  2031  195  0  2226
gimp  0  0 29183  0  29183
gimpman  0  0 13625  0  13625
gltt  0  0  151  0  151
gsview  0  123  0  0  123
gtk  0  0 11350  0  11350
gv  0  278  123  0  402
html2ps  0  36  300  0  337

Package  /  /usr/X11  /usr  /var  Σ
imagemag  0  3443  1377  0  4821
imfilm  0  0  5616  0  5616
imlib  0  0  3022  0  3022
impict  0  0  1727  0  1727
jpeg  0  97  164  0  262
libfbm  0  392  15  0  408
libgif  0  292  0  0  292
libgrout  0  0  1291  0  1291
libjpeg  0  371  0  0  371
liblug  0  344  997  0  1341
libmpeg  0  0  320  0  320
libnetpb  0  189  0  0  189
libpng  0  368  106  0  474
librle  0  295  6  0  302
libtiff  0  1502  0  0  1502
mgp  0  259  539  0  799
mlcreat  0  1349  2654  0  4003
mpegplay  0  126  4  0  131
muscan  0  23  53  0  77
netpbm  0  1309  45  0  1354
photocd  0  53  214  0  268
pictures  0  0  4014  0  4014
pixmp  0  196  0  0  196
plmagick  0  0  100  0  100
povray  0  570  5873  0  6444
pstoedit  0  0  226  0  226
sane  1  716  5605  0  6323
sced  0  495  415  0  911
smii  89  0  7  0  97
surf  0  0  1815  0  1815
tgif  0  1942  14  0  1957
tiff  0  366  332  0  698
tkdraw  0  294  3  0  297
xanim  0  542  906  0  1449
xbmbrosr  0  65  0  0  65
xgrabsc  0  80  13  0  94
xgrasp  0  56  140  0  197
xli  0  235  0  0  235
xmgr  0  1889  1893  0  3782
xmorph  0  1125  8  0  1134
xmrm  0  299  1155  0  1454
xpaint  0  469  18  0  487
xpdf  0  0  616  0  616
xray  0  8630 14434  0  23064
xrayps  0  0 115351  0  115351
xwarp  0  259  5  0  265
zgv  0  0  228  0  228

Package  /  /usr/X11  /usr  /var  Σ
Σ  90  33049 397567  0  430729

Series kde1

kadmin  887  0  0  0  887
kgames  4254  0  25  0  4279
kgoemon  10713  0  0  0  10713
kgraph  1878  0  0  0  1878
kicons  1767  0  0  0  1767
klibs  3307  0  32  0  3339
kmulti  1835  0  22  0  1857
knet  3091  0  22  0  3114
ksupp  1279  0  9  0  1288
kutils  2554  0  22  0  2576
Σ  31565  0  132  0  31698

Series kde2

kbase  6935  0  36  0  6972
Σ  6935  0  36  0  6972

Series kde3

kdir  22  0  0  0  22
Σ  22  0  0  0  22

Series kde4

kadm  1721  0  0  0  1721
kdevel  364  0  0  0  364
kmisc  3629  0  0  0  3629
knetutil  826  0  0  0  826
kscience  293  0  0  0  293
ktools  6858  0  0  0  6858
Σ  13691  0  0  0  13691

Series n1

acmennrp  0  0  1666  0  1666
amanda  1  0  1685  3  1689
amd  1  0  851  0  852
apache  39  0  3371  0  3410
apassl  22  0  4740  1  4764
archie  0  0  39  0  39
argus  0  0  574  0  575
autofs  5  0  96  0  101
bind  2  0  1685  3  1691
bind8  13  0  1248  5  1266
bing  0  0  52  0  52
bulkmail  0  0  27  0  27
cipe  26  0  104  0  130
cnews  0  0  973  30  1004
deliver  0  0  82  1  83
dhcp  5  0  523  0  528
diald  5  0  324  0  329
dochost  8  0  9  1  19

Package  /  /usr/X11  /usr  /var  Σ
elm  0  0  584  0  584
emil  0  0  342  0  342
firewall  20  0  228  0  249
fortify  0  0  3458  0  3458
ftpdir  0  0  883  0  883
gettyps  0  0  84  0  84
gup  0  0  38  5  44
htdig  2635  0  1381  0  4017
hylafax  3  0  4375  410  4789
i4l  439  19  887  4  1351
i4lfirm  0  0  644  0  644
inetcfg  0  0  38  0  38
inn  0  0  1382  95  1477
ipfwadm  25  0  10  0  35
iptraf  0  0  253  0  253
ipxrip  0  0  32  0  32
ircd  0  0  609  0  610
ircii  0  0  508  0  508
irciihlp  0  0  357  0  357
ldap  2040  0  1850  0  3890
lprold  2  0  238  1  242
lynx  0  0  1953  0  1953
mailx  0  0  126  0  126
marsnwe  32  0  597  0  630
mdomo  10  0  691  5  706
metamail  0  0  376  0  376
mgetty  20  0  1035  3  1058
mh  0  0  4093  0  4093
minicom  1  0  316  1  318
mininews  0  0  36  0  36
mirror  0  0  154  0  154
mrtg  0  0  262  0  262
mutt  0  0  725  0  725
ncftp  0  0  237  0  237
ncpfs  142  0  961  0  1104
netacct  2  0  93  0  96
netatalk  6  0  658  1  666
netcat  0  0  121  0  121
netdiag  198  0  0  0  198
nn  0  0  666  0  666
nn spool  0  0  656  0  656
nntpd  0  0  150  0  151
pgp  0  0  1849  0  1849
pine  0  0  1641  0  1641
pine4  0  0  2628  0  2628
plp  3  0  1800  0  1804
pop  0  0  2485  0  2485
ppp  1  0  130  0  131

Package  /  /usr/X11  /usr  /var  Σ
procmail  0  0  180  0  180
proftpd  1  0  343  0  345
radiusc  24  0  555  0  579
radiusd  9  0  55  2  67
rdist  0  0  108  0  108
rsync  0  0  231  0  231
rzsz  0  0  84  0  84
samba  2  0  2877  0  2880
sendfax  3  0  129  3  136
sendmail  119  0  1527  4  1651
seyon  0  186  2  0  188
sf  15  0  811  0  826
slip  5  0  195  0  200
slrn  0  0  189  18  208
smail  11  0  580  5  596
smtpd  0  0  173  4  178
snmp  2  0  425  0  428
squid  40  0  578  2  620
ssh  2  0  816  0  818
ssleay  0  0  2282  0  2282
strn  0  0  797  0  797
suck  0  0  143  0  143
susefax  0  2  445  0  447
suseppp  26  0  44  1  72
tcpdump  0  0  272  0  272
tin  0  0  406  0  406
tleds  0  10  16  0  26
tripwire  0  0  293  125  419
trn  0  0  315  0  315
trn spl  0  0  293  0  293
url get  0  0  53  0  53
uucp  0  0  1084  4  1089
uucpcfg  0  0  0  17  17
vacation  0  0  35  0  35
vbox  2  16  389  1  409
vbox2b4  1  0  416  0  417
w3mir  0  0  1071  1  1072
webalize  9  0  175  0  184
weblint  0  0  126  0  126
wget  3  0  326  0  329
whois  0  0  21  0  21
wuftpd  1  0  154  0  156
wwwoffle  20  0  423  73  518
xibod  15  1008  3  0  1026
xinetd  5  0  164  0  170
xntp  2  0  2499  0  2501
xrn  0  225  0  0  225
yaps  8  0  229  0  238

Package  /  /usr/X11  /usr  /var  Σ
ypclient  1  0  113  1  116
ypserv  1  0  249  15  266
Σ  6033  1466 83363  845  91745

Series n2

faxprint  0  0  1  1  3
Σ  0  0  1  1  3

Series pay1

abuse  0  0  6080  0  6080
ac3d dem  0  1721  2495  0  4217
acroread  0  8366  25  0  8391
adabas  1  0 93800  0  93802
adadoc  0  0 14209  0  14209
adadok  0  0 22535  0  22535
adbsmydb  0  0 41458  0  41458
antivir  0  0  1380  0  1380
arcad  132704  0 190294  0  322999
arkeia  0  0 22532  0  22533
bartels  53240  0  1  0  53241
blender  0  7208  243  0  7451
c16 demo  0  0  5717  0  5717
cactus  0  0 21676  0  21676
cint  0  0  8845  0  8845
crisp  0  1690  8559  0  10249
cysched  1633  0  880  0  2514
eloquen  9078  0  1730  5121  15929
executor  0  0  7242  0  7242
flagship  2  0 15312  0  15314
flagstry  2  0 14621  0  14623
graffy  0  0  5868  0  5868
heitml  0  0  3961  0  3961
insure  34510  0  2193  0  36703
ivs demo  0  0  548  0  548
ixware  0  0 17737  0  17737
mapledmo  0  134  4512  0  4646
mars e  0  0  252  0  252
mhitdemo  0  80 11323  2  11406
mind  0  0 15569  0  15569
netscape  18785  3  1  0  18790
not demo  0  0  3669  0  3669
rk512  223  0  257  0  480
simcity  0  0  4664  0  4664
snavig  14442  0  20  0  14463
sniff  25131  0  43  0  25174
solid  5663  0  3129  0  8793
terroid  0  0  1971  0  1971
uc  0  0  1394  0  1394
vdx  0  2580  0  0  2580

Package  /  /usr/X11  /usr  /var  Σ
vishop  1  0 44626  0  44627
vse  0  0  6018  0  6018
wingz  0  1  3647  0  3649
xcept4  0  0  1858  0  1858
xess3  0  0  4003  0  4003
xgal  0  198  165  0  364
xv  0  855  1069  0  1924
yardsql  0  0 20648  0  20648
Σ  295415  22836 638779  5123  962164

Series sgm1

docbk30  0  0  5337  2  5340
gf  0  0  463  0  463
html dtd  0  0  3221  0  3221
hyperref  0  0  588  0  588
iso ent  0  0  93  3  97
jade dsl  0  0  1856  0  1856
jadetex  0  0  1294  0  1294
linuxdoc  0  0  753  0  753
sgmltool  0  0  821  0  821
sgrep  6  67  98  0  171
sp  0  0  4075  0  4075
Σ  6  67 18599  5  18679

Series snd1

aumix  0  0  53  0  53
dap  0  0  706  0  706
gmod  0  205  188  0  394
jazz  0  345  98  0  444
mxv  0  2180  57  0  2237
rosegard  0  8  1163  0  1171
rplay  1  0  364  0  366
sapphire  0  0  1407  0  1407
slab  0  6966  214  0  7181
snd au  0  0  1513  0  1513
snd mod  0  0  827  0  827
snd wav  0  0  2692  0  2692
sox  0  0  333  0  333
speakf  0  0  307  0  307
tcd  0  0  38  0  38
tclmidi  0  0  1285  0  1285
tracker  0  0  477  0  477
wavplay  0  1319  71  0  1391
workbone  0  0  36  0  36
workman  0  0  307  0  307
xmcd  0  2151  76  0  2228
xmitsumi  0  910  2  0  912
xmix  0  14  0  0  14
xmp  0  0  182  0  182

Package  /  /usr/X11  /usr  /var  Σ
xqmixer  0  44  2  0  46
Σ  1  14142 12398  0  26547

Series spl1

dicts  0  0 10555  0  10555
iamerica  0  0  5790  0  5790
ibritish  0  0  5790  0  5790
idanish  0  0  1650  0  1650
idutch  0  0  3121  0  3121
ifrench  0  0  1503  0  1503
iitalian  0  0  1654  0  1654
inorsk  0  0  1091  0  1091
ispanish  0  0  2813  0  2813
ispell  0  0  211  0  211
iswedish  0  0  788  0  788
words  0  0  812  0  812
Σ  0  0 35778  0  35778

Series spl2

igerman  0  0  3593  0  3593
Σ  0  0  3593  0  3593

Series tcl1

blt  0  0  822  0  822
convert  0  158  5  0  164
dejagnu  0  0  424  0  424
eddi  0  270  29  0  300
expect  0  0  698  0  698
filerun  0  357  1  0  358
itcl  0  0 25301  0  25301
jstools  0  32  2548  0  2580
listexec  0  15  0  0  15
omnimoni  0  0  238  0  238
samt  1  17  212  0  232
scotty  0  1  7614  0  7615
spectcl  0  1468  38  0  1506
tcl  0  0  1338  0  1338
tcl new  0  0  1991  0  1991
tcl old  0  0  951  0  951
tclx  0  0  2667  0  2667
tix  0  5379  1510  0  6889
tix new  0  2789  1897  0  4687
tixlpq  0  12  3  0  16
tk  0  2842  286  0  3129
tk new  0  2340  745  0  3085
tk old  0  1196  1303  0  2499
tkdesk  0  2228  33  0  2262
tkfont  0  12  28  0  41
tkinfo  0  131  10  0  142
tkinspec  0  145  0  0  145

Package  /  /usr/X11  /usr  /var  Σ
tkirc  0  302  86  0  389
tkmail  0  863  19  0  883
tkman  0  219  41  0  261
tkman8  0  290  22  0  313
tkppp  0  23  9  0  32
tkrat  0  1976  95  0  2071
tkstep  0  1587  160  0  1747
tkworld  0  622  32  0  654
tkxcd  0  21  25  0  46
tuba  0  0 14075  0  14075
vtcl  0  0  419  0  419
xcdroast  0  874  86  0  960
xdmcfg  0  4  92  0  97
xf  0  243  6233  0  6476
xmclock  0  69  0  0  69
zircon  0  1  489  0  490
Σ  1  26486 72575  0  99077

Series tex1

bibview  0  219  0  0  219
chess  0  0  958  0  958
colortbl  0  0  90  0  90
detex  0  0  20  0  20
dvi2tty  0  0  46  0  46
dviutils  0  0  255  0  255
l2h  2  0  1506  223  1732
l2hhtml  0  0  635  0  635
lyx  0  1248  2928  0  4176
musictex  0  0  4231  0  4231
psutils  0  0  191  0  191
tetex  1  0 40025  3  40030
texcad  0  82  29  0  112
tktex  0  36  117  0  154
transfig  0  221  94  0  315
ts  0  406  37  0  444
xfig  0  647  1222  0  1870
xtem  0  2754  154  0  2908
Σ  3  5613 52538  226  58386

Series x1

sax  0  559  10  1663  2233
xaw3d  0  965  8  0  974
xaw3dd  0  299  0  0  300
xaw95  0  984  10  0  994
xaw95d  0  616  1  0  617
xdevel  0  5416  0  0  5416
xdmbgrd  0  12  0  0  12
xdmsc  3  0  5  0  9
xdoc  0  376  0  0  376

Package  /  /usr/X11  /usr  /var  Σ
xextra  0  5214  0  0  5214
xf86  0  8906  5  8  8920
xfnt100  0  2089  0  0  2089
xfntbig  0  2440  0  0  2440
xfntcyr  0  371  0  0  371
xfntscl  0  1599  0  0  1599
xfsetup  0  1226  0  0  1226
xlkit  0  14035  0  0  14035
xman  0  1232  0  0  1232
xpm  0  174  0  0  174
xshared  0  1921  0  0  1921
xxprt  0  2895  0  0  2895
Σ  3  51329  39  1671  53047

Series x2

oldlibs5  0  0  1150  0  1150
oldlibs6  0  0  2817  0  2817
Σ  0  0  3967  0  3967

Series x3

modegen  0  21  1  0  22
Σ  0  21  1  0  22

Series x3d1

dev3dfx  4  0  0  0  4
glide v2  0  0  720  0  720
glide vg  0  0  628  0  628
glide vr  0  0  700  0  700
glidesdk  0  0  7428  0  7428
mesa  0  0 11003  0  11003
mesa3dfx  0  0 11086  0  11086
rushlibs  0  8  1  0  9
rushsvga  0  2641  1  0  2642
Σ  4  2649 31567  0  34220

Series xap1

addrbo  0  116  330  0  446
amaya  0  0  3893  0  3893
arena  0  1202  0  0  1202
argoups  333  2940  3607  0  6881
arrow  0  9592  4  0  9596
asedit  0  1578  0  0  1578
auis ful  0  0 38371  0  38371
auis ml  0  0  5018  0  5018
auis src  0  0  1786  0  1786
auis wp  0  0  5674  0  5674
axe  0  203  28  0  232
axe3d  0  204  28  0  232
babaya  0  7  0  0  7
bitchx  0  0  1475  0  1475

Package  /  /usr/X11  /usr  /var  Σ
calctool  0  57  0  0  57
cellular  0  0  667  0  667
conv  0  6  18  0  25
coolmail  0  33  1  0  35
elem  0  17  19  0  36
fader  0  6  5  0  11
fct  0  0  24  912  936
felix  0  0  1060  0  1060
gnuplot  0  0  3517  0  3517
gvim  5  390  1  0  398
hfsutils  0  0  5066  0  5066
knews  0  406  0  0  406
kradio  30  0  1  0  31
maxwell  0  0  4452  0  4452
mfm  0  1250  143  0  1394
mfsm  0  1013  7  0  1020
mgdiff  0  1158  0  0  1158
mupad  0  0 31775  0  31775
nedit  0  2212  589  0  2801
offix  0  402  78  0  481
pan  0  1028  3  0  1031
papyrus  0  2485  25  0  2510
pcb  0  695  0  0  695
pixedit  0  2054  2  0  2057
plan  0  2154  257  0  2412
plotmtv  0  1951  1008  0  2959
procmtr  0  25  33  0  58
qps  0  1434  30  0  1464
rasmol  0  206  1292  0  1498
roottail  0  10  0  0  10
rxvt  0  168  175  0  344
scilab  0  51 25559  0  25610
snns  0  0  5532  0  5532
snnsdoc  0  0  4134  0  4134
spice  0  0  1848  0  1848
spicecad  0  0  6868  0  6868
susecurs  0  4  0  0  4
tn5250  0  0  363  0  363
tosh ut  0  0  115  0  115
vidtext  0  136  189  56  381
vnc  0  1372  1142  0  2515
x3270  0  455  282  0  738
xarchie  0  281  0  0  281
xautolck  0  18  0  0  18
xawtv  0  111  37  0  148
xbanner  0  76  163  0  240
xbarcode  0  78  2  0  81
xbench  0  98  4  0  103

Package  /  /usr/X11  /usr  /var  Σ
xcb  0  16  0  0  16
xcnf  0  34  2  0  37
xcolors  0  9  0  0  9
xcoral  0  519  6225  0  6744
xdf  0  16  4  0  20
xdtm  0  143  110  0  254
xephem  0  4596  52  0  4648
xfileman  0  99  60  0  160
xflame  0  904  1  0  905
xfm  0  583  0  0  583
xfmail  0  810  64  0  874
xformat  0  26  24  0  50
xftp  0  1448  10  0  1459
xgopher  0  260  0  0  260
xgrok  0  2  1778  0  1780
xgroups  0  60  18  0  79
xhtalk  0  58  6  0  64
xinfo  0  38  5  0  44
xirc  0  639  60  0  699
xkeycaps  0  626  0  0  626
xled  0  25  1  0  27
xless  0  37  0  0  37
xlock  0  1781  0  0  1781
xlogin  0  4  0  0  4
xlogmas  0  41  8  0  50
xlpq  0  24  23  0  47
xmail  0  135  0  0  135
xmftp  0  2446  12  0  2458
xnetload  0  13  25  0  39
xosview  0  69  22  0  91
xplot  0  0  374  0  374
xplz  0  11076  45  0  11122
xprompt  0  16  0  0  16
xrolodex  0  1170  0  0  1170
xrpm  0  0  260  1  261
xscrns  0  10606  32  0  10639
xselect  0  9  0  0  9
xsession  0  48  38  0  87
xsky  0  25378  0  0  25378
xspline  0  11  0  0  11
xspread  0  176  433  0  610
xsysinfo  0  15  9  0  24
xtail  0  1034  1  0  1036
xtalk  0  1176  660  0  1836
xtar  0  1140  7  0  1148
xtexit  0  10  3  0  13
xvhelp  0  48  46  0  94
xvi  0  0  477  0  477

Package  /  /usr/X11  /usr  /var  Σ
xvnews  0  150  0  0  150
xvtdl  0  222  0  0  222
xwpick  0  41  0  0  41
xzap  0  39  0  0  39
xzet  0  50  37  0  87
Σ  368  105558 167632  969  274558

Series xdev1

ddd  0  302 10338  0  10640
dlgedit  0  837  23  0  860
ezwgl  0  11518  186  0  11704
iv  0  0  6970  0  6970
jx  0  126647  4464  0  131112
qtdevel  0  2213  3610  0  5823
qtext  0  597  0  0  597
qtlib  0  1199  43  0  1242
vtk  0  76081  912  0  76993
vtkdata  0  0 128476  0  128476
wxwdemo  0  0  1472  0  1472
wxwdocht  0  0  1994  0  1994
wxwdocps  0  0  4533  0  4533
wxwin  0  0  4614  0  4614
wxwstat  0  0  3335  0  3335
wxxt  0  3490  3036  0  6527
xforms  0  1972  2027  0  4000
xtpanel  0  342  15  0  357
xwb  0  79  25  0  104
xwpe  0  0  570  0  570
xxgdb  0  85  0  0  85
ygl  0  891  51  0  943
Σ  0  226253 176694  0  402951

Series xsrv1

xvga16  0  2001  0  0  2001
Σ  0  2001  0  0  2001

Series xsrv2

x8514  0  1759  0  0  1759
xagx  0  1949  0  0  1949
xi128  0  2213  0  0  2213
xmach32  0  1910  0  0  1910
xmach64  0  2042  0  0  2042
xmach8  0  1772  0  0  1772
xmono  0  2075  0  0  2075
xp9k  0  1973  0  0  1973
xs3  0  2472  0  0  2472
xs3v  0  2191  0  0  2191
xsvga  0  3154  0  0  3154
xw32  0  1812  0  0  1812

Package  /  /usr/X11  /usr  /var  Σ
Σ  0  25322  0  0  25322

Series xsrv3

xcyrix  0  2446  1  5  2452
xglint  0  2285  1  5  2291
xp9x00  0  2464  1  5  2470
xrenditn  0  2476  1  5  2482
xsis  0  2483  1  5  2489
Σ  0  12154  5  25  12184

Series xsrv4

xneomag  0  2010  9  0  2020
Σ  0  2010  9  0  2020

Series xv1

ssbase  0  0  1404  0  1404
ssexmp  0  0  430  0  430
xtoolpl  0  30  0  0  30
xv32aout  0  0  3205  0  3205
xv32base  0  7  2225  1  2233
xv32dev  0  50  1160  0  1211
xv32exmp  0  0  1116  0  1116
xv32slib  0  0  1941  0  1941
Σ  0  87 11481  1  11570

Series xwm1

3dpixm  0  2425  0  0  2425
3dpixms  0  1195  0  0  1195
afterstp  0  595  157  0  753
amiwm  0  115  7  0  122
bowman  0  162  75  0  238
cdesim  0  87  119  0  207
ctwm  0  258  60  0  319
fvwm  0  1310  295  0  1605
fvwm1  0  390  234  0  624
fvwm95  0  775  171  0  947
icewm  0  833  32  0  866
icons  0  5233  2  0  5236
lxuser  0  0  0  1  1
mlvwm  0  118  44  0  163
olvwm  0  0  288  0  288
pixmaps  0  2561  1  0  2563
qvwm  0  124  58  0  183
susewm  7  496  290  2  796
wm2  0  65  7  0  72
wmaker  0  1564  211  0  1776
wmthemes  0  20728  6  0  20735
Σ  7  39034  2057  3  41114

Series xwm2

enlight  0  3285  0  0  3285

0.3. Disk usage of the series

Package  /  /usr/X11  /usr  /var  Σ
esound  0  0  299  0  299
fnlib  0  0  596  0  596
Σ  0  3285  895  0  4180

0.3 Disk usage of the series

Serie  /  /usr/X11  /usr  /var  Σ
a1  16245  193  68022  167  84646
a2  1820  0  531  0  2351
ap1  1382  6737  98372  1022  107531
ap2  0  0  2370  0  2370
beo1  0  734  20546  0  21282
d1  76  5461  374270  8  379816
d2  0  0  44  0  44
doc1  5  2442  315769 11939  330157
e1  14  60089  44821  2  104928
emu1  209  1822  20379 10629  33044
fun1  0  69719  35149  447  105340
gnm1  15019  0  69  0  15088
gnm2  11745  1  42  0  11789
gnm3  17174  0  251  0  17428
gra1  90  33049  397567  0  430729
kde1  31565  0  132  0  31698
kde2  6935  0  36  0  6972
kde3  22  0  0  0  22
kde4  13691  0  0  0  13691
n1  6033  1466  83363  845  91745
n2  0  0  1  1  3
pay1  295415  22836  638779  5123  962164
sgm1  6  67  18599  5  18679
snd1  1  14142  12398  0  26547
spl1  0  0  35778  0  35778
spl2  0  0  3593  0  3593
tcl1  1  26486  72575  0  99077
tex1  3  5613  52538  226  58386
x1  3  51329  39  1671  53047
x2  0  0  3967  0  3967
x3  0  21  1  0  22
x3d1  4  2649  31567  0  34220
xap1  368  105558  167632  969  274558
xdev1  0  226253  176694  0  402951
xsrv1  0  2001  0  0  2001
xsrv2  0  25322  0  0  25322
xsrv3  0  12154  5  25  12184
xsrv4  0  2010  9  0  2020

Serie  /  /usr/X11  /usr  /var  Σ
xv1  0  87  11481  1  11570
xwm1  7  39034  2057  3  41114
xwm2  0  3285  895  0  4180
Σ  417833k  720560k 2690341k 33083k 3862054k

0.4. Which series contains which package?

0.4 Which series contains which package?

Package  Series Page Package Series Page Package Series Page
3d chess  fun1 57 aumix  snd1 121 cdwrite  ap1  14
3dpixm  xwm1 167 autoconf  d1  30 cdwtools  ap1  15
3dpixms xwm1 167 autofs  n1  88 cellular  xap1 144
a2ps  ap1  12 autolog  ap1  13 cfengine  ap1  15
aaa base  a1  1  automake  d1  30 cgmdraw gra1  71
aaa dir  a1  1  axe  xap1 143 changes  doc1 45
abuse  pay1 107 axe3d  xap1 143 chess  tex1 133
ac3d dem pay1 107 babaya  xap1 144 chipmunk ap1  15
acmennrp  n1  86 bartels  pay1 108 cint  pay1 109
acroread  pay1 107 base  a1  1  cipe  n1  89
adabas  pay1 107 bash  a1  2  clisp  d1  31
adadoc  pay1 107 bash1  a1  2  cnews  n1  89
adadok  pay1 107 batalion  fun1 57 colortbl  tex1 133
adamem  emu1 53 battball  fun1 57 compat  a2  11
adbsmydb pay1 107 bc  ap1  14 compress  a1  3
addrbo  xap1 142 bdflush  a1  2  conv  xap1 144
afterstp  xwm1 167 beg  d1  30 convert  tcl1 126
alien  ap1  12 bibview  tex1 133 coolmail xap1 144
allman  doc1 44 bin86  d1  30 cpio  a1  3
amanda  n1  87 bind  n1  88 craft  fun1 58
amaya  xap1 142 bind8  n1  89 crisp  pay1 109
amd  n1  87 bing  n1  89 cron  a1  3
amiwm  xwm1 168 binutils  d1  30 crosfire  fun1 58
antipoli  fun1 57 bison  d1  31 ctags  ap1  15
antivir  pay1 108 bitchx  xap1 144 ctwm  xwm1 168
apache  n1  87 blender  pay1 108 cvs  d1  32
apassl  n1  88 blt  tcl1 126 cysched  pay1 110
apmd  ap1  13 books  doc1 44 daliclck  fun1 58
aps  ap1  13 booksfnt  doc1 45 dap  snd1 121
arc  ap1  13 bowman xwm1 168 db  a1  3
arcad  pay1 108 bsdgames fun1 58 ddd  xdev1 161
archie  n1  88 bttv  gra1  70 deco  ap1  15
arena  xap1 142 bttvgrab  gra1  71 dejagnu  tcl1 126
argoups  xap1 142 buffer  ap1  14 deliver  n1  90
argus  n1  88 bulkmail  n1  89 des  d1  32
arkeia  pay1 108 bwatch  beo1 29 detex  tex1 133
arrow  xap1 142 bwbasic  d1  31 dev3dfx  x3d1 141
asclock  fun1 57 bzip  ap1  14 devs  a1  3
asedit  xap1 142 c16 demo pay1 109 dhcp  n1  90
ash  ap1  13 cactus  pay1 109 diald  n1  90
astrolog  fun1 57 calctool  xap1 144 dialog  ap1  15
at  a1  1  cdb  doc1 45 dicts  spl1 124
auis ful  xap1 143 cdcons  ap1  14 diff  a1  3
auis ml  xap1 143 cdesim  xwm1 168 dlgedit  xdev1 161
auis src  xap1 143 cdparano  ap1  14 dns faq  doc1 45
auis wp  xap1 143 cdrecord  ap1  14 docbk30 sgm1 119

Package Series Page Package Series Page Package Series Page
dochost  n1  90 filmsbig  gra1  71 glide vg x3d1 141
dosemu  emu1 53 find  a1  4  glide vr x3d1 141
dosfsck  ap1  15 firewall  n1  90 glidesdk x3d1 141
duconv  ap1  16 fishtank  fun1 59 glimpse  ap1  16
dump  ap1  16 flagship  pay1 112 global  d1  34
dvi2tty  tex1 134 flagstry  pay1 113 gltt  gra1 72
dviutils  tex1 134 flex  d1  32 gmod  snd1 121
eazy  a1  3  flying  fun1 59 gmp  d1  34
ed  ap1  16 fnlib  xwm2 171 gnadmin gnm3 69
eddi  tcl1 127 fortify  n1  91 gnat  d1  34
edy  ap1  16 fortune  fun1 59 gnats  d1  34
efence  d1  32 fpk  d1  32 gncore  gnm2 69
ei  emu1 54 freeciv  fun1 59 gngames gnm3 69
elem  xap1 144 freetype  gra1  71 gngraph gnm3 70
elib  e1  50 frisk  fun1 59 gnguile gnm3 70
elispint  doc1 45 ftpdir  n1  91 gnlibs  gnm1 69
elispman doc1 45 fvwm  xwm1 168 gnmedia gnm3 70
elm n1 90 fvwm1 xwm1 169 gnobjc gnm3 70
eloquen  pay1 111 fvwm95 xwm1 169 gnuchess fun1 60
elvis  a1  3  g77  d1  32 gnugo  fun1 60
emacs w3 e1  51 gawk  a1  4  gnuhtml doc1 46
emiclock fun1 58 gawkdoc doc1 46 gnuplot  xap1 145
emil  n1  90 gcal  ap1  16 gobjc  d1  34
empire  fun1 58 gcc  d1  32 gpc  d1  35
enlight  xwm2 171 gccinfo  doc1 46 gpm  ap1  17
enscript  ap1  16 gcl  d1  33 gpp  d1  35
esound  xwm2 171 gd  gra1  72 gppinfo  doc1 46
executor  pay1 112 gdb  d1  33 gppshare  a1  4
expect  tcl1 127 gdbm  a1  4  graffy  pay1 113
ext2fs  a1  4  gdlib  d1  33 groff  a1  4
ezwgl  xdev1 161 gdtmon  ap1  16 gs both  ap1  17
f2c  d1  32 ge auc  e1  51 gs fonto  ap1  17
f2caout  d1  32 ge exec  e1  51 gs fonts  ap1  17
fader  xap1 144 ge info  doc1 46 gs lib  ap2  28
faxedit  gra1  71 ge lisp  e1  52 gs serv  ap1  17
faxprint  n2  106 ge site  e1  52 gs vga  ap1  17
fbm  gra1  71 gettext  d1  34 gs x11  ap1  17
fbm2fli  gra1  71 gettyps  n1  91 gshogi  fun1 60
fct  xap1 145 gf  sgm1 119 gsview  gra1 72
felix  xap1 145 gforth  d1  34 gtk  gra1 72
fhs  doc1 45 gif  gra1  72 guavac  d1  35
figlet  fun1 58 gimp  gra1  72 guile  d1  35
file  a1  4  gimpman gra1  72 gup  n1  91
filerun  tcl1 127 git  ap1  16 gv  gra1 72
fileutil  a1  4  glchess  fun1 60 gvim  xap1 145
films  gra1  71 glide v2 x3d1 141 gzip  a1  5

0.4. Which series contains which package?

Package  Series Page Package Series Page Package Series Page
heitml  pay1 113 ipfwadm  n1  92 kmulti  kde1 82
hello  ap1  18 iptraf  n1  92 knet  kde1 82
hextris  fun1 60 ipxrip  n1  92 knetutil  kde4 85
hfsutils  xap1 145 ircd  n1  92 knews  xap1 145
howto  doc1 46 ircii  n1  93 kradio  xap1 145
howtodvi doc1 46 irciihlp  n1  93 kscience kde4 85
howtohtm doc1 46 isapnp  ap1  19 ksupp  kde1 83
howtoinf  doc1 46 isdnpop  doc1 47 ktools  kde4 86
howtops  doc1 46 iso ent  sgm1 119 kutils  kde1 83
howtosgm doc1 46 ispanish  spl1 125 l2h  tex1 134
htdig  n1  91 ispell  spl1 125 l2hhtml  tex1 134
html2ps  gra1  73 iswedish spl1 126 lacheck  e1  52
html dtd  sgm1 119 itcl  tcl1 127 lam  beo1 29
hylafax  n1  91 iv  xdev1 161 lclint  d1  36
hyperref  sgm1 119 ivs demo pay1 113 ldap  n1  93
i4l  n1  92 ixware  pay1 114 ldp  doc1 47
i4ldoc  doc1 46 jade dsl sgm1 119 ldpman  doc1 47
i4lfirm  n1  92 jadetex  sgm1 119 ldpmande doc1 47
iamerica  spl1 124 java  d1  35 ldpmanfr doc1 48
ibcs2  emu1 54 javadoc  doc1 47 ldpmanit doc1 48
ibritish  spl1 125 javarunt  d1  35 ldso  a1  5
icewm  xwm1 169 jazz  snd1 121 leda  d1  36
icmake  d1  35 jdk115  d1  35 less  a1  5
icons  xwm1 169 jdk116  d1  36 lha  ap1  20
id utils  ap1  18 jed  ap1  19 libaout  a1  5
idanish  spl1 125 joe  ap1  19 libc  d1  37
idled  ap1  18 jove  ap1  19 libcinfo  doc1 48
idutch  spl1 125 joystick  d2  44 libd  d1  37
ifrench  spl1 125 jpeg  gra1 73 libfbm  gra1 74
igerman  spl2 126 jresep  d1  36 libgif  gra1 74
iitalian  spl1 125 jstools  tcl1 127 libgpp  d1  37
imagemag gra1  73 jx  xdev1 161 libgrout  gra1 74
imaze  fun1 60 kadm  kde4 84 libjpeg  gra1 74
imfilm  gra1  73 kadmin  kde1 81 liblug  gra1 74
imlib  gra1  73 kaffe  d1  36 libmpeg  gra1 74
impict  gra1  73 kbase  kde2 83 libnetpb  gra1 74
indent  d1  35 kbd  a1  5  libobjc  d1  38
ines  emu1 54 kdevel  kde4 84 libpng  gra1 74
inetcfg  n1  92 kdir  kde3 83 librle  gra1 75
inf2htm  doc1 47 kernmod  a1  5  libtiff  gra1 75
ingres  ap1  18 kgames  kde1 82 libz  a1  5
ingrtool  ap1  18 kgoemon kde1 82 lilo  a1  6
inlinem  ap1  18 kgraph  kde1 82 lincity  fun1 60
inn  n1  92 kicons  kde1 82 linclude  d1  38
inorsk  spl1 125 klibs  kde1 82 linux  d1  38
insure  pay1 113 kmisc  kde4 85 linuxdoc sgm1 120

Package  Series Page Package Series Page Package  Series Page
listexec  tcl1 127 mkdosfs  ap1  21 offix  xap1 147
loadlin  a1  6  mkisofs  ap1  21 oldlibs5  x2  140
localedb  a1  6  mktemp  ap1  21 oldlibs6  x2  140
logsurf  ap1  20 mlcreat  gra1  75 oldshlbs  a2  11
lprold  n1  93 mlvwm  xwm1 170 olvwm  xwm1 170
lsof  ap1  20 mmv  ap1  21 omnimoni tcl1 127
lthreads  d1  40 mocka  d1  40 oneko  fun1 61
lx hack  d1  40 modegen  x3  140 onyx  ap1  22
lx suse  d1  40 modules  a1  6  oonsoo  fun1 61
lxuser  xwm1 169 moontool fun1 61 p2c  d1  41
lynx  n1  93 mpegplay gra1  75 pacman  fun1 62
lyx  tex1 134 mpegutil  ap1  22 pan  xap1 147
maelstr  fun1 60 mrtg  n1  94 papyrus  xap1 147
mailx  n1  93 mt st  ap1  22 patch  d1  41
make  d1  40 mtc  d1  40 pcb  xap1 147
makewhat ap1  20 mtools  ap1  22 pcmcia  a1  7
man  a1  6  mupad  xap1 146 pdksh  ap1  22
man9  doc1 48 muscan  gra1  75 perl  a1  8
manix  fun1 61 musictex tex1 135 perlref  doc1 48
manyfaqs doc1 48 mutt  n1  94 pg datab  ap1  23
mapledmo pay1 114 mxv  snd1 121 pg ifa  ap1  23
mars e  pay1 114 nc  ap1  22 pg iface  ap1  23
marsnwe  n1  93 ncftp  n1  95 pgp  n1  96
maxwell  xap1 146 ncpfs  n1  95 photocd  gra1  76
mbedit  ap1  20 ncurses  a1  7  pico  ap1  23
mc  ap1  20 nedit  xap1 146 pictures  gra1  76
mdomo  n1  93 net3d  fun1 61 pine  n1  96
mdutils  ap1  21 net tool  a1  7  pine4  n1  96
meltflip  fun1 61 netacct  n1  95 pipeman  fun1 62
mesa  x3d1 141 netatalk  n1  95 pixedit  xap1 147
mesa3dfx x3d1 141 netcat  n1  95 pixmaps xwm1 170
metamail  n1  94 netcfg  a1  7  pixmp  gra1  76
mfm  xap1 146 netdiag  n1  95 plan  xap1 147
mfsm  xap1 146 nethack  fun1 61 playone  fun1 62
mgdiff  xap1 146 netpbm  gra1  75 plmagick gra1  76
mgetty  n1  94 netscape pay1 115 plotmtv  xap1 147
mgp  gra1  75 nkita  a1  7  plp  n1  96
mh  n1  94 nkitb  a1  7  pmake  d1  41
mhitdemo pay1 114 nn  n1  95 pop  n1  96
mico  d1  40 nn spool  n1  95 postgres  ap1  23
mind  pay1 115 nntpd  n1  95 povray  gra1  77
minicom  n1  94 not demo pay1 115 ppa  ap1  24
minigcc  d1  40 nvi  a1  7  ppp  n1  97
mininews  n1  94 obst  ap1  22 ppp nt  doc1 48
mirrma  fun1 61 octave  ap1  22 prcs  d1  41
mirror  n1  94 odshtml  doc1 48 procmail  n1  97

Package Series Page Package Series Page Package Series Page
procmtr  xap1 148 scilab  xap1 148 splitvt  ap1  25
proftpd  n1  97 scotty  tcl1 128 squid  n1  100
ps  a1  8  screen  ap1  25 ssbase  xv1 166
psgml  e1  52 scsiinfo  ap1  25 ssexmp  xv1 166
pstoedit  gra1  77 sdb cgi  doc1 49 ssh  n1  100
psutils  tex1 135 sdb de  doc1 49 ssleay  n1  100
puzzle  fun1 62 sdb en  doc1 49 stella  emu1 55
pvm  beo1 29 seejpeg  ap1  25 stonx  emu1 55
pyth doc doc1 49 selfhtml doc1 49 stonxsft emu1 55
python  d1  41 sendfax  n1  98 strace  d1  42
qps  xap1 148 sendmail n1  98 strn  n1  101
qtdevel  xdev1 161 seyon  n1  99 su1  ap1  26
qtext  xdev1 162 sf  n1  99 suck  n1  101
qtlib  xdev1 162 sgmltool sgm1 120 sudo  ap1  26
quota  ap1  24 sgrep  sgm1 120 surf  gra1  78
qvwm  xwm1 170 sh utils  a1  9  suse lx  doc1 49
radiusc  n1  97 shadow  a1  9  suse lxe doc1 49
radiusd  n1  97 sharutil  ap1  25 susebuch doc1 49
rasmol  xap1 148 shlibs  a1  9  susecurs xap1 150
rcs  d1  41 shlibs6  a1  9  susefax  n1  101
rdist  n1  97 simcity  pay1 115 susehilf  doc1 49
recode  ap1  24 slab  snd1 122 susepak doc1 50
rfc  doc1 49 slang  d1  41 susepake doc1 50
rk512  pay1 115 slip  n1  99 suseppp  n1  102
rman  ap1  24 slrn  n1  99 susewm xwm1 170
rockdiam fun1 62 smail  n1  100 svgaaout  d1  42
roottail  xap1 148 smaltalk  d1  42 svgalib  d1  43
rosegard  snd1 121 smeiffel  d1  42 svgatext  ap1  26
rplay  snd1 121 smii  gra1 78 swi pl  d1  43
rpm  a1  8  smtpd  n1  100 syslogd  a1  9
rpm2html ap1  24 snavig  pay1 115 sysvinit  a1  9
rpmfind  ap1  24 snd au  snd1 122 taper  ap1  26
rsync  n1  97 snd mod snd1 122 tcd  snd1 122
rushlibs  x3d1 141 snd wav snd1 122 tcl  tcl1 128
rushsvga x3d1 142 sniff  pay1 116 tcl new  tcl1 128
rxvt  xap1 148 snmp  n1  100 tcl old  tcl1 129
rzsz  n1  98 snns  xap1 148 tclmidi  snd1 123
samba  n1  98 snnsdoc xap1 149 tclx  tcl1 129
samt  tcl1 128 solid  pay1 116 tcpdump  n1  102
sane  gra1  78 sox  snd1 122 tcsh  ap1  26
sapphire  snd1 122 sp  sgm1 120 teapot  ap1  26
sastroid  fun1 62 space  fun1 62 termcap  a1  10
sax  x1  137 speakf  snd1 122 terminfo  a1  10
sc  ap1  25 spectcl  tcl1 128 terroid  pay1 116
sced  gra1  78 spice  xap1 150 tetex  tex1 135
scheme  d1  41 spicecad xap1 150 texcad  tex1 136

Package Series Page Package  Series Page Package  Series Page
texinfo  a1  10 units  ap1  27 wxwdemo xdev1 163
textutil  a1  10 unzip  ap1  27 wxwdocht xdev1 163
tgif  gra1 78 url get  n1  103 wxwdocps xdev1 163
thrust  fun1 62 util  a1  10 wxwin  xdev1 163
tiff  gra1 78 uucp  n1  103 wxwstat  xdev1 163
timezone  a1  10 uucpcfg  n1  103 wxxt  xdev1 163
tin  n1  102 uucpdoc  doc1 50 x3270  xap1 151
tix  tcl1 129 vacation  n1  103 x3d  fun1 63
tix new  tcl1 129 vbox  n1  103 x8514  xsrv2 164
tixlpq  tcl1 129 vbox2b4  n1  104 xabacus  fun1 63
tk  tcl1 129 vdx  pay1 117 xagx  xsrv2 164
tk new  tcl1 130 vgacard  fun1 63 xancur  fun1 63
tk old  tcl1 130 vgb  emu1 55 xanim  gra1 78
tkdesk  tcl1 130 vice  emu1 55 xaos  fun1 63
tkdraw  gra1 78 vidtext  xap1 150 xarchie  xap1 151
tkfont  tcl1 130 vim  a1  11 xautolck  xap1 151
tkinfo  tcl1 131 vishop  pay1 117 xaw3d  x1  137
tkinspec  tcl1 131 vnc  xap1 151 xaw3dd  x1  137
tkirc  tcl1 131 vse  pay1 117 xaw95  x1  138
tkmail  tcl1 131 vtcl  tcl1 132 xaw95d  x1  138
tkman  tcl1 131 vtk  xdev1 162 xawtv  xap1 151
tkman8  tcl1 131 vtkdata  xdev1 162 xball  fun1 63
tkppp  tcl1 131 w3mir  n1  104 xbanner  xap1 151
tkrat  tcl1 131 wavplay  snd1 123 xbarcode  xap1 152
tkstep  tcl1 132 wdiff  ap1  27 xbench  xap1 152
tktex  tex1 136 webalize  n1  104 xbill  fun1 63
tkworld  tcl1 132 weblint  n1  104 xbl  fun1 64
tkxcd  tcl1 132 wget  n1  104 xblast  fun1 64
tleds  n1  102 whois  n1  105 xbmbrosr gra1 79
tn5250  xap1 150 wie dvi  doc1 50 xboard  fun1 64
toolpack  d1  43 wie geht  doc1 50 xboing  fun1 64
toppp  doc1 50 wie html  doc1 50 xbomb  fun1 64
tosh ut  xap1 150 wie info  doc1 50 xbombs  fun1 64
tracker  snd1 123 wie pdf  doc1 50 xcb  xap1 152
transfig  tex1 136 wie ps  doc1 50 xcdroast  tcl1 132
tripwire  n1  102 wie sgml  doc1 50 xcept4  pay1 117
trn  n1  103 wine  emu1 55 xclass  d1  43
trn spl  n1  103 wingz  pay1 117 xcnf  xap1 152
ts  tex1 136 wm2  xwm1 171 xcolors  xap1 152
tuba  tcl1 132 wmaker  xwm1 171 xcoral  xap1 152
tvision  d1  43 wmthemes xwm1 171 xcyrix  xsrv3 165
tya  d1  43 words  spl1 126 xdelta  ap1  27
uae  emu1 55 workbone snd1 123 xdemine  fun1 64
uc  pay1 116 workman  snd1 123 xdevel  x1  138
umsprogs a1  10 wuftpd  n1  105 xdf  xap1 153
unarj  ap1  27 wwwoffle  n1  105 xdmbgrd  x1  138

Package Series Page Package Series Page Package Series Page
xdmcfg  tcl1 132 xled  xap1 155 xprompt xap1 157
xdmsc  x1  139 xless  xap1 155 xpuzzles fun1 67
xdoc  x1  139 xli  gra1 79 xpvm  beo1 29
xdtm  xap1 153 xlife  fun1 65 xqmixer snd1 124
xe exec  e1  52 xlkit  x1  140 xray  gra1 81
xe info  doc1 50 xlock  xap1 155 xrayps  gra1 81
xe lisp  e1  53 xlogin  xap1 155 xrenditn xsrv3 166
xearth  fun1 64 xlogmas  xap1 155 xrn  n1  106
xengine  fun1 65 xlpq  xap1 156 xroach  fun1 67
xephem  xap1 153 xlyap  fun1 65 xrobots  fun1 67
xess3  pay1 118 xmach32 xsrv2 165 xrolodex xap1 157
xextra  x1  139 xmach64 xsrv2 165 xrpm  xap1 157
xf  tcl1 133 xmach8  xsrv2 165 xs3  xsrv2 165
xf86  x1  139 xmahjong fun1 65 xs3v  xsrv2 165
xf86html doc1 50 xmail  xap1 156 xscheme  d1  44
xfig  tex1 136 xman  x1  140 xscrns  xap1 157
xfileman xap1 153 xmcd  snd1 123 xselect  xap1 157
xflame  xap1 153 xmclock  tcl1 133 xsession xap1 158
xfm  xap1 153 xmemory fun1 66 xshared  x1  140
xfmail  xap1 153 xmftp  xap1 156 xshogi  fun1 67
xfnt100  x1  139 xmgr  gra1 80 xsis  xsrv3 166
xfntbig  x1  139 xmine  fun1 66 xskat  fun1 67
xfntcyr  x1  140 xmines  fun1 66 xsky  xap1 158
xfntscl  x1  140 xmitsumi snd1 123 xsnow  fun1 67
xformat  xap1 154 xmix  snd1 124 xsok  fun1 67
xforms  xdev1 163 xmono  xsrv2 165 xsol  fun1 67
xfract  fun1 65 xmorph  gra1 80 xspline  xap1 158
xfsetup  x1  140 xmountns fun1 66 xspread  xap1 158
xftp  xap1 154 xmp  snd1 124 xspringi fun1 68
xgal  pay1 118 xmris  fun1 66 xsvga  xsrv2 165
xgammon fun1 65 xmrm  gra1 80 xsysinfo xap1 158
xgas  fun1 65 xneomag xsrv4 166 xtacy  fun1 68
xglint  xsrv3 166 xnetload  xap1 156 xtail  xap1 158
xgopher  xap1 154 xntp  n1  106 xtalk  xap1 158
xgrabsc  gra1 79 xosview  xap1 156 xtar  xap1 159
xgrasp  gra1 79 xp9k  xsrv2 165 xtartan  fun1 68
xgrok  xap1 154 xp9x00  xsrv3 166 xteddy  fun1 68
xgroups  xap1 154 xpaint  gra1 80 xtem  tex1 137
xhtalk  xap1 154 xpat2  fun1 66 xtetris  fun1 68
xi128  xsrv2 165 xpdf  gra1 80 xtexit  xap1 159
xibod  n1  105 xphoon  fun1 66 xthing  fun1 68
xinetd  n1  106 xpilot  fun1 66 xtoolpl  xv1 167
xinfo  xap1 155 xpinguin fun1 67 xtpanel xdev1 164
xirc  xap1 155 xplot  xap1 157 xtron  fun1 68
xjewel  fun1 65 xplz  xap1 157 xv  pay1 118
xkeycaps xap1 155 xpm  x1  140 xv32aout xv1 167

Package  Series Page Package Series Page Package Series Page
xv32base  xv1 167
xv32dev  xv1 167
xv32exmp xv1 167
xv32slib  xv1 167
xvga16  xsrv1 164
xvhelp  xap1 159
xvi  xap1 159
xvier  fun1 68
xvnews  xap1 159
xvtdl  xap1 159
xw32  xsrv2 165
xwarp  gra1 81
xwb  xdev1 164
xwpe  xdev1 164
xwpick  xap1 160
xxgdb  xdev1 164
xxprt x1 140
xzap  xap1 160
xzet  xap1 160
xzx  emu1 56
yacc d1 44
yahtzee  fun1 69
yamm  ap1  27
yaps  n1  106
yardsql  pay1 118
yast a1 11
ygl  xdev1 164
ypclient  n1  106
ypserv  n1  106
ytree  ap1  27
z81  emu1 57
z88 ap1 28
zgv  gra1 81
zip  ap1  28
ziptool  ap1  28
zircon  tcl1 133
zoo  ap1  28
zsh  ap1  28